A Complete Bibliography of Publications in SIAM Journal on Imaging Sciences

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 October 2023 Version 1.33

Title word cross-reference

(BV, L^{1}) [AT11]. 0 [CLL15, CHL16]. 1 [LHW^{+}15, WLYU15]. 2 [AM16, CCBB14, DSYT10, GTO14, GBFA10, GBFA12, LdGKW19, LL22, PS11, SM18, TAF^{+}20, UC13, Wol09, YY13]. 3 [DSYT10, DGH11, EST20, GS13, HMS17, HZDZ23, NTDB19, TAR^{+}19, TPM20, WCN^{+}19, WLL^{+}21, YLLY19, YLLY20, ZC20]. 4 [LHB^{+}18]. ^{q} [HW13]. α [FSV10]. B_{z} [SSL23]. βL_{q} [HCGN23]. D [AM16]. D_{2} [RS20]. ℓ^{1} [SZW14]. ℓ_{0} [CJPT13, Nik13, SBFA15, SBFA16]. ℓ_{0,∞} [PG19]. ℓ_{1} [BK15, LY13, GCN21, SX12, YOGD08]. ℓ_{2} [CJPT13, GCN21]. ℓ_{2,1} [LLC14]. ℓ_{∞,1} [CWR19]. Γ [GBFA12, Naj17, GBFA10]. H^{1} [MB16]. ∞ [ETT15]. K [DPC13, MÁS^{+}22]. L1 [GO09]. L^{1} [MGKR15, LMM17, CJK10]. L^{1}/L^{2} [HL13]. L^{1}TV [DHN09]. L_{1} [HCGN23, WLJ22]. l_{1−2} [MLH17]. L_{1}/L_{2} [WTNL21]. L_{2} [FKLS12]. l_{p} [GH23]. N [HCCS20, BHM12, ZCZL22]. p [ETT15, GWY09, HFE19]. P_{2} [FRV18]. π [HF12]. SE(2) [BDMS15]. SE(3) [ARF16]. T [LLYG14, LS19, TM18]. T V^{ϕ} [HVW15]. φ [KSS19].

-Bar [AM16]. -Based [PG19]. -Convergence [GBFA10, GBFA12, Naj17]. -D [GBFA12, GBFA10]. -Dimensional [AM16, DGH11, LdGKW19, WCN^{+}19, Wol09, ZCZL22, HCCS20]. -Distribution [DPC13]. -gon [MÁS^{+}22]. -harmonic [GWY09]. -Kernels [FSV10]. -Laplacian [ETT15, HFE19]. -Line [HF12]. -Manifolds [YLLY19, YLLY20]. -Map [LLYG14]. -mean [MGKR15]. -Minimal [KSS19]. -Minimization [YOGD08]. -Models [HW13]. -norm [Nik13, WLJ22]. -Product [TM18]. -Regularization [MB16]. -regularized [GO09]. -Space [BHM12]. -Symmetric [RS20].

/TV [FKLS12]. /TV-Image [FKLS12]. /Underexposed [HJS13].

1PI [KK08].

A-Contrario [vGPR22]. Abdominal [AKLS17]. Abel [AAD^{+}08]. Aberration [HSR^{+}23]. Absence [BH17]. Absorbing [FST20]. Absorption [ASK22, FFA11]. Absorption-Diffusion [FFA11]. Accelerated [AHL22, HMS17, HPZ16, OCLP15, STY11, SYB22, WNS^{+}22, XXYC22, YSB20]. Accelerating [Che14, DHK20, PMZ20]. Account [PS11]. Accounting [PBU^{+}22]. Accuracy [PFA^{+}19, dSO22]. Accurate [BBC11, Far19, SMA11, YJL^{+}17, ZC20]. Acoustic [AD23, DGMW23, DMZ18, JLZ19a, AGO21]. Acousto [LSYZ21, RB18]. Acousto-Electric [RB18, LSYZ21]. Acquisition [BWB14, STCB13]. Acting [Bat23]. Activation [CG19]. Active [ABG13a, ARY10, JPC12, LQZ23, MWWY21, NPJI17, SDM17, TMSP20, YSB20]. Adapted [CJ14, CFM^{+}20, DKP09, Mär11]. Adaptive [ACSW12, ADD12, BZNC16, BGH^{+}21, BPT11, Get11, HSY20, KWRC20, LLBS14, LCS^{+}16, LLLX17, LSWW22, LGL^{+}22, PH14, WZLH20, YGS^{+}19, ZZ19]. Adaptivity [LLSZ09]. Additive [HKLM21, LY12]. Adjoint [CCPS23, LQS14, MPM^{+}17]. Adjustable [LG23]. Admittivity [DHSS13]. ADMM [ACL16, CCMY15, HMY16, LG23, YY17]. ADMM/Douglas [ACL16]. Adversarial [AAG23]. Affected [SG15]. Affine [CAT08, CPRS21, FAS^{+}15, KO16, LVEB09, MY09, RDM18, SCM^{+}12, STV09, SNDP13, Zhu16]. Affine-Scaling [CPRS21]. Algebraic [APST19, TV17]. Algorithm [AM16, BT09, BPS16, BAA14, BAS15, BWB14, BLC10, CCMS13, CKL17, CMLZ18, Che14, CLY19, CLYZ21, CJT^{+}12, Col22, DFM^{+}12, DMTZ16, FR14, FK10, Gil14b, HYY14, HK19, HMXY22, HDH16, HPZ16, HH18, ISW13, KHD^{+}15, KSW20, KMDL19, KSZ12, LY13, LNPS17, LDCG14, LBM13, LLSZ09, LCS^{+}16, LZ16, LSC^{+}18, LLS^{+}20, MWWY21, MB15, MMT18, MÁS^{+}22, OCBP14, PS19, QSUZ11, SG22, STY11, SSL23, SWGL15, THC11, TBKF15, WYYZ08, WGL^{+}22, WGGX22, YYZW09, YLLY19, YLLY20, ZLD^{+}18]. Algorithms [AB10, ADGM14, ACN16, ACL16, BBK22, CTY13, CPP09, CHH^{+}12, CQ21, CCPS23, CNS10, CG19, EZC10, HY12, HPPZ19, KK08, LO17, LS11, LMT23, ODBP15, RTH21, RE15, TM18, YOGD08, ZBSZ22]. Aligned [CLL15]. Alignment [EHL17, OGL15]. Almost [BHM12]. along [Get11]. Alternating [CYY11, CTY13, CEM19, Che14, DTL^{+}21, GOSB14, KHD^{+}15, LLC14, LLS^{+}20, OCLP15, PS16, WYYZ08, YPC17]. Ambiguity [BCD^{+}12]. AMP [ET18, ET18]. Ampère [STV09]. Amplitude [Sto11]. Analysis [ALKÖP19, AGP18, BKBD16, BCP13b, BK15, BGP^{+}17, BCD^{+}12, CC14, DDPV20, DAB^{+}20, DB13, Dro14, FZ20, FH15, FAS^{+}15, GH23, GPST15, GCN21, GK14, Gil14a, GDT18, GL09, GL13, HSF^{+}19, HN17, HY12, HHK^{+}18, HW13, HQ19, HLL^{+}23, LPSS15, LSW14, Lou08, LY18, MB10, NHKD22, PMS20, QS15, RNH19, RGLB14, SDM17, SG22, SHVC19, SSL23, VF13, VF14, Wah15, WSL13, WY17, WQ20, WQ21, WCA^{+}18, WDCT09, Yin10, ZvDT^{+}17]. Analytic [MH17]. Analytical [GKQR20]. Analytics [BH15a]. Analyzing [BFJQ18]. Anatomy [DATP17]. Angle [BGL^{+}21, SZSH11, WTNL21]. Angles [BG15]. Angular [PFA^{+}19]. Anisotropic [BGM14, BP18, BPLX21, CLPS19, CFM09, CFM^{+}20, DGMW23, FSV10, LJL22, LZOX15, YGS^{+}19]. Anisotropy [LMM17]. Any [MÁS^{+}22]. Aperture [AC12, AH17, BCP13b, BMPT16, BGP^{+}17, BK18, BG20, CB11, DFM^{+}12, FSY09, GP15, KT22, LPT20b, LDS20, Voc15, WY14, WY17, YY15]. Apertures [WY12]. Appearance [CV13, NFV22]. Application [AGO21, ACL16, ABSM20, ABR10, BCP13a, BMS23, BGL^{+}21, CCR^{+}12, CHH^{+}12, CFM^{+}20, DL21, DHZ21, GTP^{+}23, HHR08, HSS21, HLST15, HQ19, JM16, Kla11, LS18a, LKW^{+}19, LPSS15, LZZ^{+}23, Lou08, MRM20, MH17, Mui09, PYW^{+}14, RL15, RB15, RDM18, RG16, ST19, Sdi13, SZW14, TPM20, WFBFA11, YY17, ZH21, MSKL09]. Applications [AARW19, BH17, BB14, BBH^{+}23, BBK22, Bel13, BLSW14, CHHN21, CKL17, CFM09, CCQY20, CV13, DDGL19, DPSV17, DB13, Dro14, ERS18, EEF23, ETT15, ELX13, GKL13, HP11, HMZZ19, HK14, JKSV20, KL18a, LY15, LLC14, LLYG14, MMM12, MB10, NK20b, OV14, PLMS20, RLL14, RW13, SHVC19, SMSY11, TM18, WSL13, WLTC12, XY13, YPC17, YOGD08, ZCO18, Zhu16]. Applied [ALKÖP19, BCMO08]. Approach [AN20, BZNC16, BLM^{+}22, BDMS15, BDM15, BCP13a, BK18, BDM^{+}20, CT17, CDP19, CCP12, CLC13, CJPT13, CJPT15, CP16, DSYT10, DDPV20, DD13, DLW16, DAG11, FLZ14, FH11, GH18, GDF15, GBFA10, GBFA12, HWZ22, HHK^{+}18, HHJ^{+}23, HNAC^{+}15, HSR^{+}23, HKLM21, JLN14, KGB15, KP13, MTWB14, MQLC16, MWBB12, MGKR15, NW13b, PG19, PCCP19, PPE^{+}09, PYA^{+}12, RDG09, RTW20, RB18, RW09, SDZ15, STY11, SV08, VDPD20, WN13a, Wan16b, YLLX20, YK16, YZL^{+}18]. Approaches [LS11, NK20b]. Approximate [Col22, GR23, PLCD20]. Approximating [BHV12, KN14]. Approximation [FRV18, GT15, Han12, HR15, JHSX11, LdGKW19, MF13, Pey15, Tsy09b, WE17]. Approximations [BCD^{+}12, FF13, SK23, ZCO18]. Arbitrary [AR20, BFJQ18, LDCG14, WDCT09]. Archaeology [HHJ^{+}23]. Area [CE12, CAT08, MÁS^{+}22, OAUC^{+}20, Yue23]. Area-Preserving [Yue23]. Arises [GSC12]. Arising [JM16, MH17, YCU19]. Array [GP14, GPST15, YBZ^{+}21]. Artifact [ZDL18]. Artifacts [AAG23, BFJQ18, HF12, PUW17, Pal16, YCF^{+}16]. Artist [HMS17]. ASIFT [MY09]. Aspects [HVW15]. Aspherical [GS16]. Assess [AKLS17]. Assignment [BHS23, HSÅS18, KMDL19, ZZPS20]. Assimilation [PM08]. Assisted [LLW23]. Associated [FSV10, HLL^{+}23, LVEB09]. Assumptions [RVCB19]. Astronomical [PPE^{+}09]. Astronomy [BCP13a]. Asymptotic [AD23, ADB^{+}21, ABSM20, CDLZN23, GK14, MHP17, TM16a]. Atlas [ADK15, DAB^{+}20, DL14]. Atrophy [AMY16]. Attenuated [LQS14]. Attenuation [LQS14]. Augmented [LY13, LLS^{+}13, MGKR15, THC11, WT10]. Autocorrelation [LMSY13]. Autoencoders [ST23]. Autoencoding [GAT22]. Automated [CLPS19]. Automatic [BAA14, CJK10, Fou10, KHD^{+}15]. Automatically [TAF^{+}20]. Averaged [LH18, XWH22]. Averaging [RW09]. Away [Mil18].

B [Sdi13]. B-Spline [Sdi13]. Back [ACN16, TG21]. Back-Projection [ACN16, TG21]. Background [YPC17]. Background/Foreground [YPC17]. Backprojection [DMZ18, HF12]. Backpropagation [AGM14]. Backscatter [GNH^{+}22, KKN19]. Backscattered [SSSW09]. Backscattering [TBKF15]. Backward [GKL13, RFP13, RL15]. Balanced [STY11]. Balancing [WC23]. Ball [CWR19]. Banach [JK23, MD15]. Banach-Like [MD15]. Band [Her19, LMT23, SM16]. Band-Limited [Her19, LMT23, SM16]. Bandwidth [SDR20]. Bar [AM16, CvG10, HRSZ16, ISW13, SG22]. Barcode [LEZX14]. Barrier [WH15]. Based [ABG^{+}13c, ACSW12, AT11, BS21b, BQ22, BAA14, BPP22, BCP13a, BS15, BH12, BH15a, BH15b, BEFL21, COS09, CCMS13, CGMP11, CLL11, CTY13, CTWY15, CLDM18, Che14, CBZ18, CGN^{+}13, DSYT10, DL18a, DD13, DPZ20, DL14, DPC13, DLW16, DL21, DLÖS23, DHP19, DMTZ16, FA09, FGS12, FGPT17, GLR18, GB11, GPST13, GSXH18, GEB15, GLQ15, GDT18, GM10, HPZ11, HDH16, HW13, HSR^{+}23, HSÅS18, ISW13, KSW20, LL14, Lan19, LMSY13, LAZ^{+}18, LS19, LPT20a, LPT21, LNZS10, LCS^{+}16, LLC14, LLS^{+}13, LWY16, LY18, MWWY21, MRM20, Mär11, MPM^{+}17, MPGMD19, NNYZ17, NS17, PKPE21, PG19, PPE^{+}09, PYA^{+}12, QSUZ11, RTW20, RKO22, SG22, SHB^{+}18, SKJ^{+}19, STY11, STA22, SEMS19, SDA15, TZS13, TM18, TG21, TPM20, VF13, Wah15, WZ17, WZLH20, WC23, WG22, YGS^{+}19, YMA22, YLH23, ZBN17, ZHW22]. based [LTW^{+}10, Mah12, NT11, RGLB14, VFPA22, MYZ13, PBU^{+}22]. Bases [HLKH14, YGLD17]. Basis [CJ14, GH18]. Bayesian [ASK22, ADK15, DDPV20, DMP18, FR14, GDF15, HPZ22, LDA^{+}22, LBM13, LKW^{+}19, LDS20, MDA^{+}23, Per17, PCP^{+}16, RPW19, ST23, SLS19, SN11, SNM17, TAR^{+}19, VDPD20, ZYZL20]. Be [SMA11]. Beam [LSC^{+}18]. Beamforming [LHLP20, SDR20]. Beams [SDR20]. Beltrami [LZ17a, LL22, LLWG13, RDSK09, WZYX13, WkZ14]. Benchmark [ELB18]. Bertozzi [CFM15]. Best [ADGM14]. Between [ADB^{+}21, AD23, BGV09, FAS^{+}15, GSGJ21, MMT18, SHS10]. Beyond [EKV23, ZTO15, GT15]. Bézier [AGSW16]. Bias [DAG11, MHP17]. Bias-Variance [DAG11]. Biclustering [TS14]. Bijective [Lip14]. Bilateral [Ang13]. Bilevel [DV22, De 23, KP13, PPRV22]. Bilinear [LS18b]. Binary [LMT23]. BinaryRelax [YZL^{+}18]. Bioinspired [BD22]. Biology [AD23]. Biperiodic [LN13]. Bistatic [WY14]. Black [Mil18]. Blind [BR15, BBFA14, BGG19, Car10, CEM19, CB18, CHM13, FM23, GS13, Gil14b, HLST15, HH18, JBS17, KGD21, LEZX14, Mar09, PBU^{+}22, QLZ20, RKT^{+}13, RB15, SX12, WSL13, WLJ22, Yan13]. Blind-Spot [CHM13]. Blob [RK19]. Block [BWB14, HLST15, LP19, LZ17b, RNH19, SMA11, XY13]. Block-Constrained [BWB14]. Blur [LEZX14, YY17, ZWN14]. Blurred [DZ13, SDZ15]. Blurring [EW15]. Blurry [CYZ14]. BM3D [ET18]. Bochner [Car10]. Bodies [HLLS14]. Boltzmann [NTDB19]. Boosted [WGL^{+}22]. Boosting [RE15]. Born [GT15]. Both [LQS14]. Bound [CN22, WGGX22]. Boundaries [APST19]. Boundary [AKL^{+}21, DHSS13, FST23, HP15, Lip14, OAUC^{+}20, ZC15]. Bounded [NPV16]. Bounds [ADGM14, ZBSZ22]. Box [GL13, Mil18]. Brain [BCD19, CLL15, DL14, GTP^{+}23, KT14, LPP^{+}09, StTBRV12]. Breaking [BBL^{+}23]. Breast [CNS10]. Bregman [BBES21, COS09, GO09, LSW14, MBBS14, WT10, YOGD08, Yin10, YY22, ZvDT^{+}17]. Bregmanized [ZBBO10]. Bridges [AvdMSS22]. Brightness [Lan19]. Budget [HNAC^{+}15]. Bundle [BHB21]. Bundles [Bat10, WN21]. Buried [KKN19, TBKF15].

Butterfly [DFM^{+}12]. BV [BP18].

Cahn [CFM15, BKSW14, BS15, BHS09, GLS18]. Calculus [RW13]. Calibration [CJK10, EST20, LS18b, WHY^{+}15]. Camera [JM16, KT16, LSZ18, ÖSB15, RVCB19, SG22]. Camera-Based [SG22]. Cameras [GSGJ21, GOF16, MH17, SXS^{+}15, YBZ^{+}21]. Can [SMA11]. Capacitance [BAA14]. Capped [PC21]. Cardiac [LKW^{+}19]. Cardinal [AR15]. Carlo [DMP18, PMZ20, dSO22]. Carrier [ST11]. Cartoon [JK15, Kut13, XXQJ20]. Cartoon-Texture [XXQJ20]. Cascadic [MRSS08]. Case [GHM23b, LRP17, YSB20]. Cauchy [BdHKU22, HW20, SDZ15]. Cavity [NK16, QYZ19]. CEL0 [SBFA15, SBFA16]. Centering [HSS21]. Centric [GF22]. Cerebral [CHPS09]. Chain [DMP18, PMZ20]. Chan [NPJI17]. Change [FH11, PCP^{+}16, WFBFA11]. Changes [BG14]. Characteristic [MWWY21]. Characteristics [XXQJ20]. Characterization [DL18a, GL09, SJD^{+}15]. Chatter [DHK20]. Cheeger [CFM09]. Chief [Sap08]. Chirps [NDM^{+}11]. Choice [CJK10, NLH^{+}16]. Choose [BBES21]. Chrominance [PAB^{+}15]. Chromoscopy [FFA11]. Circle [SC10]. Circular [CE12, HP15, OAUC^{+}20]. Class [Bat23, DHZ21, EZC10, GR23, GKL13, HL13, KK08, YPC17]. Classes [BMP13, KT14]. Classification [AZ13, ATW14, BS21a, BD22, CCFBY13, DPZ20, EEF23, GH15, LL22, MKB13, SZSH11, WMT^{+}09]. CLEAR [DPSV17]. Clifford [Bat10]. Closed [CLL15, KWRC20]. Cloud [DPH^{+}13, LZ17a, MMT18]. Clouds [CHL16, HKL20, MCL16]. Cluster [MPGMD19]. Clustered [Kut13]. Clustering [CCQY20, SX12, TV17, YLLX20]. Clutter [BPT11, YY15]. Code [SG22]. Coded [RKT^{+}13]. Codes [CvG10, ISW13, NDM^{+}11]. Coding [GTU14, PG19, WMT^{+}09]. Coefficient [ASK22, KKN^{+}18, KLNY23, KLN^{+}23]. Coefficients [MHM23]. Coherence [BBJ^{+}18, FL12, HR15, HSR^{+}23, Mär11]. Coherent [BBE^{+}21, BG21, CDA21, FH11, MNP16]. Collaborative [DMSC16, ZLD^{+}18]. Color [ABR10, BHB21, Bat23, BPP22, BPLX21, CN22, HP11, JNW19, LTKG21, LTKG23, MBBS14, SZGW18, TPM20, WN21, WYN22]. Colorization [PAB^{+}15, PABT17]. Combinatorial [CGTN11]. Combining [BHB21, Lou08]. Comet [HF12]. Cometric [MMM12]. Common [Bat10, RVCB19, RS20, SS11]. Communication [AAG23, HK23]. Communications [CCR^{+}12]. Communications-Inspired [CCR^{+}12]. Compact [MD15]. Compactification [BHS22]. Compactly [HMZZ19]. Comparative [Her19]. Comparison [MY09, RDM18]. Compartmental [CK09]. Compensated [ZCO18]. Compensation [WYN22]. Competition [LNZS10]. Complementarity [De 23]. Completion [BZNC16, BBP09, CESV13, CN22, HCCS20, LSWW22, LZZ^{+}23, XY13, ZN19, ZBSZ22]. Complex [AGO21, HZ14, HMZZ19, HW22, NAF^{+}14, PYW^{+}14, TAR^{+}19]. Complex-Valued [AGO21, HW22]. Component [BCP13b, GH23, HPPZ19]. Components [HKLM21, LS17]. Composite [GM18, HPZ16, PYW^{+}14]. Composition [AR15]. Compressed [AAG23, ACS21, ADD12, BH17, CCW20, DPVW14, FSY10, LHB^{+}18, NDM^{+}11, Poo15, RB15, YOGD08, ZH21]. Compression [LLWG13, MPGMD19, NK20b, SZGW18, TPM20]. Compressive [ADX21, CCR^{+}12, CCBB14, Fan09, FL12, GSL^{+}22, GY12, HPPZ19, KL18b, LLC14, MJC^{+}19, Mui09, PLCD20, RKT^{+}13, Rom09, STCB13, SXS^{+}15, STA22, SJD^{+}15, WHY^{+}15, YBZ^{+}21, ZLD^{+}18]. Compton [JM16, KT16, MH17, Rig17, WQ20]. Computation [DMP18, FN17, KBW13, KLYY21, MDA^{+}23, OAUC^{+}20, WkZ14]. Computational [CJT^{+}12, DATP17, HNAC^{+}15, KGC11]. Compute [MÁS^{+}22]. Computed [LSC^{+}18, MPL^{+}18, MPM^{+}17, PJS21, PLCD20, WZ17, ZE23]. Computer [ODBP15]. Computing [BHFPG21, KLS^{+}17, QLL19, StTBRV12]. Concave [PC21, Per19]. Concepts [BPP22]. Concise [KO16]. Condition [CLC13, GSGJ21, HMXY22, Sdi13]. Conditional [GL17]. Conditionally [CHPS09]. Conditioned [HK23]. Conditions [Col22, DV22, ZC15]. Conductivities [BGM14, CFdGK09, LJL22]. Conductivity [AKL^{+}21, CHKL23, DGMW23, RB18, WR14]. Cone [JM16, LSC^{+}18]. Cone-Beam [LSC^{+}18]. Confidence [Per17]. Conformal [AKZ13, CHL16, CLLGL20, HWZ22, KLYY21, LTW^{+}10, QLL19, SWGL15, WkZ14, ZCZL22, ZCL22, LL14]. Conformality [SCL20]. Conical [MH17]. Conjugate [BLM^{+}22, CZ10]. Connected [CR18, CLLGL20, ZCL22]. Connectedness [NL10]. Connection [HLW20]. Connections [NPJI17, StTBRV12]. Connectivity [HSF^{+}19, LPP^{+}09]. Consensus [BCSB18, HZDZ23, TS14]. Conservation [Lan19]. Conservative [FPM17]. Consistent [CP21, DATP17]. Consisting [JM16]. Constant [BBH^{+}23, LO17, NS14, NNZC08, OJ16]. Constrained [ATTY16, BC15, BWB14, CTY13, CPP09, CCMY15, CGN^{+}13, DGJS16, GS13, Her19, MB15, MB16, ZZ21]. Constrains [KZ14]. Constraint [HP11, HW20, LLBS14]. Constraints [AB10, AMY16, De 23, KR13, SBC22, SNB13, TV20]. Construction [CW22]. Contact [AKL^{+}21]. Content [AE08, Seg22]. Continuation [RM10]. Continuous [BLM^{+}22, CCKW14, CGTN11, Gol11, GL09, LS11, MS22, PWSU16, SBFA15, SBFA16, WY14]. Continuum [HFE19]. Contour [Get11, LQZ23, NTV10, YSB20, ZCO18]. Contours [ARY10, BBP09, JPC12, KZ18, MWWY21, SDM17]. Contraction [HY12]. Contractive [LY15]. Contrario [CCBR13, LRP17, vGPR22, SNB13]. Contrast [ALZ20, DD10, HSY20, WN13b]. Control [ATTY16, BMW09, LPP^{+}09, ZL21]. Controlled [BV16, TAF^{+}20, WR14]. Convergence [ACL16, CTWY15, CDLZN23, CCPS23, ENR20, GH18, GBFA10, GBFA12, HY12, HYY14, HMY16, HMXY22, HW22, HLL^{+}23, Ish14, JK23, LDCG14, MS22, Naj17, RB15, SSL23, TG21, YWW^{+}23]. Convergent [CMLZ18, CEM19, Col22, HW13, KLYY21, LY13, TBKF15]. Conversion [HMS17]. Convex [AR15, BR15, BLM^{+}22, BK17, CCZ13, CCP12, CPP09, CCMY15, CJPT15, CG19, CAT08, Dar15, DZ13, EZC10, GHFT23, GSC13, HHR08, HL13, HPZ16, KYW13, KLS^{+}17, LMSS19, LY15, LWM^{+}18, ÖSB15, PPRV22, PYAC13, PCBC10, PCCP19, PYA^{+}12, RPW19, SO08, SCC14, SS13, TSG^{+}11, XWH22, ZWN14]. Convexification [KKN20, KKN19, KLNY23, KLN^{+}23]. Convexity [ACL16, SCL20, ZCO18]. Convolution [ACN16, BP18, CDS17, GB18, HK14, LLLX17, Rom09, SDL22, WGGX22, YGLD17]. Convolutional [LGCWY18, PG19, SDL22, YHC18, ZE23]. Coordinate [FW10, RNH19, XY13]. Cormack [RLL14]. Cormack-Type [RLL14]. Corner [KZS14]. Corners [GB11]. Corrected [ZN19]. Correcting [MHP17].

Correction [BZNC16, CN17, DD10, HL13, HSR^{+}23, KT14, LHT^{+}21]. Correlated [AC12]. Correlation [FGPT17, GS10, GPST13, IVW16, LPT20a, LPT21, SEMS19, Voc15]. Correlation-Based [FGPT17, GPST13, LPT20a]. Correlations [GP09]. Corresponding [CFdGK09]. Corrupted [Yan13]. Could [REM17]. Coupled [PCP^{+}16, PBU^{+}22, SCC14]. Covariance [CLMT15, GGJ^{+}22, KKS15]. Covariant [BB14, FAS^{+}15, DPSV17]. Covering [RDM18]. Cradle [YCF^{+}16]. CRFs [JDA^{+}19]. Criterion [ROD15]. Critical [KZ18]. Cross [GP09, GS10]. Cryo [BGPS17, RS20, SS12, SS11, SZSH11, WSW13, HSS21, KKS15]. Cryo-Electron [RS20, SZSH11]. Cryo-EM [BGPS17, SS12, SS11, WSW13, HSS21, KKS15]. Cryogenic [DLÖS23]. Crystal [LY18]. Crystalline [HSSP09]. CT [BFJQ18, WTNL21, ZD16, ZDL18]. CUR [CHHN21]. CURE [DJLS20]. Curl [PS11]. Current [BGM14, LJL22, SSL23]. Curvature [BL14b, CLK14, CFSS16, DJLS20, EEF23, HKL20, MMM12, MGKR15, ZC12]. Curve [BG14, LBFA23, RG16]. Curvelet [EHB09, Sto11]. Curvelets [GTO14]. Curves [BBHMA17, BMP13, CAT08, Get11, KK08, NK20a, RLL14, SMSY11]. Curvilinear [FW10, GWY09]. Cut [LO17, WZLH20]. CycleGAN [SOK^{+}20]. Cycles [RG16]. Cyclic [BLSW14]. Cylinder [MH17]. Cylindrical [Hal11].

D [GBFA12, CCBB14, DSYT10, EST20, GS13, GTO14, GBFA10, HRSZ16, HMS17, HZDZ23, LHW^{+}15, LL22, LHB^{+}18, NTDB19, PS11, SM18, TAR^{+}19, TAF^{+}20, TPM20, UC13, WLL^{+}21, WLYU15, YY13, ZC20]. D-Bar [HRSZ16]. Data [ABK15, AAB^{+}11, ARYZ18a, ARYZ18b, BBE^{+}21, BDMS15, BLSW14, BT18, BdHKU22, BGMZ23, BFJQ18, BG15, BHSW18, BMS23, CCW20, CDS17, Car10, CH16, CDA21, CJK10, CFM^{+}20, CPRS21, DPH^{+}13, DL14, DLLY17, DLL19, DJLS20, EEF23, ETT15, FA09, GHFT23, GPPM15, GF22, GSZ17, GHM23b, HQ16, HP15, HL13, HPZ22, Hub13, JGM^{+}12, JLZ19a, Kla11, KKN^{+}18, KKN19, KT16, KL18b, LLW23, LHC^{+}23, SDL22, SAS17, SX12, TAR^{+}19, TM18, TP18, WDS14, XZZ19, ZD16, ZCO18, ZYZL20, ZZPS20]. Data-Assisted [LLW23]. Data-Driven [BDMS15, GHFT23, HPZ22, LHC^{+}23, ZD16]. Data-Fidelity [HL13]. Datasets [LS18a, RR15]. DC [WGL^{+}22]. Deblurring [BAS15, COS09, CTY13, Che14, CvG10, FKLS12, KL18a, LLS^{+}13, LEZX14, MYZ13, MRSS08, NK20b, OV14]. Decay [WCU13]. Decoding [ISW13, SG22]. Decolorization [JLN14, YLH23]. Decomposition [AdHW15, AT11, CHHN21, CTWY15, DM20, FKLS12, HKLM21, LRV21, MHM23, OV14, RTH21, Sto11, VFPA22, XXQJ20, YY13]. Decompositions [BGM^{+}16, HKBH13, TM18, Tii14]. Decompression [BH12, BH15a, BH15b, SYO15]. Deconvolution [BR15, BCP13a, Car10, EHB09, GS10, GS13, HH18, JBS17, Mar09, MBFG20, QLZ20, ShDC^{+}19, WLJ22, ZBBO10]. Decoupled [CFBP23]. Dedicated [DAB^{+}20]. Deep [ALKÖP19, BHB21, BGL^{+}21, CM20, EKV23, LAZ^{+}18, LHC^{+}23, RKO22, YHC18, YZL^{+}18]. Definite [CKA17]. Definition [Con17]. Deflection [SJD^{+}15]. Deformable [BGH18, BGH^{+}21, SDL22]. Deformation [ADK15, DLV23, SY14, SWGL15]. Deformations [CÖ18, LL14, Wol09]. Deforming [SMSY11]. Degeneration [SZGW18]. Degree [Lip14]. Dehazing [FLZ14, GVCPB15]. Delayed [XXYC22]. Delta [KV23]. Denoiser [LG23].

Denoising [All09, BLSW14, BL14b, BLC10, CEM21, DPN18, DL21, ET18, EKV23, FPM17, FLZ14, FQXC17, GGJ^{+}22, HMZZ19, HBD18, KYW13, KSS19, LMSY13, LMM17, LNPS17, LBM13, MGKR15, PLCD20, RE15, REM17, STA22, SSN09, TM16a, WSL13, WM13, Wan16a, ZHW22, ZC12, All08]. Dense [JDA^{+}19, KGC11, Lin18]. Densification [DNB21]. Densities [BGM14, LJL22]. Density [BJM15, BdHKU22, BWB14, CFdGK09, CCKW14, CKA17, CR18, MJC^{+}19]. Density-Equalizing [CR18]. Dependent [BDM15, BMPT16, TBKF15]. Depth [BP14]. Derivative [ABG^{+}13c, PYW^{+}14, Wah15]. Derivatives [BB14, XXYC22]. Descent [BBES21, CLYZ21, FLG23, GFB^{+}23, HPZ16, HH18, JK23, QLZ20, RNH19, TG21, VF13, WNS^{+}22, XY13]. Descent-Based [VF13]. Description [Nik13, vGPR22]. Descriptors [DL18a, DL18b, SCM^{+}12]. Design [CCR^{+}12, GPB17, GS16, HP17, KSZ12, LWY16, SN11, ZLD^{+}18]. Designs [CW18]. Despeckling [MRM20]. Detail [FPM17, GQY14]. Detail-Preserving [GQY14]. Details [ARYZ18a, ARYZ18b]. Detectable [HQ16]. Detection [BBL^{+}23, BPT11, BP14, BMW09, DKP09, Dro14, FH11, GBFA10, GBFA12, HNAC^{+}15, LRP17, LWY16, Mah12, PYW^{+}14, PCP^{+}16, RK19, SDA15, WY10, WFBFA11]. Detector [PB23]. Detectors [ES15, JM16, PFA^{+}19, ROD15, SRG10]. Determinantal [LDG21]. Determination [KHD^{+}15, SS11, WSW13]. Deterministic [NDM^{+}11]. Deviations [LSZ18, WSW13]. Device [LWY16]. Devil [FPM17]. Diagrams [AFGK23]. Dictionaries [ADD12, FF13]. Dictionary [AE08, BKBD16, LLS^{+}13, LGCWY18, NK20b, RKO22, SHB^{+}18, SHVC19, SUFU20, TMSP20, XZC^{+}12]. Dielectric [BL14a]. Diffeomorphic [ADK15, AMY16, ARY10, BJM15, CT13, CÖ18, DGSL23, HW20, MB15, MB16, Sdi13, ZL21]. Diffeomorphism [GDT18]. Diffeomorphism-Based [GDT18]. Diffeomorphisms [SYB22]. Difference [BPLX21, CLL11, LZOX15]. Differences [BLSW14, HP17, TSG^{+}11]. Different [DAMM12, Her19]. Differentiable [AGSW16]. Differential [AT11, DM20, MTWB14, PH20]. Differentiation [Bel13]. Diffraction [BBE^{+}21, BQ22, GNH^{+}22, HLST15]. Diffuse [ASK22]. Diffusion [AvdMSS22, BCGR14, Car10, CDHS13, FQXC17, FFA11, GPB17, GKL13, HR15, MRM20, QYW10, SHVC19, StTBRV12, SZGW18, SSN09, VBK13, YGS^{+}19]. Digital [ALKÖP19, ADB^{+}21, CNS10, HSR^{+}23, KV23, KSZ12, SG15, Zhu16]. Dimension [LDCG14, PYAC13]. Dimensional [AM16, CDRS16, CvG10, CFM^{+}20, DGH11, Dar15, DDPV20, DLÖS23, DH20, EEF23, EKOÅ10, FR14, Gri10, HCCS20, HWZ22, HBD18, KKN20, KGV14, KL19, KT16, LR18, LdGKW19, Lou08, MWBB12, OSZ17, SS11, SW13, SUFU20, TMP13, VDPD20, WCN^{+}19, Wol09, YCU19, ZCZL22, LR17]. Dimensions [BGM14, BLM14, BG15, BGG17, FST20, GKQR20, MBFG20, TT22, ZJ21]. DIP [BHB21]. DIP-VBTV [BHB21]. Dipoles [CV17]. Direct [ACI08, CH16, CHZ21, DMTZ16, JLZ19a, LZZ18, MTWB14, NFV22, XZZ19]. Directed [WFBFA11]. Direction [BMPT16, CYY11, CTY13, CEM19, GOSB14, OCLP15, SS12, YPC17]. Direction- [BMPT16]. Directional [CY09, HMZZ19, LCS^{+}16, PMS20, PLMS20, Sto11, WT13]. Directionality [HZ14]. Directionally [PB23]. Directions [GHM23b, SW13]. Discrepancies [CDS17]. Discrepancy [LLSV14]. Discrete [BER15, Con17, ERS18, ENR20, FPPA14, GLR18, HR15, LMT23, MC16, NPS18, RLS18, RW13, SCL20, SWGL15, WkZ14, WDCT09]. Discretization [BMS23, KK17, WR14]. Discretizations [CP21]. Discriminative [DPZ20]. Disk [KLYY21]. Dispersed [Ish14]. Dispersed-Dot [Ish14]. Displacement [BI15, GTP^{+}23, HSNS18]. Dissipative [NK16]. Distance [BBES21, CCST22, DD20, LZ17a, Mär11, VF14]. Distances [BHS23, CAT08, DAMM12, MD15, NS17, ZvDT^{+}17]. Distortion [SCGAF^{+}15, SLS22]. Distributed [AC12, WY12, XXYC22]. Distribution [CDH16, DHSS13, DPC13, HSSP09, HHJ^{+}23, PYA^{+}12]. Distribution-Based [PYA^{+}12]. Distributions [CKA17, LS19, MPL^{+}18, SNDP13]. Dithering [TSG^{+}11]. Divergence [LBFA23, PS11, YY22]. Divergences [DGSL23]. Diverse [CB11]. Diversity [PFS10]. Division [SCGAF^{+}15]. Document [BGG19, MPGMD19]. Domain [ABK15, CTWY15, EHB09, KZ15, LZ16, WZ17, ZD16, ZDL18]. Domains [APST19, AR20, CYY11, CB18, Gri10, HP15, RBLS14, Wol09]. Dot [Ish14]. Double [GLS18]. Doubly [YL21]. Douglas [ACL16, BPS16, FZ20]. Douglas-Rachford [BPS16]. Driven [BDMS15, BGMZ23, CCW20, FFA11, GHFT23, HPZ22, LHC^{+}23, RTW20, SOK^{+}20, ZD16]. Dual [CTWY15, CY09, CCPS23, Col22, CGN^{+}13, CP16, DHN09, EZC10, Gol11, HY12, HYY14, HMXY22, KZ15, LP19, MSMC15, OV14, ST11, WT10, ZZ21]. Dual-Domain [KZ15]. Duality [DL21]. Duality-Based [DL21]. Dyadic [AdHW15]. Dynamic [AARW19, BP18, CV13, CPRS21, HQ16, TA14]. Dynamical [STCB13]. Dynamics [HSH13].

Earth [Tsy09a]. Echo [FA09]. Echo-Based [FA09]. Echolocation [ATW14]. Edge [BPG08, GY12, GL13, KGB15, Mah12, PYW^{+}14, RK19, SRG10, YYZW09]. Edge-Matching [KGB15]. Edge-Preserving [YYZW09]. Edges [BMW09, DB13, GL09, HNAC^{+}15]. Editor [Sap08]. Editor-in-Chief [Sap08]. Effect [ADK15, ES15]. Effective [YK16]. Effects [DD10, KSPR17]. Efficient [ACN16, BG14, CWR19, DHN09, DMP18, Fou10, GL17, HHMT16, HDH16, JDA^{+}19, Lan19, LQZ23, MDA^{+}23, NDM^{+}11, NNZC08, QSUZ11, RB15, RDSK09, SZW14, WGL^{+}22, YLLY20]. Efficiently [PKCS18]. Eigenfunctions [AGP18]. Eigenmap [LZ17a]. Eigenstates [CDH^{+}21]. Eigenvalues [GPB17]. Eigenvectors [BHFPG21, SS11, SZSH11]. Eikonal [GLQ15]. Elastic [ACI08, DLL19, HLLS14, LLS19, MPM^{+}17, NK20a, PH20, RW09, YJL^{+}17]. Elastica [DGT19, HWC21, LTKG21, LTKG23, RLS18, THC11, YK16]. Elasticity [ABG^{+}13c, OGL15]. Elastography [BI15, BMS23, GTP^{+}23]. Electric [ABG13a, HHR08, RB18, LSYZ21]. Electrical [AKL^{+}21, AAJ^{+}16, AM16, AKLS17, DHSS13, GH18, Gri10, HRSZ16, JKSV20, PAM12, SAS17, WR14]. Electrolocation [ABG13a]. Electromagnetic [ALZ20, BLM14, CH16, LN13, LHLP20, LLW13, Wah15]. Electromagnetism [LR17, LR18]. Electron [DLÖS23, GNH^{+}22, Kla11, RS20, SZSH11]. Electrosensing [BS21a]. Element [BGH18, BMS23]. Ellipses [CF18]. Ellipsoid [LWM^{+}18]. Elliptic [AB10, ABSM20, GKQR20]. Ellipticity [AZ13]. Embedded [ALZ20]. Embedding [CM20]. Empirical [DDPV20, GTO14, VFPA22, VDPD20]. Encoded [HPZ22]. Endowed [MCL16]. Energies [SLS22]. Energy [CvG10, DL14, DB13, HLL^{+}23, Ish14, KLYY21, PKPE21, SSSW09, Yue23, BS09]. Energy-Based [PKPE21]. Engine [REM17]. Enhanced [GVCPB15, KV23]. Enhancement [GM10, GKL13, HSY20, WN13b]. Enhancing [GM15, HR15, LHB^{+}18]. Ensembles [GDT18]. Ensuring [NL10]. Entropic [EHL17, Pey15]. Entropy [MPGMD19]. Entropy-Based [MPGMD19]. Enumeration [LWM^{+}18]. Equalization [WN13b, YLH23]. Equalizing [CR18]. Equation [BS15, CFM15, DH20, Hub13, KLNY23, KLN^{+}23, MPM^{+}17, ZH21]. Equations [ABSM20, AT11, Car10, CW22, Dar15, DM20, GKL13, HW22, MS17, STV09, YGS^{+}19]. Equilibrium [BCSB18]. Equivalence [ADB^{+}21]. Equivalent [ZBSZ22]. Erratum [GBFA12, SBFA16]. Error [BGH^{+}21, CN22, ZBSZ22]. Esedoglu [CFM15]. Essential [TD17]. Estimate [BGMZ23]. Estimates [BGH^{+}21]. Estimating [AGP18]. Estimation [ADK15, AR20, BGV09, BCP13b, BDS18, CDH16, CGÖ19, CKA17, CN17, CLMT15, DSYT10, DDPV20, DAMM12, DPC13, DATP17, DLÖS23, FA09, FFA11, GPST13, HSNS18, KSZ11, KKS15, KBW13, Lan19, LS19, LPT20b, LHB^{+}18, MHP17, NFV22, ÖSB15, Per17, Per19, SCGAF^{+}15, SS12, SY14, SDA15, VDPD20, YWW^{+}23, YY22, dSO22]. Estimator [ASK22, DVFP14]. Euclidean [ZCO18]. Euler [DGT19, HP11, HWC21, RLS18, THC11, YK16]. Even [BHV12]. Evolution [BG14, PH20, SV08]. Evolutions [AvdMSS22]. Evolving [Lan19]. Exact [CJPT15, JM16, LSZ18, QLZ20, SBFA15, SBFA16, YY19]. Example [LRP17]. Examples [All08, All09]. Exchangeable [FH11]. Exemplar [CGMP11, RTW20, TPM20]. Exemplar-Based [CGMP11, RTW20, TPM20]. Existence [ERS18]. Expansion [FH15]. Expansions [ABSM20, RLL14]. Expectation [LLSZ09, LM13, YMA22]. Expectation-Maximization [LLSZ09]. Experimental [KKN^{+}18, KKN19, SN11, TBKF15]. Experiments [GPB17, LM13, SWGL15, VDPD20]. Explicit [DH20, HLST15, PMZ20]. Exposures [RVCB19]. Expression [DL14]. Extended [AGK^{+}12, BPG08, GM10, LDS20, LLS19, TMP13, TMP18]. Extended-Sampling-Bayesian [LDS20]. Extending [Naj17]. Exterior [BdHKU22, SM16]. Extraction [ALKÖP19, BAA14, WLTC12, YPC17]. Extragalactic [HHJ^{+}23]. Extrapolation [ERS18, RDSK09]. Extremal [MCL16]. Extreme [LBFA23]. Eyes [GF22].

Fabric [NNYZ17]. Fabry [Aco19]. Face [DPZ20]. Faces [LRP17]. Facet [LWM^{+}18]. Facial [HKBH13]. Factor [AKZ13]. Factorization [GS17, LN13, LWM^{+}18, PN23, PKCS18, TV20, XY13, ZH21]. Fading [FSY09]. Family [AZ13, RLL14]. Far [DLL19, GS17, GHM23b, HLLS14, JLZ19a, JLZ19b, LLW13]. Far-Field [DLL19, GHM23b, HLLS14, LLW13]. Faraday [KK17]. Fast [AB10, ACN16, BT09, BBC11, BAA14, BAS15, BGP^{+}17, BKSW14, CTY13, CHH^{+}12, FLZ14, FGPT17, GHFT23, GS13, GOSB14, HPPZ19, IVW16, KBW13, LO17, LY12, LP19, Liu21, SAS17, SW14, THC11, TM18, XWH22, YYZW09, YK16, Zhu16, CLL15]. Faster [FK10]. Fatemi [CTWY15, LP19, NPJI17]. Fatness [AKLS17]. Feasibility [LSW14]. Feature [DB10, FH15, MCL16, PFA^{+}19, RTW20, ROD15, Rig17, WLTC12]. Feature-Driven [RTW20]. Feature-Endowed [MCL16]. Feature-Preserving [DB10]. Features [CFBP23, DGH11, Far19, RDG09]. Few [OJ16]. Fiber [CDH16, GSL^{+}22]. Fibered [GH23]. Fidelity [HL13, WZYX13].

Fidelity-Beltrami-Sparsity [WZYX13]. Field [BL14a, Bat10, DLLY17, DLL19, FA09, Fou10, GS17, GHM23b, HLLS14, HSNS18, JLZ19a, JLZ19b, LLW13, MWBK14, NS17, NTDB19, NHKD22, NL10, PCP^{+}16, Sdi13, SSSW09, WCA^{+}18, XZZ19, YL21]. Fields [AC12, BLM^{+}22, CY09, CCFBY13, FW10, Her19, KvD12, NPJI17, PS11, RBB20]. Figure [PYA^{+}12]. Figure-Ground [PYA^{+}12]. Film [HMS17]. Filter [BHI11, FGPT17, LR16, LM11, Mah12, TT22]. Filtering [Ang13, Bel13, BPT11, BCMO08, GZC^{+}15, KZ15, LS18a, NT11, RDSK09, SMSY11, WDCT09]. Filters [LS19, Mil13, SSN09]. Filtrated [TV17]. Finding [ELX13, PKCS18]. Fine [ARYZ18a, ARYZ18b, Dro14]. Fingerprinting [DPVW14, DHP19, WE17]. Finite [BGH18, BMS23, CLL11, Dar15, HP17]. Finite-Difference [CLL11]. Finite-Dimensional [Dar15]. Finsler [HSF^{+}19]. First [BBC11, Col22, EZC10, GT15, LGCWY18]. First- [LGCWY18]. First-Order [BBC11, Col22]. Fish [ABG13a]. Fisher [VHO20]. Fitting [CJK10]. Fixed [CEM21, FZ20, RTH21, ZZ21]. Fixed-Point [CEM21]. Flares [SLS19]. FLASH [CLL15]. Flat [PB23]. Flexible [CLPS19, CSS08]. Flicker [DD10]. Flicker-Like [DD10]. Flickering [SG15]. Flip [SLS22]. Flip-Free [SLS22]. Flow [BGK15, BMW09, CFSS16, CGTN11, HP11, KGC11, KZ18, LDCG14, PYA^{+}12, SI23, SXS^{+}15, WkZ14, YY19]. Flows [ACDG18, AGP18, CMY10, CALG21, EEF23, FSV10, Pey15, ZZPS20, GWY09]. Fluctuations [SSSW09]. Fluid [PM08]. Fluidic [RW13]. Fluorescence [DLW16, RZ13]. Flutter [TMR13, TM16b]. fMRI [JGM^{+}12]. Focus [LEZX14]. Focusing [ES15]. Folded [PC21]. Folds [QLL19]. Force [LQZ23]. Foreground [YPC17]. Formal [GBFA10, GBFA12]. Formation [GPPM15]. Formula [FAS^{+}15]. Formulas [DH20, Hal11, HF12]. Formulation [CTWY15, CBB14, KSW20]. Formulations [BGH^{+}21, SYB22, YSB20, BS09]. Forward [GKL13, HP17, KL18a, MPM^{+}17, MHM23, RFP13, RL15, ZE23]. Forward-Adjoint [MPM^{+}17]. Forward-Backward [RFP13, RL15]. Foundation [WLYU15, Yue23]. Fourier [BCP13a, AGH14, ACS21, BGV09, GSZ17, KL18b, LMT23, MJC^{+}19, OJ16]. Fourier-Based [BCP13a]. Fourth [Dro14]. Fractal [LVEB09]. Fractional [BS15, HW20, YGS^{+}19, ZC15]. Fractional-Order [HW20, ZC15]. Frame [Bat10, COS09, CCMS13, CCW20, Che14, CBZ18, LHC^{+}23, LZD^{+}16, STY11, TZS13, ZD16]. Frame-Based [COS09, Che14, STY11]. Framelet [LCS^{+}16]. Framelet-Based [LCS^{+}16]. Framelets [HZ14, HMZZ19, YHC18, YGLD17]. Frames [GL13, PWSU16]. Framework [AKLS17, Bat10, BBHMA17, BBES21, BT18, BH15a, BH15b, CDA21, DPVW14, DDGL19, DPZ20, DMSC16, EZC10, Gil14a, GDT18, GSZ17, HSNS18, Lan19, LZ16, LYZ20, LWM^{+}18, MY09, Naj17, PABT17, SNM17, SYB22, TS14, UC13, YHC18, YSB20, YY22, YJL^{+}17, ZCZL22, ZLD^{+}18, ZTO15, vGPR22]. Fréchet [DATP17]. Fredholm [CCBB14]. Free [BCSB18, FH11, Rig17, SLS22]. Frequency [BGPS17, BdHKU22, BPT11, BMPT16, BG15, CDLZN23, FPM17, LS17, LHLP20, PS19]. Frequency-Dependent [BMPT16]. Frobenius [CHM13]. Full [AGO21, EGvL^{+}18, NHKD22]. Full-Field [NHKD22]. Full-Waveform [AGO21, EGvL^{+}18]. Fully [MY09]. Function [BHI11, DAMM12, HP15, MWWY21, Mah12, NS17, NL10, SDA15, WCN^{+}19]. Function-Based [MWWY21]. Functionals [ABG^{+}13c, AGP18, BPS16, BGM^{+}16, HK14, KR13, PPRV22, WH15]. Functions [AL15, BBH^{+}23, CG19, HSSP09, LO17, LHLP20, Mär11, PC21, TSG^{+}11, TPM20, ZCO18, ZJ21]. Fundamental [LLLX17]. Fusion [DBCS14, HBM12, JZMN21, KZS14]. Fuzzy [LNZS10].

Galaxy [AZ13]. Gamma [AC12, CYZ14, WCA^{+}18]. Gauge [KvD12]. Gauss [JKSV20]. Gaussian [DD13, CHPS09, CKA17, CJPT15, DPN18, DD20, GL17, GPB17, GGJ^{+}22, WM13, XFPA14, Yan13]. Gaussian-Impulse [DD13]. GCV [GS13]. Gene [DL14]. General [DMSC16, EHB09, EZC10, KK08, LO17, NK20a, SHS10, WGGX22, YHC18]. Generalization [BHS09, YY19]. Generalizations [LSW14, Yin10]. Generalized [ABG^{+}13b, APST19, AH17, BS09, BBH^{+}23, BKP10, BHSW18, CDHS13, Che14, CLMT15, DPN18, GR23, GB18, HMXY22, KR17, LHW^{+}15, LPT20a, LLC14, RFP13, RL15, RTH21, SSSW09, VBK13, WLYU15, WQ21, ZHW22]. Generation [ADGM14, AC12, BCC^{+}16, DAB^{+}20, TPM20]. Generic [RZ15]. Genus [CLL15, CHL16, LW14, YLLY20]. Genus- [CLL15, CHL16]. Genus-One [YLLY20]. Geodesic [BER15, CKL17, CDA21, MWWY21, Mon14, RW13, StTBRV12, SV08, YCU19]. Geodesics [BDMS15, FN17, KN14, NPV16, NPS18, ZBO14]. Geometric [ACDG18, BGL13, DSYT10, DGH11, DHZ21, DB10, FH15, GSC12, HMS17, HSÅS18, LW14, SBS23, SMSY11, SNB13, WLL^{+}21]. Geometrically [CGMP11]. Geometries [BAA14]. Geometry [AKR13, CLMT15, HSF^{+}19, LPP^{+}09]. Gesture [LWY16]. Gesture-Based [LWY16]. Gillette [CFM15]. Ginzburg [DB13]. Global [CLLGL20, CK09, EKOÅ10, KGB15, Nik13, NL10, PCBC10, SHVC19, TM16a, ZCL22]. Globally [CMLZ18, LY13, TBKF15]. gon [MÁS^{+}22]. GPU [ACN16, HMS17]. GrAdient [DVFP14, ABR10, BBES21, CMY10, CZ10, CALG21, Dro14, FLG23, HPZ11, HYY14, HPZ16, JK23, LAZ^{+}18, MSMC15, Pey15, QLZ20, RLS18, STY11, TV20, TG21, WNS^{+}22]. Gradient-Based [HPZ11, LAZ^{+}18]. Gradients [KD12, NW13b]. Graduated [YL21]. Grain [AFGK23, HSSP09]. Graph [BT18, CM20, EHL17, HLW20, LS18a, SBS23, SBC22, TMSP20]. Graph-PDE [SBS23]. Graphical [HSÅS18]. Graphs [BvGL^{+}23, CE12, CGN^{+}13, EEF23, ETT15, HFE19, KSS19, LO17, MKB13, RL15, SDM17, ZZPS20]. Gravity [HQ19]. Gray [MRM20]. Grayvalue [BHS09]. Greedy [PG19]. Green [Mah12]. Grid [AN20, CCW20, CDP19, Fou10, LBFA23, OJ16]. Grids [FL12, SC10]. Gromov [BHS23]. Ground [PYA^{+}12]. Group [BHS22, GSXH18, LZZ^{+}23, PC21]. Group-Tube [LZZ^{+}23]. Groups [Bat23, ZBO14]. Guarantee [YWW^{+}23]. Guarantees [ADX21, KBW13, RB15]. Guided [CGMP11, EB16, FL12, GY12, GGJ^{+}22, HMS17, WM13]. Guidefill [HMS17].

Haar [CSS08, LCS^{+}16]. Haar-Wavelet [CSS08]. Hadamard [BPS16, ENR20]. Half [RZ15, RYZ18]. Half-Quadratic [RZ15, RYZ18]. Halftoning [Ish14, KV23]. Hamilton [Dar15, DM20]. Hamiltonian [AvdMSS22]. Hard [GLS20]. Harmonic [AGM14, ALKÖP19, BCD19, FG23, HKLM21, LL22, SSL23, GWY09, CLL15]. HDMI [HBD18]. HDR [ADGM14, RVCB19]. Head [BP14]. Heart [SI23]. Heavy [BPT11]. Hellinger [CCST22]. Helmholtz [CDLZN23, FH11].

Hessian [LPSS15, WE17]. Heterogeneity [KKS15]. Heterogeneous [HLST15, MQLC16, PCP^{+}16]. Hidden [HLKH14]. Hierarchical [JGM^{+}12, LRV21]. High [ALZ20, BdHKU22, BG20, BGG17, BLC10, CDLZN23, DDPV20, EEF23, FPM17, FPT20, HBD18, KWRC20, KT22, LW14, MWBB12, PYAC13, PAM12, TA14, VDPD20, WZ17, WT10, ZJ21]. High-Dimension [PYAC13]. High-Dimensional [DDPV20, EEF23, HBD18, VDPD20]. High-Frequency [CDLZN23]. High-Genus [LW14]. High-Order [FPT20, PAM12]. High-Resolution [BG20, KT22]. High-Speed [WZ17]. Higher [ABSM20, DB10, FQC16, JDA^{+}19, JK15, PMS20, PLMS20, QYW10, SBS23, SNDP13, SRG10]. Higher-Order [DB10, FQC16, JDA^{+}19, JK15, PMS20, PLMS20, SBS23, SNDP13]. Highly [Far19, SAS17]. Hilbert [XWH22]. Hilliard [BKSW14, BS15, BHS09, CFM15, GLS18]. Hippocampal [WLTC12]. Histogram [SS13, WN13b, YLH23]. Holographic [BEFL21]. Holography [AGM14, CDSV18]. Homodyned [DPC13]. Homogeneous [AGP18, BGM^{+}16, CALG21, SDA15]. Homography [LLLX17]. Horn [LDCG14]. Hough [ADB^{+}21, BMP13]. Hubble [Car10]. Human [BCGR14, WLL^{+}21]. Hybrid [HRSZ16, HYY14, MSMC15, RGZ13, ZZ19]. Hydrodynamics [DAW21]. Hyperbolic [DHZ21, DMTZ16]. Hyperfields [GTU14]. Hypergraph [CCQY20]. Hypermodels [CHPS09]. Hyperspectral [PBU^{+}22, RKT^{+}13, ShDC^{+}19, XZC^{+}12]. Hypersurfaces [BHM12]. Hypoelliptic [BCGR14]. Hypothesis [Dem09].

Identifiability [BR15]. Identification [AB10, ATW14, CJ12, GM15, HK19, LQS14, SI23, SNM17]. II [All08, ARYZ18b, BH15b, DGH11, DDPV20, LR18, LHW^{+}15, NW13b]. III [All09]. Ill [GHFT23, KR13, RNH19, SKJ^{+}19]. Ill-Posed [GHFT23, KR13, RNH19, SKJ^{+}19]. Illumination [BPG08, DGH11, NMP15, SG15, WFBFA11]. Illusory [KZS14]. Image [AC09, ADGM14, AE08, All09, AT11, ABR10, ADD12, BGH18, BGH^{+}21, Bat10, BB14, BHB21, Bat23, Bel13, BBFA14, BHS22, BG14, BL14b, BAS15, BC15, BDM15, BCP13a, BH15a, BH15b, BEFL21, BPLX21, BDS18, COS09, CCZ13, CLPS19, CMY10, CCBR13, CLK14, CLL11, CYY11, CYZ14, CCQY20, CPP09, CZ10, CHH^{+}12, Che14, CÖ18, CGÖ19, CLY19, CLYZ21, CN22, CJPT13, CJPT15, CFSS16, CFM^{+}20, CG19, DGH11, DPSV17, DPN18, DD10, DGT19, DL18b, DB13, DHN09, DGJS16, DBCS14, EHB09, ERS18, EEF23, EKV23, ETT15, FPT20, FLZ14, FF13, FQC16, FQXC17, FFA11, FM23, FKLS12, FRV18, GVCPB15, GSXH18, Get11, Gol11, GNH^{+}22, GKL13, GL13, GSZ17, GGJ^{+}22, HP11, HMZZ19, HW20, HWZ22, HBM12, HWC21, HHK^{+}18, HW13, HL13, HK14, HBD18, HSÅS18, HCGN22, HCGN23, HKLM21]. Image [JNW19, JLN14, JZMN21, JK15, KD12, KGV14, KvD18, KL18a, KRW10, KT16, KSPR17, KZ18, LS18a, LMSY13, LMSS19, LDG21, LG23, LBM13, LLBS14, LNZS10, LZ16, LYZ20, LRV21, LGL^{+}22, LHC^{+}23, LLS^{+}13, LZ17b, LTKG21, LQZ23, LTKG23, LR16, LZOX15, Lou08, LY18, MYZ13, MRM20, MB15, MB16, MS22, MS17, Mär11, MKB13, MY09, MRSS08, MPGMD19, MGKR15, Nat16, NW13a, NPS18, NK20b, NW13b, NPJI17, OV14, OSZ17, PM08, PRTW21, PAB^{+}15, PKPE21, PBU^{+}22, PV14, RTW20, RTH21, RLS18, RZ15, RDM18, RE15, RDSK09, RKO22, Seg22, STY11, SZGW18, STA22, SBS23, SCL20, SDA15, SS13, TZS13, TPG16, TM12, Tii14, VZE16, VSU15, VF13, VF14, WYYZ08, WZYX13, WSL13, WM13, WN13a, WN13b, WZLH20, WN21, WGL^{+}22, XXQJ20, YYZW09, YWW^{+}23, YMA22, YK16, YGLD17, ZWJ19, ZD16, ZC15, ZBN17, ZDL18]. Image [ZC20, ZL21, ZYZL20, ZC12, Zhu16, All08, BS09, MSKL09]. Image-Driven [FFA11]. Image-Signature-Dictionary [AE08]. Image/Video [Zhu16]. Imagery [Car10, WY17]. Images [AC12, ADB^{+}21, BHI11, BPS16, BER15, BGL13, BP14, BHS09, CC14, CYZ14, CBB14, DDGL19, Dem09, DZ13, FBU15, FAS^{+}15, GTP^{+}23, GB18, GBFA10, HSS21, HSY20, KT14, LVEB09, LS17, LY12, LNPS17, LMT23, MWBB12, MB10, NPS18, NTDB19, NDM^{+}11, NNZC08, OJ16, PC21, PN23, PCCP19, PCP^{+}16, SDZ15, SM16, SZSH11, SW14, SG15, WSW13, Wan16b, WCA^{+}18, XZC^{+}12, Yan13, YLLX20, YGS^{+}19, YCF^{+}16, BGV09, GBFA12]. Imaging [ADX21, AGO21, AARW19, ACI08, AAB^{+}11, AGK^{+}12, ABG^{+}13c, AGM14, AdHW15, ACL16, BR15, BGM14, BL14a, BV16, BBE^{+}21, BBH^{+}23, BPG08, BPT11, BCP13b, BK15, BMPT16, BK17, BGP^{+}17, BK18, BG20, BG21, BCC^{+}16, BCMO08, CCMS13, CHHN21, CFdGK09, CDRS16, CV17, CMP14, CHH^{+}12, CDHS13, CH16, CGÖ19, CDH^{+}21, CK09, CPRS21, Dar15, DM20, DV22, De 23, DFM^{+}12, DLLY17, DHP19, DHZ21, Dro14, DMZ18, ES15, EZC10, Fan09, FSY09, FGPT17, GP09, GPST13, GP14, GP15, GPST15, GPB17, GT15, GPPM15, GKQR20, GH15, GM10, GSL^{+}22, GY12, GHM23b, GS16, HLST15, HPZ22, HP17, KT22, KK17, KL18b, LDA^{+}22, LKR18, LKW^{+}19, LPT20a, LPT20b, LPT21, LLSV14, LHLP20, LY15, LCS^{+}16, LZZ18, Liu21, LJL22, MDA^{+}23, MNP16, MNPT17, MB10, Mui09, NMP15, PPRV22, PLMS20, PJS21]. Imaging [PPE^{+}09, QYW10, RKT^{+}13, RB15, RPW19, RVCB19, Sap10, SLS19, StTBRV12, ST11, SSSW09, ShDC^{+}19, SDR20, SJD^{+}15, SG15, TM18, TBKF15, TA14, TMP13, TMP18, Tsy09a, Tsy09b, VBK13, Voc15, WY12, WY14, XZZ19, YWW^{+}23]. Impact [RVCB19]. Impedance [AKL^{+}21, AAJ^{+}16, AM16, AKLS17, DHSS13, GH18, Gri10, GH15, HRSZ16, HHR08, JKSV20, PAM12, SAS17, WR14]. Impedances [AKL^{+}21]. Imperfect [KL18a]. Implementation [LHW^{+}15, Mon14, ZL21]. Implementations [ACN16]. Implicit [Bel13, Fou10, HLST15, LSWW22, YWW^{+}23]. Implicit-Explicit [HLST15]. Improper [CAT08]. Improved [ADX21, BAS15, HMXY22, SRG10]. Improvement [CBB14]. Impulse [DD13, MYZ13, Yan13, ZWJ19, ZBN17]. Impulsive [CJ12]. In-Line [CDSV18]. Incidence [BZ18, BG15]. Incident [LLS19]. Inclusion [ACI08, CHKL23, HHMT16]. Inclusions [BLM14, Gri10, Wah15, YJL^{+}17]. Incompatibility [BCD19]. Incomplete [BFJQ18, CCBB14, XZC^{+}12]. Incompressible [HLKH14, Wol09]. Inconsistent [CW22]. Incorporating [LYZ20]. Increasing [HZ14]. Incremental [WMT^{+}09]. Independently [GOF16]. Indicator [AL15, MRM20]. Indicator-Based [MRM20]. Indirect [CÖ18]. Induced [LZZ^{+}23]. Inducing [LMSS19]. Induction [DGMW23, QS15]. Industrial [PJS21, ShDC^{+}19]. Inequalities [ACSW12, LLBS14]. Inequality [CW22]. Inertial [CCMY15, OCBP14, XXYC22, PS16]. Inertial-Accelerated [XXYC22]. Inexact [DL21, MB15, RYZ18]. Inf [BP18]. Inf-Convolution [BP18]. Inference [LAZ^{+}18, SN11, YL21, ZYZL20]. Infimal [CDS17, GB18, HK14, WGGX22]. Infinite [BL14a, BZ18]. Influence [SM16]. Information [AM16, BJM15, BPP22, CH16, KLS^{+}17, LYZ20, VHO20]. Information-Based [BPP22]. Information-Theoretic [KLS^{+}17]. Inhomogeneity [LYZ20]. Inhomogeneous [QYZ19]. Initial [MPL^{+}18]. Initio [RS20]. Injection [SSL23]. Inpainting [BKSW14, BS15, BHS09, CGMP11, CSS08, CYY11, CN22, DAW21, GL17, GLS18, GK14, HP11, HMZZ19, HMS17, LZ16, LHC^{+}23, Mär11, NAF^{+}14, RTW20, WSL13, Yan13, YK16, ZCO18]. Input [LWY16]. Inscribed [LWM^{+}18, MÁS^{+}22]. Inspection [NNYZ17]. Inspired [CCR^{+}12, NPS18]. Instance [SZW14]. Instantaneous [MB10]. Instruction [LWY16]. Instruction/Input [LWY16]. Integrability [CHM13, Tii14]. Integral [CCBB14, FGS12]. Integrals [OAUC^{+}20]. Integrands [PPRV22]. Integrated [DHP19]. Integration [SBS23]. Integro [AT11]. Integro-Differential [AT11]. Intended [MB10]. Intensity [BMW09, DPC13, KT14, KSPR17, LL14, LY12, LYZ20, MNPT17, NMP15, Seg22]. Intensity-Based [LL14]. Intensity-Only [MNPT17, NMP15]. Interaction [ARYZ18b]. Interactions [NL10]. Interactive [PABT17]. Interest [MÁS^{+}22, SNB13]. Interferometric [BG20, MNPT17, YY19]. Interferometry [BG21, MJC^{+}19]. Interior [HHK^{+}18, LHW^{+}15, SZGW18, WLYU15, WC23]. Interior-Point [HHK^{+}18]. Internal [BGMZ23, BMS23, QYZ19]. Interpolated [SM16]. Interpolation [CLK14, CF18, Get11, KD12, MS22, Sdi13]. Interpolations [ZCO18]. Interpretation [SO13, SSN09]. Intersection [RTH21]. Interslice [Sdi13]. Intrinsic [ST19, WLTC12]. Invariance [WFBFA11]. Invariant [AKR13, KSS19, KZ18, LZ17a, LLLX17, MY09, RR15, RDM18, SCM^{+}12, VSU15, ZBO14]. Invariants [FGS12, OAUC^{+}20]. Inverse [AKM11, AAJ^{+}16, AHL22, BZ18, BT09, BDM^{+}20, BG15, BGG17, BMS23, BGL^{+}21, Che14, CDLZN23, CPW^{+}14, CHKL23, CJT^{+}12, Col22, DDPV20, DLL19, GLS20, GAT22, GS17, Han12, HP17, HLLS14, HQ19, HSNS18, JLZ19a, JK23, KR17, KKN20, KKN^{+}18, KKN19, KLNY23, KLN^{+}23, LUZZ22, LR17, LR18, LN13, LH18, LDS20, LSYZ21, LLW23, LS18b, LZZ18, LLS19, LHT^{+}21, Mar09, QYZ19, RPW19, RYZ18, ST23, SKJ^{+}19, SO08, SOK^{+}20, TBKF15, VZE16, VDPD20, WZYX13, XWH22, XZZ19, YHC18]. Inversion [AGO21, BGMZ23, CHZ21, DLW16, DH20, DMTZ16, EGvL^{+}18, GR23, Hal11, JM16, LMT23, Mon14, MH17, YY15, YCU19, YY19]. Invertibility [SCGAF^{+}15, Sdi13]. Invisible [PB23]. Ionosphere [Tsy09a]. iPALM [PS16]. iPiano [OCBP14]. Isometric [ST19, Wol09]. Isometry [BH17, GCN21]. Isotropic [BPLX21, GPB17, LZOX15]. Issues [RVCB19]. Iterated [BHI11]. Iteration [LH18, WT10, XWH22]. Iterations [COS09, KR17]. Iterative [BT09, BAS15, CLK14, CPP09, GLS20, HN17, LR18, WY10, YOGD08]. Iteratively [GHFT23, ODBP15].

Jacobi [Dar15, DM20, LP19]. Jet [KSW20]. Jigsaw [HLW20]. Joint [BvGL^{+}23, BDS18, CGÖ19, CLY19, CBZ18, DSYT10, DL18a, DAB^{+}20, DLW16, GAT22, JHSX11, JBS17, LGL^{+}22, MPL^{+}18, OGL15, YWW^{+}23, YJL^{+}17, ZDL18]. JPEG [BH12]. Junctions [BG14].

Kaczmarz [CQ21, HHJ^{+}23, HW22, LZ18, LH18, XWH22]. Kalman [GZC^{+}15]. Kantorovich [CCST22, HQ19, LLSV14, MSKL09]. KDE [FK10]. Kendall [SUFU20]. Kernel [ACSW12, CKA17, RLL14, RG16].

Kernel-Based [ACSW12]. Kernels [FSV10]. Keypoints [TPM20]. Kinematic [HHJ^{+}23]. Kirchhoff [DL14]. Knowledge [FGS12, KZ14]. Known [KT14]. Krylov [Her19, MB15, YY17].

Labeling [GSC13, HSÅS18, LAZ^{+}18, LS11, SBS23, ZZPS20]. Lagrangian [BLM^{+}22, LLS^{+}13, MGKR15, THC11, WT10]. Lagrangian-Based [LLS^{+}13]. Lambertian [CT17]. Lamé [BdHKU22, HSNS18]. Landau [DB13]. Landmark [CKL17, CLL15, LL14, LTW^{+}10, LLYG14, MS17]. Landmark- [LL14]. Landmarks [MMM12]. Langevin [DMP18, LDA^{+}22, MS17]. Laplace [LZ17a]. Laplacian [CCQY20, CM20, ETT15, HFE19, HLW20, LS18a]. Large [CÖ18, FR14, FD20, HPZ16, LL14, LKW^{+}19, SN11, WE17]. Large-Scale [FD20, HPZ16, LKW^{+}19, WE17]. Larger [HMY16]. Lattice [HK19, NNYZ17, NTDB19]. Lattice-Based [NNYZ17]. Law [KK17]. Laws [Lan19, SV08]. Layered [GS10]. Layover [WY17]. LDDMM [Her19, SNDP13]. Learn [LO17]. Learnable [CLYZ21, LZZ^{+}23, VFPA22]. Learned [FLG23, LHT^{+}21]. Learning [BS21b, BKBD16, CP21, CDA21, DV22, De 23, EKV23, KP13, LZ18, LGCWY18, NS17, PB23, PRTW21, PFS10, PKPE21, RTH21, RKO22, SHB^{+}18, SHVC19, ST19, SKJ^{+}19, STA22, SOK^{+}20, SUFU20, TMSP20, XZC^{+}12, YHC18]. Learning-Based [RKO22, SKJ^{+}19]. LEAst [DPSV17, ELX13, LSZ18, LS18b, Nik13, SBFA15, SBFA16, WSW13]. LEAst-Square [DPSV17]. Left [SI23, ZBO14]. Left-Invariant [ZBO14]. Length [WMT^{+}09, vGPR22]. Lens [GS16, SCGAF^{+}15]. Level [AKM11, ACDG18, AHL22, EST20, FPT20, GB11, KYW13, KBW13, LAZ^{+}18, MRM20, RGLB14, SV08, SDA15, WFBFA11].

Level-Set [AHL22, FPT20, SV08]. Level-Set-based [RGLB14]. Levels [BH17]. Lidar [TAR^{+}19]. Lie [Bat23]. Lifted [KMDL19]. Lifting [BR15, BLM^{+}22, DLÖS23]. Light [HSY20, KZ14, MWBK14, SJD^{+}15, WFBFA11]. Like [DD10, DL18b, Kla11, Lan19, MD15, BPS16]. Likelihood [CJPT15, DDPV20, VDPD20]. Likely [KSPR17]. Limit [BBL^{+}23, CDLZN23]. Limited [AAB^{+}11, AH17, BGL^{+}21, Her19, Kla11, KRW10, LMT23, LDS20, SM16, WTNL21]. Limited-Angle [BGL^{+}21, WTNL21]. Limited-View [AAB^{+}11]. Limiting [HVW15]. Limits [HFE19, dSO22]. Line [BHS23, CDSV18, HF12, JM16, KGC11]. Linear [AL15, ACL16, AH17, BT09, BCMO08, DLLY17, FAS^{+}15, GHM23a, HW22, JK23, KGD21, KR13, LS18b, RNH19, RYZ18, STCB13, SN11, SZGW18, WSL13]. Linear-Quadratic [KGD21]. Linearization [BGG17]. Linearized [AAJ^{+}16, COS09, CCST22, Che14, LSW14, OCLP15, PS16, Yin10]. Linearly [BC15, CCMY15, Col22, LY13, ZZ21]. Lines [PCCP19, RS20, SS11, WFBFA11, ZCO18]. Linking [KZ18]. Links [AD23, ZJ21]. Liouville [CDLZN23]. Lipschitz [BC15, WGGX22, ZWJ19]. Lithography [CJT^{+}12]. Little [REM17]. LL1 [PBU^{+}22]. LL1-Based [PBU^{+}22]. Lobe [MQLC16]. Local [ACL16, BHM12, CFdGK09, DGH11, DD10, DLLY17, FH15, IVW16, KSZ11, KvD18, LLSZ09, LQZ23, LM11, RDG09, YGLD17, YLH23, ZTO15]. Local-Nonlocal [YGLD17]. Localization [ABG^{+}13c, CHPS09, ST19, Wah15, WCN^{+}19, dSO22]. Localized [CMP14, CDH^{+}21, Far19, LS17, AAG23]. Locally [LG23, MAP11, SNDP13, WSL13, XZZ19]. Locating [LLW13]. Location [LSZ18]. Log [DPC13, Per19, RB18]. Log-Concave [Per19]. Log-Conductivity [RB18].

Log-Moments [DPC13]. Logarithmic [Car10, CFM15]. Long [HNAC^{+}15]. Longitudinal [CDA21]. Looking [CM20]. Loop [KWRC20]. Loss [CN22, TPG16]. Lossy [TPM20]. Low [CDP19, GGJ^{+}22, HSY20, JHSX11, LHLP20, LLS^{+}20, MDA^{+}23, NNYZ17, OSZ17, PKCS18, SO13, SK23, ZN19]. Low-Frequency [LHLP20]. Low-Light [HSY20]. Low-Photon [MDA^{+}23]. Low-Rank [CDP19, GGJ^{+}22, LLS^{+}20, NNYZ17, PKCS18, SK23, ZN19]. Lower [WGGX22]. Luminance [PAB^{+}15]. Luminance-Chrominance [PAB^{+}15]. Lung [SKJ^{+}19]. LUTs [TPM20].

MacAdam [CF18]. Macrostructure [PH20]. MAGMA [HPZ16]. Magnetic [BCC^{+}16, DPVW14, DHP19, DGMW23, GTP^{+}23, HCCS20, LCS^{+}16, QS15, RB15, SAS17, WE17, GH18]. Magneto [DGMW23]. Magneto-Acoustic [DGMW23]. Magnetoacoustic [QS15]. Majorization [GM18]. Majorize [CJPT13]. Majorize-Minimize [CJPT13]. Malik [GKL13]. Manhattan [BAA14]. Manifold [BT18, BHSW18, CDA21, EHL17, GF22, HH18, LNPS17, NPS18, OSZ17, SDL22, ST19, SYB22, TD17, WDS14]. Manifold-Structured [SDL22]. Manifold-Valued [BT18, BHSW18, CDA21, LNPS17, NPS18, WDS14]. Manifolds [AGSW16, BGK15, BPS16, CC14, DL21, DLÖS23, ENR20, FAS^{+}15, MMM12, SHS10, YLLY19, YLLY20]. Many [BH17]. Map [DLÖS23, LLWG13, LLYG14]. Mapped [CBB14]. Mapping [AKZ13, BCD19, CPW^{+}14, DL14, LPP^{+}09, LLLX17, LLYG14, Nat16, PUW17, Pal16]. Mapping-Adaptive [LLLX17]. Mappings [Lip14, MCL16, Yue23, ZCZL22]. Maps [AFGK23, CR18, DNB21, LL14, SJD^{+}15, WkZ14]. Marching [BGPS17, BAA14]. Markov [BLM^{+}22, CCFBY13, DMP18, PMZ20, PCP^{+}16, WCA^{+}18, YL21]. Mask [BR15]. Mass [CLC13, HLL^{+}23, KR17, Lan19]. Match [CLC13]. Matched [FGPT17]. Matched-Filter [FGPT17]. Matches [SMA11]. Matching [BJM15, BF15, CKL17, CLC13, FH15, Far19, KGB15, LLLX17, LTW^{+}10, LLYG14, NT11, RDG09, RG16, SBC22, SNB13, WC23]. Material [DL14, MQLC16, NS14]. Materials [HSSP09, PH20]. Mathematical [AN20, AKLS17, De 23, GT15, HSR^{+}23, LWY16, VFPA22, WLYU15]. Matrices [CV13, CKA17, GPB17, TD17]. Matrix [CESV13, CN22, GGJ^{+}22, HZDZ23, JHSX11, KKS15, LSWW22, LWM^{+}18, PN23, PKCS18, TV20, VHO20, WT13, WC23]. Max [PYA^{+}12]. Max-Flow [PYA^{+}12]. Maximally [PRTW21]. Maximization [GAT22, JLN14, LLSZ09, YLH23]. Maximum [CGTN11, DDPV20, LWM^{+}18, MÁS^{+}22, Per17, Per19, VDPD20]. Maximum-a-Posteriori [Per17, Per19]. MaxPol [HP17]. MBO [MKB13]. Mean [CFSS16, DPC13, FK10, NPJI17, ZC12, MGKR15]. Means [ACSW12, DATP17, DAG11, HP15, JGKL17, LSC^{+}18]. Measure [CLMT15, DLÖS23, FR14, GSC12, Ish14, LVEB09]. Measure-Based [DLÖS23]. Measure-Theoretic [FR14]. Measure-Valued [LVEB09]. Measured [GHM23b]. Measurement [AGM14, BK18, KKN^{+}18, LLW13]. Measurements [AGH14, ACS21, AH17, CFdGK09, CCBB14, FST23, GS17, HSNS18, IVW16, KBW13, MNPT17, PS19, PV14, QYZ19, TBKF15, WHY^{+}15]. Measures [AZ13, LdGKW19]. Measuring [SG15]. Media [AGO21, ALZ20, BL14a, BK17, BG20, BG21, FSY09, GS10, YJL^{+}17]. Median [LR16]. Medical [BCMO08, CCMS13, CFM^{+}20, Dem09, ZYZL20]. Medium [AGM14, FST20, GP09]. Meets [DMP18, LDA^{+}22]. Mellin [BGV09]. Mesh [BGH^{+}21, FG23, Lip14, ZHW22]. Meshes [BBH^{+}23, CFM^{+}20, Lip14]. Meshing [CHL16]. Message [PLCD20, Sap08]. Messages [HSÅS18]. Metal [ZDL18]. Metamaterials [AD23]. Metamorphosis [ERS18, ENR20]. Method [AL15, AH17, AHL22, ABR10, BBC11, BHFPG21, CCZ13, CLL11, CYY11, CTY13, CYZ14, CEM19, CWR19, CZ10, CH16, CDH16, CHZ21, CJK10, CJ12, CPRS21, DGT19, DPC13, DL21, DLLY17, DHN09, Dro14, DMZ18, ELX13, FPT20, FKLS12, GHFT23, GHM23a, GLQ15, GS17, HRSZ16, HWC21, HLST15, HNW09, HW22, Kla11, KKN^{+}18, KKN19, KLNY23, KLN^{+}23, LY12, LR17, LR18, LN13, LP19, LZ18, LNZS10, LY15, LDS20, LLW23, LLS^{+}13, LZD^{+}16, LZ17b, LZZ18, LLS19, LSW14, LQS14, MSMC15, NLH^{+}16, OCLP15, Pal16, PMZ20, RNH19, RGZ13, SKJ^{+}19, SO08, THC11, WN13b, WG22, WT10, WE17, XY13, XZZ19, YPC17, YL21, YK16, Yin10, ZN19, ZHW22, ZZ19, ZZ21, GWY09, GO09]. Methodology [VDPD20]. Methods [AKM11, ACSW12, BGH18, CDRS16, CDSV18, CTWY15, Dar15, DGJS16, EST20, FW14, GCN21, GEB15, Gol11, GOSB14, GM10, HPZ11, HN17, HL13, JKSV20, JLZ19a, KGC11, LKR18, LPT20b, LLLX17, LGCWY18, Mar09, MRSS08, RDSK09, TV20, WC23, WT10, XXYC22, ZCO18, dSO22, BS09]. Metric [HQ19, KN14, MD15, SMSY11]. Metrics [BHM12, BBHMA17, FAS^{+}15, KvD18, NPV16, NK20a, RR15, RG16]. Micro [CDRS16, PH20]. Micro/Macrostructure [PH20]. Microlocal [WQ20, WQ21]. Micromechanical [PH20]. Microscopy [DLÖS23, RS20, SZSH11, SEMS19, dSO22]. Migration [AdHW15]. Minimal [CCP12, KLS^{+}17, KSS19]. Minimax [ACSW12]. Minimization [ADX21, CZ10, Che14, CW18, CvG10, Con17, DTL^{+}21, GH23, GCN21, GM18, GSXH18, GGJ^{+}22, HPZ16, HLL^{+}23, HCGN22, HCGN23, KHD^{+}15, KLYY21, LLC14, LLS^{+}20, LQZ23, MLH17, NW13a, NNZC08, OGL15, PS16, SX12, WYYZ08, WTNL21, YLLX20, YOGD08, YY22, Yue23, ZN19]. Minimize [CJPT13]. Minimizer [Nik13]. Minimizers [Nik13]. Minimizing [BPS16]. Minimum [Ish14, WMT^{+}09, vGPR22]. Mining [JGM^{+}12]. Mirror [GFB^{+}23, HPZ16]. Mismatch [CCPS23]. Missing [BBE^{+}21, CJ14, DJLS20]. Mixed [AAD^{+}08, ADK15, BGH18, BGH^{+}21, CDS17, CWR19, CCFBY13, DD13, HL13, Yan13]. Mixed-Norm [CWR19]. Mixed-State [CCFBY13]. Mixing [XFPA14]. Mixture [BP14, DPN18, DD20, GGJ^{+}22, HBD18, WM13]. Mixtures [KGD21]. Mobile [ABG^{+}13b]. Modality [Rig17]. Mode [VFPA22, YY13]. Model [ADK15, BHB21, BP18, BDM^{+}20, BS15, BP14, BH12, BDS18, CCZ13, CTWY15, CGÖ19, CW22, DL18a, DAB^{+}20, DPN18, DGT19, DZ13, DMTZ16, DMZ18, DBCS14, ERS18, ENR20, FW14, FQC16, GTP^{+}23, GLR18, GT15, GDF15, GNH^{+}22, GF22, GGJ^{+}22, HW20, HWC21, HDH16, HSY20, JZMN21, KYW13, LP19, LLC14, LLS^{+}20, LTKG21, LQZ23, LZOX15, LM13, MRM20, MF13, NW11, OSZ17, OGL15, PAB^{+}15, PCP^{+}16, SO08, SBC22, SX12, SNB13, THC11, WZYX13, WH15, WZLH20, WYN22, WGGX22, ZC15, ZDL18, ZWN14]. Model-Based [LLC14]. Model-Centric [GF22]. Modeling [AE08, ABG13a, CCFBY13, MAP11, SY14]. Modelling [ES15]. Models [CT17, CTY13, CK09, DM20, DD20, DL18b, DHP19, Dro14, DMSC16, FQXC17, FFA11, GDF15, HW13, HVW15, HBD18, HSÅS18, KSW20, KL18a, KP13, LY13, LAZ^{+}18, LTKG23, MLH17, NFV22, NTDB19, NPJI17, NL10, Per19, PKPE21, PCBC10, RBB20, SXS^{+}15, SCGAF^{+}15, SN11, TM16b, WT10, XFPA14, ZWJ19, ZC20, ZE23].

Modes [CDH^{+}21, CALG21]. Modification [CFdGK09]. Modified [FPT20]. Modular [GDT18]. Modulation [KV23]. Modulo [BBK22]. Modulus [DGJS16]. Molecule [dSO22]. Molecules [RK19, RS20, VHO20]. Moments [DPC13]. Momentum [SNDP13, WNS^{+}22]. Monge [MSKL09, STV09]. monochromatic [PJS21]. Monogenic [Sto11]. Monotone [PRTW21]. Monte [DMP18, PMZ20, dSO22]. Moreau [DMP18]. Morphing [NPS18, RGLB14, TAF^{+}20]. Morphological [Ang13, PFS10]. Morphology [AN20, PH14, VFPA22]. Most [KSPR17]. Motion [BCP13b, BMW09, BDS18, CGÖ19, CN17, CCFBY13, EKOÅ10, FA09, FH11, GOF16, HLKH14, KSZ11, Lan19, LPT20b, LHB^{+}18, ÖSB15, PM08, SM18, SXS^{+}15, SNM17]. Motion-Flow [SXS^{+}15]. Motions [BHS22]. Motor [CHM13]. Mouse [DL14]. Movement [DGH11]. Moving [BGK15, BGP^{+}17, CB11, DKP09, FGPT17, GOF16, GHM23b, KKN20, WY12, WY14]. Moving-Target [CB11]. MR [CHH^{+}12, KT14, LKW^{+}19, YWW^{+}23]. MR-Perfusion [LKW^{+}19]. MREIT [GH18, SSL23]. MRI [AAG23, CCW20, CBZ18, EB16, ET18, HR15, SHVC19]. Muller [NDM^{+}11]. Multi [BS21a, DM20, WZ17]. Multi-Domain [WZ17]. Multi-scale [BS21a]. Multi-time [DM20]. Multiatlas [GZC^{+}15]. Multichannel [JBS17, MB10, QLZ20, WZYX13, YYZW09]. Multiclass [LS11]. Multicontrast [EB16]. Multiconvex [XY13]. Multidimensional [BMW09]. Multienergy [LSC^{+}18]. Multifractal [WCA^{+}18]. Multifrequency [AAJ^{+}16, GH15, GS17, GHM23b, MNPT17]. Multigrid [BLC10, FD20, NTDB19]. Multilabel [PYAC13]. Multilevel [HPZ16, HPPZ19, KGC11]. Multimarginal [SK23]. Multimodal [CLY19, EST20]. Multiphase [BFG19, BPLX21, LNZS10, LQZ23, TZS13]. Multiple [ATTY16, ARF16, BK18, BG15, DVFP14, FSY09, GH15, LR17, LR18, LLW13, NTDB19, RVCB19, SC10, WHY^{+}15]. Multiple-Secret [SC10]. Multiplexing [SXS^{+}15]. Multiplicative [CYZ14, DZ13, HNW09, KYW13, LNS10, LLS^{+}20, SO08, YGS^{+}19, ZWN14]. Multipliers [CTY13, CEM19, OCLP15, YPC17]. Multiply [ZCL22]. Multiresolution [MRSS08, NLH^{+}16]. Multiscale [AdHW15, BF15, CC14, DLV23, FPM17, FAS^{+}15, FQXC17, HWZ22, HLLS14, JGM^{+}12, KRW10, LZ17a, LHW^{+}15, LLW13, LRV21, Sto11, ZH21, ZvDT^{+}17]. Multispectral [GB11]. Multistatic [AGK^{+}12]. Multitaper [AR20]. Multivalued [StTBRV12]. Multivariate [LS19, WCA^{+}18]. Multiview [PV14, YLLX20]. Multiwindow [BF15]. Mumford [BFG19, BPLX21, CCZ13, HP11, KSW20, Kla11, KR13, Mah12]. Myriad [LS19].

Nanostructures [CDRS16]. Narrowband [WY14]. Natural [FQC16, GSC12]. Near [BL14a, DLLY17, FA09, KGD21, MWBK14, TT22, XZZ19]. Near-Field [BL14a, DLLY17, FA09, XZZ19]. Near-Quadrature [TT22]. Near-Separable [KGD21]. Neighborhood [SSN09]. NESTA [BBC11]. Nested [CPP09]. Nesterov [HHJ^{+}23]. Network [FGPT17, KK17, YWW^{+}23]. Networks [ALKÖP19, BGL^{+}21, BHFPG21, HPZ22, LHC^{+}23, SDL22, Wan16a, YZL^{+}18, ZJ21]. Neumann [DH20, QSUZ11]. Neural [ALKÖP19, BGL^{+}21, BHFPG21, HPZ22, Wan16a, YZL^{+}18, ZJ21]. Neuroanatomical [DL14, RGLB14]. Newton [BS09, CWR19, CJK10, CJ12, DL21, Her19, JKSV20, MB15]. Newton-type [BS09]. NMF [RBLS14]. NMR [SX12]. Noise [CDS17, CYZ14, CLDM18, CJK10, CJ12, DD13, DZ13, FW14, FQC16, GS10, GPST13, GPST15, GPB17, HNW09, JGKL17, KHD^{+}15, KGV14, LNS10, LLS^{+}20, LCD22, MYZ13, SDZ15, SO08, SG15, SDA15, WT13, WGL^{+}22, Yan13, YGS^{+}19, ZWJ19, ZBN17, ZWN14]. Noises [AGM14]. Noisy [AS18, GP09, HSS21, LR16, SW13, Wan16b, XZC^{+}12, ZN19, BGV09]. Non [BAA14, BC15, CT17, WGGX22, ZWJ19, ZTO15]. Non-Lambertian [CT17]. Non-Lipschitz [BC15, WGGX22, ZWJ19]. Non-Local [ZTO15]. Non-Manhattan [BAA14]. Nonadditive [HL13]. Nonasymptotic [LKR18]. Nonconvex [CZ10, CW18, CLYZ21, DTL^{+}21, GM18, HW13, HVW15, KLS^{+}17, LMSS19, NNZC08, OCBP14, ODBP15, PYAC13, PKCS18, PS16, WCN^{+}19, WGL^{+}22, XXYC22, YPC17, YLLX20, ZBN17]. Nonconvex-TV [ZBN17]. Nondiffusing [FST20]. Nonhomogeneous [ZC15]. Noninvasive [CDRS16]. Nonlinear [AC09, AGM14, BHFPG21, CZ10, CFM15, CJ12, CW22, DMTZ16, DMZ18, FQXC17, GLS20, HSNS18, Mar09, OGL15, RW09, SHB^{+}18, SKJ^{+}19, SO08, ZvDT^{+}17]. Nonlocal [ACSW12, CJ14, CBB14, DL18a, DBCS14, DAG11, EEF23, HFE19, JGKL17, JPC12, LNPS17, LBM13, LPSS15, LSC^{+}18, LZ17b, LLS^{+}20, RTW20, SSN09, SBS23, YGLD17, ZBBO10]. Nonnegative [ELX13, Gil14b, NK20b, SX12, TV20, XY13]. Nonnegatively [DGJS16]. Nonoriented [CT13]. Nonoverlapping [LP19]. Nonparametric [PCP^{+}16, SHS10, SDA15]. Nonrigid [AKR13, LZ17a]. Nonseparable [LMSS19]. Nonsmooth [CZ10, CLYZ21, DTL^{+}21, HL13, KLS^{+}17, NNZC08, ODBP15, PS16, YPC17]. Nonstationary [YY15]. Nonuniform [AGH14]. Nonuniformity [KT14]. Norm [CWR19, CM20, GSXH18, KGV14, KR13, LY13, LZZ^{+}23, WLJ22, ZN19, Nik13]. Normal [RG16, YK16]. Normalized [WZLH20]. Norms [CY09]. Note [Wan16a]. Novel [CV13, GSZ17, HSR^{+}23, LL22, LWY16, TV20, YLLY19, ZHW22]. Nuclear [GSXH18, HCCS20, LY13, LZZ^{+}23, ZN19]. Nuclear-Norm [LY13]. Nullity [BKBD16]. Nullspace [BH17]. Numerical [BLM14, BBHMA17, BCD^{+}12, CFdGK09, CNS10, Dro14, FN17, FST20, FST23, GKQR20, Hub13, KKN^{+}18, KLNY23, KLN^{+}23, Lan19, Mon14, YCU19, ZC15]. Numerics [BH15b].

Object [AAD^{+}08, LL22, SY14, SZW14]. Objective [TG21]. Objects [ARYZ18a, ARYZ18b, BGP^{+}17, DKP09, FSY10, FGPT17, RW13, SMSY11, TBKF15, WZ17]. Observation [GHM23b]. Observer [NTV10]. Obstacle [BG15, DLL19, GLS18]. Occluded [NT11]. Occlusions [TCH08]. Ocean [Liu21]. Off [CCW20, CDP19, LBFA23, OJ16]. Off-the-Grid [CCW20, CDP19, LBFA23, OJ16]. One [AGP18, BGM^{+}16, CvG10, EKOÅ10, FW10, GHM23b, Han12, KGV14, LR17, LLS19, YLLY20]. One-Dimensional [EKOÅ10, KGV14]. One-Homogeneous [AGP18, BGM^{+}16]. One-Way [FW10]. Online [LGCWY18, NTV10, ShDC^{+}19]. Only [MNPT17, NMP15]. onto [CWR19]. Open [CR18]. Operations [TM18]. Operator [ACN16, DGT19, GLQ15, LHT^{+}21, MPM^{+}17, MHM23, OV14]. Operators [ACN16, BGL^{+}21, BHFPG21, EW15, PRTW21]. Optical [ASK22, BGK15, BMW09, CJT^{+}12, GSL^{+}22, HSR^{+}23, KGC11, LDCG14, SSSW09]. Optics [KWRC20]. Optimal [ASK22, ATTY16, BJM15, CCBR13, CLC13, FPPA14, GLR18, HLL^{+}23, JGKL17, KR17, KL18b, LdGKW19, MMT18, NS17, PPO14, SHB^{+}18, SOK^{+}20, SK23, TM16b, ZL21]. Optimality [ABK15, DV22, GSGJ21, SHVC19]. Optimization [AAD^{+}08, BC15, BK15, BK17, BCSB18, CDSV18, CMP14, CCMY15, DLW16, DTL^{+}21, EKOÅ10, EZC10, FD20, GM18, GOSB14, KGC11, KLS^{+}17, KP13, LWM^{+}18, OCBP14, ODBP15, PC21, RL15, RPW19, RM10, RZ15, RYZ18, RB18, Sap10, SYB22, WCN^{+}19, WE17, XY13, YSB20, ZWN14, ZZ21]. Optimization-Based [DLW16]. Optimization-Free [BCSB18]. Optimized [LTW^{+}10]. Oracle [ACSW12]. Orbiter} [GPPM15]. Order [ABSM20, BHV12, BBC11, BLSW14, BDM^{+}20, BLC10, Col22, DHZ21, Dro14, DB10, DMZ18, EZC10, FPT20, FQC16, HW20, JDA^{+}19, JK15, KSW20, LNPS17, LGCWY18, PMS20, PLMS20, PAM12, QYW10, SBS23, SNDP13, SRG10, WT10, ZC15, ZHW22]. Ordering [VZE16]. Organization [KZ18]. Orientation [CDH16, DLÖS23, HSSP09, WSW13, XXQJ20, ZC20]. Orientation-Preserving [ZC20]. Orthogonal [HZDZ23]. Orthonormal [Bat10]. Oscillation [GB18]. Oscillatory [HKLM21]. Osher [CTWY15, LP19, NPJI17]. Other [Car10]. Out-of-Focus [LEZX14]. Outer [DHSS13]. Outliers [CB18]. Over- [HJS13]. Over-/Underexposed [HJS13]. Overlapping [CTWY15]. Overparameterized [GEB15, RBB20].

Packet [YY13]. Paintings [HBM12, YCF^{+}16]. Pair [MPM^{+}17]. Pairs [BV16, TT22]. Pansharpening [DBCS14]. Paper [EKV23]. Parabolic [AdHW15, HP15, KSZ12]. Paradigms [BKBD16]. Paradox [TMR13]. Parallel [BPS16, CHH^{+}12, KL19, LCS^{+}16, MAP11, SDL22, YWW^{+}23]. Parallelizable [CLLGL20, ZCL22]. Parameter [BMS23, CLPS19, CJ12, DVFP14, FW14, FFA11, FH11, HSNS18, KP13, LS19, NLH^{+}16, SCGAF^{+}15, dSO22]. Parameterization [CLL15, CHL16, CLLGL20, ZCL22]. Parameterizations [KLYY21, YLLY19, YLLY20]. Parameterized [Bat23, Her19, MPL^{+}18]. Parameters [BdHKU22, CJK10, DDPV20, LNS10, NS14, VDPD20]. Parametric [AKM11, BG14, CPRS21, EST20, FF13, UC13]. Parametrization [KO16]. Paring [FK10]. Part [ARYZ18b, BH15a, BH15b, DDPV20, GHM23b, LR17, LR18, LHW^{+}15, VDPD20, WLYU15]. Partial [CLLGL20, DM20, FST23, GSL^{+}22, Hub13, SCL20, SBC22, ZCL22]. Partially [CHH^{+}12, LEZX14, NT11]. Particle [BBL^{+}23, DAW21, NT11]. Partitioning [SW14]. Partitions [CCBR13, CCP12, KLS^{+}17]. Parts [vGPR22]. Passing [PLCD20]. Passive [BGP^{+}17, GP09, GPST13, GP15, LKR18, WY12]. PAT [RGZ13]. Patch [AH23, BEFL21, DDGL19, DPN18, DD13, GLR18, GGJ^{+}22, NK20b, SO13, STA22, VZE16, XXQJ20, YMA22]. Patch-Based [BEFL21, DD13, STA22, YMA22]. Patch-Rank [SO13]. Patches [AC09, KZ14, TM12, YGLD17]. Path [BBES21, FG23]. Paths [BER15]. Pattern [FL12, HLLS14]. Pattern-Guided [FL12]. Patterned [NNYZ17]. Patterns [JLZ19b]. PCA [BGL13, VSU15]. PCM [GSZ17]. PDE [AB10, BS21b, BDMS15, Her19, SBS23, SYB22, WG22, YSB20]. PDE-Based [BS21b, WG22]. PDE-Constrained [Her19]. PDEs [DHZ21, LPP^{+}09]. Peaceman [CDH16, LY15]. Penalization [GLQ15]. Penalization-Regularization-Operator [GLQ15]. Penalty [GHFT23, HWC21, HR15, SBFA15, SBFA16, XWH22]. Perception [BPP22]. Perfect [CLC13].

Performance [ADGM14, GM15, KKN19, KBW13, LKR18, TM16a]. Perfusion [LKW^{+}19]. Perimeter [KR13, MÁS^{+}22]. Perimeters [EEF23]. Periodicity [TP18]. Perona [GKL13]. Perot [Aco19]. Perspective [BCD^{+}12, GZC^{+}15, GKQR20, HY12, LLLX17, MTWB14]. PET [CBZ18, CK09, CPRS21]. Peterson [FN17]. Petersson [KN14]. Pharmacokinetic [CK09]. Phase [ABFM14, AdHW15, BQ22, CESV13, CLDM18, CMLZ18, CEM19, CH16, ELB18, FZ20, FD20, GFB^{+}23, IVW16, JLZ19a, KWRC20, LY18, NTDB19, PS19, Sto11, YY22, ZZ19]. Phase-Field [NTDB19]. Phase-Space [LY18]. Phased [DLL19]. Phaseless [DLL19, HW22, JLZ19a, JLZ19b, KKN^{+}18, XZZ19]. Phases [MNP16]. Photoacoustic [Aco19, ACS21, ABSM20, CN17, ES15, FRV18, HN17, HHMT16, Kow14, LHB^{+}18, MPL^{+}18, MPM^{+}17, NS14, NK16, NHKD22, PB23, QSUZ11, RZ13, ST23, ZH21]. Photographing [HBM12]. Photographs [DAMM12, HJS13]. Photomask [CJT^{+}12]. Photometric [MF13, MTWB14, MWBK14, MQLC16, SNB13]. Photon [KRW10, MDA^{+}23, TAR^{+}19]. Photon-Limited [KRW10]. Photorealistic [CFBP23]. Physics [DHP19]. Physics-Based [DHP19]. Picking [BBL^{+}23]. Pictorial [KvD12]. Piecewise [AGSW16, BBH^{+}23, BdHKU22, CDA21, LO17, NS14, NNZC08, OJ16]. Piecewise-Bézier [AGSW16]. Piecewise-Geodesic [CDA21]. Pipeline [RVCB19]. Pixel [Bat23, GM15, LUZZ22, LAZ^{+}18]. Pixel-Level [LAZ^{+}18]. Planar [CHKL23, HSH13]. Plane [CAT08, NK20a]. Planning [FG23]. Plasmon [CDH^{+}21]. Plasmonic [ARYZ18a, ARYZ18b]. Play [BCSB18, FLG23, LDA^{+}22, LG23]. Plug [BCSB18, FLG23, LDA^{+}22, LG23]. Plug-and-Play [BCSB18, FLG23, LG23].

PnP [FLG23]. PnP-ReG [FLG23]. Point [BG21, CLC13, CHL16, CEM21, DPH^{+}13, DAMM12, FSY09, FZ20, GHM23b, HY12, HKL20, HMXY22, HHK^{+}18, JLZ19b, KKN20, LZ17a, LDG21, LBFA23, MWBK14, MCL16, MMT18, NS17, SZW14, WCN^{+}19, WLJ22, WC23, ZZ21]. Point-Set [CLC13]. Point-to-Subspace [SZW14]. Points [GBFA10, GBFA12, RTH21, SNB13, TMSP20]. Poisson [CYZ14, CLDM18, CJPT15, FQC16, GTU14, KHD^{+}15, WHY^{+}15, ZYZL20]. Poissonian [Che14]. Polarizable [CV17]. Polarization [ABFM14, ABG^{+}13b, APST19, PN23]. Polychromatic [FM23]. Polyenergetic [CNS10]. Polynomial [SCGAF^{+}15]. Population [HHJ^{+}23]. Population-Kinematic [HHJ^{+}23]. Pose [DSYT10, GSGJ21, WLL^{+}21]. Posed [GHFT23, KR13, RNH19, SKJ^{+}19]. Posedness [Aco19, CT17]. Positive [BHV12, CDHS13, CKA17, MC16, QYW10]. Possible [AKZ13]. Posterior [GAT22, LM13]. Posteriori [BGH^{+}21, Per17, Per19]. Postreconstructed [Dem09]. Postregistered [Dem09]. Potential [GLS18, NL10]. Potentials [JDA^{+}19]. Power [AH23, BV16, BHFPG21, CFdGK09, Naj17]. Practical [FAS^{+}15]. Preconditioned [CMLZ18, LG23]. Preconditioning [GM15, RL15]. Prediction [SMSY11]. Presence [AGM14, CCPS23]. Preserving [BCMO08, DB10, GQY14, Seg22, SY14, YYZW09, YLLY19, YLLY20, Yue23, ZC20]. Pressure [GTP^{+}23, MPL^{+}18]. Primal [BGH18, BGH^{+}21, CCPS23, Col22, DHN09, EZC10, Gol11, HY12, HYY14, HMXY22, MSMC15, OV14, ZZ21]. Primal-Dual [CCPS23, Col22, DHN09, EZC10, Gol11, HY12, HYY14, HMXY22, MSMC15, OV14]. Principal [BCP13b, GH23, HPPZ19, LS17]. Principle [FH11, vGPR22]. Prior [BHB21, DSYT10, FQC16, GAT22, HLKH14, LHC^{+}23, LZD^{+}16, MYZ13, PKPE21, SCL20, TCH08]. Priori [FGS12, AM16]. Priors [AH23, Bat23, DPN18, HHMT16, HPZ22, LDA^{+}22, LGL^{+}22, SS13, WCA^{+}18]. PRO [CEM21]. Probabilistic [NTV10, NPJI17, TM16b]. Probing [ST11]. Problem [AAJ^{+}16, BGPS17, BFG19, CDLZN23, CPW^{+}14, CHKL23, CHM13, DL14, GSGJ21, HY12, KSZ11, KKS15, KKN20, KKN^{+}18, KKN19, KLNY23, KLN^{+}23, LSYZ21, LLW23, LQS14, ST23, SBFA15, WQ20, MSKL09, SBFA16]. Problems [AB10, AKM11, AHL22, BGH18, BT09, BHS23, BMS23, BGL^{+}21, CPP09, CLC13, CJ12, Col22, CP16, DDPV20, DV22, De 23, DMTZ16, ELB18, EKOÅ10, ELX13, FR14, GHFT23, GEB15, GAT22, GS17, HFE19, HMXY22, HHK^{+}18, HL13, HP17, HQ19, HSNS18, JK23, KR17, KR13, KMDL19, LUZZ22, LR17, LR18, LH18, LDS20, LS18b, LLS19, LSW14, LHT^{+}21, MDA^{+}23, PYAC13, PS16, RNH19, RPW19, RYZ18, SKJ^{+}19, SOK^{+}20, SCC14, SK23, TS14, VZE16, VDPD20, WZYX13, WC23, XWH22, XXYC22, YPC17, YHC18, GO09]. Process [GPPM15]. Processes [LDG21]. Processing [AD23, BHS22, CCQY20, CFSS16, DPH^{+}13, EEF23, ETT15, FW14, Gol11, HWC21, HL13, LMSS19, LDG21, Lip14, LZOX15, MKB13, OSZ17, VZE16, WZYX13, Zhu16, BS09]. Procrustes [KvD18]. Product [HZ14, HMZZ19, TM18]. Profiles [KGV14]. Programming [CDHS13, KLS^{+}17, LY15, ÖSB15, SS11]. Programs [De 23]. Projected [HHJ^{+}23, TG21]. Projection [ACN16, CCR^{+}12, CWR19, CJT^{+}12, CEM21, DPZ20, DMTZ16, Gil14b, HSH13, KBW13, LLC14, TG21]. Projection-Based [DMTZ16]. Projections [AS18, BD22, SW13]. Promoting [CMP14]. Proof [LDCG14]. Propagation [YMA22].

Properties [KR13, LM13, WCU13].

Property [BH17, GCN21, HR15]. Provable [GFB^{+}23, WLL^{+}21]. Provably [KGD21, PKCS18]. Proximal [BHFPG21, CMLZ18, CCMY15, CG19, DTL^{+}21, DMP18, HLST15, LY15, LLS^{+}20, OCBP14, PPO14, PMZ20, PS16, STY11, TV20]. Pseudo [CDH^{+}21, PJS21]. Pseudo-monochromatic [PJS21]. Pseudodifferential [BGL^{+}21]. Ptychographic [CEM19, FD20, HLST15]. Ptychography [FZ20, FM23]. Pure [CN22, GM15]. Pure-Pixel [GM15]. Pursuit [HPPZ19, LO17]. Puzzles [HLW20, KGB15].

Quadratic

[CDHS13, KGD21, RZ15, RYZ18].

Quadrature [TT22]. Qualitative [LRV21]. Quantification [RPW19, TP18, ZYZL20]. Quantitative [AAD^{+}08, ABSM20, CPW^{+}14, DHP19, FRV18, HHMT16, KT22, LRV21, NS14, Nat16, PUW17, Pal16, RGZ13, RZ13]. Quantized [YZL^{+}18]. Quantum [BPP22]. Quasi [LL14, LLBS14, QLL19, TP18, Wol09, WkZ14, ZCZL22, ZCL22]. Quasi-Conformal [QLL19, WkZ14, ZCZL22, ZCL22, LL14]. Quasi-Isometric [Wol09]. Quasi-Variational [LLBS14]. Quasilinear [DHZ21]. Quaternion [CN22, PN23]. Query [SZW14]. Quotient [DATP17, HH18, TD17].

Rachford

[ACL16, BPS16, CDH16, FZ20, LY15]. Radar [AC12, BCP13b, BK18, CB11, DFM^{+}12, KT22, Voc15, WY14]. Radial [MHM23]. Radiative [FST20, FST23, Hub13, KLNY23, KLN^{+}23]. Radon [ACN16, ADB^{+}21, BBK22, CHZ21, GR23, GKQR20, HQ16, Hal11, MH17, RLL14, WQ21, ZD16, ZDL18]. Random [AC12, BR15, BLM^{+}22, BK17, BG21, BD22, CCFBY13, FH11, GHM23a, GPB17, LLW23, NL10, PCP^{+}16, Rom09, SC10, SW13, TM12, WCA^{+}18, YL21]. Randomized [HW22, LZ18, SDR20, TM18]. Randomly [GS10]. Range [AC09, TA14]. Rank [CDP19, GH23, GGJ^{+}22, JHSX11, LLS^{+}20, MLH17, NNYZ17, PKCS18, SO13, SK23, YLLX20, ZN19]. Ranking [ROD15]. Rapid [BGPS17]. Rate [CTWY15, TG21]. Rates [ACSW12]. Ratio [GPST13, GPST15]. Rational [Han12, KSZ12]. Raw [SG15]. Ray [AAD^{+}08, DLW16, LQS14, Mon14, PJS21, StTBRV12, YCU19, YCF^{+}16, ZE23, BFJQ18, LUZZ22]. Ray-Tracing [StTBRV12]. rays [GPPM15]. Reaction [SZGW18]. Real [BH17, BHS23]. Real-World [BH17]. Realization [Zhu16]. Receivers [FGPT17]. Reciprocity [FW10]. Reclaiming [Mil18]. Recognition [DPZ20, HKBH13, SZW14, VSU15]. Recomposition [FPM17]. Reconstructing [ARYZ18a, ARYZ18b]. Reconstruction [AAD^{+}08, AR13, AKL^{+}21, APST19, AL15, BLM14, BGPS17, BQ22, BBL^{+}23, BP18, BFJQ18, BH15a, BH15b, BDM17, BvGL^{+}23, BDS18, BCSB18, CCW20, CFdGK09, CJ14, CHH^{+}12, CGÖ19, CLY19, CLYZ21, CBZ18, CNS10, CN17, DHSS13, DGMW23, EB16, ET18, EST20, ESS16, FGS12, FRV18, FST20, FST23, GTP^{+}23, Gri10, GY12, GSZ17, HSSP09, HF12, HKL20, HHK^{+}18, HHJ^{+}23, HK14, HSH13, HCGN22, JLZ19b, KWRC20, KL19, KL18a, KRW10, KT16, LBFA23, LCS^{+}16, LSC^{+}18, Lou08, MPL^{+}18, Nat16, NW13a, NNZC08, PLCD20, PS11, PKPE21, PV14, QYZ19, RGZ13, RLL14, Rig17, RZ15, RS20, RB18, RKO22, SNM17, TAR^{+}19, TPM20, WYYZ08, WY10, WTNL21, YJL^{+}17, ZD16, ZBBO10, ZDL18, ZYZL20]. Reconstruction-Segmentation [BvGL^{+}23]. Reconstructions [AGH14, BI15]. Recovering [ALZ20, BBE^{+}21, HP15, HJS13]. Recovery [ADX21, BH17, BBC11, BdHKU22, CPP09, CW18, CQ21, CG19, DHK20, DJLS20, FM23, GB18, GOF16, HPZ11, Hub13, HCGN23, LSZ18, LQS14, MLH17, OJ16, PC21, PRTW21, PBU^{+}22, QLZ20, WHY^{+}15, WLL^{+}21, YJL^{+}17, ZJ21]. Recurrence [XXQJ20]. Recursive [BGG17]. RED [CEM21, REM17]. RED-PRO [CEM21]. Reduced [BDM^{+}20, DMZ18, GH18, ZZ21]. Reduces [ACS21]. Reducing [DHK20, PUW17]. Reduction [DMTZ16, FQC16, Pal16, SX12, ZDL18]. Redundancy [DDGL19]. Redundant [AE08]. Reed [NDM^{+}11]. Reference [JLZ19b]. Refinement [BGH^{+}21, CLK14, DLÖS23, DNB21]. Refinements [LRV21]. Refitting [DPSV17]. Reflection [ZZ19]. Reflectivities [BMPT16]. Reflectors [BPG08, TMP13, TMP18]. Reformulation [De 23]. Refoundation [BPP22]. Refractions [PS11]. ReG [FLG23]. Regime [ARYZ18b]. Region [CPRS21, DSYT10, HW13, LNZS10, MÁS^{+}22, NT11, WLJ22]. Region-Based [DSYT10, NT11]. Regions [HF12, HJS13, Per17, SDA15]. Registration [AMY16, ATTY16, BGH18, BGH^{+}21, CT13, CÖ18, DGSL23, DL18a, DAB^{+}20, DL14, DNB21, FF13, GNH^{+}22, HW20, HWZ22, KSPR17, LZ17a, LL14, LGL^{+}22, LTW^{+}10, LW14, LLYG14, MB15, MB16, MS17, OGL15, Sdi13, SNDP13, SBC22, VF13, VF14, WG22, ZC20, ZL21, MSKL09]. Regression [LKW^{+}19, PAM12, SHS10]. Regularization [AGO21, All09, ACDG18, BBJ^{+}18, Bat10, BB14, CCW20, CLPS19, CBZ18, CJPT13, CJK10, CEM21, CGN^{+}13, DDPV20, DV22, DJLS20, DB10, EGvL^{+}18, GLQ15, GQY14, HW20, HW13, JK15, KR13, LMSS19, LBFA23, LNS10, LSWW22, LTKG21, LTKG23, MB16, PCBC10, PCCP19, RLS18, REM17, SYO15, SRG10, SCC14, TA14, VZE16, VDPD20, WZ17, WZLH20, WN21, WLJ22, WDS14, YWW^{+}23, YGLD17, ZD16, ZBBO10, All08]. Regularization-Based [HW13]. Regularizations [RBLS14]. Regularized [Che14, CQ21, CPRS21, FPPA14, GHFT23, HKL20, HQ19, MRM20, NLH^{+}16, Nik13, PAM12, PPE^{+}09, SBFA15, SBFA16, TMSP20, WGGX22, XY13, GO09]. Regularizing [DLÖS23, FLG23, HR15, KR17]. Reinterpretation [GM10]. Related [AL15, ACSW12, BHS23, LUZZ22, SCM^{+}12]. Relative [GSGJ21]. Relaxation [BR15, HWC21, NTDB19, PYAC13, SCC14, YZL^{+}18, ZL21]. Relaxations [BLM^{+}22, GSC13, JDA^{+}19]. Relaxed [BPLX21, DL14, JKSV20, KYW13, YK16, ZHW22]. Relaxometry [HCCS20]. Relevance [KHD^{+}15]. Reliable [BF15]. Remote [FSY10, PCP^{+}16]. Removal [BCD19, CDS17, CLDM18, HNW09, JGKL17, LNS10, LLS^{+}20, LCD22, YGS^{+}19, ZWJ19, ZBN17, ZWN14]. Removing [DD13, YCF^{+}16]. Rendition [Mil18]. Repeatability [ROD15]. Representation [ABK15, CT13, DLV23, DPZ20, GSXH18, LL22, LLS^{+}13, LLWG13, MYZ13]. Representation-Based [DPZ20]. Representations [EW15, NNYZ17]. Representatives [MPGMD19]. Represented [ACN16]. Resistivity [KK17]. Resolution [ABG^{+}13c, BK15, BGP^{+}17, BG20, BGG17, Fou10, KWRC20, KT22, MC16, Wah15, WR14, CDH^{+}21, PBU^{+}22, SEMS19]. Resolution-Controlled [WR14]. Resonance [BCC^{+}16, DPVW14, DHP19, GTP^{+}23, GH18, HCCS20, LCS^{+}16, RB15, SAS17, WE17]. Restart [WNS^{+}22]. Restoration [ABR10, BHB21, Bat23, BBFA14, BG14, BC15, BGG19, BDM15, BCP13a, CLPS19, CZ10, CJPT15, DPSV17, DHN09, DGJS16, GSXH18, GKL13, HW13, JHSX11, JNW19, JK15, KGV14, LG23, LLBS14, LZ17b, STY11, SS13, TPG16, WN21, WGL^{+}22, Yan13, YYZW09, YMA22, YY17, ZWJ19, ZC15, ZBN17]. Restoring [DZ13, SDZ15]. Restricted [BH17, GCN21]. Result [LBFA23]. Resulting [DGH11]. Results [BZ18, CCPS23, LRV21, MS22]. Retinex [NW11, WH15, ZTO15]. Retrieval [ABFM14, BBJ^{+}18, BQ22, CESV13, CLDM18, CMLZ18, CEM19, ELB18, FZ20, FD20, GFB^{+}23, HZDZ23, IVW16, JLZ19a, PS19, WT13, YY22]. Reversal [Kow14, NK16]. Reverse [AdHW15]. Reverse-Time-Migration-Type [AdHW15]. Revisited [GTO14, HSMS22]. Revisiting [LAZ^{+}18, Per19]. Revolution [EKV23]. Reweighted [LLS^{+}20, ODBP15, PH14, ZDL18]. Rician [FSY09, LCD22, WGL^{+}22]. Ridge [RK19]. Ridgelet [EHB09]. Ridgelets [GTO14]. Riemann [HW20]. Riemannian [AGSW16, BDMS15, CC14, CFSS16, DL21, DLÖS23, FAS^{+}15, GDT18, KLN^{+}23, LPP^{+}09, MMM12, SHS10, TD17, ZBO14]. Riesz [WCU13]. Rigid [BHS22, HLLS14]. Risk [DVFP14]. Robust [BCP13b, BD22, CHHN21, CN22, CJT^{+}12, CLMT15, ESS16, GH23, Gil14b, JHSX11, KGD21, PS19, PV14, SZW14, WSL13, Wan16b, ZBSZ22]. Robustness [LKR18]. ROF [BPS16, WT10]. ROF-like [BPS16]. Role [GP14, Poo15]. Root [CWR19]. Rotating [LPT21, WCN^{+}19]. Rotation [LZ17a, Rig17, VSU15, WZ17, BGV09]. Rotation-Free [Rig17]. Rotation-Invariant [LZ17a, VSU15]. Rough [BL14a, BZ18, DLLY17, LZZ18, XZZ19]. Rubinstein [HQ19, LLSV14]. Rudin [CTWY15, LP19, NPJI17].

Saddle [HY12, HMXY22]. Saddle-Point [HY12]. Saint [DL14]. Salient [KZ18]. Sampled [ZCO18]. Samples [OJ16, SM16].

Sampling [ADX21, AL15, AH17, BWB14, BCC^{+}16, CCKW14, CHZ21, DLLY17, GHM23a, JLZ19a, LDS20, LLS19, MBFG20, MJC^{+}19, YBZ^{+}21, ZJ21]. SAR [GT15, ST11, Tsy09a, Tsy09b]. Satellite [GKL13]. Satellites [LPT20a, LPT21]. Saturation [JNW19, WN21, WYN22]. Saturation-Value [JNW19, WN21, WYN22]. Scalable [RPW19]. Scale [AKR13, FPM17, FD20, Gil14a, HPZ16, LKW^{+}19, Lin18, Mar09, PWSU16, RR15, SN11, SO08, WDCT09, WE17, BS21a]. Scale-Space [WDCT09]. Scaled [NT11]. Scaling [CPRS21, KSZ12, BGV09]. Scanner [SG22]. Scanning [GSL^{+}22]. Scatterer [JLZ19b]. Scatterers [CMP14, GM10, LLW13, Liu21]. Scattering [AHL22, BZ18, BG20, BDM^{+}20, BG15, BGG17, CH16, CDLZN23, DLL19, FST20, GP09, GP14, GLS20, Han12, HLLS14, JLZ19a, KKN20, KKN19, LR17, LR18, LN13, LDS20, LZZ18, LLS19, QYZ19, Rig17, WT13, WQ21, XZZ19]. Scenes [DKP09, NAF^{+}14, TAR^{+}19, YY15]. Scheduled [WNS^{+}22]. Scheme [FPT20, GQY14, MKB13, MJC^{+}19, RLS18, SLS22]. Schemes [BCC^{+}16, MB16, PPRV22]. Schunck [LDCG14]. Schwarzschild [FG23]. Science [EZC10]. Sciences [Dar15, DM20, Sap10]. Scientific [Car10]. Screened [GTU14]. Search [CWR19, KGC11, GWY09]. Second [BLSW14, DHZ21, KSW20, LNPS17, LGCWY18, ZHW22]. Second-Order [DHZ21, LGCWY18, ZHW22]. Secret [SC10]. Section [Sap10]. Sectional [ES15, MMM12]. Segmentation [BG14, BvGL^{+}23, BPLX21, CCZ13, CCMS13, CCBR13, CYZ14, CFM^{+}20, CCFBY13, DSYT10, DL18a, DAB^{+}20, FPT20, GB11, GZC^{+}15, HRSZ16, LNZS10, LYZ20, LGL^{+}22, LZD^{+}16, LQZ23, NFV22, NTDB19, NPJI17, OGL15, SCL20, TZS13, TCH08, Wan16b, WZLH20, ZvDT^{+}17]. Segmentation/Registration [DL18a, OGL15]. Segmenting [LY12]. Seismic [GS13]. Selection [CLPS19, CCBR13, DVFP14, FW14, GDF15, Lin18, MPGMD19, WLJ22]. Selective [TCH08, TMP13]. Self [FBU15, HLKH14, LVEB09, LS18b, ZZPS20]. Self-Assignment [ZZPS20]. Self-Calibration [LS18b]. Self-Similar [FBU15, HLKH14]. Self-Similarity [LVEB09]. Semi [GLR18]. Semi-Discrete [GLR18]. Semiblind [BCP13a]. Semiconvex [MSMC15]. Semidefinite [BHV12, CDHS13, QYW10, SS11]. Semidiscrete [BCGR14]. Semismooth [CJK10, CJ12, DL21]. Semisupervised [TMSP20]. Sensing [AAG23, ACS21, ADD12, BH17, BEFL21, CCW20, CCR^{+}12, CCBB14, DPVW14, FSY10, FL12, LLC14, LHB^{+}18, NDM^{+}11, Poo15, PCP^{+}16, RKT^{+}13, RB15, Rom09, SXS^{+}15, YOGD08, ZH21, ZLD^{+}18]. Sensitive [PB23]. Sensitivity [LR17, LR18, LHLP20, MB10, YWW^{+}23]. Sensor [FA09, GP09]. Sensors [Aco19, SG15]. Separability [WLJ22]. Separable [CCMY15, KGD21, PN23, SHVC19, ZZ21]. Separation [BGG19, CB18, Gil14b, HK19, JBS17, KGD21, KSPR17, Kut13, PH14, PYA^{+}12, SX12]. Sequence [PM08]. Sequences [DD10]. Sequential [HDH16, KLS^{+}17, dSO22]. Series [QSUZ11, RLL14]. Series-Based [QSUZ11]. Set [AKM11, ACDG18, AHL22, CLC13, EST20, FPT20, GB11, KBW13, RGLB14, SV08, WC23]. Sets [ALKÖP19, CFM09, KSS19, LLBS14, MD15]. Setup [MC16]. Shading [BCD^{+}12, CT17, KZ14, MF13]. Shah [HP11, BFG19, BPLX21, CCZ13, KSW20, Kla11, KR13, Mah12]. Shah-Like [Kla11]. Shah-Type [KR13]. Shape [ABK15, AR13, AR15, AKL^{+}21, AZ13, ATW14, AL15, AvdMSS22, BHM12, BCD^{+}12, CT17, DSYT10, DHSS13, DL18a, DAB^{+}20, EST20, FA09, FGS12, GDF15, GDT18, GTU14, LL22, LZD^{+}16, LTW^{+}10, MF13, MHP17, NPV16, OGL15, QYZ19, RGLB14, RW09, RW13, SCL20, SY14, SBC22, TCH08, WG22]. Shape-based [LTW^{+}10]. Shape-from-Shading [CT17, MF13]. Shapes [AKR13, ATTY16, CT13, KZS14, SUFU20, TAF^{+}20]. Shared [PKPE21]. Sharing [SC10]. Sharp [Sdi13]. Sharpening [Bel13, CMY10, MWBB12]. Sharpness [Col22]. Shear [Zhu16]. ShearLab [KSZ12]. Shearlet [GK14, GL09]. Shift [FK10]. Shooting [CKL17]. Short [AAG23, HK23, WLJ22]. Short-and-Sparse [WLJ22]. Shot [Han12]. Shrinkage [BT09, BAS15, CSS08]. Shrinkage-Thresholding [BT09]. Shutter [TMR13, TM16b]. SIFT [DL18b, SCM^{+}12]. SIFT-Like [DL18b]. Sigma [KV23]. Sigma-Delta [KV23]. Signal [AD23, GPST13, GPST15, GH15, GM10, HCGN23, KT22, SSN09, WY10, WHY^{+}15]. Signal-Subspace-Based [GM10]. Signal-to-Noise [GPST13]. Signals [CW18, CW22, GP09, GS10]. Signature [AE08, CE12, LL22]. Silico [GTP^{+}23]. Similar [FBU15, HLKH14, Kow14]. Similarities [FAS^{+}15]. Similarity [EHL17, LVEB09, WZLH20]. Simple [MÁS^{+}22]. Simplex [LWM^{+}18, MMT18]. Simplex-Structured [LWM^{+}18]. Simplicial [Yue23]. Simplified [FRV18]. Simplifying [NK20a]. Simply [CR18, CLLGL20]. Simply-Connected [CLLGL20]. Simulation [AC12, GL17, SKJ^{+}19]. Simultaneous [AKL^{+}21, DHSS13, JK15, LHB^{+}18, TA14]. Single [BBL^{+}23, FLZ14, HSSP09, HLLS14, KKN^{+}18, LUZZ22, LLW13, LR16, MQLC16, SSL23, TAR^{+}19, VHO20, dSO22]. Single-Grain [HSSP09]. Single-Lobe [MQLC16]. Single-Molecule [dSO22]. Single-Photon [TAR^{+}19]. Singular [KN14, MHM23]. Singularities [HQ16, Hub13]. Sinkhorn [DGSL23, KR17, KMDL19]. SISAL [HSMS22]. Size [Wan16a]. Sizes [HMY16]. Skeletons [STV09]. Sketching [LY18]. Sliced [LZ17a]. Sliced-Wasserstein [LZ17a]. Slices [BDM17, KL19]. Small [ACI08, ARYZ18a, ARYZ18b, Wah15]. Smooth [BdHKU22, CG19, HHK^{+}18, TPM20]. Smoothed [CP16, DAW21]. Smoothing [CLL11, CZ10, CY09, DGT19, FG23, Mil13, SM18]. Smoothness [Tii14]. Sobolev [BBHMA17, CMY10, KD12, KGV14, NPV16]. Solar [GPPM15, SLS19]. Solution [BGPS17, ZC15]. Solutions [BS21b, ELX13, KN14, PKCS18, PCBC10]. Solve [KKN^{+}18]. Solver [HW13, LQZ23]. Solvers [BKSW14]. Solving [CCBB14, GHFT23, GAT22, HW22, HLW20, KGB15, LR17, LR18, LLBS14, LH18, SKJ^{+}19, SK23]. Some [BMS23, CT17, DDGL19]. Sound [HN17, MPL^{+}18, NHKD22, QSUZ11]. Soup [MMT18]. Source [AB10, BV16, BGG19, CHPS09, CB18, GP14, GPST15, Gil14b, GS17, GHM23b, HHR08, Hub13, JBS17, KGD21, KKN20, LSYZ21, LLW23, LQS14, SX12, WCN^{+}19, WLJ22]. Sources [ALZ20, FST20, FST23, GHM23a, KZ14, Liu21, MWBK14, MC16]. Space [AdHW15, BHM12, BBHMA17, BBFA14, BER15, CLPS19, Car10, CV13, DD20, FN17, GLR18, Her19, KvD12, KN14, LS17, Lin18, LY18, Mar09, NK20a, RDM18, RW13, SYO15, SO08, SMSY11, TD17, WDCT09, ZZ19]. Space-Frequency [LS17]. Space-Time [Lin18, SYO15]. Space-Variant [CLPS19]. Spaceborne [ST11, Tsy09b]. Spaces [CKA17, DATP17, JK23, MD15, NPV16, NPS18, Tii14, XWH22, ZCO18]. SPARCOM [SEMS19]. Sparse [AR13, AE08, BBC11, BD22, CDP19, CW18, DAW21, DHK20, EW15, ELX13, FBU15, FSY10, FF13, FGPT17, GSXH18, GDF15, GS17, GHM23b, HPZ11, HK23, JHSX11, JDA^{+}19, LZ18, LLS^{+}13, MYZ13, MLH17, PC21, PG19, QLZ20, RBLS14, RB18, SLS19, SN11, SJD^{+}15, SX12, WY10, WY12, WLJ22, WE17, YJL^{+}17, ZBBO10, ZCO18]. Sparsifying [RB15]. Sparsity [BKBD16, CMP14, CBZ18, GEB15, JGM^{+}12, Kut13, LMSS19, NNYZ17, SEMS19, TV20, WZYX13]. Sparsity-Inducing [LMSS19]. Spatial [HZDZ23, LKW^{+}19, LSC^{+}18, LGL^{+}22, SXS^{+}15, WZLH20, WYN22, ZBO14, ZD16, ZDL18]. Spatial-Radon [ZD16, ZDL18]. Spatial-Temporal [LKW^{+}19]. Spatially [EW15, JZMN21, LNS10, YY17]. Spatiotemporal [CGÖ19, CDA21, CK09]. SPD [CV13]. Special [BMP13, DGMW23, Sap10]. Speckle [GSL^{+}22]. SPECT [CK09, LQS14]. Spectral [ABK15, ARF16, BGM^{+}16, CM20, Gil14a, KRW10, LKR18, LSC^{+}18, PBU^{+}22, YY22, ZvDT^{+}17]. Spectrometer [GPPM15]. Spectrometer/Telescope [GPPM15]. Spectrometry [MB10]. Spectroscopic [ARYZ18a, ARYZ18b]. Spectroscopy [PPE^{+}09, SX12]. Spectrum [CJ14, CM20]. Speed [HN17, MPL^{+}18, NHKD22, QSUZ11, WZ17]. Sphere [CW18, HP11, Lan19]. Sphere-Like [Lan19]. Spheres [CAT08, LCD22, GWY09]. Spherical [BP14, CW18, CLL15, CHL16, LHC^{+}23, OAUC^{+}20]. Spike [AARW19]. Spline [GL13, Sdi13]. Split [LSW14, WT10, YK16, GO09]. Splitting [CDH16, CG19, DGT19, FZ20, GLQ15, LY15, LCD22, OV14, PPO14, RFP13, RL15, SLS22]. Splittings [MSMC15]. Spot [CHM13]. Spotlight [CB11]. Spread [DAMM12, NS17, WCN^{+}19]. Square [DPSV17]. Squares [ELX13, LS18b, Nik13, SBFA15, SBFA16].

SSIM [MS22]. Stability

[ABG^{+}13c, BMS23, DGMW23, Wah15]. Stabilization [DD10, SM18]. Stabilized [PMZ20]. Stable [AGH14, HCGN22, NW13a, ÖSB15]. Stacking [SG15]. Stage [CCZ13, CYZ14, CLY19, GSZ17, LYZ20, LLW23, YLH23]. Start [Tsy09b]. Start-Stop [Tsy09b]. State [CCFBY13, LQS14]. Static [HSNS18]. Stationary [FW14, XFPA14]. Statistical [ACL16, DDGL19, Dem09, GDF15, RDG09, RGLB14]. Statistically [WLL^{+}21, YY15]. Statistics [AC09, FBU15, LNPS17]. Steepest [HH18]. Steerable [LS17, LS18a, PFA^{+}19, TT22, UC13, VSU15, WCU13]. Steering [PWSU16]. Stein [DVFP14]. Stellar [HHJ^{+}23]. Stencils [Get11]. Step [HMY16, LR17]. Stereo [BF15, MTWB14, MWBK14, MQLC16, SMA11]. Stitching [NW13b, WN13a]. STIX [GPPM15]. Stochastic [AvdMSS22, BAA14, DL18b, DTL^{+}21, HSH13, JK23, RM10, VHO20, WNS^{+}22, XXYC22, ZZ21]. Stokes [Her19]. Stop [Tsy09b]. Stored [DL14]. Strategies [ADX21, EGvL^{+}18, NMP15, SM18]. Stratified [Liu21]. Streaking [PUW17]. Stretch [HLL^{+}23, Yue23]. Strict [AN20]. Strictly [LY15]. Strong [ARYZ18b, ACL16, BS21b, CMP14]. Structural [AS18, HSF^{+}19]. Structure [EB16, EKOÅ10, HKLM21, KvD18, LRMU15, SS11]. Structure-Guided [EB16]. Structured [ELX13, GB18, JGM^{+}12, LAZ^{+}18, LWM^{+}18, PS19, SDL22]. Structures [Dro14, Fan09, LN13, RGLB14, SI23, SCL20]. Student [LS19]. Student- [LS19]. Study [ADGM14, HMY16, Her19, KK08, SKJ^{+}19]. Style [CFBP23]. Sub [BDMS15, CFSS16, GDT18]. Sub-Riemannian [BDMS15, CFSS16, GDT18]. Subaperture [HSR^{+}23]. Subaperture-Based [HSR^{+}23]. Subcellular [SNM17]. Sublinear [HNAC^{+}15]. Subordination [Car10]. Subpixel [DAMM12]. Subspace [CJPT13, GM10, HL13, KT22, SZW14, TV17, YY17]. Subwavelength [Fan09]. Successive [Gil14b]. Sufficient [GSGJ21, Sdi13]. SUGAR [DVFP14]. Super [CDH^{+}21, MC16, PBU^{+}22, SEMS19]. Super-Resolution [MC16, CDH^{+}21, PBU^{+}22, SEMS19]. Superlinearly [HW13]. Superresolution [AARW19, AH23, ALZ20, CDP19, HDH16, HK23, LR16, PCCP19, TA14]. Support [HHR08, WY10]. Supported [HMZZ19]. SURE [WM13]. Surface [AMY16, CDH^{+}21, ESS16, HKL20, HSH13, KZ14, KZ18, LTW^{+}10, WkZ14, ZC12]. Surface-Localized [CDH^{+}21]. Surfaces [AGSW16, BL14a, BZ18, CLL15, CR18, CLLGL20, DLLY17, HHK^{+}18, Lan19, LZZ18, LW14, WLTC12, WDCT09, XZZ19, ZCL22, ZJ21]. Surrogates [ZBSZ22]. Survey [EKV23]. Susceptibility [BCD19, CPW^{+}14, Nat16, PUW17, Pal16]. Symbol [ISW13]. Symbol-Based [ISW13]. Symmetric [BHV12, BPS16, CKA17, HMY16, RK19, RS20]. Symmetrizing [Mil13]. Synchronization [ARF16, GOF16, SS12]. Synchrosqueezed [LY18, YY13]. Synthesis [CJT^{+}12, GLR18, TPG16]. Synthesizing [XFPA14]. Synthetic [AC12, BCP13b, BMPT16, BGP^{+}17, BK18, BG20, CB11, DFM^{+}12, FSY09, GP15, KT22, LPT20b, Voc15, WY14, WY17, YY15]. Synthetic-Aperture [CB11]. System [CV13, KWRC20, SZGW18, WHY^{+}15]. Systems [AvdMSS22, BHI11, FW10, GK14, STCB13, ShDC^{+}19, ZLD^{+}18].

Tail [HF12]. Taken [SW13]. Takes [Mil18]. Taking [PS11]. Tale [YGLD17]. Tangent [VF14]. Tapered [BZ18]. Target [ABG^{+}13b, CB11, JLZ19b, YY15]. Targets [AGK^{+}12, FSY09, KKN19, NT11, WY12, WY14]. Taylor [KSW20]. Technique [BGG19, NK16]. Teichmuller [LLYG14, FN17, KN14, MCL16]. Telegraph [MRM20]. Telescope [GPPM15, Car10]. Template [CKL17, DATP17, MHP17]. TEMPO [MCL16]. Temporal [LKW^{+}19]. Tensor [BZNC16, CCQY20, CDHS13, CQ21, GH23, GPB17, HCCS20, HZ14, HMZZ19, HKBH13, KK17, LRMU15, LPT20b, LZZ^{+}23, NK20b, PBU^{+}22, QYW10, StTBRV12, SK23, TM18, VBK13, XY13, YLLX20, ZN19, ZBSZ22]. Tensor-Tensor [HKBH13]. Tensors [ABG^{+}13b, APST19, BHV12]. Term [LQZ23]. Terminating [TMP18]. Terms [CFM15, MMM12]. Terrain [DPH^{+}13]. Testing [DDGL19, Dem09]. Tetrahedral [FG23]. Texture [GL17, GLR18, GB18, Gil14a, JK15, KGV14, Kut13, LLWG13, MAP11, SO13, XFPA14, XXQJ20, YGS^{+}19]. Textured [Wan16b]. Textures [CCFBY13, TAF^{+}20]. TFV [GSZ17]. TGV [BH15a, BH15b]. TGV-Based [BH15a, BH15b]. Thanks [Naj17]. Their [BB14, DHZ21, Mon14, RVCB19]. Theorem [CHM13, MBFG20, SSSW09]. Theorems [FW10]. Theoretic [FR14, KLS^{+}17]. Theoretical [AGP18, DDPV20, SDM17, Yue23]. Theory [BBK22, CHHN21, CT17, CB11, De 23, GSC12, HK19, LZ18, LPP^{+}09, LLLX17, TM16b, WGGX22, ZJ21]. Thermoacoustic [QSUZ11]. Thin [Gri10]. Three [BLM14, CDRS16, CFM^{+}20, DLÖS23, GKQR20, Gri10, HWZ22, KKN20, KL19, KT16, LR17, LR18, LYZ20, SS11, TT22, YCU19]. Three-Dimensional [CDRS16, CFM^{+}20, DLÖS23, Gri10, HWZ22, KKN20, KL19, KT16, LR18, SS11, YCU19, LR17]. Three-Stage [LYZ20]. Thresholding [BT09, BAS15, CCZ13, GLS20]. Tight [CCMS13, CCW20, CBZ18, GSC13, HZ14, HMZZ19, PWSU16, ZD16]. Tight-Frame [CCMS13]. Tilts [RDM18]. Time [AdHW15, BER15, BDM15, BPT11, ERS18, ENR20, Kow14, Lin18, NK16, NTDB19, PS19, SYO15, TBKF15, YCU19, DM20]. Time-Dependent [TBKF15]. Time-Frequency [BPT11, PS19]. Tissue [Kow14]. Tomographic [AS18, CN17, DLW16, PS11, RKO22]. Tomography [ASK22, AAD^{+}08, Aco19, AKL^{+}21, AAJ^{+}16, ACS21, AM16, AKLS17, ABSM20, BS21b, BQ22, BGL^{+}21, DHSS13, DGMW23, FRV18, GH18, GLQ15, Gri10, HN17, HRSZ16, HHR08, HHMT16, HF12, HZDZ23, HSR^{+}23, JKSV20, KHD^{+}15, KSZ11, Kla11, Kow14, LHW^{+}15, LSC^{+}18, Lou08, LHB^{+}18, MPL^{+}18, MPM^{+}17, NS14, NK16, NHKD22, NLH^{+}16, PB23, PJS21, PLCD20, PAM12, PH20, QSUZ11, QS15, RZ13, RLL14, Rig17, RB18, ST23, SW13, SAS17, WZ17, WLYU15, WQ20, WQ21, WR14, YCU19, ZH21, ZE23, ZZ19]. Tomosynthesis [CNS10]. Tone [CBB14]. Top [DATP17]. Topological [ABG^{+}13c, ABR10, CDRS16, Dro14, LR17, LR18, SI23, SNM17, Wah15]. Topology [BG14, BCMO08, SY14, TP18]. Topology-Preserving [SY14]. Total [AGO21, All08, All09, BHB21, BBH^{+}23, BQ22, BKP10, BH12, BHSW18, BPLX21, CLL11, CP21, CTY13, CLDM18, CvG10, Con17, DL18a, DV22, DL21, DMSC16, EB16, EGvL^{+}18, GB18, Get11, GS13, Gil14a, GSC12, HHMT16, HNW09, HCGN22, JNW19, KSS19, KPR16, LMM17, LHW^{+}15, LRMU15, LZOX15, LM11, LM13, MYZ13, Mar09, NW13a, NW11, NNYZ17, NLH^{+}16, OGL15, PPRV22, PMS20, PLMS20, Poo15, SRG10, VBK13, WYYZ08, WN21, WYN22, WLYU15, WDS14, YY17, ZC15, ZHW22]. Trace [BBJ^{+}18]. Traces [DH20]. Tracing [StTBRV12]. Tracking [ABG^{+}13b, CV13, NT11, NTV10, SY14, SMSY11]. Tractable [BLM^{+}22]. Trading [SDR20]. Training [PPRV22, YZL^{+}18]. Trajectories [CCKW14, CDA21, HSH13, VHO20].

Transcranial [MPM^{+}17]. Transfer [CFBP23, Hub13, KLN^{+}23]. Transform [AAD^{+}08, ACN16, BHI11, BBK22, BMP13, CHZ21, GKQR20, GL09, Hal11, JM16, LUZZ22, LZ16, LZZ^{+}23, LR16, LQS14, MH17, MJC^{+}19, RLL14, Sto11, YY13, YCU19, BGV09]. Transformation [KO16, LHC^{+}23, ZBO14]. Transformed [CB18, HCGN22, RBLS14]. Transforms [ADB^{+}21, GR23, GTO14, LVEB09, LMT23, LY18, Mon14, NK20a, RB15, UC13, WCU13, WQ21, Zhu16]. Transient [AAB^{+}11]. translation [BGV09]. Transmission [CDH^{+}21, DLW16, KHD^{+}15]. Transport [BJM15, CLC13, FPPA14, GLR18, HLL^{+}23, KR17, KLNY23, LdGKW19, Mär11, MMT18, NS17, PPO14, SHB^{+}18, SDL22, SOK^{+}20, SK23]. Transport-Based [SHB^{+}18]. Travel [YCU19]. Traveltime [GLQ15, ZZ19]. Tree [KPR16]. Triangular [AN20, BBH^{+}23]. Triangulated [WDCT09]. Triangulation [Fou10]. Truncated [GSXH18, MLH17]. Trust [CPRS21, HW13]. Trust-Region [CPRS21, HW13]. Tube [LZZ^{+}23]. Tubular [Gri10]. Tumors [NTDB19]. Turbulent [HLKH14]. Turning [AFGK23]. TV [GSZ17, ADX21, CGN^{+}13, DMSC16, HHK^{+}18, HW13, HK14, LZ17b, MBBS14, WT10, YK16, ZBN17]. TV-Based [CGN^{+}13]. TV-Image [FKLS12]. TV-Type [HK14]. Tweedie [LDA^{+}22]. Twist [BCGR14]. Two [BGM14, BG21, BG15, BGG17, CCZ13, CYZ14, CLY19, DAMM12, DPC13, DH20, FST20, GSZ17, KYW13, KL19, LLW23, Lou08, MBFG20, SCGAF^{+}15, SW13, SUFU20, TMP13, YK16, YGLD17, YLH23]. Two-Dimensional [DH20, KL19, Lou08, SW13, SUFU20, TMP13]. Two-Level [KYW13]. Two-Parameter [SCGAF^{+}15]. Two-Point [BG21]. Two-Stage [CCZ13, CYZ14, CLY19, GSZ17, LLW23, YLH23]. Type [AdHW15, BP18, DD20, DGMW23, GKL13, HHMT16, HK14, PPRV22, RLL14, SRG10, BS09, KR13].

Ultra [WY14]. Ultra-Narrowband [WY14]. Ultrafast [AARW19]. Ultrasound [AARW19, BI15, LZD^{+}16, NTDB19]. Unbiased [DVFP14, ROD15]. Unbounded [LZZ18]. Uncertainty [EST20, MS17, RPW19, ZYZL20]. Underexposed [HJS13]. Undersampled [ACS21, SAS17]. Unified [CT17]. Uniform [AKZ13, BH17]. Unifying [UC13, ZCZL22, ZTO15]. Uniqueness [BZ18, JLZ19a, LMT23, MF13, Nik13, XZZ19]. Universal [FN17, GK14, KN14]. Unknown [HZDZ23, PH20, PBU^{+}22, SM16, SW13]. Unmixing [PBU^{+}22]. Unplugged [BCSB18]. Unresolved [FL12]. Unrolled [YWW^{+}23]. Unsquared [LSZ18, WSW13]. Unsupervised [BS21b, HBD18, SHB^{+}18, SOK^{+}20, ZZPS20]. Updating [LLS^{+}13]. Upwind [CLL11]. Use [MPM^{+}17, Tsy09b]. Using [AAD^{+}08, AE08, ACDG18, ABG^{+}13b, ARYZ18a, ARYZ18b, AdHW15, ALKÖP19, ADD12, BBJ^{+}18, BCD19, BHI11, BG14, BGG17, BGM^{+}16, BCSB18, CCZ13, CCMS13, CFBP23, CCQY20, CWR19, CZ10, CLC13, CDH16, CY09, CCBB14, CCFBY13, CLMT15, DSYT10, DGSL23, EHB09, EST20, FH11, FGPT17, GL17, GP09, Gri10, GL09, HSF^{+}19, HKBH13, HWC21, HSÅS18, HCGN22, KR17, KGV14, KL19, LZ17a, LDA^{+}22, LNPS17, LHW^{+}15, LPT20b, LPP^{+}09, LHLP20, LWY16, LLS19, LLWG13, NW13a, NW13b, NNYZ17, NNZC08, ÖSB15, PC21, PMZ20, RB15, RLS18, SXS^{+}15, StTBRV12, SS12, SZSH11, SCL20, SRG10, THC11, TBKF15, TCH08, Tii14, TP18, WY12, WSW13, WY14, WCN^{+}19, WYN22, WLYU15, WCA^{+}18, WkZ14, WE17, Yan13, YMA22, ZC12, ATTY16, BGV09, JLZ19b, XXQJ20]. Utilizing [ST23].

Value [JNW19, LY12, MHM23, WN21, WYN22]. Valued [AGO21, BT18, BS15, BHSW18, CDA21, GSC13, HW22, LVEB09, LNPS17, LTKG21, NPS18, SW14, WDS14]. Values [Bat23, BPS16, SM16]. Variability [AS18, PBU^{+}22]. Variable [AAD^{+}08, BWB14, CCKW14, HN17, MHM23, MJC^{+}19, NHKD22, QSUZ11, Tii14]. Variables [FH11]. Variance [DAG11, JLN14, LQZ23, YLH23, ZZ21]. Variant [BBFA14, CCZ13, CLPS19, YY17]. Variants [Her19]. Variation [AGO21, All09, BHB21, BBH^{+}23, BQ22, BKP10, BH12, BHSW18, BPLX21, CLL11, CP21, CTY13, CLDM18, CvG10, Con17, DL18a, DV22, DL21, DMSC16, EB16, EGvL^{+}18, GB18, Get11, GS13, Gil14a, GSC12, HHMT16, HNW09, HCGN22, JNW19, KSS19, KPR16, LMM17, LHW^{+}15, LRMU15, LZOX15, LM11, LM13, Mar09, NPV16, NW13a, NW11, NNYZ17, NLH^{+}16, OGL15, PPRV22, PMS20, PLMS20, Poo15, SRG10, VBK13, WYYZ08, WN21, WYN22, WLYU15, WDS14, YY17, ZC15, ZHW22, All08, MYZ13]. Variation-Based [BQ22, BH12, CLL11, CLDM18]. Variation-Type [PPRV22, SRG10]. Variational [BLSW14, BDM15, BH15a, BH15b, BDS18, CMLZ18, CGÖ19, CW22, CP16, CBB14, Dar15, DAB^{+}20, DZ13, DBCS14, DNB21, FLZ14, FW14, FR14, GVCPB15, GEB15, HFE19, HL13, HSY20, HKLM21, JLN14, JZMN21, KYW13, KP13, LMSY13, LLBS14, LYZ20, LCD22, MWBB12, NW13b, PM08, PABT17, PCBC10, RLS18, ST23, SDZ15, SBC22, SS13, WN13a, WN13b, WH15, WZLH20, YYZW09, ZC20]. Varifold [CT13]. Varying [EW15, LNS10]. VBTV [BHB21]. Vector [Bat10, BHB21, BK18, BS15, CY09, Fou10, GSC13, Her19, LTKG21, PS11, RDSK09, RBB20, Sdi13, SW14, WN21].

Vector-Valued [BS15, GSC13, LTKG21, SW14]. Vectorial [DMSC16, GSC12, SCC14, WT10]. Vehicles [CHM13]. Velocity [TM16b]. Venant [DL14]. Ventricle [SI23]. Version [HMY16]. Vese [NPJI17]. Vessel [CCMS13]. Via [DMZ18, BCP13b, BMW09, CCW20, CMY10, CESV13, CEM19, CDHS13, CN22, CLLGL20, CvG10, CEM21, DMTZ16, FPT20, FQXC17, GH23, GOF16, GK14, GS13, GM15, Han12, Hub13, KZS14, KK17, KV23, LZ17a, LL14, LWM^{+}18, LS18b, LCD22, LSW14, OAUC^{+}20, OGL15, PKCS18, QLZ20, RGLB14, RPW19, RDSK09, STY11, WY10, WTNL21, WMT^{+}09, YLLX20, YY22, ZvDT^{+}17, ZCL22, dSO22]. Vibrations [SG15]. Video [DKP09, HDH16, HK14, JHSX11, LZD^{+}16, LLWG13, NAF^{+}14, PABT17, SM18, SXS^{+}15, SYO15, TP18, Zhu16]. View [AAB^{+}11, HZDZ23]. Viewing [MTWB14, SS12, SZSH11]. Views [ARF16]. Virtual [GP14, GPST15]. Visco [AGO21]. Visco-acoustic [AGO21]. Viscous [RW13]. Visibility [PPE^{+}09]. Visibility-Based [PPE^{+}09]. Visible [BBP09]. Vision [BCGR14, ODBP15, SMA11]. Visual [NTV10, SC10]. Volume [BDM17, KL19, LWM^{+}18, OAUC^{+}20, YLLY19, YLLY20]. Volume-Preserving [YLLY19, YLLY20]. Volumetric [HLL^{+}23, RR15]. Voronoi [CLMT15]. Vortical [SI23]. Voxel [KT14].

Walk [TM12]. Warp [Seg22]. WARPd [Col22]. Warping [KSPR17]. Wasserstein [AH23, BHS23, CP16, DD20, HSÅS18, LZ17a, Pey15, SHB^{+}18, TPG16]. Wasserstein-Type [DD20]. Water [Kow14]. Watershed [Naj17]. Wave [AAB^{+}11, BZ18, BGMZ23, DH20, FW10, HP15, LWY16, LLS19, MPM^{+}17, MHM23, YY13, ZH21]. Waveform [AGO21, BGMZ23, CB11, EGvL^{+}18]. Waveform-Diverse [CB11]. Waveforms [WY14]. Wavefront [ALKÖP19]. Waveguide [Liu21, TMP18]. Waveguides [TMP13]. Wavelet [CSS08, CYY11, DB13, EW15, FBU15, FKLS12, GL13, HLKH14, LZD^{+}16, PWSU16, TZS13, UC13]. Wavelets [WCU13, GTO14]. Waves [DLL19, DMZ18, SSSW09]. Way [FW10]. Weak [HK23]. Weak* [Ish14]. Weak*-Convergence [Ish14]. Weakly [ABG13a, Far19]. Weighted [BPLX21, CN22, DL18a, KV23, LO17, LLC14, LZOX15, OGL15, YK16]. Weighted- [LLC14]. Weighting [JZMN21]. Weights [JGKL17, YZL^{+}18]. Weil [FN17, KN14]. Welding [CLLGL20, ZCL22]. Well [Aco19, CT17, HK23]. Well-Conditioned [HK23]. Well-Posedness [Aco19, CT17]. Which [GSC12]. Whiteness [LMSY13]. Whole [BCD19, vGPR22]. Windows [DKP09]. Wirtinger [YY19]. Wisely [BBES21]. Within [PABT17, BMW09]. without [ACL16, BBL^{+}23, CH16, KZ14, MNP16, TA14]. World [BH17]. WPPFlows [AH23]. WPPNets [AH23].

X [GPPM15, AAD^{+}08, BFJQ18, DLW16, LUZZ22, LQS14, Mon14, PJS21, YCU19, YCF^{+}16, ZE23]. X-Ray [AAD^{+}08, DLW16, LQS14, Mon14, PJS21, YCU19, YCF^{+}16, ZE23, BFJQ18, LUZZ22]. X-rays [GPPM15].

Zooming [BH15a, BH15b, CLK14].

References

Ammari:2011:TWI

[AAB^{+}11] Habib Ammari, Mark Asch, Lili Guadarrama Bustos, Vincent Jugnon, and Hyeonbae Kang. Transient wave imaging with limited-view data. SIAM Journal on Imaging Sciences, 4(4):1097–1121, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i4/p1097_s1.

Abramson:2008:QOR

[AAD^{+}08] Mark A. Abramson, Thomas J.

Asaki, J. E. Dennis, Jr., Kevin R. O’Reilly, and Rachael L. Pingel. Quantitative object reconstruction using Abel transform X-ray tomography and mixed variable optimization. SIAM Journal on Imaging Sciences, 1(3):322–342, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Alaifari:2023:SCL

[AAG23] Rima Alaifari, Giovanni S.

Alberti, and Tandri Gauksson. Short communication: Localized adversarial artifacts for compressed sensing MRI. SIAM Journal on Imaging Sciences, 16(4):SC14– SC26, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1503221.

Alberti:2016:LIP

[AAJ^{+}16] Giovanni S. Alberti, Habib Ammari, Bangti Jin, Jin-Keun Seo, and Wenlong Zhang. The linearized inverse problem in multifrequency electrical impedance tomography. SIAM Journal on Imaging Sciences, 9 (4):1525–1551, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Alberti:2019:DSS

[AARW19] Giovanni S. Alberti, Habib Ammari, Francisco Romero, and Timothée Wintz. Dynamic spike superresolution and applications to ultrafast ultrasound imaging. SIAM Journal on Imaging Sciences, 12(3):1501– 1527, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Adavani:2010:FAS

[AB10] Santi S. Adavani and George Biros. Fast algorithms for source identification problems with elliptic PDE constraints. SIAM Journal on Imaging Sciences, 3(4):791–808, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Alexeev:2014:PRP

[ABFM14] Boris Alexeev, Afonso S. Bandeira, Matthew Fickus, and Dustin G. Mixon. Phase retrieval with polarization. SIAM Journal on Imaging Sciences, 7 (1):35–66, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ammari:2013:MAE

[ABG13a] Habib Ammari, Thomas Boulier, and Josselin Garnier. Modeling active electrolocation in weakly electric fish. SIAM Journal on Imaging Sciences, 6(1):285–321, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Ammari:2013:TMT

[ABG^{+}13b] Habib Ammari, Thomas Boulier, Josselin Garnier, Hyeonbae Kang, and Han Wang. Tracking of a mobile target using generalized polarization tensors. SIAM Journal on Imaging Sciences, 6 (3):1477–1498, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Ammari:2013:LSR

[ABG^{+}13c] Habib Ammari, Elie Bretin, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, and Abdul Wahab. Localization, stability, and resolution of topological derivative based imaging functionals in elasticity. SIAM Journal on Imaging Sciences, 6(4): 2174–2212, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Aflalo:2015:OSD

[ABK15] Yonathan Aflalo, Haim Brezis, and Ron Kimmel. On the optimality of shape and data representation in the spectral domain. SIAM Journal on Imaging Sciences, 8(2):1141–1160, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Auroux:2010:ATG

[ABR10] Didier Auroux, Lamia Jaafar Belaid, and Badreddine Rjaibi. Application of the topological gradient method to color image restoration. SIAM Journal on Imaging Sciences, 3(2):153–175, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Aspri:2020:AEH

[ABSM20] Andrea Aspri, Elena Beretta, Otmar Scherzer, and Monika Muszkieta. Asymptotic expansions for higher order elliptic equations with an application to quantitative photoacoustic tomography. SIAM Journal on Imaging Sciences, 13(4):1781– 1833, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Adams:2009:NSR

[AC09] Henry Adams and Gunnar Carlsson. On the nonlinear statistics of range image patches. SIAM Journal on Imaging Sciences, 2(1):110–117, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Allan:2012:GCG

[AC12] Jeremy M. Allan and Michael J. Collins. The generation of correlated gamma distributed random fields for the simulation of synthetic aperture radar images. SIAM Journal on Imaging Sciences, 5(4):1261–1290, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Alvarez:2018:LSR

[ACDG18] Luis Alvarez, Carmelo Cuenca, Jesús Ildefonso Díaz, and Esther González. Level set regularization using geometric flows. SIAM Journal on Imaging Sciences, 11(2):1493–1523, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Ammari:2008:DEI

[ACI08] Habib Ammari, Pierre Calmon, and Ekaterina Iakovleva. Direct elastic imaging of a small inclusion. SIAM Journal on Imaging Sciences, 1(2):169–187, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Aspelmeier:2016:LLC

[ACL16] Timo Aspelmeier, C. Charitha, and D. Russell Luke. Local linear convergence of the ADMM/Douglas–Rachford algorithms without strong convexity and application to statistical imaging. SIAM Journal on Imaging Sciences, 9(2):842–868, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Andersson:2016:FAE

[ACN16] Fredrik Andersson, Marcus Carlsson, and Viktor V. Nikitin. Fast algorithms and efficient GPU implementations for the Radon transform and the backprojection operator represented as convolution operators. SIAM Journal on Imaging Sciences, 9 (2):637–664, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Acosta:2019:WPP

[Aco19]  Sebastian Acosta.  Well-

posedness for photoacoustic tomography with Fabry–Perot sensors. SIAM Journal on Imaging Sciences, 12(4):1669– 1685, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Alberti:2021:CSP

[ACS21] Giovanni S. Alberti, Paolo Campodonico, and Matteo Santacesaria. Compressed sensing photoacoustic tomography reduces to compressed sensing for undersampled Fourier measurements. SIAM Journal on Imaging Sciences, 14(3):1039–1077, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Arias-Castro:2012:OIM

[ACSW12] Ery Arias-Castro, Joseph Salmon, and Rebecca Willett. Oracle inequalities and minimax rates for nonlocal means and related adaptive kernel-based methods. SIAM Journal on Imaging Sciences, 5(3):944–992, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Ammari:2023:ALB

[AD23] Habib Ammari and Bryn Davies. Asymptotic links between signal processing, acoustic metamaterials, and biology. SIAM Journal on Imaging Sciences, 16(1):64– 88, ???? 2023. CODEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1510352.

Aramini:2021:AEB

[ADB^{+}21] Riccardo Aramini, Fabrice Delbary, Mauro C. Beltrametti, Claudio Estatico, Michele Piana, and Anna Maria Massone. On the asymptotic equivalence between the Radon and the Hough transforms of digital images. SIAM Journal on Imaging Sciences, 14(2):506–529, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Averbuch:2012:ACI

[ADD12] Amir Averbuch, Shai Dekel, and Shay Deutsch. Adaptive compressed image sensing using dictionaries. SIAM Journal on Imaging Sciences, 5(1):57– 89, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Aguerrebere:2014:BAH

[ADGM14] Cecilia Aguerrebere, Julie Delon, Yann Gousseau, and Pablo Musé. Best algorithms for HDR image generation. a study of performance bounds. SIAM Journal on Imaging Sciences, 7 (1):1–34, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Andersson:2015:MRT

[AdHW15] Fredrik Andersson, Maarten V.

de Hoop, and Herwig Wendt. Multiscale reverse-time-migration-type imaging using the dyadic parabolic decomposition of phase space. SIAM Journal on Imaging Sciences, 8(4):2383– 2411, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Allassonniere:2015:BME

[ADK15] S. Allassonnière, S. Durrleman, and E. Kuhn. Bayesian mixed effect atlas estimation with a diffeomorphic deformation model. SIAM Journal on Imaging Sciences, 8(3):1367– 1395, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Adcock:2021:IRG

[ADX21] Ben Adcock, Nick Dexter, and Qinghong Xu. Improved recovery guarantees and sampling strategies for TV minimization in compressive imaging. SIAM Journal on Imaging Sciences, 14(3):1149–1183, ???? 2021.

CODEN SJISBI. ISSN 1936-4954.

Aharon:2008:SRM

[AE08] Michal Aharon and Michael Elad. Sparse and redundant modeling of image content using an image-signature-dictionary. SIAM Journal on Imaging Sciences, 1(3):228–247, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Alpers:2023:TGM

[AFGK23] Andreas Alpers, Maximilian Fiedler, Peter Gritzmann, and Fabian Klemm. Turning grain maps into diagrams. SIAM Journal on Imaging Sciences, 16(1):223–249, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1491988.

Adcock:2014:SRN

[AGH14] Ben Adcock, Milana Gataric, and Anders Hansen. On stable reconstructions from nonuniform Fourier measurements. SIAM Journal on Imaging Sciences, 7(3):1690–1723, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ammari:2012:MIE

[AGK^{+}12] Habib Ammari, Josselin Garnier, Hyeonbae Kang, Mikyoung Lim, and Knut Sølna. Multistatic imaging of extended targets. SIAM Journal on Imaging Sciences, 5(2):564–600, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Ammari:2014:BIN

[AGM14] Habib Ammari, Josselin Garnier, and Pierre Millien. Backpropagation imaging in nonlinear harmonic holography in the presence of measurement and medium noises. SIAM Journal on Imaging Sciences, 7(1): 239–276, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Aghamiry:2021:CVI

[AGO21] Hossein S. Aghamiry, Ali Gholami, and Stéphane Operto. Complex-valued imaging with total variation regularization: an application to full-waveform inversion in visco-acoustic media. SIAM Journal on Imaging Sciences, 14(1):58–91, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Aujol:2018:TAF

[AGP18] Jean-Franois Aujol, Guy Gilboa, and Nicolas Papadakis. Theoretical analysis of flows estimating eigenfunctions of one-homogeneous functionals. SIAM Journal on Imaging Sciences, 11(2):1416–1440, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Absil:2016:DPB

[AGSW16] P.-A. Absil, Pierre-Yves Gousenbourger, Paul Striewski, and Benedikt Wirth. Differentiable piecewise-Bézier surfaces on Riemannian manifolds. SIAM Journal on Imaging Sciences, 9 (4):1788–1828, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Audibert:2017:GLS

[AH17] Lorenzo Audibert and Houssem Haddar. The generalized linear sampling method for limited aperture measurements. SIAM Journal on Imaging Sciences, 10(2):845–870, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Altekruger:2023:WWP

[AH23] Fabian Altekrüger and Johannes Hertrich. WPPNets and WPPFlows: The power of Wasserstein patch priors for superresolution. SIAM Journal on Imaging Sciences, 16(3): 1033–1067, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1496542.

Audibert:2022:ALS

[AHL22] Lorenzo Audibert, Houssem Haddar, and Xiaoli Liu. An accelerated level-set method for inverse scattering problems. SIAM Journal on Imaging Sciences, 15(3):1576– 1600, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1457783.

Agnelli:2021:SRC

[AKL^{+}21] Juan P. Agnelli, Ville Kolehmainen, Matti J. Lassas, Petri Ola, and Samuli Siltanen. Simultaneous reconstruction of conductivity, boundary shape, and contact impedances in electrical impedance tomography. SIAM Journal on Imaging Sciences, 14(4):1407–1438, ???? 2021.

CODEN SJISBI. ISSN 1936-4954.

Ammari:2017:MFA

[AKLS17] Habib Ammari, Hyeuknam Kwon, Seungri Lee, and Jin Keun Seo. Mathematical framework for abdominal electrical impedance tomography to assess fatness. SIAM Journal on Imaging Sciences, 10(2): 900–919, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Aghasi:2011:PLS

[AKM11] Alireza Aghasi, Misha Kilmer, and Eric L. Miller. Parametric level set methods for inverse problems. SIAM Journal on Imaging Sciences, 4(2): 618–650, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i2/p618_s1.

Aflalo:2013:SIG

[AKR13] Yonathan Aflalo, Ron Kimmel, and Dan Raviv. Scale invariant geometry for nonrigid shapes. SIAM Journal on Imaging Sciences, 6(3):1579–1597, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Aflalo:2013:CMU

[AKZ13] Yonathan Aflalo, Ron Kimmel, and Michael Zibulevsky. Conformal mapping with as uniform as possible conformal factor. SIAM Journal on Imaging Sciences, 6(1):78–101, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Arens:2015:IFS

[AL15] T. Arens and A. Lechleiter. Indicator functions for shape reconstruction related to the linear sampling method. SIAM Journal on Imaging Sciences, 8 (1):513–535, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Andrade-Loarca:2019:EDW

[ALKÖP19] Hector Andrade-Loarca, Gitta Kutyniok, Ozan Öktem, and Philipp C. Petersen. Extraction of digital wavefront sets using applied harmonic analysis and deep neural networks. SIAM Journal on Imaging Sciences, 12(4):1936–1966, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Allard:2008:TVR

[All08] William K. Allard. Total variation regularization for image denoising. II. Examples. SIAM Journal on Imaging Sciences, 1 (4):400–417, ???? 2008. CO-DEN SJISBI. ISSN 1936-4954.

Allard:2009:TVR

[All09] William K. Allard. Total variation regularization for image denoising, III. Examples. SIAM Journal on Imaging Sciences, 2 (2):532–568, ???? 2009. CO-DEN SJISBI. ISSN 1936-4954.

Ammari:2020:SRE

[ALZ20] Habib Ammari, Bowen Li, and Jun Zou. Superresolution in recovering embedded electromagnetic sources in high contrast media. SIAM Journal on Imaging Sciences, 13(3):1467–1510, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Alsaker:2016:BAP

[AM16] Melody Alsaker and Jennifer L.

Mueller. A D-bar algorithm with a Priori information for 2-dimensional electrical impedance tomography. SIAM Journal on Imaging Sciences, 9 (4):1619–1654, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Arguillere:2016:DSR

[AMY16] Sylvain Arguillère, Michael I.

Miller, and Laurent Younes. Diffeomorphic surface registration with atrophy constraints. SIAM Journal on Imaging Sciences, 9(3):975–1003, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Abdalla:2020:MMT

[AN20] Mohsen Abdalla and Benedek Nagy. Mathematical morphology on the triangular grid: The strict approach. SIAM Journal on Imaging Sciences, 13(3): 1367–1385, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Angulo:2013:MBF

[Ang13] Jesús Angulo. Morphological bilateral filtering. SIAM Journal on Imaging Sciences, 6(3): 1790–1822, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Ammari:2019:RDA

[APST19] Habib Ammari, Mihai Putinar, Andries Steenkamp, and Faouzi Triki. Reconstruction of domains with algebraic boundaries from generalized polarization tensors. SIAM Journal on Imaging Sciences, 12(4):2097– 2118, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Aghasi:2013:SSR

[AR13] Alireza Aghasi and Justin Romberg. Sparse shape reconstruction. SIAM Journal on Imaging Sciences, 6(4):2075– 2108, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Aghasi:2015:CCS

[AR15] Alireza Aghasi and Justin Romberg. Convex cardinal shape composition. SIAM Journal on Imaging Sciences, 8(4): 2887–2950, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Anden:2020:MEA

[AR20] Joakim Andén and José Luis Romero. Multitaper estimation on arbitrary domains. SIAM Journal on Imaging Sciences, 13(3):1565–1594, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Arrigoni:2016:SSM

[ARF16] Federica Arrigoni, Beatrice Rossi, and Andrea Fusiello. Spectral synchronization of multiple views in SE(3). SIAM Journal on Imaging Sciences, 9 (4):1963–1990, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Arrate:2010:DAC

[ARY10] Felipe Arrate, J. Tilak Ratnanather, and Laurent Younes.

Diffeomorphic active contours. SIAM Journal on Imaging Sciences, 3(2):176–198, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Ammari:2018:RFDa

[ARYZ18a] Habib Ammari, Matias Ruiz, Sanghyeon Yu, and Hai Zhang. Reconstructing fine details of small objects by using plasmonic spectroscopic data. SIAM Journal on Imaging Sciences, 11(1):1–23, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Ammari:2018:RFDb

[ARYZ18b] Habib Ammari, Matias Ruiz, Sanghyeon Yu, and Hai Zhang. Reconstructing fine details of small objects by using plasmonic spectroscopic data. Part II: The strong interaction regime. SIAM Journal on Imaging Sciences, 11(3):1931–1953, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Anden:2018:SVN

[AS18] Joakim Andén and Amit Singer.

Structural variability from noisy tomographic projections. SIAM Journal on Imaging Sciences, 11(2):1441–1492, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Abhishek:2022:OBE

[ASK22] Anuj Abhishek, Thilo Strauss, and Taufiquar Khan. An optimal Bayesian estimator for absorption coefficient in diffuse optical tomography. SIAM Journal on Imaging Sciences, 15(2):797–821, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1462842.

Athavale:2011:IDE

[AT11] Prashant Athavale and Eitan Tadmor. Integro-differential equations based on (BV, L^{1}) image decomposition. SIAM Journal on Imaging Sciences, 4 (1):300–312, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p300_s1.

Arguillere:2016:RMS

[ATTY16] Sylvain Arguillère, Emmanuel Trélat, Alain Trouvé, and Laurent Younes. Registration of multiple shapes using constrained optimal control. SIAM Journal on Imaging Sciences, 9 (1):344–385, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Ammari:2014:SIC

[ATW14] Habib Ammari, Minh Phuong Tran, and Han Wang. Shape identification and classification in echolocation. SIAM Journal on Imaging Sciences, 7(3): 1883–1905, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Arnaudon:2022:DBS

[AvdMSS22] Alexis Arnaudon, Frank van der Meulen, Moritz Schauer, and Stefan Sommer. Diffusion bridges for stochastic Hamiltonian systems and shape evolutions. SIAM Journal on Imaging Sciences, 15(1):293– 323, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1406283.

Aktas:2013:FSE

[AZ13] Mehmet Ali Aktas and Jovisa Zunić. A family of shape ellipticity measures for galaxy classification. SIAM Journal on Imaging Sciences, 6(2):765–781, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Bernal:2014:SAB

[BAA14] Francisco Bernal, Juan A.

Acebrón, and Immanuel Anjam. A stochastic algorithm based on fast marching for automatic capacitance extraction in non-Manhattan geometries. SIAM Journal on Imaging Sciences, 7 (4):2657–2674, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Bhotto:2015:IFI

[BAS15] Md. Zulfiquar Ali Bhotto, M. Omair Ahmad, and M. N. S. Swamy. An improved fast iterative shrinkage thresholding algorithm for image deblurring. SIAM Journal on Imaging Sciences, 8(3):1640–1657, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Batard:2010:CBC

[Bat10] Thomas Batard. Clifford bundles: a common framework for image, vector field, and orthonormal frame field regularization. SIAM Journal on Imaging Sciences, 3(3):670–701, ????

2010. CODEN SJISBI. ISSN 1936-4954.

Batard:2023:CPC

[Bat23]  Thomas Batard.  A class

of priors for color image restoration parameterized by Lie groups acting on pixel values. SIAM Journal on Imaging Sciences, 16(3):1235– 1280, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1504664.

Batard:2014:CDT

[BB14] Thomas Batard and Marcelo Bertalmío. On covariant derivatives and their applications to image regularization. SIAM Journal on Imaging Sciences, 7 (4):2393–2422, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Becker:2011:NFA

[BBC11] Stephen Becker, Jérôme Bobin, and Emmanuel J. Candès. NESTA: a fast and accurate first-order method for sparse recovery. SIAM Journal on Imaging Sciences, 4 (1):1–39, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p1_s1.

Barmherzig:2021:RMD

[BBE^{+}21] David A. Barmherzig, Alex H. Barnett, Charles L. Epstein, Leslie F. Greengard, Jeremy F. Magland, and Manas Rachh. Recovering missing data in coherent diffraction imaging. SIAM Journal on Imaging Sciences, 14(2):620–644, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Benning:2021:CYP

[BBES21] Martin Benning, Marta M.

Betcke, Matthias J. Ehrhardt, and Carola-Bibiane Schönlieb. Choose your path wisely: Gradient descent in a Bregman distance framework. SIAM Journal on Imaging Sciences, 14(2): 814–843, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

BenHadj:2014:SVB

[BBFA14] Saima Ben Hadj, Laure Blanc-Féraud, and Gilles Aubert. Space variant blind image restoration. SIAM Journal on Imaging Sciences, 7(4):2196– 2225, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Baumgartner:2023:TGV

[BBH^{+}23] Lukas Baumgärtner, Ronny Bergmann, Roland Herzog, Stephan Schmidt, and José Vidal-Núnez. Total generalized variation for piecewise constant functions on triangular meshes with applications in imaging. SIAM Journal on Imaging Sciences, 16(1): 313–339, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1505281.

Bauer:2017:NFS

[BBHMA17] Martin Bauer, Martins Bruveris, Philipp Harms, and Jakob Møller-Andersen. A numerical framework for Sobolev metrics on the space of curves. SIAM Journal on Imaging Sciences, 10(1):47–73, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Bao:2018:CRU

[BBJ^{+}18] Chenglong Bao, George Barbastathis, Hui Ji, Zuowei Shen, and Zhengyun Zhang. Coherence retrieval using trace regularization. SIAM Journal on Imaging Sciences, 11(1):679–706, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Beckmann:2022:MRT

[BBK22] Matthias Beckmann, Ayush Bhandari, and Felix Krahmer. The modulo Radon transform: Theory, algorithms, and applications. SIAM Journal on Imaging Sciences, 15(2): 455–490, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1424615.

Bendory:2023:TSP

[BBL^{+}23] Tamir Bendory, Nicolas Boumal, William Leeb, Eitan Levin, and Amit Singer. Toward single particle reconstruction without particle picking: Breaking the detection limit. SIAM Journal on Imaging Sciences, 16 (2):886–910, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1503828.

Bellettini:2009:CVC

[BBP09] Giovanni Bellettini, Valentina Beorchia, and Maurizio Paolini. Completion of visible contours. SIAM Journal on Imaging Sciences, 2(3):777–799, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Bian:2015:LCN

[BC15] Wei Bian and Xiaojun Chen.

Linearly constrained non-Lipschitz optimization for image restoration. SIAM Journal on Imaging Sciences, 8(4):2294–2322, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Boyer:2016:GSS

[BCC^{+}16] Claire Boyer, Nicolas Chauffert, Philippe Ciuciu, Jonas Kahn, and Pierre Weiss. On the generation of sampling schemes for magnetic resonance imaging. SIAM Journal on Imaging Sciences, 9(4):2039–2072, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Breuss:2012:PSS

[BCD^{+}12] Michael Breuß, Emiliano Cristiani, Jean-Denis Durou, Maurizio Falcone, and Oliver Vogel. Perspective shape from shading: Ambiguity analysis and numerical approximations. SIAM Journal on Imaging Sciences, 5 (1):311–342, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Bao:2019:WBS

[BCD19] Chenglong Bao, Jae Kyu Choi, and Bin Dong. Whole brain susceptibility mapping using harmonic incompatibility removal. SIAM Journal on Imaging Sciences, 12(1):492–520, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Boscain:2014:HDH

[BCGR14] U. Boscain, R. A. Chertovskih, J. P. Gauthier, and A. O. Remizov. Hypoelliptic diffusion and human vision: a semidiscrete new twist. SIAM Journal on Imaging Sciences, 7(2):669–695, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Buades:2008:TPL

[BCMO08] A. Buades, A. Chien, J. M. Morel, and S. Osher. Topology preserving linear filtering applied to medical imaging. SIAM Journal on Imaging Sciences, 1 (1):26–50, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Bonettini:2013:NSD

[BCP13a] S. Bonettini, A. Cornelio, and M. Prato. A new semiblind deconvolution approach for fourier-based image restoration: an application in astronomy. SIAM Journal on Imaging Sciences, 6(3):1736–1757, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Borcea:2013:SAR

[BCP13b] Liliana Borcea, Thomas Callaghan, and George Papanicolaou. Synthetic aperture radar imaging and motion estimation via robust principal component analysis. SIAM Journal on Imaging Sciences, 6(3):1445–1476, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Buzzard:2018:PPU

[BCSB18] Gregery T. Buzzard, Stanley H.

Chan, Suhas Sreehari, and Charles A. Bouman. Plug-and-play unplugged: Optimizationfree reconstruction using consensus equilibrium. SIAM Journal on Imaging Sciences, 11(3): 2001–2020, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Bruhin:2022:BRP

[BD22] Nina Dekoninck Bruhin and Bryn Davies. Bioinspired random projections for robust, sparse classification. SIAM Journal on Imaging Sciences, 15(4):1833–1850, ???? 2022. CODEN SJISBI. ISSN 1936-

4954. URL https://epubs.
siam.org/doi/10.1137/22M1503579.

Bhattacharyya:2022:RPS

[BdHKU22] Sombuddha Bhattacharyya, Maarten V. de Hoop, Vitaly Katsnelson, and Gunther Uhlmann. Recovery of piecewise smooth density and Lamé parameters from high frequency exterior Cauchy data. SIAM Journal on Imaging Sciences, 15(4):1910– 1943, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1480951.

Bogelein:2015:TDV

[BDM15] Verena Bögelein, Frank Duzaar, and Paolo Marcellini. A time dependent variational approach to image restoration. SIAM Journal on Imaging Sciences, 8 (2):968–1006, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Bretin:2017:VRS

[BDM17] Elie Bretin, François Dayrens, and Simon Masnou. Volume reconstruction from slices. SIAM Journal on Imaging Sciences, 10(4):2326–2358, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Borcea:2020:ROM

[BDM^{+}20] Liliana Borcea, Vladimir Druskin, Alexander V. Mamonov, Mikhail Zaslavsky, and Jörn Zimmerling. Reduced order model approach to inverse scattering. SIAM Journal on Imaging Sciences, 13(2):685–723, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Bekkers:2015:PAD

[BDMS15] E. J. Bekkers, R. Duits, A. Mashtakov, and G. R. Sanguinetti. A PDE approach to data-driven sub-Riemannian geodesics in SE(2). SIAM Journal on Imaging Sciences, 8(4): 2740–2770, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Burger:2018:VMJ

[BDS18] Martin Burger, Hendrik Dirks, and Carola-Bibiane Schönlieb. A variational model for joint motion estimation and image reconstruction. SIAM Journal on Imaging Sciences, 11(1):

94–128, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Bruckstein:2021:PBH

[BEFL21] Alfred Marcel Bruckstein, Martianus Frederic Ezerman, Adamas Aqsa Fahreza, and San Ling. Patchbased holographic image sensing. SIAM Journal on Imaging Sciences, 14(1):198–223, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Belyaev:2013:IID

[Bel13] Alexander Belyaev. Implicit image differentiation and filtering with applications to image sharpening. SIAM Journal on Imaging Sciences, 6(1):660–679, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Berkels:2015:TDG

[BER15] B. Berkels, A. Effland, and M. Rumpf. Time discrete geodesic paths in the space of images. SIAM Journal on Imaging Sciences, 8(3):1457–1488, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Buades:2015:RMM

[BF15] Antoni Buades and Gabriele Facciolo. Reliable multiscale and multiwindow stereo matching. SIAM Journal on Imaging Sciences, 8(2):888–915, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Bucur:2019:MMS

[BFG19] Dorin Bucur, Ilaria Fragalà, and Alessandro Giacomini. The multiphase Mumford–Shah problem. SIAM Journal on Imaging Sciences, 12(3):1561– 1583, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Borg:2018:ARA

[BFJQ18] Leise Borg, Jürgen Frikel, Jakob Sauer Jørgensen, and Eric Todd Quinto. Analyzing reconstruction artifacts from arbitrary incomplete X-ray CT data. SIAM Journal on Imaging Sciences, 11(4):2786–2814, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Benninghoff:2014:EIS

[BG14] Heike Benninghoff and Harald Garcke. Efficient image segmentation and restoration using parametric curve evolution with junctions and topology changes. SIAM Journal on Imaging Sciences, 7(3):1451– 1483, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Borges:2015:IOS

[BG15] Carlos Borges and Leslie Greengard. Inverse obstacle scattering in two dimensions with multiple frequency data and multiple angles of incidence. SIAM Journal on Imaging Sciences, 8(1): 280–298, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Borcea:2020:HRI

[BG20] Liliana Borcea and Josselin Garnier. High-resolution interferometric synthetic aperture imaging in scattering media. SIAM Journal on Imaging Sciences, 13(1):291–316, ????

2020. CODEN SJISBI. ISSN 1936-4954.

Borcea:2021:IRM

[BG21] Liliana Borcea and Josselin Garnier. Imaging in random media by two-point coherent interferometry. SIAM Journal on Imaging Sciences, 14(4):1635– 1668, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Borges:2017:HRI

[BGG17] Carlos Borges, Adrianna Gillman, and Leslie Greengard. High resolution inverse scattering in two dimensions using recursive linearization. SIAM Journal on Imaging Sciences, 10(2):641–664, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Boccuto:2019:BSS

[BGG19] Antonio Boccuto, Ivan Gerace, and Valentina Giorgetti. A blind source separation technique for document restoration. SIAM Journal on Imaging Sciences, 12(2):1135–1162, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Barnafi:2018:PMF

[BGH18] Nicolás Barnafi, Gabriel N. Gatica, and Daniel E. Hurtado. Primal and mixed finite element methods for deformable image registration problems. SIAM Journal on Imaging Sciences, 11(4):2529–2567, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Barnafi:2021:AMR

[BGH^{+}21] Nicolas Barnafi, Gabriel N. Gatica, Daniel E. Hurtado, Willian Miranda, and Ricardo Ruiz-Baier. Adaptive mesh refinement in deformable image registration: a posteriori error estimates for primal and mixed formulations. SIAM Journal on Imaging Sciences, 14(3):1238– 1272, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Bauer:2015:OFM

[BGK15] Martin Bauer, Markus Grasmair, and Clemens Kirisits. Optical flow on moving manifolds. SIAM Journal on Imaging Sciences, 8(1):484–512, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Bigot:2013:GPI

[BGL13] J. Bigot, R. Gouet, and A. López. Geometric PCA of images. SIAM Journal on Imaging Sciences, 6(4):1851–1879, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Bubba:2021:DNN

[BGL^{+}21] Tatiana A. Bubba, Mathilde Galinier, Matti Lassas, Marco Prato, Luca Ratti, and Samuli Siltanen. Deep neural networks for inverse problems with pseudodifferential operators: an application to limited-angle tomography. SIAM Journal on Imaging Sciences, 14(2):470– 505, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Bal:2014:IAC

[BGM14] Guillaume Bal, Chenxi Guo, and François Monard. Imaging of anisotropic conductivities from current densities in two dimensions. SIAM Journal on Imaging Sciences, 7(4): 2538–2557, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Burger:2016:SDU

[BGM^{+}16] Martin Burger, Guy Gilboa, Michael Moeller, Lina Eckardt, and Daniel Cremers. Spectral decompositions using onehomogeneous functionals. SIAM Journal on Imaging Sciences, 9 (3):1374–1408, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Borcea:2023:WID

[BGMZ23] Liliana Borcea, Josselin Garnier, Alexander V. Mamonov, and Jörn Zimmerling. Waveform inversion with a data driven estimate of the internal wave. SIAM Journal on Imaging Sciences, 16(1): 280–312, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1517342.

Borcea:2017:RAP

[BGP^{+}17] L. Borcea, J. Garnier, G. Papanicolaou, K. Solna, and C. Tsogka. Resolution analysis of passive synthetic aperture imaging of fast moving objects. SIAM Journal on Imaging Sciences, 10(2):665–710, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Barnett:2017:RSC

[BGPS17] Alex Barnett, Leslie Greengard, Andras Pataki, and Marina Spivak. Rapid solution of the cryo-EM reconstruction problem by frequency marching. SIAM Journal on Imaging Sciences, 10(3):1170–1195, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Bigot:2009:ETR

[BGV09] Jérémie Bigot, Fabrice Gamboa, and Myriam Vimond. Estimation of translation, rotation, and scaling between noisy images using the Fourier–Mellin transform. SIAM Journal on Imaging Sciences, 2(2):614–645, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Bredies:2012:TVB

[BH12] K. Bredies and M. Holler. A total variation-based JPEG decompression model. SIAM Journal on Imaging Sciences, 5 (1):366–393, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Bredies:2015:TBFa

[BH15a] Kristian Bredies and Martin Holler. A TGV-based framework for variational image decompression, zooming, and reconstruction. Part I: Analytics. SIAM Journal on Imaging Sciences, 8(4):2814–2850, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Bredies:2015:TBFb

[BH15b] Kristian Bredies and Martin Holler. A TGV-based framework for variational image decompression, zooming, and reconstruction. Part II: Numerics. SIAM Journal on Imaging Sciences, 8(4):2851–2886, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Bastounis:2017:AUR

[BH17] Alexander Bastounis and Anders C. Hansen. On the absence of uniform recovery in many real-world applications of compressed sensing and the restricted isometry property and nullspace property in levels. SIAM Journal on Imaging Sciences, 10(1):335–371, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Batard:2021:DVC

[BHB21] Thomas Batard, Gloria Haro, and Coloma Ballester. DIP-VBTV: a color image restoration model combining a deep image prior and a vector bundle total variation. SIAM Journal on Imaging Sciences, 14(4): 1816–1847, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Bungert:2021:NPM

[BHFPG21] Leon Bungert, Ester Hait-Fraenkel, Nicolas Papadakis, and Guy Gilboa. Nonlinear power method for computing eigenvectors of proximal operators and neural networks. SIAM Journal on Imaging Sciences, 14(3):1114–1148, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Barnsley:2011:HTF

[BHI11] Michael F. Barnsley, Brendan Harding, and Konstantin Igudesman. How to transform and filter images using iterated function systems. SIAM Journal on Imaging Sciences, 4(4):1001–1028, ???? 2011. CODEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i4/p1001_s1.

Bauer:2012:ALM

[BHM12] Martin Bauer, Philipp Harms, and Peter W. Michor. Almost local metrics on shape space of hypersurfaces in n-space. SIAM Journal on Imaging Sciences, 5 (1):244–310, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Burger:2009:CHI

[BHS09] Martin Burger, Lin He, and Carola-Bibiane Schönlieb. Cahn– Hilliard inpainting and a generalization for grayvalue images. SIAM Journal on Imaging Sciences, 2(4):1129–1167, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Bendory:2022:CRM

[BHS22] Tamir Bendory, Ido Hadi, and Nir Sharon. Compactification of the rigid motions group in image processing. SIAM Journal on Imaging Sciences, 15(3): 1041–1078, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1429448.

Beinert:2023:APR

[BHS23] Robert Beinert, Cosmas Heiss, and Gabriele Steidl. On assignment problems related to Gromov–Wasserstein distances on the real line. SIAM Journal on Imaging Sciences, 16(2): 1028–1032, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1497808.

Bredies:2018:TGV

[BHSW18] K. Bredies, M. Holler, M. Storath, and A. Weinmann. Total generalized variation for manifoldvalued data. SIAM Journal on Imaging Sciences, 11(3):1785– 1848, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Barmpoutis:2012:ASP

[BHV12] Angelos Barmpoutis, Jeffrey Ho, and Baba C. Vemuri. Approximating symmetric positive semidefinite tensors of even order. SIAM Journal on Imaging Sciences, 5(1):434–464, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Bal:2015:DRU

[BI15] Guillaume Bal and Sébastien Imperiale. Displacement reconstructions in ultrasound elastography. SIAM Journal on Imaging Sciences, 8(2):1070–1089, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Bauer:2015:DDM

[BJM15] Martin Bauer, Sarang Joshi, and Klas Modin. Diffeomorphic density matching by optimal information transport. SIAM Journal on Imaging Sciences, 8 (3):1718–1751, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Borcea:2015:RAI

[BK15] Liliana Borcea and Ilker Kocyigit. Resolution analysis of imaging with ℓ_{1} optimization. SIAM Journal on Imaging Sciences, 8(4):3015–3050, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Borcea:2017:IRM

[BK17] Liliana Borcea and Ilker Kocyigit. Imaging in random media with convex optimization. SIAM Journal on Imaging Sciences, 10(1):147–190, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Borcea:2018:MMV

[BK18] Liliana Borcea and Ilker Kocyigit. A multiple measurement vector approach to synthetic aperture radar imaging. SIAM Journal on Imaging Sciences, 11(1):770–801, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Bian:2016:SNP

[BKBD16] Xiao Bian, Hamid Krim, Alex Bronstein, and Liyi Dai. Sparsity and nullity: Paradigms for analysis dictionary learning. SIAM Journal on Imaging Sciences, 9(3):1107–1126, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Bredies:2010:TGV

[BKP10] Kristian Bredies, Karl Kunisch, and Thomas Pock. Total generalized variation. SIAM Journal on Imaging Sciences, 3(3): 492–526, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Bosch:2014:FSC

[BKSW14] Jessica Bosch, David Kay, Martin Stoll, and Andrew J. Wathen. Fast solvers for Cahn– Hilliard inpainting. SIAM Journal on Imaging Sciences, 7(1): 67–97, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Bao:2014:NFI

[BL14a] Gang Bao and Peijun Li. Nearfield imaging of infinite rough surfaces in dielectric media. SIAM Journal on Imaging Sciences, 7(2):867–899, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Bertalmio:2014:DID

[BL14b] Marcelo Bertalmío and Stacey Levine. Denoising an image by denoising its curvature image. SIAM Journal on Imaging Sciences, 7(1):187–211, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Brito-Loeza:2010:MAH

[BLC10] Carlos Brito-Loeza and Ke Chen.

Multigrid algorithm for high order denoising. SIAM Journal on Imaging Sciences, 3(3):363–389, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Bao:2014:NRE

[BLM14] Gang Bao, Junshan Lin, and Séraphin M. Mefire. Numerical reconstruction of electromagnetic inclusions in three dimensions. SIAM Journal on Imaging Sciences, 7(1):558–577, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Bauermeister:2022:LCC

[BLM^{+}22] Hartmut Bauermeister, Emanuel Laude, Thomas Möllenhoff, Michael Moeller, and Daniel Cremers. Lifting the convex conjugate in Lagrangian relaxations: A tractable approach for continuous Markov random fields. SIAM Journal on Imaging Sciences, 15(3): 1253–1281, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1433241.

Bergmann:2014:SOD

[BLSW14] Ronny Bergmann, Friederike Laus, Gabriele Steidl, and Andreas Weinmann. Second order differences of cyclic data and applications in variational denoising. SIAM Journal on Imaging Sciences, 7(4):2916–2953, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Beltrametti:2013:HTS

[BMP13] M. C. Beltrametti, A. M. Massone, and M. Piana. Hough transform of special classes of curves. SIAM Journal on Imaging Sciences, 6(1):391–412, ????

2013. CODEN SJISBI. ISSN 1936-4954.

Borcea:2016:SAI

[BMPT16] Liliana Borcea, Miguel Moscoso, George Papanicolaou, and Chrysoula Tsogka. Synthetic aperture imaging of directionand frequency-dependent reflectivities. SIAM Journal on Imaging Sciences, 9(1):52–81, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Bretin:2023:SFE

[BMS23] Elie Bretin, Pierre Millien, and Laurent Seppecher. Stability for finite element discretization of some inverse parameter problems from internal data: Application to elastography. SIAM Journal on Imaging Sciences, 16(1):340–367, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1428522.

Brune:2009:DIM

[BMW09] Christoph Brune, Helmut Maurer, and Marcus Wagner. Detection of intensity and motion edges within optical flow via multidimensional control. SIAM Journal on Imaging Sciences, 2(4):1190–1210, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Brazey:2014:NSM

[BP14] Denis Brazey and Bruno Portier. A new spherical mixture model for head detection in depth images. SIAM Journal on Imaging Sciences, 7(4):

2423–2447, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Bergounioux:2018:AIC

[BP18] Maïtine Bergounioux and E. Papoutsellis. An anisotropic infconvolution BV type model for dynamic reconstruction. SIAM Journal on Imaging Sciences, 11(1):129–163, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Borcea:2008:EII

[BPG08] Liliana Borcea, George Papanicolaou, and Fernando Guevara Vasquez. Edge illumination and imaging of extended reflectors. SIAM Journal on Imaging Sciences, 1(1):75–114, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Bui:2021:WDA

[BPLX21] Kevin Bui, Fredrick Park, Yifei Lou, and Jack Xin. A weighted difference of anisotropic and isotropic total variation for relaxed Mumford–Shah color and multiphase image segmentation. SIAM Journal on Imaging Sciences, 14(3):1078–1113, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Berthier:2022:QIB

[BPP22] Michel Berthier, Nicoletta Prencipe, and Edoardo Provenzi. A quantum information-based refoundation of color perception concepts. SIAM Journal on Imaging Sciences, 15(4): 1944–1976, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1476071.

Bergmann:2016:PDR

[BPS16] Ronny Bergmann, Johannes Persch, and Gabriele Steidl. A parallel Douglas-Rachford algorithm for minimizing ROFlike functionals on images with values in symmetric Hadamard manifolds. SIAM Journal on Imaging Sciences, 9(3):901–937, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Borcea:2011:ATF

[BPT11] L. Borcea, G. Papanicolaou, and C. Tsogka. Adaptive timefrequency detection and filtering for imaging in heavy clutter. SIAM Journal on Imaging Sciences, 4(3):827–849, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p827_s1.

Beinert:2022:TVB

[BQ22] Robert Beinert and Michael Quellmalz. Total variationbased reconstruction and phase retrieval for diffraction tomography. SIAM Journal on Imaging Sciences, 15(3):1373– 1399, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1474382.

Bahmani:2015:LBD

[BR15] Sohail Bahmani and Justin Romberg. Lifting for blind deconvolution in random mask imaging: Identifiability and convex relaxation. SIAM Journal on Imaging Sciences, 8(4): 2203–2238, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Bar:2009:GNT

[BS09] Leah Bar and Guillermo Sapiro. Generalized Newton-type methods for energy formulations in image processing. SIAM Journal on Imaging Sciences, 2(2): 508–531, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Bosch:2015:FIM

[BS15] Jessica Bosch and Martin Stoll. A fractional inpainting model based on the vector-valued Cahn–Hilliard equation. SIAM Journal on Imaging Sciences, 8 (4):2352–2382, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Baldassari:2021:MSC

[BS21a] Lorenzo Baldassari and Andrea Scapin. Multi-scale classification for electrosensing. SIAM Journal on Imaging Sciences, 14(1):26–57, ???? 2021. CO-DEN SJISBI. ISSN 1936-4954.

Bar:2021:SSP

[BS21b] Leah Bar and Nir Sochen. Strong solutions for PDE-based tomography by unsupervised learning. SIAM Journal on Imaging Sciences, 14(1):128– 155, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Beck:2009:FIS

[BT09] Amir Beck and Marc Teboulle. A fast iterative shrinkagethresholding algorithm for linear inverse problems. SIAM Journal on Imaging Sciences, 2 (1):183–202, ???? 2009. CO-DEN SJISBI. ISSN 1936-4954.

Bergmann:2018:GFM

[BT18] Ronny Bergmann and Daniel Tenbrinck. A graph framework for manifold-valued data. SIAM Journal on Imaging Sciences, 11(1):325–360, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Bardsley:2016:IPC

[BV16] Patrick Bardsley and Fernando Guevara Vasquez. Imaging with power controlled source pairs. SIAM Journal on Imaging Sciences, 9(1):185–211, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Budd:2023:JRS

[BvGL^{+}23] Jeremy M. Budd, Yves van Gennip, Jonas Latz, Simone Parisotto, and Carola-Bibiane Schönlieb. Joint reconstructionsegmentation on graphs. SIAM Journal on Imaging Sciences, 16(2):911–947, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M151546X.

Boyer:2014:AVD

[BWB14] Claire Boyer, Pierre Weiss, and Jérémie Bigot. An algorithm for variable density sampling with block-constrained acquisition. SIAM Journal on Imaging Sciences, 7(2):1080–1107, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Bao:2018:URS

[BZ18] Gang Bao and Lei Zhang. Uniqueness results for scattering and inverse scattering by infinite rough surfaces with tapered wave incidence. SIAM Journal on Imaging Sciences, 11(1):361–375, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Bai:2016:ACA

[BZNC16] Minru Bai, Xiongjun Zhang, Guyan Ni, and Chunfeng Cui. An adaptive correction approach for tensor completion. SIAM Journal on Imaging Sciences, 9(3):1298–1323, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Cohen:2021:MHG

[CALG21] Ido Cohen, Omri Azencot, Pavel Lifshits, and Guy Gilboa. Modes of homogeneous gradient flows. SIAM Journal on Imaging Sciences, 14(3):913– 945, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Carasso:2010:BSL

[Car10] Alfred S. Carasso. Bochner subordination, logarithmic diffusion equations, and blind deconvolution of Hubble Space Telescope imagery and other scientific data. SIAM Journal on Imaging Sciences, 3(4): 954–980, ???? 2010. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v3/i4/p954_s1.

Craizer:2008:ADC

[CAT08] Marcos Craizer, Moacyr Alvim, and Ralph Teixeira. Area distances of convex plane curves and improper affine spheres. SIAM Journal on Imaging Sciences, 1(3):209–227, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Cheney:2011:TWD

[CB11] Margaret Cheney and Brett Borden. Theory of waveformdiverse moving-target spotlight synthetic-aperture radar. SIAM Journal on Imaging Sciences, 4(4):1180–1199, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i4/p1180_s1.

Chenot:2018:BSS

[CB18] Cécile Chenot and Jérôme Bobin. Blind source separation with outliers in transformed domains. SIAM Journal on Imaging Sciences, 11(2):1524–1559, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Cyriac:2014:NVF

[CBB14] Praveen Cyriac, Thomas Batard, and Marcelo Bertalmío. A nonlocal variational formulation for the improvement of tone mapped images. SIAM Journal on Imaging Sciences, 7(4): 2340–2363, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Choi:2018:PMJ

[CBZ18] Jae Kyu Choi, Chenglong Bao, and Xiaoqun Zhang. PET–MRI joint reconstruction by joint sparsity based tight frame regularization. SIAM Journal on Imaging Sciences, 11(2):1179– 1204, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Calderero:2014:MAI

[CC14] Felipe Calderero and Vicent Caselles. Multiscale analysis for images on Riemannian manifolds. SIAM Journal on Imaging Sciences, 7(2):1108–1170, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Cloninger:2014:SFI

[CCBB14] Alexander Cloninger, Wojciech Czaja, Ruiliang Bai, and Peter J. Basser. Solving 2D Fredholm integral from incomplete measurements using compressive sensing. SIAM Journal on Imaging Sciences, 7(3):1775– 1798, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Cardelino:2013:CSO

[CCBR13] Juan Cardelino, Vicent Caselles, Marcelo Bertalmío, and Gregory Randall. A contrario selection of optimal partitions for image segmentation. SIAM Journal on Imaging Sciences, 6 (3):1274–1317, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Crivelli:2013:MTM

[CCFBY13] Tomás Crivelli, Bruno Cernuschi-Frias, Patrick Bouthemy, and Jian-Feng Yao. Motion textures: Modeling, classification, and segmentation using mixedstate Markov random fields. SIAM Journal on Imaging Sciences, 6(4):2484–2520, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Chauffert:2014:VDS

[CCKW14] Nicolas Chauffert, Philippe Ciuciu, Jonas Kahn, and Pierre Weiss. Variable density sampling with continuous trajectories. SIAM Journal on Imaging Sciences, 7(4):1962–1992, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Cai:2013:VSM

[CCMS13] Xiaohao Cai, Raymond Chan, Serena Morigi, and Fiorella Sgallari. Vessel segmentation in medical imaging using a tightframe–based algorithm. SIAM Journal on Imaging Sciences, 6 (1):464–486, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Chen:2015:IPA

[CCMY15] Caihua Chen, Raymond H.

Chan, Shiqian Ma, and Junfeng Yang. Inertial proximal ADMM for linearly constrained separable convex optimization. SIAM Journal on Imaging Sciences, 8 (4):2239–2267, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Chambolle:2012:CAM

[CCP12] Antonin Chambolle, Daniel Cremers, and Thomas Pock. A convex approach to minimal partitions. SIAM Journal on Imaging Sciences, 5(4): 1113–1158, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Chouzenoux:2023:CRP

[CCPS23] Emilie Chouzenoux, Andrés Contreras, Jean-Christophe Pesquet, and Marion Savanier. Convergence results for primaldual algorithms in the presence of adjoint mismatch. SIAM Journal on Imaging Sciences, 16(1):1–34, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1490223.

Chang:2020:HCU

[CCQY20] Jingya Chang, Yannan Chen, Liqun Qi, and Hong Yan. Hypergraph clustering using a new Laplacian tensor with applications in image processing. SIAM Journal on Imaging Sciences, 13(3):1157–1178, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Carson:2012:CIP

[CCR^{+}12] William R. Carson, Minhua Chen, Miguel R. D. Rodrigues, Robert Calderbank, and Lawrence Carin. Communications-inspired projection design with application to compressive sensing. SIAM Journal on Imaging Sciences, 5 (4):1185–1212, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Cai:2022:LHK

[CCST22] Tianji Cai, Junyi Cheng, Bernhard Schmitzer, and Matthew Thorpe. The linearized Hellinger–Kantorovich distance. SIAM Journal on Imaging Sciences, 15(1):45–83, ???? 2022. CODEN SJISBI. ISSN 1936-

4954. URL https://epubs.
siam.org/doi/10.1137/21M1400080.

Cai:2020:DDT

[CCW20] Jian-Feng Cai, Jae Kyu Choi, and Ke Wei. Data driven tight frame for compressed sensing MRI reconstruction via off-the-grid regularization. SIAM Journal on Imaging Sciences, 13(3): 1272–1301, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Cai:2013:TSI

[CCZ13] Xiaohao Cai, Raymond Chan, and Tieyong Zeng. A two-stage image segmentation method using a convex variant of the Mumford–Shah model and thresholding. SIAM Journal on Imaging Sciences, 6(1):368–390, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Chevallier:2021:CFL

[CDA21] Juliette Chevallier, Vianney Debavelaere, and Stéphanie Allassonnière. A coherent framework for learning spatiotemporal piecewise-geodesic trajectories from longitudinal manifoldvalued data. SIAM Journal on Imaging Sciences, 14(1):349– 388, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Chen:2016:FOD

[CDH16] Yannan Chen, Yu-Hong Dai, and Deren Han. Fiber orientation distribution estimation using a Peaceman–Rachford splitting method. SIAM Journal on Imaging Sciences, 9(2):573–604, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Chow:2021:SLT

[CDH^{+}21] Yat Tin Chow, Youjun Deng, Youzi He, Hongyu Liu, and Xianchao Wang. Surface-localized transmission eigenstates, superresolution imaging, and pseudo surface plasmon modes. SIAM Journal on Imaging Sciences, 14(3):946–975, ???? 2021. CO-DEN SJISBI. ISSN 1936-4954.

Chen:2013:PSG

[CDHS13] Yannan Chen, Yuhong Dai, Deren Han, and Wenyu Sun. Positive semidefinite generalized diffusion tensor imaging via quadratic semidefinite programming. SIAM Journal on Imaging Sciences, 6(3):1531–1552, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Chen:2023:HFL

[CDLZN23] Shi Chen, Zhiyan Ding, Qin Li, and Leonardo Zepeda-Núñez. High-frequency limit of the inverse scattering problem: Asymptotic convergence from inverse Helmholtz to inverse Liouville. SIAM Journal on Imaging Sciences, 16 (1):111–143, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M147075X.

Catala:2019:LRA

[CDP19] Paul Catala, Vincent Duval, and Gabriel Peyré. A low-rank approach to off-the-grid sparse superresolution. SIAM Journal on Imaging Sciences, 12(3): 1464–1500, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Carpio:2016:NIT

[CDRS16] A. Carpio, T. G. Dimiduk, M. L. Rapún, and V. Selgas. Noninvasive imaging of threedimensional micro and nanostructures by topological methods. SIAM Journal on Imaging Sciences, 9(3):1324–1354, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Calatroni:2017:ICD

[CDS17] Luca Calatroni, Juan Carlos De Los Reyes, and Carola-Bibiane Schönlieb. Infimal convolution of data discrepancies for mixed noise removal. SIAM Journal on Imaging Sciences, 10(3): 1196–1233, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Carpio:2018:OML

[CDSV18] A. Carpio, T. G. Dimiduk, V. Selgas, and P. Vidal. Optimization methods for in-line holography. SIAM Journal on Imaging Sciences, 11(2):923– 956, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Calder:2012:CAS

[CE12] Jeff Calder and Selim Esedoğlu.

On the circular area signature for graphs. SIAM Journal on Imaging Sciences, 5(4):1355– 1379, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Chang:2019:BPP

[CEM19] Huibin Chang, Pablo Enfedaque, and Stefano Marchesini. Blind ptychographic phase retrieval via convergent alternating direction method of multipliers. SIAM Journal on Imaging Sciences, 12(1): 153–185, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Cohen:2021:RDF

[CEM21] Regev Cohen, Michael Elad, and Peyman Milanfar. Regularization by denoising via fixedpoint projection (RED-PRO). SIAM Journal on Imaging Sciences, 14(3):1374–1406, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Candes:2013:PRM

[CESV13] Emmanuel J. Candès, Yonina C. Eldar, Thomas Strohmer, and Vladislav Voroninski. Phase retrieval via matrix completion. SIAM Journal on Imaging Sciences, 6(1):199–225, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Chevallier:2018:IME

[CF18] Emmanuel Chevallier and Ivar Farup. Interpolation of the MacAdam ellipses. SIAM Journal on Imaging Sciences, 11(3): 1979–2000, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Canham:2023:UDF

[CFBP23] Trevor Canham, Adrián Martín Fernández, Marcelo Bertalmío, and Javier Portilla. Using decoupled features for photorealistic style transfer. SIAM Journal on Imaging Sciences, 16(3): 1687–1726, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1512491.

Capdeboscq:2009:IMN

[CFdGK09] Y. Capdeboscq, J. Fehrenbach, F. de Gournay, and O. Kavian. Imaging by modification: Numerical reconstruction of local conductivities from corresponding power density measurements. SIAM Journal on Imaging Sciences, 2(4):1003– 1030, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Caselles:2009:ACS

[CFM09] Vicent Caselles, Gabriele Facciolo, and Enric Meinhardt. Anisotropic Cheeger sets and applications. SIAM Journal on Imaging Sciences, 2(4):1211– 1254, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Cherfils:2015:BEG

[CFM15] Laurence Cherfils, Hussein Fakih, and Alain Miranville. On the Bertozzi–Esedoğlu– Gillette–Cahn–Hilliard equation with logarithmic nonlinear terms. SIAM Journal on Imaging Sciences, 8(2):1123–1140, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Clerici:2020:AAM

[CFM^{+}20] Francesco Clerici, Nicola Ferro, Stefania Marconi, Stefano Micheletti, Erika Negrello, and Simona Perotto. Anisotropic adapted meshes for image segmentation: Application to three-dimensional medical data. SIAM Journal on Imaging Sciences, 13(4):2189–2212, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Citti:2016:SRM

[CFSS16] G. Citti, B. Franceschiello, G. Sanguinetti, and A. Sarti. Sub-Riemannian mean curvature flow for image processing. SIAM Journal on Imaging Sciences, 9(1):212–237, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Combettes:2019:PAS

[CG19] Patrick L. Combettes and Lilian E. Glaudin. Proximal activation of smooth functions in splitting algorithms for convex image recovery. SIAM Journal on Imaging Sciences, 12(4): 1905–1935, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Cao:2011:GGE

[CGMP11] Frédéric Cao, Yann Gousseau, Simon Masnou, and Patrick Pérez. Geometrically guided exemplar-based inpainting. SIAM Journal on Imaging Sciences, 4(4):1143–1179, ???? 2011. CODEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.

siam.org/siims/resource/1/
sjisbi/v4/i4/p1143_s1.

Couprie:2013:DCT

[CGN^{+}13] Camille Couprie, Leo Grady, Laurent Najman, Jean-Christophe Pesquet, and Hugues Talbot. Dual constrained TV-based regularization on graphs. SIAM Journal on Imaging Sciences, 6 (3):1246–1273, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Chen:2019:NVM

[CGÖ19] Chong Chen, Barbara Gris, and Ozan Öktem. A new variational model for joint image reconstruction and motion estimation in spatiotemporal imaging. SIAM Journal on Imaging Sciences, 12(4):1686–1719, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Couprie:2011:CCM

[CGTN11] Camille Couprie, Leo Grady, Hugues Talbot, and Laurent Najman. Combinatorial continuous maximum flow. SIAM Journal on Imaging Sciences, 4 (3):905–930, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p905_s1.

Chen:2016:DIM

[CH16] Zhiming Chen and Guanghui Huang. A direct imaging method for electromagnetic scattering data without phase information. SIAM Journal on Imaging Sciences, 9(3):1273– 1297, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Chen:2014:RGI

[Che14] Dai-Qiang Chen. Regularized generalized inverse accelerating linearized alternating minimization algorithm for framebased Poissonian image deblurring. SIAM Journal on Imaging Sciences, 7(2):716–739, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Chen:2012:FAI

[CHH^{+}12] Yunmei Chen, William Hager, Feng Huang, Dzung Phan, Xiaojing Ye, and Wotao Yin. Fast algorithms for image reconstruction with application to partially parallel MR imaging. SIAM Journal on Imaging Sciences, 5(1):90–118, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Cai:2021:RCD

[CHHN21] HanQin Cai, Keaton Hamm, Longxiu Huang, and Deanna Needell. Robust CUR decomposition: Theory and imaging applications. SIAM Journal on Imaging Sciences, 14(4):1472– 1503, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Choi:2023:IPP

[CHKL23] Doosung Choi, Johan Helsing, Sangwoo Kang, and Mikyoung Lim. Inverse problem for a planar conductivity inclusion. SIAM Journal on Imaging Sciences, 16 (2):969–995, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1522395.

Choi:2016:SCP

[CHL16] Gary Pui-Tung Choi, Kin Tat Ho, and Lok Ming Lui. Spherical conformal parameterization of genus-0 point clouds for meshing. SIAM Journal on Imaging Sciences, 9(4):1582– 1618, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Coletta:2013:FIT

[CHM13] Meredith L. Coletta, R. Andrew Hicks, and Shari Moskow. The Frobenius integrability theorem and the blind-spot problem for motor vehicles. SIAM Journal on Imaging Sciences, 6(3): 1367–1384, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Calvetti:2009:CGH

[CHPS09] Daniela Calvetti, Harri Hakula, Sampsa Pursiainen, and Erkki Somersalo. Conditionally Gaussian hypermodels for cerebral source localization. SIAM Journal on Imaging Sciences, 2(3): 879–909, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Chow:2021:DSM

[CHZ21] Yat Tin Chow, Fuqun Han, and Jun Zou. A direct sampling method for the inversion of the Radon transform. SIAM Journal on Imaging Sciences, 14(3): 1004–1038, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Clason:2012:SNM

[CJ12] Christian Clason and Bangti Jin. A semismooth Newton method for nonlinear parameter identification problems with impulsive noise. SIAM Journal on Imaging Sciences, 5(2):505–536, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Chambolle:2014:ABN

[CJ14] Antonin Chambolle and Khalid Jalalzai. Adapted basis for nonlocal reconstruction of missing spectrum. SIAM Journal on Imaging Sciences, 7(3):1484– 1502, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Clason:2010:SNM

[CJK10] Christian Clason, Bangti Jin, and Karl Kunisch. A semismooth Newton method for L^{1} data fitting with automatic choice of regularization parameters and noise calibration. SIAM Journal on Imaging Sciences, 3(2):199–231, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Chouzenoux:2013:MMS

[CJPT13] Emilie Chouzenoux, Anna Jezierska, Jean-Christophe Pesquet, and Hugues Talbot. A majorize-minimize subspace approach for ℓ_{2}-ℓ_{0} image regularization. SIAM Journal on Imaging Sciences, 6(1):563–591, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Chouzenoux:2015:CAI

[CJPT15] Emilie Chouzenoux, Anna Jezierska, Jean-Christophe Pesquet, and Hugues Talbot. A convex approach for image restoration with exact Poisson– Gaussian likelihood. SIAM Journal on Imaging Sciences, 8 (4):2662–2682, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Choy:2012:RCA

[CJT^{+}12] Siu Kai Choy, Ningning Jia, Chong Sze Tong, Man Lai Tang, and Edmund Y. Lam. A robust computational algorithm for inverse photomask synthesis in optical projection lithography. SIAM Journal on Imaging Sciences, 5(2):625–651, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Clarkson:2009:GCP

[CK09] Eric Clarkson and Matthew A.

Kupinski. Global compartmental pharmacokinetic models for spatiotemporal SPECT and PET imaging. SIAM Journal on Imaging Sciences, 2(1):203–225, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Chevallier:2017:KDE

[CKA17] Emmanuel Chevallier, Emmanuel Kalunga, and Jesús Angulo. Kernel density estimation on spaces of Gaussian distributions and symmetric positive definite matrices. SIAM Journal on Imaging Sciences, 10(1): 191–215, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Camassa:2017:GLS

[CKL17] Roberto Camassa, Dongyang Kuang, and Long Lee. A geodesic landmark shooting algorithm for template matching and its applications. SIAM Journal on Imaging Sciences, 10(1):303–334, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Chen:2013:PMC

[CLC13] Pengwen Chen, Ching-Long Lin, and I-Liang Chern. A perfect match condition for pointset matching problems using the optimal mass transport approach. SIAM Journal on Imaging Sciences, 6(2):730–764, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Chang:2018:TVB

[CLDM18] Huibin Chang, Yifei Lou, Yuping Duan, and Stefano Marchesini. Total variation-based phase retrieval for Poisson noise removal. SIAM Journal on Imaging Sciences, 11(1):24–55, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Cha:2014:IZC

[CLK14] Youngjoon Cha, Gi Yun Lee, and Seongjai Kim. Image zooming by curvature interpolation and iterative refinement. SIAM Journal on Imaging Sciences, 7 (2):1284–1308, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Chambolle:2011:UFD

[CLL11] Antonin Chambolle, Stacey E. Levine, and Bradley J. Lucier.

An upwind finite-difference method for total variationbased image smoothing. SIAM Journal on Imaging Sciences, 4 (1):277–299, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p277_s1.

Choi:2015:FFL

[CLL15] Pui Tung Choi, Ka Chun Lam, and Lok Ming Lui. FLASH: Fast Landmark Aligned Spherical Harmonic parameterization for genus-0 closed brain surfaces. SIAM Journal on Imaging Sciences, 8(1):67–94, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Choi:2020:PGC

[CLLGL20] Gary P. T. Choi, Yusan Leung-Liu, Xianfeng Gu, and Lok Ming Lui. Parallelizable global conformal parameterization of simply-connected surfaces via partial welding. SIAM Journal on Imaging Sciences, 13(3):1049–1083, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Cuel:2015:RGE

[CLMT15] Louis Cuel, Jacques-Olivier Lachaud, Quentin Mérigot, and Boris Thibert. Robust geometry estimation using the generalized Voronoi covariance measure. SIAM Journal on Imaging Sciences, 8(2):1293–1314, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Calatroni:2019:FSV

[CLPS19] Luca Calatroni, Alessandro Lanza, Monica Pragliola, and Fiorella Sgallari. A flexible space-variant anisotropic regularization for image restoration with automated parameter selection. SIAM Journal on Imaging Sciences, 12(2):1001–1037, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Chen:2019:TSA

[CLY19] Yunmei Chen, Bin Li, and Xiaojing Ye. A two-stage algorithm for joint multimodal image reconstruction. SIAM Journal on Imaging Sciences, 12(3): 1425–1463, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Chen:2021:LDA

[CLYZ21] Yunmei Chen, Hongcheng Liu, Xiaojing Ye, and Qingchao Zhang. Learnable descent algorithm for nonsmooth nonconvex image reconstruction. SIAM Journal on Imaging Sciences, 14(4):1532–1564, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Cheng:2020:SEN

[CM20] Xiuyuan Cheng and Gal Mishne. Spectral embedding norm: Looking deep into the spectrum of the graph Laplacian. SIAM Journal on Imaging Sciences, 13(2):1015–1048, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Chang:2018:VPR

[CMLZ18] Huibin Chang, Stefano Marchesini, Yifei Lou, and Tieyong Zeng. Variational phase retrieval with globally convergent preconditioned proximal algorithm. SIAM Journal on Imaging Sciences, 11(1):56–93, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Chai:2014:ISL

[CMP14] Anwei Chai, Miguel Moscoso, and George Papanicolaou. Imaging strong localized scatterers with sparsity promoting optimization. SIAM Journal on Imaging Sciences, 7(2):1358– 1387, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Calder:2010:ISS

[CMY10] J. Calder, A. Mansouri, and A. Yezzi. Image sharpening via Sobolev gradient flows. SIAM Journal on Imaging Sciences, 3 (4):981–1014, ???? 2010. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v3/i4/p981_s1.

Chung:2017:MEC

[CN17] Julianne Chung and Linh Nguyen. Motion estimation and correction in photoacoustic tomographic reconstruction. SIAM Journal on Imaging Sciences, 10(1):216–242, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Chen:2022:CII

[CN22] Junren Chen and Michael K. Ng. Color image inpainting via robust pure quaternion matrix completion: Error bound and weighted loss. SIAM Journal on Imaging Sciences, 15(3): 1469–1498, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1476897.

Chung:2010:NAP

[CNS10] Julianne Chung, James G. Nagy, and Ioannis Sechopoulos. Numerical algorithms for polyenergetic digital breast tomosynthesis reconstruction. SIAM Journal on Imaging Sciences, 3(1):133–152, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Chen:2018:IIR

[CÖ18] Chong Chen and Ozan Öktem. Indirect image registration with large diffeomorphic deformations. SIAM Journal on Imaging Sciences, 11(1):575–617, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Colbrook:2022:WLC

[Col22] Matthew J. Colbrook. WARPd: A linearly convergent firstorder primal-dual algorithm for inverse problems with approximate sharpness conditions. SIAM Journal on Imaging Sciences, 15(3):1539– 1575, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1455000.

Condat:2017:DTV

[Con17] Laurent Condat. Discrete total variation: New definition and minimization. SIAM Journal on Imaging Sciences, 10(3): 1258–1290, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Cai:2009:LBI

[COS09] Jian-Feng Cai, Stanley Osher, and Zuowei Shen. Linearized Bregman iterations for framebased image deblurring. SIAM Journal on Imaging Sciences, 2 (1):226–252, ???? 2009. CO-DEN SJISBI. ISSN 1936-4954.

Cuturi:2016:SDA

[CP16] Marco Cuturi and Gabriel Peyré. A smoothed dual approach for variational Wasserstein problems. SIAM Journal on Imaging Sciences, 9(1): 320–343, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Chambolle:2021:LCD

[CP21] Antonin Chambolle and Thomas Pock. Learning consistent discretizations of the total variation. SIAM Journal on Imaging Sciences, 14(2):778–813, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Chaux:2009:NIA

[CPP09] Caroline Chaux, Jean-Christophe Pesquet, and Nelly Pustelnik. Nested iterative algorithms for convex constrained image recovery problems. SIAM Journal on Imaging Sciences, 2(2):730–762, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Crisci:2021:RAS

[CPRS21] S. Crisci, M. Piana, V. Ruggiero, and M. Scussolini. A regularized affine-scaling trustregion method for parametric imaging of dynamic PET data. SIAM Journal on Imaging Sciences, 14(1):418–439, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Choi:2014:IPQ

[CPW^{+}14] Jae Kyu Choi, Hyoung Suk Park, Shuai Wang, Yi Wang, and Jin Keun Seo. Inverse problem in quantitative susceptibility mapping. SIAM Journal on Imaging Sciences, 7(3):1669– 1689, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Chen:2021:RKA

[CQ21] Xuemei Chen and Jing Qin. Regularized Kaczmarz algorithms for tensor recovery. SIAM Journal on Imaging Sciences, 14(4):1439–1471, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Choi:2018:DEM

[CR18] Gary P. T. Choi and Chris H. Rycroft. Density-equalizing maps for simply connected open surfaces. SIAM Journal on Imaging Sciences, 11(2):1134– 1178, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Chan:2008:IFH

[CSS08] R. H. Chan, S. Setzer, and G. Steidl. Inpainting by flexible Haar-wavelet shrinkage. SIAM Journal on Imaging Sciences, 1 (3):273–293, ???? 2008. CO-DEN SJISBI. ISSN 1936-4954.

Charon:2013:VRN

[CT13] Nicolas Charon and Alain Trouvé. The varifold representation of nonoriented shapes for diffeomorphic registration. SIAM Journal on Imaging Sciences, 6(4):2547–2580, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Camilli:2017:UAW

[CT17] Fabio Camilli and Silvia Tozza.

A unified approach to the well-posedness of some non-Lambertian models in shapefrom-shading theory. SIAM Journal on Imaging Sciences, 10(1):26–46, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Chang:2015:CRO

[CTWY15] Huibin Chang, Xue-Cheng Tai, Li-Lian Wang, and Danping Yang. Convergence rate of overlapping domain decomposition methods for the Rudin–Osher– Fatemi model based on a dual formulation. SIAM Journal on Imaging Sciences, 8(1):564–591, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Chan:2013:CTV

[CTY13] Raymond H. Chan, Min Tao, and Xiaoming Yuan. Constrained total variation deblurring models and fast algorithms based on alternating direction method of multipliers. SIAM Journal on Imaging Sciences, 6 (1):680–697, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Cheng:2013:NDS

[CV13] Guang Cheng and Baba C. Vemuri. A novel dynamic system in the space of SPD matrices with applications to appearance tracking. SIAM Journal on Imaging Sciences, 6(1): 592–615, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Cassier:2017:IPD

[CV17] Maxence Cassier and Fernando Guevara Vasquez. Imaging polarizable dipoles. SIAM Journal on Imaging Sciences, 10(3):1381–1415, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Choksi:2010:DOD

[CvG10] Rustum Choksi and Yves van Gennip. Deblurring of one dimensional bar codes via total variation energy minimization. SIAM Journal on Imaging Sciences, 3(4):735–764, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Chen:2018:SDN

[CW18] Xiaojun Chen and Robert S. Womersley. Spherical designs and nonconvex minimization for recovery of sparse signals on the sphere. SIAM Journal on Imaging Sciences, 11(2):1390–1415, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Combettes:2022:VIM

[CW22] Patrick L. Combettes and Zev C. Woodstock. A variational inequality model for the construction of signals from inconsistent nonlinear equations. SIAM Journal on Imaging Sciences, 15(1):84– 109, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1420368.

Chau:2019:EPM

[CWR19] Gustavo Chau, Brendt Wohlberg, and Paul Rodriguez. Efficient projection onto the ℓ_{∞,1} mixednorm ball using a Newton root search method. SIAM Journal on Imaging Sciences, 12(1): 604–623, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Chou:2009:SDV

[CY09] Hsiao-Fang Chou and Laurent Younes. Smoothing directional vector fields using dual norms. SIAM Journal on Imaging Sciences, 2(1):41–63, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Chan:2011:ADM

[CYY11] Raymond H. Chan, Junfeng Yang, and Xiaoming Yuan. Alternating direction method for image inpainting in wavelet domains. SIAM Journal on Imaging Sciences, 4(3):807– 826, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i3/p807_s1.

Chan:2014:TSI

[CYZ14] Raymond Chan, Hongfei Yang, and Tieyong Zeng. A two-stage image segmentation method for blurry images with Poisson or multiplicative gamma noise. SIAM Journal on Imaging Sciences, 7(1):98–127, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Chen:2010:SNC

[CZ10] Xiaojun Chen and Weijun Zhou. Smoothing nonlinear conjugate gradient method for image restoration using nonsmooth nonconvex minimization. SIAM Journal on Imaging Sciences, 3(4):765–790, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Debroux:2020:VMD

[DAB^{+}20] Noémie Debroux, John Aston, Fabien Bonardi, Alistair Forbes, Carole Le Guyader, Marina Romanchikova, and Carola-Bibiane Schönlieb. A variational model dedicated to joint segmentation, registration, and atlas generation for shape analysis. SIAM Journal on Imaging Sciences, 13(1):351–380, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Duval:2011:BVA

[DAG11] Vincent Duval, Jean-François Aujol, and Yann Gousseau. A bias-variance approach for the nonlocal means. SIAM Journal on Imaging Sciences, 4(2): 760–788, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i2/p760_s1.

Delbracio:2012:SPS

[DAMM12] Mauricio Delbracio, Andrés Almansa, Jean-Michel Morel, and Pablo Musé. Subpixel point spread function estimation from two photographs at different distances. SIAM Journal on Imaging Sciences, 5(4):1234– 1260, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Darbon:2015:CFD

[Dar15] Jérôme Darbon. On convex finite-dimensional variational methods in imaging sciences and Hamilton–Jacobi equations. SIAM Journal on Imaging Sciences, 8(4):2268–2293, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Devilliers:2017:TEC

[DATP17] Loïc Devilliers, Stéphanie Allassonnière, Alain Trouvé, and Xavier Pennec. Template estimation in computational anatomy: Fréchet means top and quotient spaces are not consistent. SIAM Journal on Imaging Sciences, 10(3):1139–1169, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Daropoulos:2021:SIS

[DAW21] Viktor Daropoulos, Matthias Augustin, and Joachim Weickert. Sparse inpainting with smoothed particle hydrodynamics. SIAM Journal on Imaging Sciences, 14(4):1669–1705, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Droske:2010:HOF

[DB10] Marc Droske and Andrea Bertozzi. Higher-order featurepreserving geometric regularization. SIAM Journal on Imaging Sciences, 3(1):21–51, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Dobrosotskaya:2013:AWG

[DB13] J. A. Dobrosotskaya and A. L. Bertozzi. Analysis of the wavelet Ginzburg–Landau energy in image applications with edges. SIAM Journal on Imaging Sciences, 6(1):698–729, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Duran:2014:NVM

[DBCS14] J. Duran, A. Buades, B. Coll, and C. Sbert. A nonlocal variational model for pansharpening image fusion. SIAM Journal on Imaging Sciences, 7(2):761–796, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Delon:2010:SFL

[DD10] Julie Delon and Agnès Desolneux. Stabilization of flickerlike effects in image sequences through local contrast correction. SIAM Journal on Imaging Sciences, 3(4):703–734, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Delon:2013:PBA

[DD13] Julie Delon and Agnès Desolneux. A patch-based approach for removing impulse or mixed gaussian-impulse noise. SIAM Journal on Imaging Sciences, 6 (2):1140–1174, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Delon:2020:WTD

[DD20] Julie Delon and Agnès Desolneux. A Wasserstein-type distance in the space of Gaussian mixture models. SIAM Journal on Imaging Sciences, 13(2): 936–970, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

DeBortoli:2019:PRI

[DDGL19] Valentin De Bortoli, Agnès Desolneux, Bruno Galerne, and Arthur Leclaire. Patch redundancy in images: a statistical testing framework and some applications. SIAM Journal on Imaging Sciences, 12(2):893– 926, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

DeBortoli:2020:MLE

[DDPV20] Valentin De Bortoli, Alain Durmus, Marcelo Pereyra, and Ana Fernandez Vidal. Maximum likelihood estimation of regularization parameters in high-dimensional inverse problems: an empirical Bayesian approach. Part II: Theoretical analysis. SIAM Journal on Imaging Sciences, 13(4):1990– 2028, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

DelosReyes:2023:BIL

[De 23] Juan Carlos De los Reyes. Bilevel imaging learning problems as mathematical programs with complementarity constraints: Reformulation and theory. SIAM Journal on Imaging Sciences, 16(3):1655– 1686, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1450744.

Demidenko:2009:SHT

[Dem09] Eugene Demidenko. Statistical hypothesis testing for postreconstructed and postregistered medical images. SIAM Journal on Imaging Sciences, 2(4): 1049–1067, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Demanet:2012:BAS

[DFM^{+}12] Laurent Demanet, Matthew Ferrara, Nicholas Maxwell, Jack Poulson, and Lexing Ying. A butterfly algorithm for synthetic aperture radar imaging. SIAM Journal on Imaging Sciences, 5(1):203–243, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Damon:2011:LIF

[DGH11] James Damon, Peter Giblin, and Gareth Haslinger. Local image features resulting from 3dimensional geometric features, illumination, and movement: II. SIAM Journal on Imaging Sciences, 4(1):386–412, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.

siam.org/siims/resource/1/
sjisbi/v4/i1/p386_s1.

Dong:2016:MMN

[DGJS16] Jun-Liang Dong, Junbin Gao, Fujiao Ju, and Jinghua Shen. Modulus methods for nonnegatively constrained image restoration. SIAM Journal on Imaging Sciences, 9(3):1226– 1246, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Donlon:2023:SRS

[DGMW23] Niall Donlon, Romina Gaburro, Shari Moskow, and Isaac Woods. Stability and reconstruction of a special type of anisotropic conductivity in magneto-acoustic tomography with magnetic induction. SIAM Journal on Imaging Sciences, 16(2):614–639, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1512260.

DeLara:2023:DRU

[DGSL23] Lucas De Lara, Alberto González-Sanz, and Jean-Michel Loubes. Diffeomorphic registration using Sinkhorn divergences. SIAM Journal on Imaging Sciences, 16(1):250–279, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1493562.

Deng:2019:NOS

[DGT19] Liang-Jian Deng, Roland Glowinski, and Xue-Cheng Tai. A new operator splitting method for the Euler elastica model for image smoothing. SIAM Journal on Imaging Sciences, 12(2): 1190–1230, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Dreier:2020:EIF

[DH20] Florian Dreier and Markus Haltmeier. Explicit inversion formulas for the twodimensional wave equation from Neumann traces. SIAM Journal on Imaging Sciences, 13(2): 589–608, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Daskalakis:2020:ASR

[DHK20] Emmanouil Daskalakis, Felix J. Herrmann, and Rachel Kuske. Accelerating sparse recovery by reducing chatter. SIAM Journal on Imaging Sciences, 13(3): 1211–1239, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Dong:2009:EPD

[DHN09] Yiqiu Dong, Michael Hintermüller, and Marrick Neri. An efficient primal-dual method for L^{1}TV image restoration. SIAM Journal on Imaging Sciences, 2 (4):1168–1189, ???? 2009. CO-DEN SJISBI. ISSN 1936-4954.

Dong:2019:QMR

[DHP19] Guozhi Dong, Michael Hintermüller, and Kostas Papafitsoros. Quantitative magnetic resonance imaging: From fingerprinting to integrated physics-based models. SIAM Journal on Imaging Sciences, 12(2):927–971, ???? 2019. CO-DEN SJISBI. ISSN 1936-4954.

Darde:2013:SRO

[DHSS13] J. Dardé, N. Hyvönen, A. Seppänen, and S. Staboulis. Simultaneous reconstruction of outer boundary shape and admittivity distribution in electrical impedance tomography. SIAM Journal on Imaging Sciences, 6 (1):176–198, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Dong:2021:CSO

[DHZ21] Guozhi Dong, Michael Hintermueller, and Ye Zhang. A class of second-order geometric quasilinear hyperbolic PDEs and their application in imaging. SIAM Journal on Imaging Sciences, 14(2):645–688, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Dong:2020:CCR

[DJLS20] Bin Dong, Haocheng Ju, Yiping Lu, and Zuoqiang Shi. CURE: Curvature regularization for missing data recovery. SIAM Journal on Imaging Sciences, 13(4):2169–2188, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Dibos:2009:AWD

[DKP09] Françoise Dibos, Georges Koepfler, and Sylvain Pelletier. Adapted windows detection of moving objects in video scenes. SIAM Journal on Imaging Sciences, 2 (1):1–19, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Derfoul:2014:RPR

[DL14] Ratiba Derfoul and Carole Le Guyader. A relaxed problem of registration based on the Saint Venant–Kirchhoff material stored energy for the mapping of mouse brain gene expression data to a neuroanatomical mouse atlas. SIAM Journal on Imaging Sciences, 7(4): 2175–2195, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Debroux:2018:JSR

[DL18a] Noémie Debroux and Carole Le Guyader. A joint segmentation/registration model based on a nonlocal characterization of weighted total variation and nonlocal shape descriptors. SIAM Journal on Imaging Sciences, 11(2):957–990, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Desolneux:2018:SIM

[DL18b] A. Desolneux and A. Leclaire. Stochastic image models from SIFT-like descriptors. SIAM Journal on Imaging Sciences, 11(4):2305–2338, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Diepeveen:2021:ISN

[DL21] Willem Diepeveen and Jan Lellmann. An inexact semismooth Newton method on Riemannian manifolds with application to duality-based total variation denoising. SIAM Journal on Imaging Sciences, 14(4):1565– 1600, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Dong:2019:IOS

[DLL19] Heping Dong, Jun Lai, and Peijun Li. Inverse obstacle scattering for elastic waves with Phased or phaseless far-field data. SIAM Journal on Imaging Sciences, 12(2):809–838, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Ding:2017:ILR

[DLLY17] Meng Ding, Jianliang Li, Keji Liu, and Jiaqing Yang. Imaging of local rough surfaces by the linear sampling method with near-field data. SIAM Journal on Imaging Sciences, 10(3): 1579–1602, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Diepeveen:2023:ROE

[DLÖS23] Willem Diepeveen, Jan Lellmann, Ozan Öktem, and Carola-Bibiane Schönlieb. Regularizing orientation estimation in cryogenic electron microscopy three-dimensional map refinement through measurebased lifting over Riemannian manifolds. SIAM Journal on Imaging Sciences, 16(3): 1440–1490, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1520773.

Debroux:2023:MDR

[DLV23] Noémie Debroux, Carole Le Guyader, and Luminita A. Vese. A multiscale deformation representation. SIAM Journal on Imaging Sciences, 16(2):

802–841, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1510200.

Di:2016:OBA

[DLW16] Zichao (Wendy) Di, Sven Leyffer, and Stefan M. Wild. Optimization-based approach for joint X-ray fluorescence and transmission tomographic inversion. SIAM Journal on Imaging Sciences, 9(1):1–23, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Darbon:2020:DMI

[DM20] Jérôme Darbon and Tingwei Meng. On decomposition models in imaging sciences and multi-time Hamilton–Jacobi partial differential equations. SIAM Journal on Imaging Sciences, 13(2):971–1014, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Durmus:2018:EBC

[DMP18] Alain Durmus, Éric Moulines, and Marcelo Pereyra. Efficient Bayesian computation by proximal Markov Chain Monte Carlo: When Langevin meets Moreau. SIAM Journal on Imaging Sciences, 11(1):473– 506, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Duran:2016:CTV

[DMSC16] J. Duran, M. Moeller, C. Sbert, and D. Cremers. Collaborative total variation: a general framework for vectorial TV models.

SIAM Journal on Imaging Sciences, 9(1):116–151, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Druskin:2016:DNI

[DMTZ16] Vladimir Druskin, Alexander V.

Mamonov, Andrew E. Thaler, and Mikhail Zaslavsky. Direct, nonlinear inversion algorithm for hyperbolic problems via projection-based model reduction. SIAM Journal on Imaging Sciences, 9(2):684–747, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Druskin:2018:NMI

[DMZ18] Vladimir Druskin, Alexander V.

Mamonov, and Mikhail Zaslavsky. A nonlinear method for imaging with acoustic waves via reduced order model backprojection. SIAM Journal on Imaging Sciences, 11(1):164– 196, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Duran:2021:VDR

[DNB21] Joan Duran, Julia Navarro, and Antoni Buades. Variational densification and refinement of registration maps. SIAM Journal on Imaging Sciences, 14(3): 879–912, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Destrempes:2013:EMH

[DPC13] François Destrempes, Jonathan Porée, and Guy Cloutier. Estimation method of the homodyned K-distribution based on the mean intensity and two log-moments. SIAM Journal on Imaging Sciences, 6(3):1499– 1530, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

DeVore:2013:PTP

[DPH^{+}13] Ronald DeVore, Guergana Petrova, Matthew Hielsberg, Luke Owens, Billy Clack, and Alok Sood. Processing terrain point cloud data. SIAM Journal on Imaging Sciences, 6(1):1–31, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Deledalle:2018:IDG

[DPN18] Charles-Alban Deledalle, Shibin Parameswaran, and Truong Q. Nguyen. Image denoising with generalized Gaussian mixture model patch priors. SIAM Journal on Imaging Sciences, 11(4): 2568–2609, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Deledalle:2017:CCL

[DPSV17] Charles-Alban Deledalle, Nicolas Papadakis, Joseph Salmon, and Samuel Vaiter. CLEAR: Covariant LEAst-Square Refitting with applications to image restoration. SIAM Journal on Imaging Sciences, 10(1): 243–284, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Davies:2014:CSF

[DPVW14] Mike Davies, Gilles Puy, Pierre Vandergheynst, and Yves Wiaux. A compressed sensing framework for magnetic resonance fingerprinting. SIAM Journal on Imaging Sciences, 7 (4):2623–2656, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Deng:2020:DPR

[DPZ20] Kangkang Deng, Zheng Peng, and Wenxing Zhu. A discriminative projection and representation-based classification framework for face recognition. SIAM Journal on Imaging Sciences, 13(3):1446–1466, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Drogoul:2014:NAT

[Dro14] Audric Drogoul. Numerical analysis of the topological gradient method for fourth order models and applications to the detection of fine structures in imaging. SIAM Journal on Imaging Sciences, 7(4):2700– 2731, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

dAvigneau:2022:LAP

[dSO22] A. Marie d’Avigneau, Sumeetpal S. Singh, and Raimund J. Ober. Limits of accuracy for parameter estimation and localization in single-molecule microscopy via sequential Monte Carlo methods. SIAM Journal on Imaging Sciences, 15 (1):139–171, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1422823.

Dambreville:2010:GAJ

[DSYT10] Samuel Dambreville, Romeil Sandhu, Anthony Yezzi, and Allen Tannenbaum. A geometric approach to joint 2D region-based segmentation and 3D pose estimation using a 3D shape prior. SIAM Journal on Imaging Sciences, 3(1):110–132, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Driggs:2021:SPA

[DTL^{+}21] Derek Driggs, Junqi Tang, Jingwei Liang, Mike Davies, and Carola-Bibiane Schönlieb. A stochastic proximal alternating minimization for nonsmooth and nonconvex optimization. SIAM Journal on Imaging Sciences, 14(4):1932–1970, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

DelosReyes:2022:OCB

[DV22] Juan Carlos De los Reyes and David Villacís. Optimality conditions for bilevel imaging learning problems with total variation regularization. SIAM Journal on Imaging Sciences, 15(4): 1646–1689, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M143412X.

Deledalle:2014:SUG

[DVFP14] Charles-Alban Deledalle, Samuel Vaiter, Jalal Fadili, and Gabriel Peyré. Stein Unbiased GrAdient estimator of the Risk (SUGAR) for multiple parameter selection. SIAM Journal on Imaging Sciences, 7(4):2448–2487, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Dong:2013:CVM

[DZ13] Yiqiu Dong and Tieyong Zeng. A convex variational model for restoring blurred images with multiplicative noise. SIAM Journal on Imaging Sciences, 6 (3):1598–1625, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Ehrhardt:2016:MMR

[EB16] Matthias J. Ehrhardt and Marta M. Betcke. Multicontrast MRI reconstruction with structure-guided total variation. SIAM Journal on Imaging Sciences, 9(3):1084–1106, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

ElBouchairi:2023:NPC

[EEF23] Imad El Bouchairi, Abderrahim Elmoataz, and Jalal Fadili. Nonlocal perimeters and curvature flows on graphs with applications in image processing and high-dimensional data classification. SIAM Journal on Imaging Sciences, 16(1): 368–392, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M148598X.

Esser:2018:TVR

[EGvL^{+}18] Ernie Esser, Lluis Guasch, Tristan van Leeuwen, Aleksandr Y. Aravkin, and Felix J. Herrmann. Total variation regularization strategies in fullwaveform inversion. SIAM Journal on Imaging Sciences, 11(1):376–406, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Easley:2009:IDU

[EHB09] Glenn R. Easley, Dennis M.

Healy, Jr., and Carlos A. Berenstein. Image deconvolution using a general ridgelet and curvelet domain. SIAM Journal on Imaging Sciences, 2(1): 253–283, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Escolano:2017:GST

[EHL17] Francisco Escolano, Edwin R. Hancock, and Miguel A. Lozano. Graph similarity through entropic manifold alignment. SIAM Journal on Imaging Sciences, 10(2):942– 978, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Enqvist:2010:GOO

[EKOÅ10] Olof Enqvist, Fredrik Kahl, Carl Olsson, and Kalle Åström. Global optimization for onedimensional structure and motion problems. SIAM Journal on Imaging Sciences, 3(4): 1075–1095, ???? 2010. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v3/i4/p1075_s1.

Elad:2023:IDD

[EKV23] Michael Elad, Bahjat Kawar, and Gregory Vaksman. Image denoising: The deep learning revolution and beyond — a survey paper. SIAM Journal on Imaging Sciences, 16(3): 1594–1654, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/23M1545859.

Elser:2018:BPP

[ELB18] Veit Elser, Ti-Yen Lan, and Tamir Bendory. Benchmark problems for phase retrieval. SIAM Journal on Imaging Sciences, 11(4):2429–2455, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Esser:2013:MFS

[ELX13] Ernie Esser, Yifei Lou, and Jack Xin. A method for finding structured sparse solutions to nonnegative least squares problems with applications. SIAM Journal on Imaging Sciences, 6(4): 2010–2046, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Effland:2020:CTD

[ENR20] Alexander Effland, Sebastian Neumayer, and Martin Rumpf. Convergence of the time discrete metamorphosis model on Hadamard manifolds. SIAM Journal on Imaging Sciences, 13(2):557–588, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Effland:2018:IET

[ERS18] Alexander Effland, Martin Rumpf, and Florian Schäfer. Image extrapolation for the time discrete metamorphosis model: Existence and applications. SIAM Journal on Imaging Sciences, 11(1):834–862, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Elbau:2015:MEF

[ES15] P. Elbau and O. Scherzer. Modelling the effect of focusing detectors in photoacoustic sectional imaging. SIAM Journal on Imaging Sciences, 8(1):1–18, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Estellers:2016:RSR

[ESS16] V. Estellers, M. A. Scott, and S. Soatto. Robust surface reconstruction. SIAM Journal on Imaging Sciences, 9(4):2073– 2098, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Eliasof:2020:MSR

[EST20] Moshe Eliasof, Andrei Sharf, and Eran Treister. Multimodal 3D shape reconstruction under calibration uncertainty using parametric level set methods. SIAM Journal on Imaging Sciences, 13(1):265–290, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Eksioglu:2018:DAM

[ET18] Ender M. Eksioglu and A. Korhan Tanc. Denoising AMP for MRI reconstruction: BM3D– AMP–MRI. SIAM Journal on Imaging Sciences, 11(3):2090– 2109, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Elmoataz:2015:LLG

[ETT15] Abderrahim Elmoataz, Matthieu Toutain, and Daniel Tenbrinck. On the p-Laplacian and ∞- Laplacian on graphs with applications in image and data processing. SIAM Journal on Imaging Sciences, 8(4):2412– 2451, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Escande:2015:SWR

[EW15] Paul Escande and Pierre Weiss. Sparse wavelet representations of spatially varying blurring operators. SIAM Journal on Imaging Sciences, 8(4):2976– 3014, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Esser:2010:GFC

[EZC10] Ernie Esser, Xiaoqun Zhang, and Tony F. Chan. A general framework for a class of first order primal-dual algorithms for convex optimization in imaging science. SIAM Journal on Imaging Sciences, 3(4): 1015–1046, ???? 2010. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v3/i4/p1015_s1.

Ferrara:2009:SME

[FA09] Matthew Ferrara and Gregory Arnold. Shape and motion estimation from near-field echobased sensor data. SIAM Journal on Imaging Sciences, 2(2): 710–729, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Fannjiang:2009:CIS

[Fan09] Albert C. Fannjiang. Compressive imaging of subwavelength structures. SIAM Journal on Imaging Sciences, 2(4): 1277–1291, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Farhan:2019:HAM

[Far19] Erez Farhan. Highly accurate matching of weakly localized features. SIAM Journal on Imaging Sciences, 12(4):1833– 1863, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Fedorov:2015:LMA

[FAS^{+}15] Vadim Fedorov, Pablo Arias, Rida Sadek, Gabriele Facciolo, and Coloma Ballester. Linear multiscale analysis of similarities between images on Riemannian manifolds: Practical formula and affine covariant metrics. SIAM Journal on Imaging Sciences, 8(3):2021–2069, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Fageot:2015:WSS

[FBU15] Julien Fageot, Emrah Bostan, and Michael Unser. Wavelet statistics of sparse and selfsimilar images. SIAM Journal on Imaging Sciences, 8(4): 2951–2975, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Fung:2020:MOL

[FD20] Samy Wu Fung and Zichao Wendy Di. Multigrid optimization for large-scale ptychographic phase retrieval. SIAM Journal on Imaging Sciences, 13(1):214– 233, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Fawzi:2013:IRS

[FF13] Alhussein Fawzi and Pascal Frossard. Image registration with sparse approximations in parametric dictionaries. SIAM Journal on Imaging Sciences, 6 (4):2370–2403, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Figueiredo:2011:IDP

[FFA11] Isabel N. Figueiredo, Pedro N. Figueiredo, and Nuno Almeida.

Image-driven parameter estimation in absorption-diffusion models of chromoscopy. SIAM Journal on Imaging Sciences, 4 (3):884–904, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p884_s1.

Fan:2023:SHP

[FG23] Xu-Qian Fan and Wenyong Gong. Schwarzschild harmonic path planning and tetrahedral mesh smoothing. SIAM Journal on Imaging Sciences, 16 (1):568–584, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1493446.

Fournier:2017:MFC

[FGPT17] J. Fournier, J. Garnier, G. Papanicolaou, and C. Tsogka. Matched-filter and correlationbased imaging for fast moving objects using a sparse network of receivers. SIAM Journal on Imaging Sciences, 10(4):2165– 2216, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Fidler:2012:SRP

[FGS12] Thomas Fidler, Markus Grasmair, and Otmar Scherzer. Shape reconstruction with a priori knowledge based on integral invariants. SIAM Journal on Imaging Sciences, 5(2):726–745, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Flenner:2011:HPA

[FH11] Arjuna Flenner and Gary Hewer. A Helmholtz principle approach to parameter free change detection and coherent motion using exchangeable random variables. SIAM Journal on Imaging Sciences, 4(1): 243–276, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p243_s1.

Farhan:2015:GEL

[FH15] Erez Farhan and Rami Hagege. Geometric expansion for local feature analysis and matching. SIAM Journal on Imaging Sciences, 8(4):2771–2813, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Freedman:2010:KPF

[FK10] Daniel Freedman and Pavel Kisilev. KDE paring and a faster mean shift algorithm. SIAM Journal on Imaging Sciences, 3(4):878–903, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Fornasier:2012:WDM

[FKLS12] M. Fornasier, Y. Kim, A. Langer, and C.-B. Schönlieb. Wavelet decomposition method for L_{2}/ TV-image deblurring. SIAM Journal on Imaging Sciences, 5 (3):857–885, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Fannjiang:2012:CPG

[FL12] Albert Fannjiang and Wenjing Liao. Coherence pattern-guided compressive sensing with unresolved grids. SIAM Journal on Imaging Sciences, 5(1):179–202, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Fermanian:2023:PRL

[FLG23] Rita Fermanian, Mikael Le Pendu, and Christine Guillemot. PnP-ReG: Learned regularizing gradient for plug-and-play gradient descent. SIAM Journal on Imaging Sciences, 16(2):585–613, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1490843.

Fang:2014:SID

[FLZ14] Faming Fang, Fang Li, and Tieyong Zeng. Single image dehazing and denoising: a fast variational approach. SIAM Journal on Imaging Sciences, 7 (2):969–996, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Filbir:2023:IRB

[FM23] Frank Filbir and Oleh Melnyk. Image recovery for blind polychromatic ptychography. SIAM Journal on Imaging Sciences, 16(3):1308– 1337, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1527155.

Feiszli:2017:NCW

[FN17] Matt Feiszli and Akil Narayan. Numerical computation of Weil–Peterson geodesics in the universal Teichmüller space. SIAM Journal on Imaging Sciences, 10(3):1322–1345, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Fournier:2010:AGR

[Fou10] Marc Fournier. Automatic grid resolution and efficient triangulation of implicit vector field. SIAM Journal on Imaging Sciences, 3(3):564–577, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Facciolo:2017:CSR

[FPM17] Gabriele Facciolo, Nicola Pierazzo, and Jean-Michel Morel. Conservative scale recomposition for multiscale denoising (the devil is in the high frequency detail). SIAM Journal on Imaging Sciences, 10(3): 1603–1626, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Ferradans:2014:RDO

[FPPA14] Sira Ferradans, Nicolas Papadakis, Gabriel Peyré, and Jean-François Aujol. Regularized discrete optimal transport. SIAM Journal on Imaging Sciences, 7(3):1853–1882, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Falcone:2020:HOS

[FPT20] Maurizio Falcone, Giulio Paolucci, and Silvia Tozza. A highorder scheme for image segmentation via a modified levelset method. SIAM Journal on Imaging Sciences, 13(1):497– 534, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Feng:2016:PNR

[FQC16] Wensen Feng, Hong Qiao, and Yunjin Chen. Poisson noise reduction with higher-order natural image prior model. SIAM Journal on Imaging Sciences, 9 (3):1502–1524, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Feng:2017:IDM

[FQXC17] Wensen Feng, Peng Qiao, Xuanyang Xi, and Yunjin Chen. Image denoising via multiscale nonlinear diffusion models. SIAM Journal on Imaging Sciences, 10(3):1234–1257, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Fraysse:2014:MTV

[FR14] A. Fraysse and T. Rodet. A measure-theoretic variational Bayesian algorithm for large dimensional problems. SIAM Journal on Imaging Sciences, 7 (4):2591–2622, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Frederick:2018:IRQ

[FRV18] Christina Frederick, Kui Ren, and Sarah Vallélian. Image reconstruction in quantitative photoacoustic tomography with the simplified P_{2} approximation. SIAM Journal on Imaging Sciences, 11(4):2847–2876, ????

2018. CODEN SJISBI. ISSN 1936-4954.

Fujiwara:2020:NRR

[FST20] Hiroshi Fujiwara, Kamran Sadiq, and Alexandru Tamasan. Numerical reconstruction of radiative sources in an absorbing and nondiffusing scattering medium in two dimensions. SIAM Journal on Imaging Sciences, 13(1):535–555, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Fujiwara:2023:NRR

[FST23] Hiroshi Fujiwara, Kamran Sadiq, and Alexandru Tamasan. Numerical reconstruction of radiative sources from partial boundary measurements. SIAM Journal on Imaging Sciences, 16(2):948–968, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1507449.

Feigin:2010:AKA

[FSV10] Micha Feigin, Nir Sochen, and Baba C. Vemuri. Anisotropic α-kernels and associated flows. SIAM Journal on Imaging Sciences, 3(4):904–925, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Fannjiang:2009:SAI

[FSY09] Albert C. Fannjiang, Knut Solna, and Pengchong Yan. Synthetic aperture imaging of multiple point targets in Rician fading media. SIAM Journal on Imaging Sciences, 2(2):344–366, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Fannjiang:2010:CRS

[FSY10] Albert C. Fannjiang, Thomas Strohmer, and Pengchong Yan. Compressed remote sensing of sparse objects. SIAM Journal on Imaging Sciences, 3(3): 595–618, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Frijlink:2010:RTO

[FW10] Martijn Frijlink and Kees Wapenaar. Reciprocity theorems for one-way wave fields in curvilinear coordinate systems. SIAM Journal on Imaging Sciences, 3(3):390–415, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Fehrenbach:2014:PSN

[FW14] Jérôme Fehrenbach and Pierre Weiss. Processing stationary noise: Model and parameter selection in variational methods. SIAM Journal on Imaging Sciences, 7(2):613–640, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Fannjiang:2020:FPA

[FZ20] Albert Fannjiang and Zheqing Zhang. Fixed point analysis of Douglas–Rachford splitting for ptychography and phase retrieval. SIAM Journal on Imaging Sciences, 13(2):609– 650, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Gonzalez:2022:SIP

[GAT22] Mario González, Andrés Almansa, and Pauline Tan. Solving inverse problems by joint posterior maximization with autoencoding prior. SIAM Journal on Imaging Sciences, 15 (2):822–859, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M140225X.

Gao:2011:LSB

[GB11] Wenhua Gao and Andrea Bertozzi. Level set based multispectral segmentation with corners. SIAM Journal on Imaging Sciences, 4(2):597– 617, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i2/p597_s1.

Gao:2018:ICO

[GB18] Yiming Gao and Kristian Bredies. Infimal convolution of oscillation total generalized variation for the recovery of images with structured texture. SIAM Journal on Imaging Sciences, 11(3):2021–2063, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Graziani:2010:FCA

[GBFA10] Daniele Graziani, Laure Blanc-Féraud, and Gilles Aubert. A formal Γ-convergence approach for the detection of points in 2-D images. SIAM Journal on Imaging Sciences, 3(3):578–594, ???? 2010. CODEN SJISBI. ISSN 1936-4954. See erratum [GBFA12].

Graziani:2012:EFC

[GBFA12] Daniele Graziani, Laure Blanc-Féraud, and Gilles Aubert. Erratum: A Formal Γ-Convergence Approach for the Detection of Points in 2-D Images. SIAM Journal on Imaging Sciences, 5 (3):1111–1112, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954. See [GBFA10].

Ge:2021:NRI

[GCN21] Huanmin Ge, Wengu Chen, and Michael K. Ng. New restricted isometry property analysis for ℓ_{1}–ℓ_{2} minimization methods. SIAM Journal on Imaging Sciences, 14(2):530– 557, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Gooya:2015:BAS

[GDF15] Ali Gooya, Christos Davatzikos, and Alejandro F. Frangi. A Bayesian approach to sparse model selection in statistical shape models. SIAM Journal on Imaging Sciences, 8(2):858–887, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Gris:2018:SRM

[GDT18] Barbara Gris, Stanley Durrleman, and Alain Trouvé. A sub-Riemannian modular framework for diffeomorphismbased analysis of shape ensembles. SIAM Journal on Imaging Sciences, 11(1):802–833, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Giryes:2015:SBM

[GEB15] R. Giryes, M. Elad, and A. M. Bruckstein. Sparsity based methods for overparameterized variational problems. SIAM Journal on Imaging Sciences, 8 (3):2133–2159, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Getreuer:2011:CST

[Get11] Pascal Getreuer. Contour stencils: Total variation along curves for adaptive image interpolation. SIAM Journal on Imaging Sciences, 4(3): 954–979, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p954_s1.

Grementieri:2022:MCD

[GF22] Luca Grementieri and Rita Fioresi. Model-centric data manifold: The data through the eyes of the model. SIAM Journal on Imaging Sciences, 15(3): 1140–1156, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1437056.

Godeme:2023:PPR

[GFB^{+}23] Jean-Jacques Godeme, Jalal Fadili, Xavier Buet, Myriam Zerrad, Michel Lequime, and Claude Amra. Provable phase retrieval with mirror descent. SIAM Journal on Imaging Sciences, 16(3): 1106–1141, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1528896.

Guo:2022:GPM

[GGJ^{+}22] Jing Guo, Yu Guo, Qiyu Jin, Michael Kwok-Po Ng, and Shuping Wang. Gaussian patch mixture model guided low-rank covariance matrix minimization for image denoising. SIAM Journal on Imaging Sciences, 15(4):1601– 1622, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1454262.

Griesmaier:2015:MII

[GH15] Roland Griesmaier and Martin Hanke. Multifrequency impedance imaging with multiple signal classification. SIAM Journal on Imaging Sciences, 8 (2):939–967, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Garmatter:2018:MRE

[GH18] Dominik Garmatter and Bastian Harrach. Magnetic Resonance Electrical Impedance Tomography (MREIT): Convergence and reduced basis approach. SIAM Journal on Imaging Sciences, 11(1):863– 887, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Gao:2023:TRP

[GH23] Kaixin Gao and Zheng-Hai Huang. Tensor robust principal component analysis via tensor fibered rank and l_{p} minimization. SIAM Journal on Imaging Sciences, 16 (1):423–460, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1473236.

Gao:2023:FDD

[GHFT23] Guangyu Gao, Bo Han, Zhenwu Fu, and Shanshan Tong. A fast data-driven iteratively regularized method with convex penalty for solving illposed problems. SIAM Journal on Imaging Sciences, 16 (2):640–670, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1506778.

Garnier:2023:LSM

[GHM23a] Josselin Garnier, Houssem Haddar, and Hadrien Montanelli. The linear sampling method for random sources. SIAM Journal on Imaging Sciences, 16(3): 1572–1593, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1531336.

Guo:2023:IMP

[GHM23b] Hongxia Guo, Guanghui Hu, and Guanqiu Ma. Imaging a moving point source from multifrequency data measured at one and sparse observation directions (Part I): Farfield case. SIAM Journal on Imaging Sciences, 16(3): 1535–1571, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/23M1545045.

Gilboa:2014:TVS

[Gil14a] Guy Gilboa. A total variation spectral framework for scale and texture analysis. SIAM Journal on Imaging Sciences, 7(4): 1937–1961, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Gillis:2014:SNP

[Gil14b] Nicolas Gillis. Successive nonnegative projection algorithm for robust nonnegative blind source separation. SIAM Journal on Imaging Sciences, 7(2): 1420–1450, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Genzel:2014:AAI

[GK14] Martin Genzel and Gitta Kutyniok. Asymptotic analysis of inpainting via universal shearlet systems. SIAM Journal on Imaging Sciences, 7(4):2301– 2339, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Guidotti:2013:IRN

[GKL13] Patrick Guidotti, Yunho Kim, and James Lambers. Image restoration with a new class of forward–backward–forward diffusion equations of Perona– Malik type with applications to satellite image enhancement. SIAM Journal on Imaging Sciences, 6(3):1416–1444, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Grathwohl:2020:IER

[GKQR20] Christine Grathwohl, Peer Christian Kunstmann, Eric Todd Quinto, and Andreas Rieder.

Imaging with the elliptic Radon transform in three dimensions from an analytical and numerical perspective. SIAM Journal on Imaging Sciences, 13(4): 2250–2280, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Guo:2009:CAE

[GL09] Kanghui Guo and Demetrio Labate. Characterization and analysis of edges using the continuous shearlet transform. SIAM Journal on Imaging Sciences, 2(3):959–986, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Guo:2013:BSW

[GL13] Weihong Guo and Ming-Jun Lai. Box spline wavelet frames for image edge analysis. SIAM Journal on Imaging Sciences, 6 (3):1553–1578, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Galerne:2017:TIU

[GL17] Bruno Galerne and Arthur Leclaire. Texture inpainting using efficient Gaussian conditional simulation. SIAM Journal on Imaging Sciences, 10(3): 1446–1474, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Glowinski:2015:PRO

[GLQ15] Roland Glowinski, Shingyu Leung, and Jianliang Qian. A penalization-regularization-operator splitting method for eikonal based traveltime tomography. SIAM Journal on Imaging Sciences, 8(2):1263–1292, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Galerne:2018:TSM

[GLR18] B. Galerne, A. Leclaire, and J. Rabin. A texture synthesis model based on semidiscrete optimal transport in patch space. SIAM Journal on Imaging Sciences, 11(4):2456– 2493, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Garcke:2018:CHI

[GLS18] Harald Garcke, Kei Fong Lam, and Vanessa Styles. Cahn–Hilliard inpainting with the double obstacle potential. SIAM Journal on Imaging Sciences, 11(3):2064–2089, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Gilbert:2020:NIH

[GLS20] Anna C. Gilbert, Howard W.

Levinson, and John C. Schotland. Nonlinear iterative hard thresholding for inverse scattering. SIAM Journal on Imaging Sciences, 13(1):108–140, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Gruber:2010:RES

[GM10] Fred K. Gruber and Edwin A.

Marengo. Reinterpretation and enhancement of signalsubspace-based imaging methods for extended scatterers. SIAM Journal on Imaging Sciences, 3(3):434–461, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Gillis:2015:EPP

[GM15] Nicolas Gillis and Wing-Kin Ma. Enhancing pure-pixel identification performance via preconditioning. SIAM Journal on Imaging Sciences, 8(2):1161– 1186, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Geiping:2018:CON

[GM18] Jonas Geiping and Michael Moeller. Composite optimization by nonconvex majorization– minimization. SIAM Journal on Imaging Sciences, 11(4):2494– 2528, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Graf:2022:IRM

[GNH^{+}22] Manuel Gräf, Sebastian Neumayer, Ralf Hielscher, Gabriele Steidl, Moritz Liesegang, and Tilmann Beck. An image registration model in electron backscatter diffraction. SIAM Journal on Imaging Sciences, 15(1):228–260, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1426353.

Goldstein:2009:SBM

[GO09] Tom Goldstein and Stanley Osher. The split Bregman method for L1-regularized problems. SIAM Journal on Imaging Sciences, 2(2):323–343, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Gaspar:2016:SIM

[GOF16] Tiago Gaspar, Paulo Oliveira, and Paolo Favaro. Synchronization of independently moving cameras via motion recovery. SIAM Journal on Imaging Sciences, 9(3):869–900, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Goldman:2011:CPD

[Gol11]  M. Goldman.  Continuous

primal-dual methods for image processing. SIAM Journal on Imaging Sciences, 4(1): 366–385, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p366_s1.

Goldstein:2014:FAD

[GOSB14] Tom Goldstein, Brendan O’Donoghue, Simon Setzer, and Richard Baraniuk. Fast alternating direction optimization methods. SIAM Journal on Imaging Sciences, 7(3):1588–1623, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Garnier:2009:PSI

[GP09] Josselin Garnier and George Papanicolaou. Passive sensor imaging using cross correlations of noisy signals in a scattering medium. SIAM Journal on Imaging Sciences, 2(2):396–437, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Garnier:2014:RSV

[GP14] Josselin Garnier and George Papanicolaou. Role of scattering in virtual source array imaging. SIAM Journal on Imaging Sciences, 7(2):1210–1236, ????

2014. CODEN SJISBI. ISSN 1936-4954.

Garnier:2015:PSA

[GP15] Josselin Garnier and George Papanicolaou. Passive synthetic aperture imaging. SIAM Journal on Imaging Sciences, 8(4): 2683–2705, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Gasbarra:2017:ERM

[GPB17] Dario Gasbarra, Sinisa Pajevic, and Peter J. Basser. Eigenvalues of random matrices with isotropic Gaussian noise and the design of diffusion tensor imaging experiments. SIAM Journal on Imaging Sciences, 10(3): 1511–1548, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Giordano:2015:PDF

[GPPM15] S. Giordano, N. Pinamonti, M. Piana, and A. M. Massone. The process of data formation for the Spectrometer/Telescope for Imaging X-rays (STIX) in Solar Orbiter. SIAM Journal on Imaging Sciences, 8(2):1315– 1331, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Garnier:2013:SNR

[GPST13] Josselin Garnier, George Papanicolaou, Adrien Semin, and Chrysoula Tsogka. Signal-to-noise ratio estimation in passive correlation-based imaging. SIAM Journal on Imaging Sciences, 6(2):1092–1110, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Garnier:2015:SNR

[GPST15] Josselin Garnier, George Papanicolaou, Adrien Semin, and Chrysoula Tsogka. Signal to noise ratio analysis in virtual source array imaging. SIAM Journal on Imaging Sciences, 8 (1):248–279, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Guo:2014:NDP

[GQY14] Weihong Guo, Jing Qin, and Wotao Yin. A new detail-preserving regularization scheme. SIAM Journal on Imaging Sciences, 7(2):1309– 1334, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ganster:2023:AIC

[GR23] Kevin Ganster and Andreas Rieder. Approximate inversion of a class of generalized Radon transforms. SIAM Journal on Imaging Sciences, 16 (2):842–866, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1512417.

Griesmaier:2010:RTT

[Gri10] Roland Griesmaier. Reconstruction of thin tubular inclusions in three-dimensional domains using electrical impedance tomography. SIAM Journal on Imaging Sciences, 3(3):340–362, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Garnier:2010:CCD

[GS10] Josselin Garnier and Knut Sølna. Cross correlation and deconvolution of noise signals in randomly layered media. SIAM Journal on Imaging Sciences, 3 (4):809–834, ???? 2010. CO-DEN SJISBI. ISSN 1936-4954.

Gholami:2013:FBS

[GS13] Ali Gholami and Mauricio D. Sacchi. Fast 3D blind seismic deconvolution via constrained total variation and GCV. SIAM Journal on Imaging Sciences, 6 (4):2350–2369, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Gutierrez:2016:ALD

[GS16] Cristian E. Gutiérrez and Ahmad Sabra. Aspherical lens design and imaging. SIAM Journal on Imaging Sciences, 9(1): 386–411, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Griesmaier:2017:FMM

[GS17] Roland Griesmaier and Christian Schmiedecke. A factorization method for multifrequency inverse source problems with sparse far field measurements. SIAM Journal on Imaging Sciences, 10(4):2119–2139, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Goldluecke:2012:NVT

[GSC12] Bastian Goldluecke, Evgeny Strekalovskiy, and Daniel Cremers. The natural vectorial total variation which arises from geometric measure theory. SIAM Journal on Imaging Sciences, 5(2):537–563, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Goldluecke:2013:TCR

[GSC13] Bastian Goldluecke, Evgeny Strekalovskiy, and Daniel Cremers. Tight convex relaxations for vector-valued labeling. SIAM Journal on Imaging Sciences, 6(3):1626–1664, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Garcia-Salguero:2021:SCO

[GSGJ21] Mercedes Garcia-Salguero and Javier Gonzalez-Jimenez. A sufficient condition of optimality for the relative pose problem between cameras. SIAM Journal on Imaging Sciences, 14(4): 1617–1634, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Guerit:2022:CIT

[GSL^{+}22] Stéphanie Guérit, Siddharth Sivankutty, John Lee, Hervé Rigneault, and Laurent Jacques. Compressive imaging through optical fiber with partial speckle scanning. SIAM Journal on Imaging Sciences, 15(2):387– 423, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1407586.

Geng:2018:TNN

[GSXH18] Tianyu Geng, Guiling Sun, Yi Xu, and Jingfei He. Truncated nuclear norm minimization based group sparse representation for image restoration. SIAM Journal on Imaging Sciences, 11(3):1878–1897, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Guo:2017:PTT

[GSZ17] Weihong Guo, Guohui Song, and Yue Zhang. PCM–TV– TFV: a novel two-stage framework for image reconstruction from Fourier data. SIAM Journal on Imaging Sciences, 10(4): 2250–2274, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Gilman:2015:MMS

[GT15] Mikhail Gilman and Semyon Tsynkov. A mathematical model for SAR imaging beyond the first Born approximation. SIAM Journal on Imaging Sciences, 8(1):186–225, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Gilles:2014:ETW

[GTO14] Jérôme Gilles, Giang Tran, and Stanley Osher. 2D empirical transforms. Wavelets, ridgelets, and curvelets revisited. SIAM Journal on Imaging Sciences, 7 (1):157–186, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Galarce:2023:DPR

[GTP^{+}23] Felipe Galarce, Karsten Tabelow, Jörg Polzehl, Christos Panagiotis Papanikas, Vasileios Vavourakis, Ledia Lilaj, Ingolf Sack, and Alfonso Caiazzo. Displacement and pressure reconstruction from magnetic resonance elastography images: Application to an in silico brain model. SIAM Journal on Imaging Sciences, 16(2): 996–1027, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M149363X.

Guler:2014:SPH

[GTU14] R. A. Guler, S. Tari, and G. Unal. Screened Poisson hyperfields for shape coding. SIAM Journal on Imaging Sciences, 7(4):2558–2590, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Galdran:2015:EVI

[GVCPB15] Adrian Galdran, Javier Vazquez-Corral, David Pardo, and Marcelo Bertalmío. Enhanced variational image dehazing. SIAM Journal on Imaging Sciences, 8(3):1519–1546, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Goldfarb:2009:CSM

[GWY09] Donald Goldfarb, Zaiwen Wen, and Wotao Yin. A curvilinear search method for p-harmonic flows on spheres. SIAM Journal on Imaging Sciences, 2(1): 84–109, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Guo:2012:EGR

[GY12] Weihong Guo and Wotao Yin.

Edge guided reconstruction for compressive imaging. SIAM Journal on Imaging Sciences, 5 (3):809–834, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Gao:2015:KFP

[GZC^{+}15] Yi Gao, Liangjia Zhu, Joshua Cates, Rob S. MacLeod, Sylvain Bouix, and Allen Tannenbaum. A Kalman filtering perspective for multiatlas segmentation. SIAM Journal on Imaging Sciences, 8(2):1007–1029, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Haltmeier:2011:IFC

[Hal11]  Markus Haltmeier.  Inver-

sion formulas for a cylindrical Radon transform. SIAM Journal on Imaging Sciences, 4 (3):789–806, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p789_s1.

Hanke:2012:OSI

[Han12] Martin Hanke. One shot inverse scattering via rational approximation. SIAM Journal on Imaging Sciences, 5(1):465–482, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Houdard:2018:HDM

[HBD18] Antoine Houdard, Charles Bouveyron, and Julie Delon. Highdimensional mixture models for unsupervised image denoising (HDMI). SIAM Journal on Imaging Sciences, 11(4):2815– 2846, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Haro:2012:PPI

[HBM12] Gloria Haro, Antoni Buades, and Jean-Michel Morel. Photographing paintings by image fusion. SIAM Journal on Imaging Sciences, 5(3):1055–1087, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Hafftka:2020:DTC

[HCCS20] Ariel Hafftka, Wojciech Czaja, Hasan Celik, and Richard G. Spencer. N-dimensional tensor completion for nuclear magnetic resonance relaxometry. SIAM Journal on Imaging Sciences, 13(1):176–213, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Huo:2022:SIR

[HCGN22] Limei Huo, Wengu Chen, Huanmin Ge, and Michael K. Ng. Stable image reconstruction using transformed total variation minimization. SIAM Journal on Imaging Sciences, 15(3): 1104–1139, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1438566.

Huo:2023:LLM

[HCGN23] Limei Huo, Wengu Chen, Huanmin Ge, and Michael K. Ng. L_{1}– βL_{q} minimization for signal and image recovery. SIAM Journal on Imaging Sciences, 16(4): 1886–1928, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1525363.

Heas:2016:EAV

[HDH16] P. Héas, A. Drémeau, and C. Herzet. An efficient algorithm for video superresolution based on a sequential model. SIAM Journal on Imaging Sciences, 9(2):537–572, ???? 2016. CODEN SJISBI. ISSN 1936-

Hernandez:2019:CSD

[Her19] Monica Hernandez. A comparative study of different variants of Newton–Krylov PDEconstrained Stokes–LDDMM parameterized in the space of band-limited vector fields. SIAM Journal on Imaging Sciences, 12(2):1038–1070, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Hass:2012:RBC

[HF12] Ryan Hass and Adel Faridani. Regions of backprojection and comet tail artifacts for π-line reconstruction formulas in tomography. SIAM Journal on Imaging Sciences, 5(4):1159–1184, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Hafiene:2019:CLN

[HFE19] Yosra Hafiene, Jalal M. Fadili, and Abderrahim Elmoataz. Continuum limits of nonlocal p-Laplacian variational problems on graphs. SIAM Journal on Imaging Sciences, 12(4):1772– 1807, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Huang:2018:BDS

[HH18] Wen Huang and Paul Hand. Blind deconvolution by a steepest descent algorithm on a quotient manifold. SIAM Journal on Imaging Sciences, 11(4): 2757–2785, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Hinterer:2023:PNK

[HHJ^{+}23] Fabian Hinterer, Simon Hubmer, Prashin Jethwa, Kirk M.

Soodhalter, Glenn van de Ven, and Ronny Ramlau. A projected Nesterov–Kaczmarz approach to stellar populationkinematic distribution reconstruction in extragalactic archaeology. SIAM Journal on Imaging Sciences, 16(1): 192–222, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1503002.

Herrmann:2018:AIP

[HHK^{+}18] Marc Herrmann, Roland Herzog, Heiko Kröner, Stephan Schmidt, and José Vidal. Analysis and an interior-point approach for TV image reconstruction problems on smooth surfaces. SIAM Journal on Imaging Sciences, 11(2):889– 922, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Hannukainen:2016:EIT

[HHMT16] A. Hannukainen, N. Hyvönen, H. Majander, and T. Tarvainen. Efficient inclusion of total variation type priors in quantitative photoacoustic tomography. SIAM Journal on Imaging Sciences, 9(3):1132–1153, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Hanke:2008:CSS

[HHR08] Martin Hanke, Nuutti Hyvönen, and Stefanie Reusswig. Convex source support and its application to electric impedance tomography. SIAM Journal on Imaging Sciences, 1(4):364–378, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Hou:2013:RSU

[HJS13] Likun Hou, Hui Ji, and Zuowei Shen. Recovering over-/ underexposed regions in photographs. SIAM Journal on Imaging Sciences, 6(4):2213– 2235, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Holler:2014:ICT

[HK14] Martin Holler and Karl Kunisch. On infimal convolution of TV-type functionals and applications to video and image reconstruction. SIAM Journal on Imaging Sciences, 7(4):2258– 2300, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

He:2019:LIS

[HK19] Yuchen He and Sung H. Kang. Lattice identification and separation: Theory and algorithm. SIAM Journal on Imaging Sciences, 12(4):2063–2096, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Hockmann:2023:SCW

[HK23] Mathias Hockmann and Stefan Kunis. Short communication: Weak sparse superresolution is well-conditioned. SIAM Journal on Imaging Sciences, 16(1): SC1–SC13, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1521353.

Hao:2013:FRU

[HKBH13] Ning Hao, Misha E. Kilmer, Karen Braman, and Randy C. Hoover. Facial recognition using tensor-tensor decompositions. SIAM Journal on Imaging Sciences, 6(1):437–463, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

He:2020:CRS

[HKL20] Yuchen He, Sung Ha Kang, and Hao Liu. Curvature regularized surface reconstruction from point clouds. SIAM Journal on Imaging Sciences, 13(4):1834– 1859, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Huska:2021:VAA

[HKLM21] Martin Huska, Sung H. Kang, Alessandro Lanza, and Serena Morigi. A variational approach to additive image decomposition into structure, harmonic, and oscillatory components. SIAM Journal on Imaging Sciences, 14(4):1749–1789, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Hintermuller:2013:SCM

[HL13] Michael Hintermüller and Andreas Langer. Subspace correction methods for a class of nonsmooth and nonadditive convex variational problems with mixed L^{1}/L^{2} data-fidelity in image processing. SIAM Journal on Imaging Sciences, 6(4): 2134–2173, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Heas:2014:SSP

[HLKH14] P. Héas, F. Lavancier, and S. Kadri-Harouna. Self-similar prior and wavelet bases for hidden incompressible turbulent motion. SIAM Journal on Imaging Sciences, 7(2):1171– 1209, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Huang:2023:CAV

[HLL^{+}23] Tsung-Ming Huang, Wei-Hung Liao, Wen-Wei Lin, Mei-Heng Yueh, and Shing-Tung Yau. Convergence analysis of volumetric stretch energy minimization and its associated optimal mass transport. SIAM Journal on Imaging Sciences, 16(3): 1825–1855, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1528756.

Hu:2014:IES

[HLLS14] Guanghui Hu, Jingzhi Li, Hongyu Liu, and Hongpeng Sun. Inverse elastic scattering for multiscale rigid bodies with a single far-field pattern. SIAM Journal on Imaging Sciences, 7 (3):1799–1825, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Hesse:2015:PHB

[HLST15] Robert Hesse, D. Russell Luke, Shoham Sabach, and Matthew K. Tam. Proximal heterogeneous block implicitexplicit method and application to blind ptychographic diffraction imaging. SIAM Journal on Imaging Sciences, 8(1):426–457, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Huroyan:2020:SJP

[HLW20] Vahan Huroyan, Gilad Lerman, and Hau-Tieng Wu. Solving jigsaw puzzles by the graph connection Laplacian. SIAM Journal on Imaging Sciences, 13(4): 1717–1753, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Hocking:2017:GGA

[HMS17] L. Robert Hocking, Russell MacKenzie, and Carola-Bibiane Schönlieb. Guidefill: GPU accelerated, artist guided geometric inpainting for 3D conversion of film. SIAM Journal on Imaging Sciences, 10(4):2049–2090, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

He:2022:GPD

[HMXY22] Bingsheng He, Feng Ma, Shengjie Xu, and Xiaoming Yuan. A generalized primaldual algorithm with improved convergence condition for saddle point problems. SIAM Journal on Imaging Sciences, 15(3): 1157–1183, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1453463.

He:2016:CSS

[HMY16] Bingsheng He, Feng Ma, and Xiaoming Yuan. Convergence study on the symmetric version of ADMM with larger step sizes. SIAM Journal on Imaging Sciences, 9(3):1467–1501, ????

2016. CODEN SJISBI. ISSN 1936-4954.

Han:2019:DCS

[HMZZ19] Bin Han, Qun Mo, Zhenpeng Zhao, and Xiaosheng Zhuang. Directional compactly supported tensor product complex tight framelets with applications to image denoising and inpainting. SIAM Journal on Imaging Sciences, 12(4): 1739–1771, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Haltmeier:2017:AIM

[HN17] Markus Haltmeier and Linh V. Nguyen. Analysis of iterative methods in photoacoustic tomography with variable sound speed. SIAM Journal on Imaging Sciences, 10(2):751– 781, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Horev:2015:DLE

[HNAC^{+}15] Inbal Horev, Boaz Nadler, Ery Arias-Castro, Meirav Galun, and Ronen Basri. Detection of long edges on a computational budget: a sublinear approach. SIAM Journal on Imaging Sciences, 8(1):458–483, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Huang:2009:NTV

[HNW09] Yu-Mei Huang, Michael K. Ng, and You-Wei Wen. A new total variation method for multiplicative noise removal. SIAM Journal on Imaging Sciences, 2 (1):20–40, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Haehnle:2011:MSE

[HP11] Jonas Haehnle and Andreas Prohl. Mumford–Shah–Euler flow with sphere constraint and applications to color image inpainting. SIAM Journal on Imaging Sciences, 4(4): 1200–1233, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i4/p1200_s1.

Haltmeier:2015:RFC

[HP15] Markus Haltmeier and Sergiy Pereverzyev, Jr. Recovering a function from circular means or wave data on the boundary of parabolic domains. SIAM Journal on Imaging Sciences, 8(1): 592–610, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Hosseini:2017:FDF

[HP17] Mahdi S. Hosseini and Konstantinos N. Plataniotis. Finite differences in forward and inverse imaging problems: Max-Pol design. SIAM Journal on Imaging Sciences, 10(4):1963– 1996, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Hovhannisyan:2019:FMA

[HPPZ19] Vahan Hovhannisyan, Yannis Panagakis, Panos Parpas, and Stefanos Zafeiriou. Fast multilevel algorithms for compressive principal component pursuit. SIAM Journal on Imaging Sciences, 12(1):624–649, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Hager:2011:GBM

[HPZ11] William W. Hager, Dzung T. Phan, and Hongchao Zhang. Gradient-based methods for sparse recovery. SIAM Journal on Imaging Sciences, 4(1): 146–165, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p146_s1.

Hovhannisyan:2016:MMA

[HPZ16] Vahan Hovhannisyan, Panos Parpas, and Stefanos Zafeiriou. MAGMA: Multilevel accelerated gradient mirror descent algorithm for large-scale convex composite minimization. SIAM Journal on Imaging Sciences, 9 (4):1829–1857, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Holden:2022:BID

[HPZ22] Matthew Holden, Marcelo Pereyra, and Konstantinos C. Zygalakis. Bayesian imaging with data-driven priors encoded by neural networks. SIAM Journal on Imaging Sciences, 15(2):892–924, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1406313.

Hahn:2016:DSD

[HQ16] B. N. Hahn and E. T. Quinto. Detectable singularities from dynamic Radon data. SIAM Journal on Imaging Sciences, 9 (3):1195–1225, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Huang:2019:ARK

[HQ19] Guanghui Huang and Jianliang Qian. Analysis of regularized Kantorovich–Rubinstein metric and its application to inverse gravity problems. SIAM Journal on Imaging Sciences, 12(3): 1528–1560, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Hohage:2015:CEP

[HR15] T. Hohage and C. Rügge. A coherence enhancing penalty for diffusion MRI: Regularizing property and discrete approximation. SIAM Journal on Imaging Sciences, 8(3):1874– 1893, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Hamilton:2016:HSD

[HRSZ16] S. J. Hamilton, J. M. Reyes, S. Siltanen, and X. Zhang. A hybrid segmentation and D-bar method for electrical impedance tomography. SIAM Journal on Imaging Sciences, 9(2):770–793, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Huhnerbein:2018:ILB

[HSÅS18] Ruben Hühnerbein, Fabrizio Savarino, Freddie Åström, and Christoph Schnörr. Image labeling based on graphical models using Wasserstein messages and geometric assignment. SIAM Journal on Imaging Sciences, 11(2):1317–1362, ???? 2018. CODEN SJISBI. ISSN 1936-

Haije:2019:SCA

[HSF^{+}19] Tom Dela Haije, Peter Savadjiev, Andrea Fuster, Robert T. Schultz, Ragini Verma, Luc Florack, and Carl-Fredrik Westin. Structural connectivity analysis using Finsler geometry. SIAM Journal on Imaging Sciences, 12(1):551–575, ???? 2019. CO-DEN SJISBI. ISSN 1936-4954.

Hoze:2013:RSS

[HSH13] N. Hoze, Z. Schuss, and D. Holcman. Reconstruction of surface and stochastic dynamics from a planar projection of trajectories. SIAM Journal on Imaging Sciences, 6(4):2430–2449, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Huang:2022:SR

[HSMS22] Chujun Huang, Mingjie Shao, Wing-Kin Ma, and Anthony Man-Cho So. SISAL revisited. SIAM Journal on Imaging Sciences, 15(2):591– 624, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1430613.

Hubmer:2018:LPE

[HSNS18] Simon Hubmer, Ekaterina Sherina, Andreas Neubauer, and Otmar Scherzer. Lamé parameter estimation from static displacement field measurements in the framework of nonlinear inverse problems. SIAM Journal on Imaging Sciences, 11(2): 1268–1293, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Hubmer:2023:SBD

[HSR^{+}23] Simon Hubmer, Ekaterina Sherina, Ronny Ramlau, Michael Pircher, and Rainer Leitgeb. Subaperture-based digital aberration correction for optical coherence tomography: a novel mathematical approach. SIAM Journal on Imaging Sciences, 16(4):1857– 1885, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1543240.

Heimowitz:2021:CNI

[HSS21] Ayelet Heimowitz, Nir Sharon, and Amit Singer. Centering noisy images with application to Cryo-EM. SIAM Journal on Imaging Sciences, 14(2):689– 716, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Hansen:2009:RSG

[HSSP09] Per Christian Hansen, Henning Osholm Sørensen, Zsuzsanna Sükösd, and Henning Friis Poulsen. Reconstruction of single-grain orientation distribution functions for crystalline materials. SIAM Journal on Imaging Sciences, 2(2):593–613, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Hsieh:2020:AVM

[HSY20] Po-Wen Hsieh, Pei-Chiang Shao, and Suh-Yuh Yang. Adaptive variational model for contrast enhancement of lowlight images. SIAM Journal on Imaging Sciences, 13(1):1–28, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Hubenthal:2013:NRS

[Hub13] Mark Hubenthal. Numerical recovery of source singularities via the radiative transfer equation with partial data. SIAM Journal on Imaging Sciences, 6(3): 1175–1198, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Hintermuller:2015:LAN

[HVW15] Michael Hintermüller, Tuomo Valkonen, and Tao Wu. Limiting aspects of nonconvex T V^{ϕ} models. SIAM Journal on Imaging Sciences, 8(4):2581– 2621, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Hintermuller:2013:NTM

[HW13] Michael Hintermüller and Tao Wu. Nonconvex TV^{q}-models in image restoration: Analysis and a trust-region regularizationbased superlinearly convergent solver. SIAM Journal on Imaging Sciences, 6(3):1385–1415, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Han:2020:DIR

[HW20] Huan Han and Zhengping Wang. A diffeomorphic image registration model with fractional-order regularization and Cauchy–Riemann constraint. SIAM Journal on Imaging Sciences, 13(3):1240–1271, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Huang:2022:LCR

[HW22] Meng Huang and Yang Wang.

Linear convergence of randomized Kaczmarz method for solving complex-valued phaseless equations. SIAM Journal on Imaging Sciences, 15(2): 989–1016, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1450537.

He:2021:PRM

[HWC21] Fang He, Xiao Wang, and Xiaojun Chen. A penalty relaxation method for image processing using Euler’s elastica model. SIAM Journal on Imaging Sciences, 14(1):389– 417, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Han:2022:MAT

[HWZ22] Huan Han, Zhengping Wang, and Yimin Zhang. Multiscale approach for threedimensional conformal image registration. SIAM Journal on Imaging Sciences, 15(3): 1431–1468, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1455929.

He:2012:CAP

[HY12] Bingsheng He and Xiaoming Yuan. Convergence analysis of primal-dual algorithms for a saddle-point problem: From contraction perspective. SIAM Journal on Imaging Sciences, 5 (1):119–149, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

He:2014:CPD

[HYY14] Bingsheng He, Yanfei You, and Xiaoming Yuan. On the convergence of primal-dual hybrid gradient algorithm. SIAM Journal on Imaging Sciences, 7(4): 2526–2537, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Han:2014:TPC

[HZ14] Bin Han and Zhenpeng Zhao. Tensor product complex tight framelets with increasing directionality. SIAM Journal on Imaging Sciences, 7(2):997– 1034, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Huang:2023:OMR

[HZDZ23] Shuai Huang, Mona Zehni, Ivan Dokmanić, and Zhizhen Zhao. Orthogonal matrix retrieval with spatial consensus for 3D unknown view tomography. SIAM Journal on Imaging Sciences, 16(3): 1398–1439, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1498218.

Ishizaka:2014:WCM

[Ish14]  Kanya Ishizaka.  Weak*-

convergence to minimum energy measure and disperseddot halftoning. SIAM Journal on Imaging Sciences, 7(2): 1035–1079, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Iwen:2013:SBA

[ISW13] Mark A. Iwen, Fadil Santosa, and Rachel Ward. A symbolbased algorithm for decoding bar codes. SIAM Journal on Imaging Sciences, 6(1):56–77, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Iwen:2016:FPR

[IVW16] Mark A. Iwen, Aditya Viswanathan, and Yang Wang. Fast phase retrieval from local correlation measurements. SIAM Journal on Imaging Sciences, 9(4):1655– 1688, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Jiang:2017:JMD

[JBS17] Ming Jiang, Jérôme Bobin, and Jean-Luc Starck. Joint multichannel deconvolution and blind source separation. SIAM Journal on Imaging Sciences, 10(4):1997–2021, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Joy:2019:ERD

[JDA^{+}19] Thomas Joy, Alban Desmaison, Thalaiyasingam Ajanthan, Rudy Bunel, Mathieu Salzmann, Pushmeet Kohli, Philip H. S. Torr, and M. Pawan Kumar. Efficient relaxations for dense CRFs with sparse higherorder potentials. SIAM Journal on Imaging Sciences, 12(1): 287–318, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Jin:2017:NMO

[JGKL17] Qiyu Jin, Ion Grama, Charles Kervrann, and Quansheng Liu. Nonlocal means and optimal weights for noise removal. SIAM Journal on Imaging Sciences, 10(4):1878–1920, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Jenatton:2012:MMF

[JGM^{+}12] Rodolphe Jenatton, Alexandre Gramfort, Vincent Michel, Guillaume Obozinski, Evelyn Eger, Francis Bach, and Bertrand Thirion. Multiscale mining of fMRI data with hierarchical structured sparsity. SIAM Journal on Imaging Sciences, 5(3):835–856, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Ji:2011:RVR

[JHSX11] Hui Ji, Sibin Huang, Zuowei Shen, and Yuhong Xu. Robust video restoration by joint sparse and low rank matrix approximation. SIAM Journal on Imaging Sciences, 4(4): 1122–1142, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i4/p1122_s1.

Jung:2015:SCT

[JK15] Miyoun Jung and Myungjoo Kang. Simultaneous cartoon and texture image restoration with higher-order regularization. SIAM Journal on Imaging Sciences, 8(1):721–756, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Jin:2023:CSG

[JK23] Bangti Jin and Zeljko Kereta. On the convergence of stochastic gradient descent for linear inverse problems in Banach spaces. SIAM Journal on Imaging Sciences, 16(2): 671–705, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1518542.

Jauhiainen:2020:RGN

[JKSV20] Jyrki Jauhiainen, Petri Kuusela, Aku Seppänen, and Tuomo Valkonen. Relaxed Gauss–Newton methods with applications to electrical impedance tomography. SIAM Journal on Imaging Sciences, 13(3):1415– 1445, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Jin:2014:VAI

[JLN14] Zhengmeng Jin, Fang Li, and Michael K. Ng. A variational approach for image decolorization by variance maximization. SIAM Journal on Imaging Sciences, 7(2):944–968, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ji:2019:IAS

[JLZ19a] Xia Ji, Xiaodong Liu, and Bo Zhang. Inverse acoustic scattering with phaseless far field data: Uniqueness, phase retrieval, and direct sampling methods. SIAM Journal on Imaging Sciences, 12(2):1163– 1189, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Ji:2019:TRR

[JLZ19b] Xia Ji, Xiaodong Liu, and Bo Zhang. Target reconstruction with a reference point scatterer using phaseless far field patterns. SIAM Journal on Imaging Sciences, 12(1):372– 391, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Jung:2016:EIC

[JM16] Chang-Yeol Jung and Sunghwan Moon. Exact inversion of the cone transform arising in an application of a Compton camera consisting of line detectors. SIAM Journal on Imaging Sciences, 9(2):520–536, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Jia:2019:CIR

[JNW19] Zhigang Jia, Michael K. Ng, and Wei Wang. Color image restoration by saturation-value total variation. SIAM Journal on Imaging Sciences, 12(2): 972–1000, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Jung:2012:NAC

[JPC12] Miyoun Jung, Gabriel Peyré, and Laurent D. Cohen. Nonlocal active contours. SIAM Journal on Imaging Sciences, 5(3): 1022–1054, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Jin:2021:VMS

[JZMN21] Zhengmeng Jin, Junkang Zhang, Lihua Min, and Michael K. Ng. A variational model for spatially weighting in image fusion. SIAM Journal on Imaging Sciences, 14(2):441–469, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Krishnamurthy:2013:LSE

[KBW13] Kalyani Krishnamurthy, Waheed U. Bajwa, and Rebecca Willett. Level set estimation from projection measurements: Performance guarantees and fast computation. SIAM Journal on Imaging Sciences, 6 (4):2047–2074, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Kazemi:2012:SGI

[KD12] Parimah Kazemi and Ionut Danaila. Sobolev gradients and image interpolation. SIAM Journal on Imaging Sciences, 5 (2):601–624, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Kovalsky:2015:GAS

[KGB15] S. Z. Kovalsky, D. Glasner, and R. Basri. A global approach for solving edge-matching puzzles. SIAM Journal on Imaging Sciences, 8(2):916–938, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Kalmoun:2011:LSM

[KGC11] El Mostafa Kalmoun, Luis Garrido, and Vicent Caselles. Line search multilevel optimization as computational methods for dense optical flow. SIAM Journal on Imaging Sciences, 4(2): 695–722, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i2/p695_s1.

Kervazo:2021:PRB

[KGD21] Christophe Kervazo, Nicolas Gillis, and Nicolas Dobigeon. Provably robust blind source separation of linear-quadratic near-separable mixtures. SIAM Journal on Imaging Sciences, 14(4):1848–1889, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Kim:2014:IRU

[KGV14] Yunho Kim, John B. Garnett, and Luminita A. Vese. Image restoration using onedimensional Sobolev norm profiles of noise and texture. SIAM Journal on Imaging Sciences, 7 (1):366–390, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Kaganovsky:2015:AMA

[KHD^{+}15] Yan Kaganovsky, Shaobo Han, Soysal Degirmenci, David G. Politte, David J. Brady, Joseph A. O’Sullivan, and Lawrence Carin. Alternating minimization algorithm with automatic relevance determination for transmission tomography under Poisson noise. SIAM Journal on Imaging Sciences, 8(3): 2087–2132, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Kapralov:2008:SAG

[KK08] M. Kapralov and A. Katsevich. A study of 1PI algorithms for a general class of curves. SIAM Journal on Imaging Sciences, 1 (4):418–459, ???? 2008. CO-DEN SJISBI. ISSN 1936-4954.

Ko:2017:RTI

[KK17] Min-Su Ko and Yong-Jung Kim. Resistivity tensor imaging via network discretization of Faraday’s Law. SIAM Journal on Imaging Sciences, 10(1):1– 25, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Klibanov:2018:NMS

[KKN^{+}18] Michael V. Klibanov, Nikolay A. Koshev, Dinh-Liem Nguyen, Loc H. Nguyen, Aaron Brettin, and Vasily N. Astratov. A numerical method to solve a phaseless coefficient inverse problem from a single measurement of experimental data. SIAM Journal on Imaging Sciences, 11(4):2339–2367, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Klibanov:2019:CMI

[KKN19] Michael V. Klibanov, Aleksandr E. Kolesov, and Dinh-Liem Nguyen. Convexification method for an inverse scattering problem and its performance for experimental backscatter data for buried targets. SIAM Journal on Imaging Sciences, 12(1): 576–603, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Khoa:2020:CTD

[KKN20] Vo Anh Khoa, Michael Victor Klibanov, and Loc Hoang Nguyen. Convexification for a three-dimensional inverse scattering problem with the moving point source. SIAM Journal on Imaging Sciences, 13(2):

871–904, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Katsevich:2015:CME

[KKS15] E. Katsevich, A. Katsevich, and A. Singer. Covariance matrix estimation for the Cryo-EM heterogeneity problem. SIAM Journal on Imaging Sciences, 8 (1):126–185, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Korolev:2018:IRI

[KL18a] Yury Korolev and Jan Lellmann. Image reconstruction with imperfect forward models and applications in deblurring. SIAM Journal on Imaging Sciences, 11(1):197–218, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Kutyniok:2018:OCI

[KL18b] Gitta Kutyniok and Wang-Q Lim. Optimal compressive imaging of Fourier data. SIAM Journal on Imaging Sciences, 11(1):507–546, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Kim:2019:TDV

[KL19] Junwoo Kim and Chang-Ock Lee. Three-dimensional volume reconstruction using twodimensional parallel slices. SIAM Journal on Imaging Sciences, 12(1):1–27, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Klann:2011:MSL

[Kla11] Esther Klann. A Mumford– Shah-like method for limited data tomography with an application to electron tomography. SIAM Journal on Imaging Sciences, 4(4):1029– 1048, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i4/p1029_s1.

Klibanov:2023:CNMb

[KLN^{+}23] Michael V. Klibanov, Jingzhi Li, Loc H. Nguyen, Vladimir Romanov, and Zhipeng Yang. Convexification numerical method for a coefficient inverse problem for the Riemannian radiative transfer equation. SIAM Journal on Imaging Sciences, 16(3): 1762–1790, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/23M1565449.

Klibanov:2023:CNMa

[KLNY23] Michael V. Klibanov, Jingzhi Li, Loc H. Nguyen, and Zhipeng Yang. Convexification numerical method for a coefficient inverse problem for the radiative transport equation. SIAM Journal on Imaging Sciences, 16(1):35–63, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1509837.

Kee:2017:SCP

[KLS^{+}17] Youngwook Kee, Yegang Lee, Mohamed Souiai, Daniel Cremers, and Junmo Kim. Sequential convex programming for computing informationtheoretic minimal partitions: Nonconvex nonsmooth optimization. SIAM Journal on Imaging Sciences, 10(4):1845– 1877, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Kuo:2021:CCE

[KLYY21] Yueh-Cheng Kuo, Wen-Wei Lin, Mei-Heng Yueh, and Shing-Tung Yau. Convergent conformal energy minimization for the computation of disk parameterizations. SIAM Journal on Imaging Sciences, 14(4):1790– 1815, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Kushinsky:2019:SAL

[KMDL19] Yam Kushinsky, Haggai Maron, Nadav Dym, and Yaron Lipman. Sinkhorn algorithm for lifted assignment problems. SIAM Journal on Imaging Sciences, 12(2):716–735, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Kushnarev:2014:AWP

[KN14] Sergey Kushnarev and Akil Narayan. Approximating the Weil–Petersson metric geodesics on the universal Teichmüller space by singular solutions. SIAM Journal on Imaging Sciences, 7(2):900–923, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Kaji:2016:CPA

[KO16] Shizuo Kaji and Hiroyuki Ochiai. A concise parametrization of affine transformation.

SIAM Journal on Imaging Sciences, 9(3):1355–1373, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Kowar:2014:TRP

[Kow14] Richard Kowar. On time reversal in photoacoustic tomography for tissue similar to water. SIAM Journal on Imaging Sciences, 7(1):509–527, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Kunisch:2013:BOA

[KP13] Karl Kunisch and Thomas Pock. A bilevel optimization approach for parameter learning in variational models. SIAM Journal on Imaging Sciences, 6(2): 938–983, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Kolmogorov:2016:TVT

[KPR16] Vladimir Kolmogorov, Thomas Pock, and Michal Rolinek. Total variation on a tree. SIAM Journal on Imaging Sciences, 9(2): 605–636, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Klann:2013:RPM

[KR13] Esther Klann and Ronny Ramlau. Regularization properties of Mumford–Shah-Type functionals with perimeter and norm constraints for linear illposed problems. SIAM Journal on Imaging Sciences, 6(1): 413–436, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Karlsson:2017:GSI

[KR17] Johan Karlsson and Axel Ringh. Generalized Sinkhorn iterations for regularizing inverse problems using optimal mass transport. SIAM Journal on Imaging Sciences, 10(4): 1935–1962, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Krishnamurthy:2010:MPL

[KRW10] Kalyani Krishnamurthy, Maxim Raginsky, and Rebecca Willett. Multiscale photon-limited spectral image reconstruction. SIAM Journal on Imaging Sciences, 3(3):619–645, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Kuhnel:2017:MLS

[KSPR17] Line Kühnel, Stefan Sommer, Akshay Pai, and Lars Lau Raket. Most likely separation of intensity and warping effects in image registration. SIAM Journal on Imaging Sciences, 10(2): 578–601, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Kirisits:2019:IMS

[KSS19] Clemens Kirisits, Otmar Scherzer, and Eric Setterqvist. Invariant φ-minimal sets and total variation denoising on graphs. SIAM Journal on Imaging Sciences, 12(4):1643–1668, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Kiefer:2020:ASO

[KSW20] Lukas Kiefer, Martin Storath, and Andreas Weinmann. An algorithm for second order Mumford–Shah models based on a Taylor jet formulation.

SIAM Journal on Imaging Sciences, 13(4):2307–2360, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Katsevich:2011:LTM

[KSZ11] A. Katsevich, M. Silver, and A. Zamyatin. Local tomography and the motion estimation problem. SIAM Journal on Imaging Sciences, 4(1): 200–219, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p200_s1.

Kutyniok:2012:SRD

[KSZ12] Gitta Kutyniok, Morteza Shahram, and Xiaosheng Zhuang. Shear-Lab: a rational design of a digital parabolic scaling algorithm. SIAM Journal on Imaging Sciences, 5(4):1291–1332, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Kim:2014:INC

[KT14] Yunho Kim and Hemant D.

Tagare. Intensity nonuniformity correction for brain MR images with known voxel classes. SIAM Journal on Imaging Sciences, 7 (1):528–557, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Kuchment:2016:TDI

[KT16] Peter Kuchment and Fatma Terzioglu. Three-dimensional image reconstruction from Compton camera data. SIAM Journal on Imaging Sciences, 9 (4):1708–1725, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Kim:2022:HRQ

[KT22] Arnold D. Kim and Chrysoula Tsogka. High-resolution, quantitative signal subspace imaging for synthetic aperture radar. SIAM Journal on Imaging Sciences, 15(3):1229– 1252, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1467109.

Kutyniok:2013:CSS

[Kut13] Gitta Kutyniok. Clustered sparsity and separation of cartoon and texture. SIAM Journal on Imaging Sciences, 6(2):848–874, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Krahmer:2023:EDH

[KV23] Felix Krahmer and Anna Veselovska. Enhanced digital halftoning via weighted sigmadelta modulation. SIAM Journal on Imaging Sciences, 16(3): 1727–1761, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M151786X.

Koenderink:2012:GFP

[KvD12] Jan Koenderink and Andrea van Doorn. Gauge fields in pictorial space. SIAM Journal on Imaging Sciences, 5(4): 1213–1233, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Koenderink:2018:LIS

[KvD18] Jan Koenderink and Andrea van Doorn. Local image structure and Procrustes metrics. SIAM Journal on Imaging Sciences, 11(1):293–324, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Ke:2020:RHR

[KWRC20] Rihuan Ke, Roland Wagner, Ronny Ramlau, and Raymond Chan. Reconstruction of the high resolution phase in a closed loop adaptive optics system. SIAM Journal on Imaging Sciences, 13(2):775–806, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Kang:2013:TLC

[KYW13] Myungjoo Kang, Sangwoon Yun, and Hyenkyun Woo. Twolevel convex relaxed variational model for multiplicative denoising. SIAM Journal on Imaging Sciences, 6(2):875–903, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Kunsberg:2014:HSC

[KZ14] Benjamin Kunsberg and Steven W.

Zucker. How shading constrains surface patches without knowledge of light sources. SIAM Journal on Imaging Sciences, 7 (2):641–668, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Knaus:2015:DDF

[KZ15] C. Knaus and M. Zwicker. Dualdomain filtering. SIAM Journal on Imaging Sciences, 8(3): 1396–1420, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Kunsberg:2018:CCI

[KZ18] Benjamin Kunsberg and Steven W. Zucker. Critical contours: an invariant linking image flow with salient surface organization. SIAM Journal on Imaging Sciences, 11(3):1849–1877, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Kang:2014:ISC

[KZS14] Sung Ha Kang, Wei Zhu, and Jackie Jianhong Shen. Illusory shapes via corner fusion. SIAM Journal on Imaging Sciences, 7 (4):1907–1936, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Lang:2019:NFE

[Lan19] Lukas F. Lang. A numerical framework for efficient motion estimation on evolving spherelike surfaces based on brightness and mass conservation laws. SIAM Journal on Imaging Sciences, 12(1):459–491, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Larsson:2018:RDS

[LAZ^{+}18] Måns Larsson, Anurag Arnab, Shuai Zheng, Philip Torr, and Fredrik Kahl. Revisiting deep structured models for pixel-level labeling with gradient-based inference. SIAM Journal on Imaging Sciences, 11(4):2610– 2628, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Laville:2023:GCR

[LBFA23] Bastien Laville, Laure Blanc-Féraud, and Gilles Aubert.

Off-the-grid curve reconstruction through divergence regularization: an extreme point result. SIAM Journal on Imaging Sciences, 16(2):867– 885, ???? 2023. CODEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1494373.

Lebrun:2013:NBI

[LBM13] M. Lebrun, A. Buades, and J. M. Morel. A nonlocal Bayesian image denoising algorithm. SIAM Journal on Imaging Sciences, 6(3):1665–1688, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Liu:2022:VRN

[LCD22] Zhifang Liu, Huibin Chang, and Yuping Duan. Variational Rician noise removal via splitting on spheres. SIAM Journal on Imaging Sciences, 15 (2):521–549, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1452792.

Li:2016:ADH

[LCS^{+}16] Yan-Ran Li, Raymond H. Chan, Lixin Shen, Yung-Chin Hsu, and Wen-Yih Isaac Tseng. An adaptive directional Haar framelet-based reconstruction algorithm for parallel magnetic resonance imaging. SIAM Journal on Imaging Sciences, 9(2): 794–821, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Laumont:2022:BIU

[LDA^{+}22] Rémi Laumont, Valentin De Bortoli, Andrés Almansa, Julie Delon, Alain Durmus, and Marcelo Pereyra. Bayesian imaging using plug & play priors: When Langevin meets Tweedie. SIAM Journal on Imaging Sciences, 15(2):701– 737, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1406349.

LeTarnec:2014:PCH

[LDCG14] Louis Le Tarnec, François Destrempes, Guy Cloutier, and Damien Garcia. A proof of convergence of the Horn–Schunck optical flow algorithm in arbitrary dimension. SIAM Journal on Imaging Sciences, 7(1): 277–293, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Launay:2021:DPP

[LDG21] Claire Launay, Agnès Desolneux, and Bruno Galerne. Determinantal point processes for image processing. SIAM Journal on Imaging Sciences, 14(1): 304–348, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Lebrat:2019:OTA

[LdGKW19] Léo Lebrat, Frédéric de Gournay, Jonas Kahn, and Pierre Weiss. Optimal transport approximation of 2-dimensional measures. SIAM Journal on Imaging Sciences, 12(2):762– 787, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Li:2020:ESB

[LDS20] Zhaoxing Li, Zhiliang Deng, and Jiguang Sun. Extendedsampling-Bayesian method for limited aperture inverse scattering problems. SIAM Journal on Imaging Sciences, 13(1): 422–444, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Lou:2014:PBD

[LEZX14] Yifei Lou, Ernie Esser, Hongkai Zhao, and Jack Xin. Partially blind deblurring of barcode from out-of-focus blur. SIAM Journal on Imaging Sciences, 7(2):740–760, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

LePendu:2023:PPP

[LG23] Mikael Le Pendu and Christine Guillemot. Preconditioned plug-and-play ADMM with locally adjustable denoiser for image restoration. SIAM Journal on Imaging Sciences, 16 (1):393–422, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1504809.

Liu:2018:FSO

[LGCWY18] Jialin Liu, Cristina Garcia-Cardona, Brendt Wohlberg, and Wotao Yin. First- and second-order methods for online convolutional dictionary learning. SIAM Journal on Imaging Sciences, 11(2):1589–1628, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Li:2022:ISA

[LGL^{+}22] Haifeng Li, Weihong Guo, Jun Liu, Li Cui, and Dongxing Xie. Image segmentation with adaptive spatial priors from joint registration. SIAM Journal on Imaging Sciences, 15(3): 1314–1344, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1444874.

Li:2018:AKI

[LH18] Housen Li and Markus Haltmeier. The averaged Kaczmarz iteration for solving inverse problems. SIAM Journal on Imaging Sciences, 11(1): 618–642, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Lucka:2018:ECS

[LHB^{+}18] Felix Lucka, Nam Huynh, Marta Betcke, Edward Zhang, Paul Beard, Ben Cox, and Simon Arridge. Enhancing compressed sensing 4D photoacoustic tomography by simultaneous motion estimation. SIAM Journal on Imaging Sciences, 11(4): 2224–2253, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Li:2023:SII

[LHC^{+}23] Jianfei Li, Chaoyan Huang, Raymond Chan, Han Feng, Michael K. Ng, and Tieyong Zeng. Spherical image inpainting with frame transformation and data-driven prior deep networks. SIAM Journal on Imaging Sciences, 16(3):

1177–1194, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M152462X.

Letourneau:2020:LFE

[LHLP20] Pierre-David Letourneau, Mitchell Tong Harris, Matthew Harper Langston, and George Papanicolaou. Low-frequency electromagnetic imaging using sensitivity functions and beamforming. SIAM Journal on Imaging Sciences, 13(2):807–843, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Lunz:2021:LOC

[LHT^{+}21] Sebastian Lunz, Andreas Hauptmann, Tanja Tarvainen, Carola-Bibiane Schönlieb, and Simon Arridge. On learned operator correction in inverse problems. SIAM Journal on Imaging Sciences, 14(1):92–127, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Lee:2015:ITU

[LHW^{+}15] Minji Lee, Yoseob Han, John Paul Ward, Michael Unser, and Jong Chul Ye. Interior tomography using 1D generalized total variation. Part II: Multiscale implementation. SIAM Journal on Imaging Sciences, 8(4): 2452–2486, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Lindeberg:2018:DSS

[Lin18] Tony Lindeberg. Dense scale selection over space, time, and space-time. SIAM Journal on Imaging Sciences, 11(1):407– 441, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Lipman:2014:BMM

[Lip14] Yaron Lipman. Bijective mappings of meshes with boundary and the degree in mesh processing. SIAM Journal on Imaging Sciences, 7(2):1263–1283, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Liu:2021:FIS

[Liu21] Keji Liu. Fast imaging of sources and scatterers in a stratified ocean waveguide. SIAM Journal on Imaging Sciences, 14(1):224–245, ???? 2021. CO-DEN SJISBI. ISSN 1936-4954.

Liu:2022:IAC

[LJL22] Huan Liu, Bangti Jin, and Xiliang Lu. Imaging anisotropic conductivities from current densities. SIAM Journal on Imaging Sciences, 15(2):860– 891, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1437810.

Lee:2018:SMP

[LKR18] Kiryung Lee, Felix Krahmer, and Justin Romberg. Spectral methods for passive imaging: Nonasymptotic performance and robustness. SIAM Journal on Imaging Sciences, 11(3):2110–2164, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Lehnert:2019:LSB

[LKW^{+}19] Judith Lehnert, Christoph Kolbitsch, Gerd Wübbeler, Amedeo Chiribiri, Tobias Schaeffter, and Clemens Elster. Large-scale Bayesian spatialtemporal regression with application to cardiac MR-perfusion imaging. SIAM Journal on Imaging Sciences, 12(4):2035– 2062, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Lam:2014:LIB

[LL14] Ka Chun Lam and Lok Ming Lui. Landmark- and intensitybased registration with large deformations via quasi-conformal maps. SIAM Journal on Imaging Sciences, 7(4):2364–2392, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Lin:2022:HBS

[LL22] Chenran Lin and Lok Ming Lui. Harmonic Beltrami signature: A novel 2D shape representation for object classification. SIAM Journal on Imaging Sciences, 15(4):1851– 1893, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1470852.

Lenzen:2014:SQV

[LLBS14] F. Lenzen, J. Lellmann, F. Becker, and C. Schnörr. Solving quasivariational inequalities for image restoration with adaptive constraint sets. SIAM Journal on Imaging Sciences, 7(4):

2139–2174, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Liao:2014:GAP

[LLC14] Xuejun Liao, Hui Li, and Lawrence Carin. Generalized alternating projection for weighted-ℓ_{2,1} minimization with applications to model-based compressive sensing. SIAM Journal on Imaging Sciences, 7 (2):797–823, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Li:2017:MAC

[LLLX17] Zheng Li, Yiguang Liu, Jipeng Li, and Wenzheng Xu. The mapping-adaptive convolution: a fundamental theory for homography or perspective invariant matching methods. SIAM Journal on Imaging Sciences, 10(4):1767–1803, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Liu:2013:ALB

[LLS^{+}13] Qiegen Liu, Dong Liang, Ying Song, Jianhua Luo, Yuemin Zhu, and Wenshu Li. Augmented Lagrangian-based sparse representation method with dictionary updating for image deblurring. SIAM Journal on Imaging Sciences, 6(3): 1689–1718, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Liu:2019:ESM

[LLS19] J. Liu, X. Liu, and J. Sun. Extended sampling method for inverse elastic scattering problems using one incident wave. SIAM Journal on Imaging Sciences, 12(2):874–892, ???? 2019. CO-DEN SJISBI. ISSN 1936-4954.

Liu:2020:MNR

[LLS^{+}20] Xiaoxia Liu, Jian Lu, Lixin Shen, Chen Xu, and Yuesheng Xu. Multiplicative noise removal: Nonlocal low-rank model and its proximal alternating reweighted minimization algorithm. SIAM Journal on Imaging Sciences, 13(3):1595– 1629, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Lellmann:2014:IKR

[LLSV14] Jan Lellmann, Dirk A. Lorenz, Carola Schönlieb, and Tuomo Valkonen. Imaging with Kantorovich–Rubinstein discrepancy. SIAM Journal on Imaging Sciences, 7(4):2833– 2859, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Leung:2009:EMA

[LLSZ09] Shingyu Leung, Gang Liang, Knut Solna, and Hongkai Zhao. Expectation-maximization algorithm with local adaptivity. SIAM Journal on Imaging Sciences, 2(3):834–857, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Li:2013:LMM

[LLW13] Jingzhi Li, Hongyu Liu, and Qi Wang. Locating multiple multiscale electromagnetic scatterers by a single far-field measurement. SIAM Journal on Imaging Sciences, 6(4):2285– 2309, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Li:2023:DAT

[LLW23] Peijun Li, Ying Liang, and Yuliang Wang. A dataassisted two-stage method for the inverse random source problem. SIAM Journal on Imaging Sciences, 16(4):1929– 1952, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/23M1562561.

Lui:2013:TMV

[LLWG13] Lok Ming Lui, Ka Chun Lam, Tsz Wai Wong, and Xianfeng Gu. Texture map and video compression using Beltrami representation. SIAM Journal on Imaging Sciences, 6(4):1880– 1902, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Lui:2014:TMM

[LLYG14] Lok Ming Lui, Ka Chun Lam, Shing-Tung Yau, and Xianfeng Gu. Teichmuller mapping (T -map) and its applications to landmark matching registration. SIAM Journal on Imaging Sciences, 7(1):391–426, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Louchet:2011:TVL

[LM11] Cécile Louchet and Lionel Moisan. Total variation as a local filter. SIAM Journal on Imaging Sciences, 4(2): 651–694, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i2/p651_s1.

Louchet:2013:PET

[LM13] Cécile Louchet and Lionel Moisan. Posterior expectation of the total variation model: Properties and experiments. SIAM Journal on Imaging Sciences, 6(4):2640–2684, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Lasica:2017:TVD

[LMM17] Michal Lasica, Salvador Moll, and Piotr B. Mucha. Total variation denoising in l^{1} anisotropy. SIAM Journal on Imaging Sciences, 10(4):1691– 1723, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Lanza:2019:SIN

[LMSS19] Alessandro Lanza, Serena Morigi, Ivan W. Selesnick, and Fiorella Sgallari. Sparsityinducing nonconvex nonseparable regularization for convex image processing. SIAM Journal on Imaging Sciences, 12(2): 1099–1134, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Lanza:2013:VID

[LMSY13] Alessandro Lanza, Serena Morigi, Fiorella Sgallari, and Anthony J. Yezzi. Variational image denoising based on autocorrelation whiteness. SIAM Journal on Imaging Sciences, 6(4): 1931–1955, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Levinson:2023:IBL

[LMT23] Howard W. Levinson, Vadim Markel, and Nicholas Triantafillou. Inversion of band-limited discrete Fourier transforms of binary images: Uniqueness and algorithms. SIAM Journal on Imaging Sciences, 16(3): 1338–1369, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1540442.

Lechleiter:2013:FME

[LN13] Armin Lechleiter and Dinh-Liem Nguyen. Factorization method for electromagnetic inverse scattering from biperiodic structures. SIAM Journal on Imaging Sciences, 6(2): 1111–1139, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Laus:2017:NDA

[LNPS17] Friederike Laus, Mila Nikolova, Johannes Persch, and Gabriele Steidl. A nonlocal denoising algorithm for manifold-valued images using second order statistics. SIAM Journal on Imaging Sciences, 10(1):416–448, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Li:2010:MNR

[LNS10] Fang Li, Michael K. Ng, and Chaomin Shen. Multiplicative noise removal with spatially varying regularization parameters. SIAM Journal on Imaging Sciences, 3(1):1–20, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Li:2010:MIS

[LNZS10] Fang Li, Michael K. Ng, Tie Yong Zeng, and Chunli Shen. A multiphase image segmentation method based on fuzzy region competition. SIAM Journal on Imaging Sciences, 3 (3):277–299, ???? 2010. CO-DEN SJISBI. ISSN 1936-4954.

Landrieu:2017:CPF

[LO17] Loic Landrieu and Guillaume Obozinski. Cut pursuit: Fast algorithms to learn piecewise constant functions on general weighted graphs. SIAM Journal on Imaging Sciences, 10(4): 1724–1766, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Louis:2008:CIR

[Lou08] Alfred K. Louis. Combining image reconstruction and image analysis with an application to two-dimensional tomography. SIAM Journal on Imaging Sciences, 1(2):188–208, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Lee:2019:FNB

[LP19] Chang-Ock Lee and Jongho Park. Fast nonoverlapping block Jacobi method for the dual Rudin–Osher–Fatemi model. SIAM Journal on Imaging Sciences, 12(4):2009–2034, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Lenglet:2009:BCM

[LPP^{+}09] Christophe Lenglet, Emmanuel Prados, Jean-Philippe Pons, Rachid Deriche, and Olivier Faugeras. Brain connectivity mapping using Riemannian geometry, control theory, and PDEs. SIAM Journal on Imaging Sciences, 2(2):285–322, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Lellmann:2015:AAN

[LPSS15] Jan Lellmann, Konstantinos Papafitsoros, Carola Schönlieb, and Daniel Spector. Analysis and application of a nonlocal Hessian. SIAM Journal on Imaging Sciences, 8(4):2161– 2202, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Leibovich:2020:GCB

[LPT20a] Matan Leibovich, George Papanicolaou, and Chrysoula Tsogka. Generalized correlationbased imaging for satellites. SIAM Journal on Imaging Sciences, 13(3):1331–1366, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Leibovich:2020:SAI

[LPT20b] Matan Leibovich, George Papanicolaou, and Chrysoula Tsogka. Synthetic aperture imaging and motion estimation using tensor methods. SIAM Journal on Imaging Sciences, 13(4):2213–2249, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Leibovich:2021:CBI

[LPT21] Matan Leibovich, George Papanicolaou, and Chrysoula Tsogka. Correlation based imaging for rotating satellites. SIAM Journal on Imaging Sciences, 14(1):271–303, ????

2021. CODEN SJISBI. ISSN 1936-4954.

Luo:2014:ASM

[LQS14] Songting Luo, Jianliang Qian, and Plamen Stefanov. Adjoint state method for the identification problem in SPECT: Recovery of both the source and the attenuation in the attenuated X-ray transform. SIAM Journal on Imaging Sciences, 7(2): 696–715, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Liu:2023:ACM

[LQZ23] Chaoyu Liu, Zhonghua Qiao, and Qian Zhang. An active contour model with local variance force term and its efficient minimization solver for multiphase image segmentation. SIAM Journal on Imaging Sciences, 16(1):144–168, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1483645.

Lopez-Rubio:2016:SSN

[LR16] Ezequiel López-Rubio. Superresolution from a single noisy image by the median filter transform. SIAM Journal on Imaging Sciences, 9(1):82–115, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

LeLouer:2017:TSS

[LR17] Frédérique Le Louër and María-Luisa Rapún. Topological sensitivity for solving inverse multiple scattering problems in threedimensional electromagnetism. Part I: One step method. SIAM Journal on Imaging Sciences, 10(3):1291–1321, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

LeLouer:2018:TSS

[LR18] Frédérique Le Louër and María-Luisa Rapún. Topological sensitivity for solving inverse multiple scattering problems in threedimensional electromagnetism. Part II: Iterative method. SIAM Journal on Imaging Sciences, 11(1):734–769, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Lefkimmiatis:2015:STT

[LRMU15] Stamatios Lefkimmiatis, Anastasios Roussos, Petros Maragos, and Michael Unser. Structure tensor total variation. SIAM Journal on Imaging Sciences, 8 (2):1090–1122, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Lisani:2017:CDF

[LRP17] Jose-Luis Lisani, Silvia Ramis, and Francisco J. Perales. A contrario detection of faces: a case example. SIAM Journal on Imaging Sciences, 10(4):2091– 2118, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Li:2021:MHI

[LRV21] Wen Li, Elena Resmerita, and Luminita A. Vese. Multiscale hierarchical image decomposition and refinements: Qualitative and quantitative results. SIAM Journal on Imaging Sciences, 14(2):844–877, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Lellmann:2011:CML

[LS11] J. Lellmann and C. Schnörr.

Continuous multiclass labeling approaches and algorithms. SIAM Journal on Imaging Sciences, 4(4):1049– 1096, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i4/p1049_s1.

Landa:2017:SPC

[LS17] Boris Landa and Yoel Shkolnisky. Steerable principal components for space-frequency localized images. SIAM Journal on Imaging Sciences, 10(2): 508–534, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Landa:2018:SGL

[LS18a] Boris Landa and Yoel Shkolnisky. The steerable graph Laplacian and its application to filtering image datasets. SIAM Journal on Imaging Sciences, 11(4):2254–2304, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Ling:2018:SCB

[LS18b] Shuyang Ling and Thomas Strohmer. Self-calibration and bilinear inverse problems via linear least squares. SIAM Journal on Imaging Sciences, 11(1):252–292, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Laus:2019:MMF

[LS19] Friederike Laus and Gabriele Steidl. Multivariate Myriad filters based on parameter estimation of Student-t distributions. SIAM Journal on Imaging Sciences, 12(4):1864–1904, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Li:2018:MCB

[LSC^{+}18] Bin Li, Chenyang Shen, Yujie Chi, Ming Yang, Yifei Lou, Linghong Zhou, and Xun Jia. Multienergy cone-beam computed tomography reconstruction with a spatial spectral nonlocal means algorithm. SIAM Journal on Imaging Sciences, 11(2):1205–1229, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Lorenz:2014:LBM

[LSW14] Dirk A. Lorenz, Frank Schöpfer, and Stephan Wenger. The linearized Bregman method via split feasibility problems: Analysis and generalizations. SIAM Journal on Imaging Sciences, 7 (2):1237–1262, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Li:2022:AIR

[LSWW22] Zhemin Li, Tao Sun, Hongxia Wang, and Bao Wang. Adaptive and implicit regularization for matrix completion. SIAM Journal on Imaging Sciences, 15(4): 2000–2022, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1489228.

Li:2021:AEI

[LSYZ21] Wei Li, John C. Schotland, Yang Yang, and Yimin Zhong.

An acousto-electric inverse source problem. SIAM Journal on Imaging Sciences, 14(4): 1601–1616, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Lerman:2018:ECL

[LSZ18] Gilad Lerman, Yunpeng Shi, and Teng Zhang. Exact camera location recovery by least unsquared deviations. SIAM Journal on Imaging Sciences, 11(4): 2692–2721, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Liu:2021:CEM

[LTKG21] Hao Liu, Xue-Cheng Tai, Ron Kimmel, and Roland Glowinski. A color elastica model for vector-valued image regularization. SIAM Journal on Imaging Sciences, 14(2):717–748, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Liu:2023:EMC

[LTKG23] Hao Liu, Xue-Cheng Tai, Ron Kimmel, and Roland Glowinski. Elastica models for color image regularization. SIAM Journal on Imaging Sciences, 16(1):461–500, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M147935X.

Lui:2010:OCS

[LTW^{+}10] Lok Ming Lui, Sheshadri Thiruvenkadam, Yalin Wang, Paul M. Thompson, and Tony F. Chan. Optimized conformal surface registration with shape-based landmark matching. SIAM Journal on Imaging Sciences, 3 (1):52–78, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Lai:2022:SPX

[LUZZ22] Ru-Yu Lai, Gunther Uhlmann, Jian Zhai, and Hanming Zhou. Single pixel X-ray transform and related inverse problems. SIAM Journal on Imaging Sciences, 15(4):1894– 1909, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1468103.

LaTorre:2009:MVI

[LVEB09] D. La Torre, E. R. Vrscay, M. Ebrahimi, and M. F. Barnsley. Measure-valued images, associated fractal transforms, and the affine self-similarity of images. SIAM Journal on Imaging Sciences, 2(2):470–507, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Lui:2014:GRH

[LW14] Lok Ming Lui and Chengfeng Wen. Geometric registration of high-genus surfaces. SIAM Journal on Imaging Sciences, 7 (1):337–365, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Lin:2018:MVI

[LWM^{+}18] Chia-Hsiang Lin, Ruiyuan Wu, Wing-Kin Ma, Chong-Yung Chi, and Yue Wang. Maximum volume inscribed ellipsoid: a new simplex-structured matrix factorization framework via facet enumeration and convex optimization. SIAM Journal on Imaging Sciences, 11(2):

1651–1679, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Liu:2016:MDN

[LWY16] Hongyu Liu, Yuliang Wang, and Can Yang. Mathematical design of a novel gesture-based instruction/input device using wave detection. SIAM Journal on Imaging Sciences, 9(2): 822–841, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Lau:2012:FMS

[LY12] Tze Siong Lau and Andy M.

Yip. A fast method for segmenting images with additive intensity value. SIAM Journal on Imaging Sciences, 5(3): 993–1021, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Lai:2013:ANN

[LY13] Ming-Jun Lai and Wotao Yin.

Augmented ℓ_{1} and nuclearnorm models with a globally linearly convergent algorithm. SIAM Journal on Imaging Sciences, 6(2):1059–1091, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Li:2015:PSC

[LY15] Xinxin Li and Xiaoming Yuan.

A proximal strictly contractive Peaceman–Rachford splitting method for convex programming with applications to imaging. SIAM Journal on Imaging Sciences, 8(2):1332– 1365, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Lu:2018:PSS

[LY18] Jianfeng Lu and Haizhao Yang. Phase-space sketching for crystal image analysis based on synchrosqueezed transforms. SIAM Journal on Imaging Sciences, 11(3):1954–1978, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Li:2020:TSV

[LYZ20] Xu Li, Xiaoping Yang, and Tieyong Zeng. A three-stage variational image segmentation framework incorporating intensity inhomogeneity information. SIAM Journal on Imaging Sciences, 13(3):1692–1715, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Li:2016:NAF

[LZ16] Fang Li and Tieyong Zeng. A new algorithm framework for image inpainting in transform domain. SIAM Journal on Imaging Sciences, 9(1):24–51, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Lai:2017:MNP

[LZ17a] Rongjie Lai and Hongkai Zhao. Multiscale nonrigid point cloud registration using rotationinvariant sliced-Wasserstein distance via Laplace–Beltrami eigenmap. SIAM Journal on Imaging Sciences, 10(2):449– 483, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Liu:2017:BNT

[LZ17b] Jun Liu and Xiaojun Zheng. A block nonlocal TV method for image restoration. SIAM Journal on Imaging Sciences, 10(2): 920–941, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Lei:2018:LTR

[LZ18] Yunwen Lei and Ding-Xuan Zhou. Learning theory of randomized sparse Kaczmarz method. SIAM Journal on Imaging Sciences, 11(1):547– 574, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Liu:2016:WFM

[LZD^{+}16] Jiulong Liu, Xiaoqun Zhang, Bin Dong, Zuowei Shen, and Lixu Gu. A wavelet frame method with shape prior for ultrasound video segmentation. SIAM Journal on Imaging Sciences, 9(2):495–519, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Lou:2015:WDA

[LZOX15] Yifei Lou, Tieyong Zeng, Stanley Osher, and Jack Xin. A weighted difference of anisotropic and isotropic total variation model for image processing. SIAM Journal on Imaging Sciences, 8(3):1798– 1823, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Liu:2018:DIM

[LZZ18] Xiaoli Liu, Bo Zhang, and Haiwen Zhang. A direct imaging method for inverse scattering by unbounded rough surfaces. SIAM Journal on Imaging Sciences, 11(2):1629–1650, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Li:2023:LGT

[LZZ^{+}23] Ben-Zheng Li, Xi-Le Zhao, Xiongjun Zhang, Teng-Yu Ji, Xinyu Chen, and Michael K. Ng. A learnable grouptube transform induced tensor nuclear norm and its application for tensor completion. SIAM Journal on Imaging Sciences, 16(3):1370– 1397, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1531907.

Mahmoodi:2012:EDF

[Mah12] Sasan Mahmoodi. Edge detection filter based on Mumford– Shah Green function. SIAM Journal on Imaging Sciences, 5 (1):343–365, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Maurel:2011:LPT

[MAP11] Pierre Maurel, Jean-François Aujol, and Gabriel Peyré. Locally parallel texture modeling. SIAM Journal on Imaging Sciences, 4(1):413–447, ???? 2011. CODEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p413_s1.

Marquina:2009:NIS

[Mar09] Antonio Marquina. Nonlinear inverse scale space methods for total variation blind deconvolution. SIAM Journal on Imaging Sciences, 2(1):64–83, ???? 2009.

CODEN SJISBI. ISSN 1936-4954.

Marz:2011:IIB

[Mär11]  Thomas März.  Image in-

painting based on coherence transport with adapted distance functions. SIAM Journal on Imaging Sciences, 4(4): 981–1000, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i4/p981_s1.

Molano:2022:ACS

[MÁS^{+}22] Rubén Molano, Mar Ávila, José Carlos Sancho, Pablo G. Rodríguez, and Andres Caro. An algorithm to compute any simple k-gon of a maximum area or perimeter inscribed in a region of interest. SIAM Journal on Imaging Sciences, 15(4): 1808–1832, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1482676.

Muhammed:2010:SAM

[MB10] Hamed Hamid Muhammed and Fredrik Bergholm. Sensitivity analysis of multichannel images intended for instantaneous imaging spectrometry applications. SIAM Journal on Imaging Sciences, 3(1):79–109, ???? 2010. CODEN ????

Mang:2015:INK

[MB15] Andreas Mang and George Biros. An inexact Newton– Krylov algorithm for constrained diffeomorphic image registration. SIAM Journal on Imaging Sciences, 8(2):1030– 1069, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Mang:2016:CRS

[MB16] Andreas Mang and George Biros. Constrained H^{1}- regularization schemes for diffeomorphic image registration. SIAM Journal on Imaging Sciences, 9(3):1154–1194, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Moeller:2014:CBT

[MBBS14] Michael Moeller, Eva-Maria Brinkmann, Martin Burger, and Tamara Seybold. Color Bregman TV. SIAM Journal on Imaging Sciences, 7(4):2771– 2806, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

McDonald:2020:STD

[MBFG20] Joseph McDonald, Brett Bernstein, and Carlos Fernandez-Granda. A sampling theorem for deconvolution in two dimensions. SIAM Journal on Imaging Sciences, 13(4):1754–1780, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Morgenshtern:2016:SRP

[MC16] Veniamin I. Morgenshtern and Emmanuel J. Candès. Superresolution of positive sources: The discrete setup. SIAM Journal on Imaging Sciences, 9(1): 412–444, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Meng:2016:TFE

[MCL16] Ting Wei Meng, Gary Pui-Tung Choi, and Lok Ming Lui. TEMPO: Feature-endowed Teichmüller extremal mappings of point clouds. SIAM Journal on Imaging Sciences, 9(4):1922– 1962, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Mennucci:2015:BLD

[MD15] A. C. G. Mennucci and A. Duci.

Banach-like distances and metric spaces of compact sets. SIAM Journal on Imaging Sciences, 8(1):19–66, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Melidonis:2023:EBC

[MDA^{+}23] Savvas Melidonis, Paul Dobson, Yoann Altmann, Marcelo Pereyra, and Konstantinos Zygalakis. Efficient Bayesian computation for low-photon imaging problems. SIAM Journal on Imaging Sciences, 16(3): 1195–1234, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1502240.

Mecca:2013:UAP

[MF13] Roberto Mecca and Maurizio Falcone. Uniqueness and approximation of a photometric shape-from-shading model. SIAM Journal on Imaging Sciences, 6(1):616–659, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Myllykoski:2015:NAL

[MGKR15] M. Myllykoski, R. Glowinski, T. Kärkkäinen, and T. Rossi. A new augmented Lagrangian approach for L^{1}-mean curvature image denoising. SIAM Journal on Imaging Sciences, 8(1): 95–125, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Moon:2017:AIC

[MH17] Sunghwan Moon and Markus Haltmeier. Analytic inversion of a conical Radon transform arising in application of Compton cameras on the cylinder. SIAM Journal on Imaging Sciences, 10(2):535–557, ???? 2017. CO-DEN SJISBI. ISSN 1936-4954.

Moon:2023:SVD

[MHM23] Minam Moon, Injo Hur, and Sunghwan Moon. Singular value decomposition of the wave forward operator with radial variable coefficients. SIAM Journal on Imaging Sciences, 16(3): 1520–1534, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1511643.

Miolane:2017:TSE

[MHP17] Nina Miolane, Susan Holmes, and Xavier Pennec. Template shape estimation: Correcting an asymptotic bias. SIAM Journal on Imaging Sciences, 10(2): 808–844, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Milanfar:2013:SSF

[Mil13] Peyman Milanfar. Symmetrizing smoothing filters. SIAM Journal on Imaging Sciences, 6 (1):263–284, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Milanfar:2018:RRW

[Mil18] Peyman Milanfar. Rendition:

Reclaiming what a black box takes away. SIAM Journal on Imaging Sciences, 11(4):2722– 2756, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Moshtaghpour:2019:VDS

[MJC^{+}19] A. Moshtaghpour, L. Jacques, V. Cambareri, P. Antoine, and M. Roblin. A variable density sampling scheme for compressive Fourier transform interferometry. SIAM Journal on Imaging Sciences, 12(2):671– 715, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Merkurjev:2013:MSG

[MKB13] Ekaterina Merkurjev, Tijana Kostić, and Andrea L. Bertozzi. An MBO scheme on graphs for classification and image processing. SIAM Journal on Imaging Sciences, 6(4):1903– 1930, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Ma:2017:TMS

[MLH17] Tian-Hui Ma, Yifei Lou, and Ting-Zhu Huang. Truncated l_{1−2} models for sparse recovery and rank minimization. SIAM Journal on Imaging Sciences, 10(3):1346–1380, ???? 2017.

CODEN SJISBI. ISSN 1936-4954.

Micheli:2012:SCT

[MMM12] Mario Micheli, Peter W. Michor, and David Mumford. Sectional curvature in terms of the cometric, with applications to the Riemannian manifolds of landmarks. SIAM Journal on Imaging Sciences, 5(1):394–433, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Merigot:2018:AOT

[MMT18] Quentin Mérigot, Jocelyn Meyron, and Boris Thibert. An algorithm for optimal transport between a simplex soup and a point cloud. SIAM Journal on Imaging Sciences, 11(2):1363– 1389, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Moscoso:2016:CIP

[MNP16] Miguel Moscoso, Alexei Novikov, and George Papanicolaou. Coherent imaging without phases. SIAM Journal on Imaging Sciences, 9(4):1689–1707, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Moscoso:2017:MII

[MNPT17] Miguel Moscoso, Alexei Novikov, George Papanicolaou, and Chrysoula Tsogka. Multifrequency interferometric imaging with intensity-only measurements. SIAM Journal on Imaging Sciences, 10(3):1005–1032, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Monard:2014:NIG

[Mon14] François Monard. Numerical implementation of geodesic Xray transforms and their inversion. SIAM Journal on Imaging Sciences, 7(2):1335–1357, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Munoz-Perez:2019:EBS

[MPGMD19] Luis F. Muñoz-Pérez, José A.

Guerrero, and Jorge E. Macías-Díaz. Entropy-based selection of cluster representatives for document image compression. SIAM Journal on Imaging Sciences, 12(4):1720–1738, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Matthews:2018:PJR

[MPL^{+}18] Thomas P. Matthews, Joemini Poudel, Lei Li, Lihong V. Wang, and Mark A. Anastasio. Parameterized joint reconstruction of the initial pressure and sound speed distributions for photoacoustic computed tomography. SIAM Journal on Imaging Sciences, 11(2):1560–1588, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Mitsuhashi:2017:FAO

[MPM^{+}17] Kenji Mitsuhashi, Joemini Poudel, Thomas P. Matthews, Alejandro Garcia-Uribe, Lihong V. Wang, and Mark A. Anastasio. A forward-adjoint operator pair based on the elastic wave equation for use in transcranial photoacoustic computed tomography. SIAM Journal on Imaging Sciences, 10(4): 2022–2048, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Mecca:2016:SLP

[MQLC16] Roberto Mecca, Yvain Quéau, Fotios Logothetis, and Roberto Cipolla. A single-lobe photometric stereo approach for heterogeneous material. SIAM Journal on Imaging Sciences, 9 (4):1858–1888, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Majee:2020:GLI

[MRM20] Sudeb Majee, Rajendra K. Ray, and Ananta K. Majee. A gray level indicator-based regularized telegraph diffusion model: Application to image despeckling. SIAM Journal on Imaging Sciences, 13(2):844–870, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Morigi:2008:CMM

[MRSS08] Serena Morigi, Lothar Reichel, Fiorella Sgallari, and Andriy Shyshkov. Cascadic multiresolution methods for image deblurring. SIAM Journal on Imaging Sciences, 1(1):51–74, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Marsland:2017:LEL

[MS17] Stephen Marsland and Tony Shardlow. Langevin equations for landmark image registration with uncertainty. SIAM Journal on Imaging Sciences, 10(2): 782–807, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Marchetti:2022:CRI

[MS22] Francesco Marchetti and Gabriele Santin. Convergence results in image interpolation with the continuous SSIM. SIAM Journal on Imaging Sciences, 15(4): 1977–1999, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M147637X.

Museyko:2009:AMK

[MSKL09] O. Museyko, M. Stiglmayr, K. Klamroth, and G. Leugering. On the application of the Monge–Kantorovich problem to image registration. SIAM Journal on Imaging Sciences, 2(4): 1068–1097, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Mollenhoff:2015:PDH

[MSMC15] Thomas Möllenhoff, Evgeny Strekalovskiy, Michael Moeller, and Daniel Cremers. The primal-dual hybrid gradient method for semiconvex splittings. SIAM Journal on Imaging Sciences, 8(2):827–857, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Mecca:2014:DDA

[MTWB14] Roberto Mecca, Ariel Tankus, Aaron Wetzler, and Alfred M. Bruckstein. A direct differential approach to photometric stereo with perspective viewing. SIAM Journal on Imaging Sciences, 7 (2):579–612, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Muise:2009:CIA

[Mui09] Robert Muise. Compressive imaging: An application. SIAM Journal on Imaging Sciences, 2 (4):1255–1276, ???? 2009. CO-DEN SJISBI. ISSN 1936-4954.

Moller:2012:VAS

[MWBB12] Michael Möller, Todd Wittman, Andrea L. Bertozzi, and Martin Burger. A variational approach for sharpening high dimensional images. SIAM Journal on Imaging Sciences, 5(1):150–178, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Mecca:2014:NFP

[MWBK14] Roberto Mecca, Aaron Wetzler, Alfred M. Bruckstein, and Ron Kimmel. Near field photometric stereo with point light sources. SIAM Journal on Imaging Sciences, 7(4):2732– 2770, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ma:2021:CFB

[MWWY21] Jun Ma, Dong Wang, Xiao-Ping Wang, and Xiaoping Yang. A characteristic function-based algorithm for geodesic active contours. SIAM Journal on Imaging Sciences, 14(3):1184–1205, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Morel:2009:ANF

[MY09] Jean-Michel Morel and Guoshen Yu. ASIFT: a new framework for fully affine invariant image comparison. SIAM Journal on Imaging Sciences, 2(2):

438–469, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Ma:2013:SRP

[MYZ13] Liyan Ma, Jian Yu, and Tieyong Zeng. Sparse representation prior and total Variation– Based image deblurring under impulse noise. SIAM Journal on Imaging Sciences, 6(4):2258– 2284, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Newson:2014:VIC

[NAF^{+}14] Alasdair Newson, Andrés Almansa, Matthieu Fradet, Yann Gousseau, and Patrick Pérez. Video inpainting of complex scenes. SIAM Journal on Imaging Sciences, 7(4):1993–2019, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Najman:2017:EPW

[Naj17] Laurent Najman. Extending the power watershed framework thanks to Γ-convergence. SIAM Journal on Imaging Sciences, 10(4):2275–2292, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Natterer:2016:IRQ

[Nat16] Frank Natterer. Image reconstruction in quantitative susceptibility mapping. SIAM Journal on Imaging Sciences, 9(3): 1127–1131, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Ni:2011:EDC

[NDM^{+}11] Kangyu Ni, Somantika Datta, Prasun Mahanti, Svetlana Roudenko, and Douglas Cochran.

Efficient deterministic compressed sensing for images with chirps and Reed–Muller codes. SIAM Journal on Imaging Sciences, 4(3):931–953, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p931_s1.

Neto:2022:DEA

[NFV22] Jeova F. S. Rocha Neto, Pedro Felzenszwalb, and Marilyn Vazquez. Direct estimation of appearance models for segmentation. SIAM Journal on Imaging Sciences, 15 (1):172–191, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1400729.

Nguyen:2022:AFF

[NHKD22] Linh Nguyen, Markus Haltmeier, Richard Kowar, and Ngoc Do. Analysis for fullfield photoacoustic tomography with variable sound speed. SIAM Journal on Imaging Sciences, 15(3):1213– 1228, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1463409.

Nikolova:2013:DML

[Nik13] Mila Nikolova. Description of the minimizers of least squares regularized with ℓ_{0}- norm. uniqueness of the global minimizer. SIAM Journal on Imaging Sciences, 6(2):904–937, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Nguyen:2016:DTR

[NK16] Linh V. Nguyen and Leonid A.

Kunyansky. A dissipative time reversal technique for photoacoustic tomography in a cavity. SIAM Journal on Imaging Sciences, 9(2):748–769, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Needham:2020:STG

[NK20a] Tom Needham and Sebastian Kurtek. Simplifying transforms for general elastic metrics on the space of plane curves. SIAM Journal on Imaging Sciences, 13(1):445–473, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Newman:2020:NTP

[NK20b] Elizabeth Newman and Misha E.

Kilmer. Nonnegative tensor patch dictionary approaches for image compression and deblurring applications. SIAM Journal on Imaging Sciences, 13(3): 1084–1112, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Nowozin:2010:GIR

[NL10] Sebastian Nowozin and Christoph H.

Lampert. Global interactions in random field models: a potential function ensuring connectedness. SIAM Journal on Imaging Sciences, 3(4): 1048–1074, ???? 2010. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v3/i4/p1048_s1.

Niinimaki:2016:MPC

[NLH^{+}16] Kati Niinimäki, Matti Lassas, Keijo Hämäläinen, Aki Kallonen, Ville Kolehmainen, Esa Niemi, and Samuli Siltanen. Multiresolution parameter choice method for total variation regularized tomography. SIAM Journal on Imaging Sciences, 9(3):938–974, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Novikov:2015:ISI

[NMP15] Alexei Novikov, Miguel Moscoso, and George Papanicolaou. Illumination strategies for intensity-only imaging. SIAM Journal on Imaging Sciences, 8 (3):1547–1573, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Ng:2017:LBP

[NNYZ17] Michael K. Ng, Henry Y. T. Ngan, Xiaoming Yuan, and Wenxing Zhang. Lattice-based patterned fabric inspection by using total variation with sparsity and low-rank representations. SIAM Journal on Imaging Sciences, 10(4):2140–2164, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Nikolova:2008:ERP

[NNZC08] Mila Nikolova, Michael K. Ng, Shuqin Zhang, and Wai-Ki Ching. Efficient reconstruction of piecewise constant images using nonsmooth nonconvex minimization. SIAM Journal on Imaging Sciences, 1(1):

2–25, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Niethammer:2017:AMF

[NPJI17] Marc Niethammer, Kilian M.

Pohl, Firdaus Janoos, and William M. Wells III. Active mean fields for probabilistic image segmentation: Connections with Chan–Vese and Rudin– Osher–Fatemi models. SIAM Journal on Imaging Sciences, 10(3):1069–1103, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Neumayer:2018:MMV

[NPS18] Sebastian Neumayer, Johannes Persch, and Gabriele Steidl. Morphing of manifold-valued images inspired by discrete geodesics in image spaces. SIAM Journal on Imaging Sciences, 11(3):1898–1930, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Nardi:2016:GSS

[NPV16] Giacomo Nardi, Gabriel Peyré, and François-Xavier Vialard. Geodesics on shape spaces with bounded variation and Sobolev metrics. SIAM Journal on Imaging Sciences, 9(1):238–274, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Naetar:2014:QPT

[NS14] W. Naetar and O. Scherzer.

Quantitative photoacoustic tomography with piecewise constant material parameters. SIAM Journal on Imaging Sciences, 7(3):1755–1774, ????

2014. CODEN SJISBI. ISSN 1936-4954.

Ngole:2017:PSF

[NS17] F. Ngolé and J.-L. Starck. Point spread function field learning based on optimal transport distances. SIAM Journal on Imaging Sciences, 10(3):1549–1578, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Nakhmani:2011:PFR

[NT11] Arie Nakhmani and Allen Tannenbaum. Particle filtering with region-based matching for tracking of partially occluded and scaled targets. SIAM Journal on Imaging Sciences, 4(1): 220–242, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p220_s1.

Nguyen:2019:MRT

[NTDB19] Khac L. Nguyen, Mohamed M. Tekitek, Philippe Delachartre, and Michel Berthier. Multiple relaxation time lattice Boltzmann models for multigrid phase-field segmentation of tumors in 3D ultrasound images. SIAM Journal on Imaging Sciences, 12(3):1324–1346, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Ndiour:2010:PCO

[NTV10] Ibrahima J. Ndiour, Jochen Teizer, and Patricio A. Vela. A probabilistic contour observer for online visual tracking. SIAM Journal on Imaging Sciences, 3 (4):835–855, ???? 2010. CO-DEN SJISBI. ISSN 1936-4954.

Ng:2011:TVM

[NW11] Michael K. Ng and Wei Wang.

A total variation model for Retinex. SIAM Journal on Imaging Sciences, 4(1):345– 365, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p345_s1.

Needell:2013:SIR

[NW13a] Deanna Needell and Rachel Ward. Stable image reconstruction using total variation minimization. SIAM Journal on Imaging Sciences, 6(2):1035– 1058, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Ng:2013:VAI

[NW13b] Michael K. Ng and Wei Wang.

A variational approach for image stitching II: Using image gradients. SIAM Journal on Imaging Sciences, 6(3):1345– 1366, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

ONeill:2020:CCA

[OAUC^{+}20] Riley C. W. O’Neill, Pedro Angulo-Umaña, Jeff Calder, Bo Hessburg, Peter J. Olver, Chehrzad Shakiban, and Katrina Yezzi-Woodley. Computation of circular area and spherical volume invariants via boundary integrals. SIAM Journal on Imaging Sciences, 13(1):53–77, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Ochs:2014:IIP

[OCBP14] Peter Ochs, Yunjin Chen, Thomas Brox, and Thomas Pock. iPiano: Inertial proximal algorithm for nonconvex optimization. SIAM Journal on Imaging Sciences, 7(2):1388– 1419, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ouyang:2015:ALA

[OCLP15] Yuyuan Ouyang, Yunmei Chen, Guanghui Lan, and Eduardo Pasiliao, Jr. An accelerated linearized alternating direction method of multipliers. SIAM Journal on Imaging Sciences, 8 (1):644–681, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Ochs:2015:IRA

[ODBP15] Peter Ochs, Alexey Dosovitskiy, Thomas Brox, and Thomas Pock. On iteratively reweighted algorithms for nonsmooth nonconvex optimization in computer vision. SIAM Journal on Imaging Sciences, 8(1):331–372, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Ozere:2015:JSR

[OGL15] Solène Ozeré, Christian Gout, and Carole Le Guyader. Joint segmentation/registration model by shape alignment via weighted total variation minimization and nonlinear elasticity. SIAM Journal on Imaging Sciences, 8 (3):1981–2020, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Ongie:2016:GRP

[OJ16] Greg Ongie and Mathews Jacob. Off-the-grid recovery of piecewise constant images from few Fourier samples. SIAM Journal on Imaging Sciences, 9 (3):1004–1041, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Ozyesil:2015:SCM

[ÖSB15] Onur Özyesil, Amit Singer, and Ronen Basri. Stable camera motion estimation using convex programming. SIAM Journal on Imaging Sciences, 8(2):1220– 1262, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Osher:2017:LDM

[OSZ17] Stanley Osher, Zuoqiang Shi, and Wei Zhu. Low dimensional manifold model for image processing. SIAM Journal on Imaging Sciences, 10(4): 1669–1690, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

OConnor:2014:PDD

[OV14] Daniel O’Connor and Lieven Vandenberghe. Primal-dual decomposition by operator splitting and applications to image deblurring. SIAM Journal on Imaging Sciences, 7(3): 1724–1754, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Pierre:2015:LCM

[PAB^{+}15] F. Pierre, J.-F. Aujol, A. Bugeau, N. Papadakis, and V.-T. Ta. Luminance-chrominance model for image colorization. SIAM Journal on Imaging Sciences, 8 (1):536–563, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Pierre:2017:IVC

[PABT17] F. Pierre, J.-F. Aujol, A. Bugeau, and V.-T. Ta. Interactive video colorization within a variational framework. SIAM Journal on Imaging Sciences, 10(4):2293– 2325, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Palamodov:2016:MRA

[Pal16] V. Palamodov. A method of reduction of artifacts of quantitative susceptibility mapping. SIAM Journal on Imaging Sciences, 9(1):481–489, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Polydorides:2012:HOR

[PAM12] Nick Polydorides, Alireza Aghasi, and Eric L. Miller. High-order regularized regression in electrical impedance tomography. SIAM Journal on Imaging Sciences, 5(3):912–943, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Pan:2023:LIP

[PB23] Bolin Pan and Marta M. Betcke. On learning the invisible in photoacoustic tomography with flat directionally sensitive detector. SIAM Journal on Imaging Sciences, 16 (2):770–801, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M148793X.

Prevost:2022:HSR

[PBU^{+}22] Clémence Prévost, Ricardo A.

Borsoi, Konstantin Usevich, David Brie, José C. M. Bermudez, and Cédric Richard. Hyperspectral super-resolution accounting for spectral variability: Coupled tensor LL1-Based recovery and blind unmixing of the unknown superresolution image. SIAM Journal on Imaging Sciences, 15 (1):110–138, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1409354.

Pan:2021:GSO

[PC21] Lili Pan and Xiaojun Chen.

Group sparse optimization for images recovery using capped folded concave functions. SIAM Journal on Imaging Sciences, 14(1):1–25, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Pock:2010:GSV

[PCBC10] Thomas Pock, Daniel Cremers, Horst Bischof, and Antonin Chambolle. Global solutions of variational models with convex regularization. SIAM Journal on Imaging Sciences, 3(4): 1122–1145, ???? 2010. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v3/i4/p1122_s1.

Polisano:2019:CAS

[PCCP19] Kévin Polisano, Laurent Condat, Marianne Clausel, and Valérie Perrier. A convex approach to superresolution and regularization of lines in images. SIAM Journal on Imaging Sciences, 12(1):211–258, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Prendes:2016:BNM

[PCP^{+}16] Jorge Prendes, Marie Chabert, Frédéric Pascal, Alain Giros, and Jean-Yves Tourneret. A Bayesian nonparametric model coupled with a Markov random field for change detection in heterogeneous remote sensing images. SIAM Journal on Imaging Sciences, 9(4):1889–1921, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Pereyra:2017:MPE

[Per17] Marcelo Pereyra. Maximuma-posteriori estimation with Bayesian confidence regions. SIAM Journal on Imaging Sciences, 10(1):285–302, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Pereyra:2019:RMP

[Per19]  Marcelo Pereyra.  Revisit-

ing maximum-a-posteriori estimation in log-concave models. SIAM Journal on Imaging Sciences, 12(1):650–670, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Peyre:2015:EAW

[Pey15] Gabriel Peyré. Entropic approximation of Wasserstein gradient flows. SIAM Journal on Imaging Sciences, 8(4):2323– 2351, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Puspoki:2019:AAS

[PFA^{+}19] Zsuzsanna Püspöki, Julien Fageot, Arash Amini, John Paul Ward, and Michael Unser. Angular accuracy of steerable feature detectors. SIAM Journal on Imaging Sciences, 12(1): 344–371, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Peyre:2010:LMD

[PFS10] Gabriel Peyré, Jalal Fadili, and Jean-Luc Starck. Learning the morphological diversity. SIAM Journal on Imaging Sciences, 3 (3):646–669, ???? 2010. CO-DEN SJISBI. ISSN 1936-4954.

Plaut:2019:GAB

[PG19] Elad Plaut and Raja Giryes.

A greedy approach to ℓ_{0,∞}- based convolutional sparse coding. SIAM Journal on Imaging Sciences, 12(1):186–210, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Peng:2014:RAM

[PH14] Guan-Ju Peng and Wen-Liang Hwang. Reweighted and adaptive morphology separation. SIAM Journal on Imaging Sciences, 7(4):2078–2104, ???? 2014. CODEN SJISBI. ISSN 1936-4954. Pourahmadian:2020:DTM

[PH20] Fatemeh Pourahmadian and Houssem Haddar. Differential tomography of micromechanical evolution in elastic materials of unknown micro/ macrostructure. SIAM Journal on Imaging Sciences, 13(3): 1302–1330, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Park:2021:PMI

[PJS21] Hyoung Suk Park, Junhwa Jung, and Jin Keun Seo. Pseudo-monochromatic imaging in industrial X-ray computed tomography. SIAM Journal on Imaging Sciences, 14(3): 1306–1325, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Park:2018:FLR

[PKCS18] Dohyung Park, Anastasios Kyrillidis, Constantine Caramanis, and Sujay Sanghavi. Finding low-rank solutions via nonconvex matrix factorization, efficiently and provably. SIAM Journal on Imaging Sciences, 11(4):2165–2204, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Pinetz:2021:SPL

[PKPE21] Thomas Pinetz, Erich Kobler, Thomas Pock, and Alexander Effland. Shared prior learning of energy-based models for image reconstruction. SIAM Journal on Imaging Sciences, 14(4): 1706–1748, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Perelli:2020:CCT

[PLCD20] Alessandro Perelli, Michael Lexa, Ali Can, and Mike E. Davies. Compressive computed tomography reconstruction through denoising approximate message passing. SIAM Journal on Imaging Sciences, 13(4):1860–1897, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Parisotto:2020:HOTb

[PLMS20] Simone Parisotto, Jan Lellmann, Simon Masnou, and Carola-Bibiane Schönlieb. Higherorder total directional variation: Imaging applications. SIAM Journal on Imaging Sciences, 13(4):2063–2104, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Papadakis:2008:VAF

[PM08] Nicolas Papadakis and Étienne Mémin. Variational assimilation of fluid motion from image sequence. SIAM Journal on Imaging Sciences, 1(4):343–363, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Parisotto:2020:HOTa

[PMS20] Simone Parisotto, Simon Masnou, and Carola-Bibiane Schönlieb. Higher-order total directional variation: Analysis. SIAM Journal on Imaging Sciences, 13(1):474–496, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Pereyra:2020:APM

[PMZ20] Marcelo Pereyra, Luis Vargas Mieles, and Konstantinos C. Zygalakis. Accelerating proximal Markov Chain Monte Carlo by using an explicit stabilized method. SIAM Journal on Imaging Sciences, 13(2):905– 935, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Pan:2023:SQM

[PN23] Junjun Pan and Michael K. Ng. Separable quaternion matrix factorization for polarization images. SIAM Journal on Imaging Sciences, 16(3): 1281–1307, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M151248X.

Poon:2015:RTV

[Poo15] Clarice Poon. On the role of total variation in compressed sensing. SIAM Journal on Imaging Sciences, 8(1):682–720, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Prato:2009:RVB

[PPE^{+}09] M. Prato, M. Piana, A. G. Emslie, G. J. Hurford, E. P. Kontar, and A. M. Massone. A regularized visibility-based approach to astronomical imaging spectroscopy. SIAM Journal on Imaging Sciences, 2(3):910–930, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Papadakis:2014:OTP

[PPO14] Nicolas Papadakis, Gabriel Peyré, and Edouard Oudet. Optimal transport with proximal splitting. SIAM Journal on Imaging Sciences, 7(1):212–238, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Pagliari:2022:BTS

[PPRV22] Valerio Pagliari, Kostas Papafitsoros, Bogdan Raiba̧ta, and Andreas Vikelis. Bilevel training schemes in imaging for total variation-type functionals with convex integrands. SIAM Journal on Imaging Sciences, 15(4):1690– 1728, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1467328.

Pesquet:2021:LMM

[PRTW21] Jean-Christophe Pesquet, Audrey Repetti, Matthieu Terris, and Yves Wiaux. Learning maximally monotone operators for image recovery. SIAM Journal on Imaging Sciences, 14(3): 1206–1237, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Pfitzenreiter:2011:TRC

[PS11] Tim Pfitzenreiter and Thomas Schuster. Tomographic reconstruction of the curl and divergence of 2D vector fields taking refractions into account. SIAM Journal on Imaging Sciences, 4(1):40–56, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p40_s1.

Pock:2016:IPA

[PS16] Thomas Pock and Shoham Sabach. Inertial Proximal Alternating Linearized Minimization (iPALM) for nonconvex and nonsmooth problems. SIAM Journal on Imaging Sciences, 9 (4):1756–1787, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Pfander:2019:RPR

[PS19] Götz E. Pfander and Palina Salanevich. Robust phase retrieval algorithm for timefrequency structured measurements. SIAM Journal on Imaging Sciences, 12(2):736– 761, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Palacios:2017:RSA

[PUW17] Benjamin Palacios, Gunther Uhlmann, and Yiran Wang. Reducing streaking artifacts in quantitative susceptibility mapping. SIAM Journal on Imaging Sciences, 10(4):1921–1934, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Puy:2014:RIR

[PV14] Gilles Puy and P. Vandergheynst. Robust image reconstruction from multiview measurements. SIAM Journal on Imaging Sciences, 7(1): 128–156, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Puspoki:2016:CSS

[PWSU16] Zsuzsanna Püspöki, John Paul Ward, Daniel Sage, and Michael Unser. On the continuous steering of the scale of tight wavelet frames. SIAM Journal on Imaging Sciences, 9(3):1042–1062, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Punithakumar:2012:CMF

[PYA^{+}12] Kumaradevan Punithakumar, Jing Yuan, Ismail Ben Ayed, Shuo Li, and Yuri Boykov. A convex max-flow approach to distribution-based figureground separation. SIAM Journal on Imaging Sciences, 5(4): 1333–1354, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Papadakis:2013:HDM

[PYAC13] Nicolas Papadakis, Romain Yildizoglu, Jean-François Aujol, and Vicent Caselles. Highdimension multilabel problems: Convex or nonconvex relaxation? SIAM Journal on Imaging Sciences, 6(4):2603–2639, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Pan:2014:CCD

[PYW^{+}14] Xiang Pan, Yongqiang Ye, Jianhong Wang, Xudong Gao, Chun He, Danwei Wang, Bin Jiang, and Lihua Li. Complex composite derivative and its application to edge detection. SIAM Journal on Imaging Sciences, 7(4): 2807–2832, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Qiu:2019:CQC

[QLL19] Di Qiu, Ka-Chun Lam, and Lok-Ming Lui. Computing quasi-conformal folds. SIAM Journal on Imaging Sciences, 12(3):1392–1424, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Qu:2020:ERM

[QLZ20] Qing Qu, Xiao Li, and Zhihui Zhu. Exact recovery of multichannel sparse blind deconvolution via gradient descent. SIAM Journal on Imaging Sciences, 13(3):1630–1652, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Qiu:2015:AMT

[QS15] Lingyun Qiu and Fadil Santosa. Analysis of the magnetoacoustic tomography with magnetic induction. SIAM Journal on Imaging Sciences, 8(3): 2070–2086, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Qian:2011:ENS

[QSUZ11] Jianliang Qian, Plamen Stefanov, Gunther Uhlmann, and Hongkai Zhao. An efficient Neumann series-based algorithm for thermoacoustic and photoacoustic tomography with variable sound speed. SIAM Journal on Imaging Sciences, 4(3): 850–883, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i3/p850_s1.

Qi:2010:HOP

[QYW10] Liqun Qi, Gaohang Yu, and Ed X. Wu. Higher order positive semidefinite diffusion tensor imaging. SIAM Journal on Imaging Sciences, 3(3):416–433, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Qu:2019:SRI

[QYZ19] Fenglong Qu, Jiaqing Yang, and Haiwen Zhang. Shape reconstruction in inverse scattering by an inhomogeneous cavity with internal measurements. SIAM Journal on Imaging Sciences, 12(2):788– 808, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Ravishankar:2015:EBC

[RB15] Saiprasad Ravishankar and Yoram Bresler. Efficient blind compressed sensing using sparsifying transforms with convergence guarantees and application to magnetic resonance imaging. SIAM Journal on Imaging Sciences, 8(4):2519– 2557, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Roy:2018:NOA

[RB18] Souvik Roy and Alfio Borzì.

A new optimization approach to sparse reconstruction of logconductivity in acousto-electric tomography. SIAM Journal on Imaging Sciences, 11(2):1759– 1784, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Rotker:2020:OMV

[RBB20] Keren Rotker, Dafna Ben Bashat, and Alex M. Bronstein. Overparameterized models for vector fields. SIAM Journal on Imaging Sciences, 13(3):1386– 1414, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Rapin:2014:NSR

[RBLS14] Jérémy Rapin, Jérôme Bobin, Anthony Larue, and Jean-Luc Starck. NMF with sparse regularizations in transformed domains. SIAM Journal on Imaging Sciences, 7(4):2020–2047, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Rabin:2009:SAM

[RDG09] J. Rabin, J. Delon, and Y. Gousseau. A statistical approach to the matching of local features. SIAM Journal on Imaging Sciences, 2(3):931–958, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Rodriguez:2018:CST

[RDM18] Mariano Rodríguez, Julie Delon, and Jean-Michel Morel. Covering the space of tilts. application to affine invariant image comparison. SIAM Journal on Imaging Sciences, 11(2): 1230–1267, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Rosman:2009:EBI

[RDSK09] Guy Rosman, Lorina Dascal, Avram Sidi, and Ron Kimmel. Efficient Beltrami image filtering via vector extrapolation methods. SIAM Journal on Imaging Sciences, 2(3):858–878, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Romano:2015:BID

[RE15] Yaniv Romano and Michael Elad. Boosting of image denoising algorithms. SIAM Journal on Imaging Sciences, 8(2):

1187–1219, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Romano:2017:LEC

[REM17] Yaniv Romano, Michael Elad, and Peyman Milanfar. The little engine that could: Regularization by denoising (RED). SIAM Journal on Imaging Sciences, 10(4):1804–1844, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Raguet:2013:GFB

[RFP13] Hugo Raguet, Jalal Fadili, and Gabriel Peyré. A generalized forward-backward splitting. SIAM Journal on Imaging Sciences, 6(3):1199–1226, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Roussillon:2016:KMN

[RG16] Pierre Roussillon and Joan Alexis Glaunès. Kernel metrics on normal cycles and application to curve matching. SIAM Journal on Imaging Sciences, 9(4): 1991–2038, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Raviv:2014:SSA

[RGLB14] Tammy Riklin Raviv, Yi Gao, James J. Levitt, and Sylvain Bouix. Statistical shape analysis of neuroanatomical structures via level-set-based shape morphing. SIAM Journal on Imaging Sciences, 7(3):1645– 1668, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ren:2013:HRM

[RGZ13] Kui Ren, Hao Gao, and Hongkai Zhao. A hybrid reconstruction method for quantitative PAT. SIAM Journal on Imaging Sciences, 6(1):32–55, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Rigaud:2017:CST

[Rig17] Gaël Rigaud. Compton scattering tomography: Feature reconstruction and rotation-free modality. SIAM Journal on Imaging Sciences, 10(4):2217– 2249, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Reisenhofer:2019:ERB

[RK19] Rafael Reisenhofer and Emily J. King. Edge, ridge, and blob detection with symmetric molecules. SIAM Journal on Imaging Sciences, 12(4):1585– 1626, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Rudzusika:2022:DLB

[RKO22] Jevgenija Rudzusika, Thomas Koehler, and Ozan Oktem. Deep learning-based dictionary learning and tomographic image reconstruction. SIAM Journal on Imaging Sciences, 15(4): 1729–1764, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1445697.

Rajwade:2013:CHI

[RKT^{+}13] Ajit Rajwade, David Kittle, Tsung-Han Tsai, David Brady, and Lawrence Carin. Coded hyperspectral imaging and blind compressive sensing. SIAM Journal on Imaging Sciences, 6 (2):782–812, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Raguet:2015:PGF

[RL15] Hugo Raguet and Loïc Landrieu. Preconditioning of a generalized forward-backward splitting and application to optimization on graphs. SIAM Journal on Imaging Sciences, 8 (4):2706–2739, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Rigaud:2014:SER

[RLL14] G. Rigaud, A. Lakhal, and A. K. Louis. Series expansions of the reconstruction kernel of the Radon transform over a Cormack-type family of curves with applications in tomography. SIAM Journal on Imaging Sciences, 7(2):924–943, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Ringholm:2018:VIR

[RLS18] Torbjørn Ringholm, Jasmina Lazić, and Carola-Bibiane Schönlieb. Variational image regularization with Euler’s elastica using a discrete gradient scheme. SIAM Journal on Imaging Sciences, 11(4):2665– 2691, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Robini:2010:OSC

[RM10] Marc C. Robini and Isabelle E.

Magnin. Optimization by stochastic continuation. SIAM Journal on Imaging Sciences, 3(4):1096–1121, ???? 2010. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.

                                        siam.org/siims/resource/1/
                                        sjisbi/v3/i4/p1096_s1.

Rabanser:2019:ABC

[RNH19] Simon Rabanser, Lukas Neumann, and Markus Haltmeier. Analysis of the block coordinate descent method for linear illposed problems. SIAM Journal on Imaging Sciences, 12(4): 1808–1832, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Rey-Otero:2015:RUC

[ROD15] Ives Rey-Otero and Mauricio Delbracio. Is repeatability an unbiased criterion for ranking feature detectors? SIAM Journal on Imaging Sciences, 8(4): 2558–2580, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Romberg:2009:CSR

[Rom09] Justin Romberg. Compressive sensing by random convolution. SIAM Journal on Imaging Sciences, 2(4):1098–1128, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Repetti:2019:SBU

[RPW19] Audrey Repetti, Marcelo Pereyra, and Yves Wiaux. Scalable Bayesian uncertainty quantification in imaging inverse problems via convex optimization. SIAM Journal on Imaging Sciences, 12(1):87–118, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Raviv:2015:SIM

[RR15] Dan Raviv and Ramesh Raskar. Scale invariant metrics of volumetric datasets. SIAM Journal on Imaging Sciences, 8(1): 403–425, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Rosen:2020:CLI

[RS20] Eitan Rosen and Yoel Shkolnisky. Common lines ab initio reconstruction of D_{2}-symmetric molecules in cryo-electron microscopy. SIAM Journal on Imaging Sciences, 13(4):1898– 1944, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Richter:2021:GIA

[RTH21] Robin Richter, Duy H. Thai, and Stephan Huckemann. Generalized intersection algorithms with fixed points for image decomposition learning. SIAM Journal on Imaging Sciences, 14(3):1273–1305, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Reshniak:2020:NFD

[RTW20] Viktor Reshniak, Jeremy Trageser, and Clayton G. Webster. A nonlocal feature-driven exemplarbased approach for image inpainting. SIAM Journal on Imaging Sciences, 13(4):2140– 2168, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Rodriguez:2019:ICA

[RVCB19] R. Gil Rodríguez, J. Vazquez-Corral, and M. Bertalmío. Issues with common assumptions about the camera pipeline and their impact in HDR imaging from multiple exposures. SIAM Journal on Imaging Sciences, 12(4):1627–1642, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Rumpf:2009:NES

[RW09] Martin Rumpf and Benedikt Wirth. A nonlinear elastic shape averaging approach. SIAM Journal on Imaging Sciences, 2(3):800–833, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Rumpf:2013:DGC

[RW13] Martin Rumpf and Benedikt Wirth. Discrete geodesic calculus in shape space and applications in the space of viscous fluidic objects. SIAM Journal on Imaging Sciences, 6(4):2581– 2602, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Robini:2018:IHQ

[RYZ18] Marc C. Robini, Feng Yang, and Yuemin Zhu. Inexact halfquadratic optimization for linear inverse problems. SIAM Journal on Imaging Sciences, 11(2):1078–1133, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Ren:2013:QFP

[RZ13] Kui Ren and Hongkai Zhao. Quantitative fluorescence photoacoustic tomography. SIAM Journal on Imaging Sciences, 6 (4):2404–2429, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Robini:2015:GHQ

[RZ15] Marc C. Robini and Yuemin Zhu. Generic half-quadratic optimization for image reconstruction. SIAM Journal on Imaging Sciences, 8(3):1752–1797, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Sapiro:2008:MEC

[Sap08] Guillermo Sapiro. Message from the Editor-in-Chief. SIAM Journal on Imaging Sciences, 1 (1):1, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Sapiro:2010:SSO

[Sap10] Guillermo Sapiro. Special section on optimization in imaging sciences. SIAM Journal on Imaging Sciences, 3 (4):1047, ???? 2010. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v3/i4/p1047_s1.

Song:2017:FMR

[SAS17] Yizhuang Song, Habib Ammari, and Jin Keun Seo. Fast magnetic resonance electrical impedance tomography with highly undersampled data. SIAM Journal on Imaging Sciences, 10(2):558–577, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Sukurdeep:2022:NVM

[SBC22] Yashil Sukurdeep, Martin Bauer, and Nicolas Charon. A new variational model for shape graph registration with partial matching constraints. SIAM Journal on Imaging Sciences, 15(1):261–292, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1418587.

Soubies:2015:CEP

[SBFA15] Emmanuel Soubies, Laure Blanc-Féraud, and Gilles Aubert. A continuous exact ℓ_{0} penalty (CEL0) for least squares regularized problem. SIAM Journal on Imaging Sciences, 8(3): 1607–1639, ???? 2015. CODEN SJISBI. ISSN 1936-4954. See erratum [SBFA16].

Soubies:2016:ECE

[SBFA16] Emmanuel Soubies, Laure Blanc-Féraud, and Gilles Aubert. Erratum: A Continuous Exact ℓ_{0} Penalty (CEL0) for Least Squares Regularized Problem. SIAM Journal on Imaging Sciences, 9(1):490–494, ???? 2016. CODEN SJISBI. ISSN 1936-4954. See [SBFA15].

Sitenko:2023:NGP

[SBS23] Dmitrij Sitenko, Bastian Boll, and Christoph Schnörr. A nonlocal graph-PDE and higherorder geometric integration for image labeling. SIAM Journal on Imaging Sciences, 16 (1):501–567, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1496141.

Shyu:2010:VMS

[SC10] Shyong Jian Shyu and Kun Chen. Visual multiple-secret sharing by circle random grids. SIAM Journal on Imaging Sciences, 3(4):926–953, ???? 2010.

CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v3/i4/p926_s1.

Strekalovskiy:2014:CRV

[SCC14] Evgeny Strekalovskiy, Antonin Chambolle, and Daniel Cremers. Convex relaxation of vectorial problems with coupled regularization. SIAM Journal on Imaging Sciences, 7(1):294– 336, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Santana-Cedres:2015:IET

[SCGAF^{+}15] Daniel Santana-Cedrés, Luis Gomez, Miguel Alemán-Flores, Agustín Salgado, Julio Esclarín, Luis Mazorra, and Luis Alvarez. Invertibility and estimation of two-parameter polynomial and division lens distortion models. SIAM Journal on Imaging Sciences, 8(3):1574–1606, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Siu:2020:ISP

[SCL20] Chun Yin Siu, Hei Long Chan, and Ronald Lok Ming Lui. Image segmentation with partial convexity shape prior using discrete conformality structures. SIAM Journal on Imaging Sciences, 13(4):2105–2139, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Sadek:2012:AID

[SCM^{+}12] R. Sadek, C. Constantinopoulos, E. Meinhardt, C. Ballester, and V. Caselles. On affine invariant descriptors related to SIFT. SIAM Journal on Imaging Sciences, 5(2):652–687, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Sutour:2015:ENL

[SDA15] Camille Sutour, Charles-Alban Deledalle, and Jean-François Aujol. Estimation of the noise level function based on a nonparametric detection of homogeneous image regions. SIAM Journal on Imaging Sciences, 8 (4):2622–2661, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Sdika:2013:SSC

[Sdi13] Michaël Sdika. A sharp sufficient condition for B-spline vector field invertibility. application to diffeomorphic registration and interslice interpolation. SIAM Journal on Imaging Sciences, 6(4):2236–2257, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Schonsheck:2022:PTC

[SDL22] Stefan C. Schonsheck, Bin Dong, and Rongjie Lai. Parallel transport convolution: Deformable convolutional networks on manifold-structured data. SIAM Journal on Imaging Sciences, 15(1):367– 386, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1407616.

Sakaridis:2017:TAA

[SDM17] Christos Sakaridis, Kimon Drakopoulos, and Petros Maragos. Theoretical analysis of active contours on graphs. SIAM Journal on Imaging Sciences, 10(3):1475–1510, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Srinivasa:2020:TBB

[SDR20] Rakshith Sharma Srinivasa, Mark A. Davenport, and Justin Romberg. Trading beams for bandwidth: Imaging with randomized beamforming. SIAM Journal on Imaging Sciences, 13(1):317–350, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Sciacchitano:2015:VAR

[SDZ15] Federica Sciacchitano, Yiqiu Dong, and Tieyong Zeng. Variational approach for restoring blurred images with Cauchy noise. SIAM Journal on Imaging Sciences, 8(3):1894–1922, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Segre:2022:IWP

[Seg22] Enrico Segre. Image warp preserving content intensity. SIAM Journal on Imaging Sciences, 15(4):1623–1645, ???? 2022. CODEN SJISBI. ISSN 1936-

4954. URL https://epubs.
siam.org/doi/10.1137/21M1452688.

Solomon:2019:SSB

[SEMS19] Oren Solomon, Yonina C.

Eldar, Maor Mutzafi, and Mordechai Segev. SPARCOM: Sparsity based super-resolution correlation microscopy. SIAM Journal on Imaging Sciences, 12(1):392–419, ???? 2019. CO-DEN SJISBI. ISSN 1936-4954.

Sur:2015:MND

[SG15] Frédéric Sur and Michel Grédiac. Measuring the noise of digital imaging sensors by stacking raw images affected by vibrations and illumination flickering. SIAM Journal on Imaging Sciences, 8(1):611–643, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Santosa:2022:BCD

[SG22] Fadil Santosa and Madeline Goh. Bar code decoding in a camera-based scanner: Analysis and algorithm. SIAM Journal on Imaging Sciences, 15(3): 1017–1040, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1449658.

Schmitz:2018:WDL

[SHB^{+}18] Morgan A. Schmitz, Matthieu Heitz, Nicolas Bonneel, Fred Ngolè, David Coeurjolly, Marco Cuturi, Gabriel Peyré, and Jean-Luc Starck. Wasserstein dictionary learning: Optimal transport-based unsupervised nonlinear dictionary learning. SIAM Journal on Imaging Sciences, 11(1):643–678, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Song:2019:ODI

[ShDC^{+}19] Yingying Song, El hadi Djermoune, Jie Chen, Cédric Richard, and David Brie. Online deconvolution for industrial hyperspectral imaging systems.

SIAM Journal on Imaging Sciences, 12(1):54–86, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Steinke:2010:NRB

[SHS10] Florian Steinke, Matthias Hein, and Bernhard Schölkopf. Nonparametric regression between general Riemannian manifolds. SIAM Journal on Imaging Sciences, 3(3):527–563, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Schwab:2019:GOS

[SHVC19] Evan Schwab, Benjamin D. Haeffele, René Vidal, and Nicolas Charon. Global optimality in separable dictionary learning with applications to the analysis of diffusion MRI. SIAM Journal on Imaging Sciences, 12(4): 1967–2008, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Sakajo:2023:TIV

[SI23] Takashi Sakajo and Keiichi Itatani. Topological identification of vortical flow structures in the left ventricle of the heart. SIAM Journal on Imaging Sciences, 16(3): 1491–1519, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1536923.

Sudhakar:2015:CIC

[SJD^{+}15] Prasad Sudhakar, Laurent Jacques, Xavier Dubois, Philippe Antoine, and Luc Joannes. Compressive imaging and characterization of sparse light deflection maps. SIAM Journal on Imaging Sciences, 8(3):1824– 1856, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Strossner:2023:LRT

[SK23] Christoph Strössner and Daniel Kressner. Low-rank tensor approximations for solving multimarginal optimal transport problems. SIAM Journal on Imaging Sciences, 16(1):169– 191, ???? 2023. CODEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1478355.

Seo:2019:LBM

[SKJ^{+}19] Jin Keun Seo, Kang Cheol Kim, Ariungerel Jargal, Kyounghun Lee, and Bastian Harrach. A learning-based method for solving ill-posed nonlinear inverse problems: a simulation study of lung EIT. SIAM Journal on Imaging Sciences, 12(3):1275– 1295, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Sciacchitano:2019:SBI

[SLS19] Federica Sciacchitano, Silvio Lugaro, and Alberto Sorrentino. Sparse Bayesian imaging of solar flares. SIAM Journal on Imaging Sciences, 12(1):319– 343, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Stein:2022:SSF

[SLS22] Oded Stein, Jiajin Li, and Justin Solomon. A splitting scheme for flip-free distortion energies. SIAM Journal on Imaging Sciences, 15 (2):925–959, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1433058.

Simon:2016:IUE

[SM16] Loïc Simon and Jean-Michel Morel. Influence of unknown exterior samples on interpolated values for band-limited images. SIAM Journal on Imaging Sciences, 9(1):152–184, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Sanchez:2018:MSS

[SM18] Javier Sánchez and Jean-Michel Morel. Motion smoothing strategies for 2D video stabilization. SIAM Journal on Imaging Sciences, 11(1):219–251, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Sabater:2011:HAC

[SMA11] N. Sabater, J.-M. Morel, and A. Almansa. How accurate can block matches be in stereo vision? SIAM Journal on Imaging Sciences, 4(1): 472–500, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p472_s1.

Sundaramoorthi:2011:NGM

[SMSY11] Ganesh Sundaramoorthi, Andrea Mennucci, Stefano Soatto, and Anthony Yezzi. A new geometric metric in the space of curves, and applications to tracking deforming objects by prediction and filtering. SIAM Journal on Imaging Sciences, 4 (1):109–145, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p109_s1.

Seeger:2011:LSB

[SN11] Matthias W. Seeger and Hannes Nickisch. Large scale Bayesian inference and experimental design for sparse linear models. SIAM Journal on Imaging Sciences, 4(1):166–199, ???? 2011. CODEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p166_s1.

Sur:2013:CMM

[SNB13] Frédéric Sur, Nicolas Noury, and Marie-Odile Berger. An A Contrario model for matching interest points under geometric and photometric constraints. SIAM Journal on Imaging Sciences, 6(4):1956– 1978, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Sommer:2013:HOM

[SNDP13] Stefan Sommer, Mads Nielsen, Sune Darkner, and Xavier Pennec. Higher-order momentum distributions and locally affine LDDMM registration. SIAM Journal on Imaging Sciences, 6 (1):341–367, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Sgouralis:2017:BTF

[SNM17] Ioannis Sgouralis, Andreas Nebenführ, and Vasileios Maroulas.

A Bayesian topological framework for the identification and reconstruction of subcellular motion. SIAM Journal on Imaging Sciences, 10(2):871– 899, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Shi:2008:NIS

[SO08] Jianing Shi and Stanley Osher.

A nonlinear inverse scale space method for a convex multiplicative noise model. SIAM Journal on Imaging Sciences, 1(3): 294–321, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Schaeffer:2013:LPR

[SO13] Hayden Schaeffer and Stanley Osher. A low patch-rank interpretation of texture. SIAM Journal on Imaging Sciences, 6 (1):226–262, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Sim:2020:OTD

[SOK^{+}20] Byeongsu Sim, Gyutaek Oh, Jeongsol Kim, Chanyong Jung, and Jong Chul Ye. Optimal transport driven CycleGAN for unsupervised learning in inverse problems. SIAM Journal on Imaging Sciences, 13(4):2281– 2306, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Stefan:2010:ITV

[SRG10] W. Stefan, R. A. Renaut, and A. Gelb. Improved total variation-type regularization using higher order edge detectors. SIAM Journal on Imaging Sciences, 3(2):232–251, ????

2010. CODEN SJISBI. ISSN 1936-4954.

Singer:2011:TDS

[SS11] A. Singer and Y. Shkolnisky. Three-dimensional structure determination from common lines in cryo-EM by eigenvectors and semidefinite programming. SIAM Journal on Imaging Sciences, 4(2):543–572, ???? 2011. CODEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i2/p543_s1.

Shkolnisky:2012:VDE

[SS12] Yoel Shkolnisky and Amit Singer. Viewing direction estimation in cryo-EM using synchronization. SIAM Journal on Imaging Sciences, 5(3):1088– 1110, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Swoboda:2013:CVI

[SS13] Paul Swoboda and Christoph Schnörr. Convex variational image restoration with histogram priors. SIAM Journal on Imaging Sciences, 6(3):1719–1735, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Song:2023:CAH

[SSL23] Yizhuang Song, Rosalind Sadleir, and Jijun Liu. Convergence analysis of the harmonic B_{z} algorithm with single injection current in MREIT. SIAM Journal on Imaging Sciences, 16 (2):706–739, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1505438.

Singer:2009:DIN

[SSN09] Amit Singer, Yoel Shkolnisky, and Boaz Nadler. Diffusion interpretation of nonlocal neighborhood filters for signal denoising. SIAM Journal on Imaging Sciences, 2(1):118–139, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Snieder:2009:FFI

[SSSW09] Roel Snieder, Francisco J.

Sánchez-Sesma, and Kees Wapenaar. Field fluctuations, imaging with backscattered waves, a generalized energy theorem, and the optical theorem. SIAM Journal on Imaging Sciences, 2 (2):763–776, ???? 2009. CO-DEN SJISBI. ISSN 1936-4954.

Smith:2011:DCP

[ST11] E. M. Smith and S. V. Tsynkov.

Dual carrier probing for spaceborne SAR imaging. SIAM Journal on Imaging Sciences, 4 (2):501–542, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i2/p501_s1.

Schwartz:2019:IIM

[ST19] Ariel Schwartz and Ronen Talmon. Intrinsic isometric manifold learning with application to localization. SIAM Journal on Imaging Sciences, 12(3): 1347–1391, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Sahlstrom:2023:UVA

[ST23] Teemu Sahlström and Tanja Tarvainen. Utilizing variational autoencoders in the Bayesian inverse problem of photoacoustic tomography. SIAM Journal on Imaging Sciences, 16 (1):89–110, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1489897.

Shi:2022:CLP

[STA22] Hui Shi, Yann Traonmilin, and Jean-François Aujol. Compressive learning for patch-based image denoising. SIAM Journal on Imaging Sciences, 15(3): 1184–1212, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1459812.

Sankaranarayanan:2013:CAL

[STCB13] Aswin C. Sankaranarayanan, Pavan K. Turaga, Rama Chellappa, and Richard G. Baraniuk. Compressive acquisition of linear dynamical systems. SIAM Journal on Imaging Sciences, 6(4):2109–2133, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Storath:2011:DMA

[Sto11] Martin Storath. Directional multiscale amplitude and phase decomposition by the monogenic curvelet transform. SIAM Journal on Imaging Sciences, 4(1):57–78, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.

siam.org/siims/resource/1/
sjisbi/v4/i1/p57_s1.

Sepasian:2012:MGR

[StTBRV12] N. Sepasian, J. H. M. ten Thije Boonkkamp, B. M. Ter Haar Romeny, and A. Vilanova. Multivalued geodesic ray-tracing for computing brain connections using diffusion tensor imaging. SIAM Journal on Imaging Sciences, 5(2):483–504, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Silva:2009:ASM

[STV09] Moacyr Alvim Silva, Ralph Teixeira, and Luiz Velho. Affine skeletons and Monge–Ampère equations. SIAM Journal on Imaging Sciences, 2(3):987– 1001, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Shen:2011:APG

[STY11] Zuowei Shen, Kim-Chuan Toh, and Sangwoon Yun. An accelerated proximal gradient algorithm for frame-based image restoration via the balanced approach. SIAM Journal on Imaging Sciences, 4(2): 573–596, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i2/p573_s1.

Song:2020:DLT

[SUFU20] Anna Song, Virginie Uhlmann, Julien Fageot, and Michael Unser. Dictionary learning for two-dimensional Kendall shapes. SIAM Journal on Imaging Sciences, 13(1):141–175, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Stocker:2008:GEL

[SV08] Christina Stöcker and Axel Voigt. Geodesic evolution laws — a level-set approach. SIAM Journal on Imaging Sciences, 1 (4):379–399, ???? 2008. CO-DEN SJISBI. ISSN 1936-4954.

Singer:2013:TDT

[SW13] A. Singer and H.-T. Wu. Two-dimensional tomography from noisy projections taken at unknown random directions. SIAM Journal on Imaging Sciences, 6(1):136–175, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Storath:2014:FPV

[SW14] Martin Storath and Andreas Weinmann. Fast partitioning of vector-valued images. SIAM Journal on Imaging Sciences, 7 (3):1826–1852, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Sun:2015:DCD

[SWGL15] Jian Sun, Tianqi Wu, Xianfeng Gu, and Feng Luo. Discrete conformal deformation: Algorithm and experiments. SIAM Journal on Imaging Sciences, 8(3): 1421–1456, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Sun:2012:NSB

[SX12] Yuanchang Sun and Jack Xin. Nonnegative sparse blind source separation for NMR spectroscopy by data clustering, model reduction, and ℓ_{1} minimization. SIAM Journal on Imaging Sciences, 5(3):886–911, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Sankaranarayanan:2015:VCS

[SXS^{+}15] Aswin C. Sankaranarayanan, Lina Xu, Christoph Studer, Yun Li, Kevin F. Kelly, and Richard G. Baraniuk. Video compressive sensing for spatial multiplexing cameras using motion-flow models. SIAM Journal on Imaging Sciences, 8 (3):1489–1518, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Staneva:2014:MES

[SY14] Valentina Staneva and Laurent Younes. Modeling and estimation of shape deformation for topology-preserving object tracking. SIAM Journal on Imaging Sciences, 7(1):427–455, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Sundaramoorthi:2022:AOP

[SYB22] Ganesh Sundaramoorthi, Anthony Yezzi, and Minas Benyamin. Accelerated optimization in the PDE framework: Formulations for the manifold of diffeomorphisms. SIAM Journal on Imaging Sciences, 15 (1):324–366, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/20M1381927.

Schaeffer:2015:STR

[SYO15] Hayden Schaeffer, Yi Yang, and Stanley Osher. Space-time regularization for video decompression. SIAM Journal on Imaging Sciences, 8(1):373–402, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Shi:2018:LRD

[SZGW18] Kehan Shi, Dazhi Zhang, Zhichang Guo, and Boying Wu. A linear reaction–diffusion system with interior degeneration for color image compression. SIAM Journal on Imaging Sciences, 11(1):442–472, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Singer:2011:VAC

[SZSH11] A. Singer, Z. Zhao, Y. Shkolnisky, and R. Hadani. Viewing angle classification of cryoelectron microscopy images using eigenvectors. SIAM Journal on Imaging Sciences, 4(2): 723–759, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i2/p723_s1.

Sun:2014:EPS

[SZW14] Ju Sun, Yuqian Zhang, and John Wright. Efficient point-to-subspace query in ℓ^{1} with application to robust object instance recognition. SIAM Journal on Imaging Sciences, 7(4): 2105–2138, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Traonmilin:2014:SHD

[TA14] Yann Traonmilin and Cecilia Aguerrebere. Simultaneous high dynamic range and superresolution imaging without regularization. SIAM Journal on Imaging Sciences, 7(3):1624–1644, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Tereshin:2020:ACM

[TAF^{+}20] Alexander Tereshin, Valery Adzhiev, Oleg Fryazinov, Felix Marrington-Reeve, and Alexander Pasko. Automatically controlled morphing of 2D shapes with textures. SIAM Journal on Imaging Sciences, 13(1):78–107, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Tachella:2019:BRC

[TAR^{+}19] Julián Tachella, Yoann Altmann, Ximing Ren, Aongus McCarthy, Gerald S. Buller, Stephen McLaughlin, and Jean-Yves Tourneret. Bayesian 3D reconstruction of complex scenes from single-photon lidar data. SIAM Journal on Imaging Sciences, 12(1):521–550, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Thanh:2015:IBO

[TBKF15] Nguyen Trung Thành, Larisa Beilina, Michael V. Klibanov, and Michael A. Fiddy. Imaging of buried objects from experimental backscattering timedependent measurements using a globally convergent inverse algorithm. SIAM Journal on Imaging Sciences, 8(1):757–786, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Thiruvenkadam:2008:SUO

[TCH08] Sheshadri R. Thiruvenkadam, Tony F. Chan, and Byung-Woo Hong. Segmentation under occlusions using selective shape prior. SIAM Journal on Imaging Sciences, 1(1):115–142, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Tron:2017:SEM

[TD17] Roberto Tron and Kostas Daniilidis. The space of essential matrices as a Riemannian quotient manifold. SIAM Journal on Imaging Sciences, 10(3): 1416–1445, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Tirer:2021:CRP

[TG21] Tom Tirer and Raja Giryes. On the convergence rate of projected gradient descent for a back-projection based objective. SIAM Journal on Imaging Sciences, 14(4):1504–1531, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Tai:2011:FAE

[THC11] Xue-Cheng Tai, Jooyoung Hahn, and Ginmo Jason Chung. A fast algorithm for Euler’s elastica model using augmented Lagrangian method. SIAM Journal on Imaging Sciences, 4 (1):313–344, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.

siam.org/siims/resource/1/
sjisbi/v4/i1/p313_s1.

Tiirola:2014:IDU

[Tii14] Juha Tiirola. Image decompositions using spaces of variable smoothness and integrability. SIAM Journal on Imaging Sciences, 7(3):1558–1587, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Taylor:2012:RWI

[TM12] Kye M. Taylor and François G.

Meyer. A random walk on image patches. SIAM Journal on Imaging Sciences, 5(2):688–725, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Talebi:2016:APG

[TM16a] Hossein Talebi and Peyman Milanfar. Asymptotic performance of global denoising. SIAM Journal on Imaging Sciences, 9(2): 665–683, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Tendero:2016:TOF

[TM16b] Yohann Tendero and Jean-Michel Morel. A theory of optimal flutter shutter for probabilistic velocity models. SIAM Journal on Imaging Sciences, 9 (1):445–480, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Tarzanagh:2018:FRA

[TM18] Davoud Ataee Tarzanagh and George Michailidis. Fast randomized algorithms for tproduct based tensor operations and decompositions with applications to imaging data. SIAM Journal on Imaging Sciences, 11(4):2629–2664, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Tsogka:2013:SIE

[TMP13] Chrysoula Tsogka, Dimitrios A. Mitsoudis, and Symeon Papadimitropoulos. Selective imaging of extended reflectors in two-dimensional waveguides. SIAM Journal on Imaging Sciences, 6(4):2714–2739, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Tsogka:2018:IER

[TMP18] Chrysoula Tsogka, Dimitrios A. Mitsoudis, and Symeon Papadimitropoulos. Imaging extended reflectors in a terminating waveguide. SIAM Journal on Imaging Sciences, 11(2): 1680–1716, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Tendero:2013:FSP

[TMR13] Yohann Tendero, Jean-Michel Morel, and Bernard Rougé. The flutter shutter paradox. SIAM Journal on Imaging Sciences, 6 (2):813–847, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Tran:2020:SDL

[TMSP20] K. H. Tran, F. M. Ngolè Mboula, J. L. Starck, and V. Prost. Semisupervised dictionary learning with graph regularized and active points. SIAM Journal on Imaging Sciences, 13(2):724–745, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Tralie:2018:QPQ

[TP18] Christopher J. Tralie and Jose A. Perea. (quasi)periodicity quantification in video data, using topology. SIAM Journal on Imaging Sciences, 11(2):1049– 1077, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Tartavel:2016:WLI

[TPG16] Guillaume Tartavel, Gabriel Peyré, and Yann Gousseau. Wasserstein loss for image synthesis and restoration. SIAM Journal on Imaging Sciences, 9 (4):1726–1755, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Tschumperle:2020:RSC

[TPM20] David Tschumperlé, Christine Porquet, and Amal Mahboubi. Reconstruction of smooth 3D color functions from keypoints: Application to lossy compression and exemplar-based generation of color LUTs. SIAM Journal on Imaging Sciences, 13(3):1511–1535, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Tepper:2014:BFC

[TS14] Mariano Tepper and Guillermo Sapiro. A biclustering framework for consensus problems. SIAM Journal on Imaging Sciences, 7(4):2488–2525, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Teuber:2011:DDC

[TSG^{+}11] T. Teuber, G. Steidl, P. Gwosdek, C. Schmaltz, and J. Weickert. Dithering by differences of convex functions. SIAM Journal on Imaging Sciences, 4(1):79–108, ???? 2011. CO-DEN SJISBI. ISSN 1936-

4954. URL http://epubs.
siam.org/siims/resource/1/
sjisbi/v4/i1/p79_s1.

Tsynkov:2009:SIT

[Tsy09a] S. V. Tsynkov. On SAR imaging through the Earth’s ionosphere. SIAM Journal on Imaging Sciences, 2(1):140–182, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Tsynkov:2009:USS

[Tsy09b]  S. V. Tsynkov.  On the

use of start-stop approximation for spaceborne SAR imaging. SIAM Journal on Imaging Sciences, 2(2):646–669, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Tang:2022:SNQ

[TT22] Tommy M. Tang and Hemant D. Tagare. Steerable near-quadrature filter pairs in three dimensions. SIAM Journal on Imaging Sciences, 15 (2):670–700, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M143529X.

Tsakiris:2017:FAS

[TV17] Manolis C. Tsakiris and René Vidal. Filtrated algebraic subspace clustering. SIAM Journal on Imaging Sciences, 10(1): 372–415, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Teboulle:2020:NPG

[TV20] Marc Teboulle and Yakov Vaisbourd. Novel proximal gradient methods for nonnegative matrix factorization with sparsity constraints. SIAM Journal on Imaging Sciences, 13(1): 381–421, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Tai:2013:WFB

[TZS13] Cheng Tai, Xiaoqun Zhang, and Zuowei Shen. Wavelet frame based multiphase image segmentation. SIAM Journal on Imaging Sciences, 6(4):2521– 2546, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Unser:2013:UPF

[UC13] Michael Unser and Nicolas Chenouard. A unifying parametric framework for 2D steerable wavelet transforms. SIAM Journal on Imaging Sciences, 6 (1):102–135, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Valkonen:2013:TGV

[VBK13] Tuomo Valkonen, Kristian Bredies, and Florian Knoll. Total generalized variation in diffusion tensor imaging. SIAM Journal on Imaging Sciences, 6 (1):487–525, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Vidal:2020:MLE

[VDPD20] Ana Fernandez Vidal, Valentin De Bortoli, Marcelo Pereyra, and Alain Durmus. Maximum likelihood estimation of regularization parameters in high-dimensional inverse problems: an empirical Bayesian approach. Part I: Methodology and experiments. SIAM Journal on Imaging Sciences, 13(4): 1945–1989, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Vural:2013:ADB

[VF13] Elif Vural and Pascal Frossard. Analysis of descent-based image registration. SIAM Journal on Imaging Sciences, 6(4): 2310–2349, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Vural:2014:AIR

[VF14] Elif Vural and Pascal Frossard. Analysis of image registration with tangent distance. SIAM Journal on Imaging Sciences, 7 (4):2860–2915, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Velasco-Forero:2022:LEM

[VFPA22] Santiago Velasco-Forero, R. Pagès, and Jesus Angulo. Learnable empirical mode decomposition based on mathematical morphology. SIAM Journal on Imaging Sciences, 15 (1):23–44, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1417867.

vonGioi:2022:WPM

[vGPR22] Rafael Grompone von Gioi, Ignacio Ramírez Paulino, and Gregory Randall. The whole and the parts: The minimum description length principle and the a-contrario framework. SIAM Journal on Imaging Sciences, 15(3):1282– 1313, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M145745X.

Vahid:2020:FIM

[VHO20] Milad R. Vahid, Bernard Hanzon, and Raimund J. Ober. Fisher information matrix for single molecules with stochastic trajectories. SIAM Journal on Imaging Sciences, 13(1): 234–264, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Voccola:2015:SAR

[Voc15] Kaitlyn Voccola. Synthetic aperture radar correlation imaging. SIAM Journal on Imaging Sciences, 8(1):299–330, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Vonesch:2015:SPR

[VSU15] Cédric Vonesch, Frédéric Stauber, and Michael Unser. Steerable PCA for rotation-invariant image recognition. SIAM Journal on Imaging Sciences, 8(3): 1857–1873, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Vaksman:2016:POR

[VZE16] Gregory Vaksman, Michael Zibulevsky, and Michael Elad. Patch ordering as a regularization for inverse problems in image processing. SIAM Journal on Imaging Sciences, 9(1): 287–319, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Wahab:2015:SRA

[Wah15] Abdul Wahab. Stability and resolution analysis of topological derivative based localization of small electromagnetic inclusions. SIAM Journal on Imaging Sciences, 8(3):1687–1717, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Wang:2016:NSD

[Wan16a] Yi-Qing Wang. A note on the size of denoising neural networks. SIAM Journal on Imaging Sciences, 9(1):275–286, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Wang:2016:NAR

[Wan16b] Zhenzhou Wang. A new approach for robust segmentation of the noisy or textured images. SIAM Journal on Imaging Sciences, 9(3):1409–1436, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Wijesinghe:2023:MBB

[WC23] Janith Wijesinghe and Pengwen Chen. Matrix balancing based interior point methods for point set matching problems. SIAM Journal on Imaging Sciences, 16(3):1068– 1105, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/22M1479476.

Wendt:2018:MAM

[WCA^{+}18] Herwig Wendt, Sébastien Combrexelle, Yoann Altmann, Jean-Yves Tourneret, Stephen McLaughlin, and Patrice Abry. Multifractal analysis of multivariate images using gamma Markov random field priors. SIAM Journal on Imaging Sciences, 11(2):1294–1316, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Wang:2019:NOD

[WCN^{+}19] Chao Wang, Raymond Chan, Mila Nikolova, Robert Plemmons, and Sudhakar Prasad. Nonconvex optimization for 3dimensional point source localization using a rotating point spread function. SIAM Journal on Imaging Sciences, 12(1): 259–286, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Ward:2013:DPR

[WCU13] John Paul Ward, Kunal Narayan Chaudhury, and Michael Unser. Decay properties of Riesz transforms and steerable wavelets. SIAM Journal on Imaging Sciences, 6(2):984–998, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Wu:2009:SSA

[WDCT09] Chunlin Wu, Jiansong Deng, Falai Chen, and Xuecheng Tai. Scale-space analysis of discrete filtering over arbitrary triangulated surfaces. SIAM Journal on Imaging Sciences, 2(2): 670–709, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Weinmann:2014:TVR

[WDS14] Andreas Weinmann, Laurent Demaret, and Martin Storath.

Total variation regularization for manifold-valued data. SIAM Journal on Imaging Sciences, 7 (4):2226–2257, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Wubbeler:2017:LSO

[WE17] Gerd Wübbeler and Clemens Elster. A large-scale optimization method using a sparse approximation of the Hessian for magnetic resonance fingerprinting. SIAM Journal on Imaging Sciences, 10(3):979–1004, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Weiss:2011:IIL

[WFBFA11] Pierre Weiss, Alexandre Fournier, Laure Blanc-Féraud, and Gilles Aubert. On the illumination invariance of the level lines under directed light: Application to change detection. SIAM Journal on Imaging Sciences, 4 (1):448–471, ???? 2011. CO-DEN SJISBI. ISSN 1936-

                                        4954. URL http://epubs.
                                        siam.org/siims/resource/1/
                                        sjisbi/v4/i1/p448_s1.

Woien:2022:PBM

[WG22] Esten Nicolai Wøien and Markus Grasmair. A PDEbased method for shape registration. SIAM Journal on Imaging Sciences, 15(2):762– 796, ???? 2022. CODEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1408932.

Wu:2022:GNL

[WGGX22] Chunlin Wu, Xueyan Guo, Yiming Gao, and Yunhua Xue. A general non-Lipschitz infimal convolution regularized model: Lower bound theory and algorithm. SIAM Journal on Imaging Sciences, 15(3): 1499–1538, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/20M1356634.

Wu:2022:EBD

[WGL^{+}22] Tingting Wu, Xiaoyu Gu, Zeyu Li, Zhi Li, Jianwei Niu, and Tieyong Zeng. Efficient boosted DC algorithm for nonconvex image restoration with Rician noise. SIAM Journal on Imaging Sciences, 15(2): 424–454, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1421660.

Wang:2015:VMB

[WH15] Wei Wang and Chuanjiang He.

A variational model with barrier functionals for Retinex. SIAM Journal on Imaging Sciences, 8(3):1955–1980, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Wang:2015:SRS

[WHY^{+}15] Liming Wang, Jiaji Huang, Xin Yuan, Kalyani Krishnamurthy, Joel Greenberg, Volkan Cevher, Miguel R. D. Rodrigues, David Brady, Robert Calderbank, and Lawrence Carin. Signal recovery and system calibration from multiple compressive Poisson measurements. SIAM Journal on Imaging Sciences, 8(3): 1923–1954, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Wong:2014:CQC

[WkZ14] Tsz Wai Wong and Hong kai Zhao. Computation of quasiconformal surface maps using discrete Beltrami flow. SIAM Journal on Imaging Sciences, 7 (4):2675–2699, ???? 2014. CO-DEN SJISBI. ISSN 1936-4954.

Wang:2022:NRS

[WLJ22] Weixi Wang, Ji Li, and Hui Ji. L_{1}-norm regularization for short-and-sparse blind deconvolution: Point source separability and region selection. SIAM Journal on Imaging Sciences, 15(3):1345– 1372, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M144904X.

Wangni:2021:TSP

[WLL^{+}21] Jianqiao Wangni, Dahua Lin, Ji Liu, Kostas Daniilidis, and Jianbo Shi. Towards statistically provable geometric 3D human pose recovery. SIAM Journal on Imaging Sciences, 14(1): 246–270, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Wong:2012:IFE

[WLTC12] Tsz Wai Wong, Lok Ming Lui, Paul M. Thompson, and Tony F. Chan. Intrinsic feature extraction on hippocampal surfaces and its applications. SIAM Journal on Imaging Sciences, 5 (2):746–768, ???? 2012. CO-DEN SJISBI. ISSN 1936-4954.

Ward:2015:ITU

[WLYU15] John Paul Ward, Minji Lee, Jong Chul Ye, and Michael Unser. Interior tomography using 1D generalized total variation. Part I: Mathematical foundation. SIAM Journal on Imaging Sciences, 8(1):226–247, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Wang:2013:SGG

[WM13] Yi-Qing Wang and Jean-Michel Morel. SURE guided Gaussian mixture image denoising. SIAM Journal on Imaging Sciences, 6 (2):999–1034, ???? 2013. CO-DEN SJISBI. ISSN 1936-4954.

Wright:2009:CMI

[WMT^{+}09] John Wright, Yi Ma, Yangyu Tao, Zhouchen Lin, and Heung-Yeung Shum. Classification via minimum incremental coding length. SIAM Journal on Imaging Sciences, 2(2):367–395, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Wang:2013:VAI

[WN13a] Wei Wang and Michael K. Ng.

A variational approach for image stitching I. SIAM Journal on Imaging Sciences, 6(3): 1318–1344, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Wang:2013:VHE

[WN13b] Wei Wang and Michael K. Ng.

A variational histogram equalization method for image contrast enhancement. SIAM Journal on Imaging Sciences, 6(3): 1823–1849, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Wang:2021:CIR

[WN21] Wei Wang and Michael K. Ng. Color image restoration by saturation-value total variation regularization on vector bundles. SIAM Journal on Imaging Sciences, 14(1):178–197, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Wang:2022:SRM

[WNS^{+}22] Bao Wang, Tan Nguyen, Tao Sun, Andrea L. Bertozzi, Richard G. Baraniuk, and Stanley J. Osher. Scheduled restart momentum for accelerated stochastic gradient descent. SIAM Journal on Imaging Sciences, 15(2): 738–761, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1453311.

Wolansky:2009:IQI

[Wol09] Gershon Wolansky. Incompressible, quasi-isometric deformations of 2-dimensional domains. SIAM Journal on Imaging Sciences, 2(4):1031–1048, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Webber:2020:MAC

[WQ20] James W. Webber and Eric Todd Quinto. Microlocal analysis of a Compton tomography problem. SIAM Journal on Imaging Sciences, 13(2):746–774, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Webber:2021:MAG

[WQ21] James W. Webber and Eric Todd Quinto. Microlocal analysis of generalized Radon transforms from scattering tomography. SIAM Journal on Imaging Sciences, 14(3):976–1003, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Winkler:2014:RCC

[WR14] Robert Winkler and Andreas Rieder. Resolution-controlled conductivity discretization in electrical impedance tomography. SIAM Journal on Imaging Sciences, 7(4):2048–2077, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Wang:2013:RLL

[WSL13] Yi Wang, Arthur Szlam, and Gilad Lerman. Robust locally linear analysis with applications to image denoising and blind inpainting. SIAM Journal on Imaging Sciences, 6(1):526–562, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Wang:2013:ODC

[WSW13] Lanhui Wang, Amit Singer, and Zaiwen Wen. Orientation determination of cryo-EM images using least unsquared deviations. SIAM Journal on Imaging Sciences, 6(4):2450–2483, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Wu:2010:ALM

[WT10] Chunlin Wu and Xue-Cheng Tai. Augmented Lagrangian method, dual methods, and split Bregman iteration for ROF, vectorial TV, and high order models. SIAM Journal on Imaging Sciences, 3(3):300–339, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Wapenaar:2013:RDS

[WT13] Kees Wapenaar and Jan Thorbecke. On the retrieval of the directional scattering matrix from directional noise. SIAM Journal on Imaging Sciences, 6(1): 322–340, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Wang:2021:LAC

[WTNL21] Chao Wang, Min Tao, James G. Nagy, and Yifei Lou. Limitedangle CT reconstruction via the L_{1}/L_{2} minimization. SIAM Journal on Imaging Sciences, 14(2):749–777, ???? 2021. CO-DEN SJISBI. ISSN 1936-4954.

Wang:2010:SSR

[WY10] Yilun Wang and Wotao Yin. Sparse signal reconstruction via iterative support detection. SIAM Journal on Imaging Sciences, 3(3):462–491, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Wang:2012:PIM

[WY12] Ling Wang and Birsen Yazici. Passive imaging of moving targets using sparse distributed apertures. SIAM Journal on Imaging Sciences, 5(3):769–808, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Wang:2014:BSA

[WY14] Ling Wang and Birsen Yazici.

Bistatic synthetic aperture radar imaging of moving targets using ultra-narrowband continuous waveforms. SIAM Journal on Imaging Sciences, 7(2): 824–866, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Wang:2017:LAS

[WY17] Ling Wang and Birsen Yazici.

Layover analysis in synthetic aperture imagery. SIAM Journal on Imaging Sciences, 10(3): 1033–1068, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Wang:2022:SCC

[WYN22] Wei Wang, Yuming Yang, and Michael K. Ng. A spatial color compensation model using saturation-value total variation. SIAM Journal on Imaging Sciences, 15(3):1400– 1430, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1453773.

Wang:2008:NAM

[WYYZ08] Yilun Wang, Junfeng Yang, Wotao Yin, and Yin Zhang. A new alternating minimization algorithm for total variation image reconstruction. SIAM Journal on Imaging Sciences, 1(3): 248–272, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Wang:2017:MDR

[WZ17] Qian Wang and Yining Zhu. Multi-domain regularization based computed tomography for high-speed rotation objects. SIAM Journal on Imaging Sciences, 10(2):602–640, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Wang:2020:VIS

[WZLH20] Faqiang Wang, Cuicui Zhao, Jun Liu, and Haiyang Huang. A variational image segmentation model based on normalized cut with adaptive similarity and spatial regularization. SIAM Journal on Imaging Sciences, 13(2):651–684, ???? 2020. CO-DEN SJISBI. ISSN 1936-4954.

Wang:2013:FBS

[WZYX13] Zelong Wang, Jubo Zhu, Fengxia Yan, and Meihua Xie. Fidelity-Beltrami-sparsity model for inverse problems in multichannel image processing. SIAM Journal on Imaging Sciences, 6(4):2685–2713, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Xia:2014:SMS

[XFPA14] Gui-Song Xia, Sira Ferradans, Gabriel Peyré, and Jean-François Aujol. Synthesizing and mixing stationary Gaussian texture models. SIAM Journal on Imaging Sciences, 7(1): 476–508, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Xia:2022:FAK

[XWH22] Yuxin Xia, Wei Wang, and Bo Han. A fast averaged Kaczmarz iteration with convex penalty for inverse problems in Hilbert spaces. SIAM Journal on Imaging Sciences, 15(3): 1079–1103, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1445181.

Xu:2020:CTI

[XXQJ20] Ruotao Xu, Yong Xu, Yuhui Quan, and Hui Ji. Cartoontexture image decomposition using orientation characteristics in patch recurrence. SIAM Journal on Imaging Sciences, 13(3):1179–1210, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Xu:2022:DSI

[XXYC22] Yangyang Xu, Yibo Xu, Yonggui Yan, and Jie Chen. Distributed stochastic inertialaccelerated methods with delayed derivatives for nonconvex problems. SIAM Journal on Imaging Sciences, 15 (2):550–590, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1435719.

Xu:2013:BCD

[XY13] Yangyang Xu and Wotao Yin.

A block coordinate descent method for regularized multiconvex optimization with applications to nonnegative tensor factorization and completion. SIAM Journal on Imaging Sciences, 6(3):1758–1789, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Xing:2012:DLN

[XZC^{+}12] Zhengming Xing, Mingyuan Zhou, Alexey Castrodad, Guillermo Sapiro, and Lawrence Carin. Dictionary learning for noisy and incomplete hyperspectral images. SIAM Journal on Imaging Sciences, 5(1):33–56, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Xu:2019:UDI

[XZZ19] Xiaoxu Xu, Bo Zhang, and Haiwen Zhang. Uniqueness and direct imaging method for inverse scattering by locally rough surfaces with phaseless near-field data. SIAM Journal on Imaging Sciences, 12(1):119–152, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Yan:2013:RIC

[Yan13] Ming Yan. Restoration of images corrupted by impulse noise and mixed Gaussian impulse noise using blind inpainting. SIAM Journal on Imaging Sciences, 6(3):1227–1245, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Yan:2021:CSA

[YBZ^{+}21] Xuefei Yan, David J. Brady, Weiping Zhang, Changzhi Yu, Yulin Jiang, Jianqiang Wang, Chao Huang, Zian Li, and Zhan Ma. Compressive sampling for array cameras. SIAM Journal on Imaging Sciences, 14(1): 156–177, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Yin:2016:RCA

[YCF^{+}16] Rujie Yin, Bruno Cornelis, Gabor Fodor, Noelle Ocon, David Dunson, and Ingrid Daubechies. Removing cradle artifacts in Xray images of paintings. SIAM Journal on Imaging Sciences, 9 (3):1247–1272, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Yeung:2019:NIT

[YCU19] T. S. Au Yeung, Eric T.

Chung, and Gunther Uhlmann. Numerical inversion of threedimensional geodesic X-ray transform arising from travel time tomography. SIAM Journal on Imaging Sciences, 12(3): 1296–1323, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Yin:2017:TTB

[YGLD17] Rujie Yin, Tingran Gao, Yue M.

Lu, and Ingrid Daubechies. A tale of two bases: Localnonlocal regularization on image patches with convolution framelets. SIAM Journal on Imaging Sciences, 10(2):711– 750, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Yao:2019:MNR

[YGS^{+}19] Wenjuan Yao, Zhichang Guo, Jiebao Sun, Boying Wu, and Huijun Gao. Multiplicative noise removal for texture images based on adaptive anisotropic fractional diffusion equations.

SIAM Journal on Imaging Sciences, 12(2):839–873, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Ye:2018:DCF

[YHC18] Jong Chul Ye, Yoseob Han, and Eunju Cha. Deep convolutional framelets: a general deep learning framework for inverse problems. SIAM Journal on Imaging Sciences, 11(2):991–1048, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Yin:2010:AGL

[Yin10] Wotao Yin. Analysis and generalizations of the linearized Bregman method. SIAM Journal on Imaging Sciences, 3(4):856–877, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Yoo:2017:JSR

[YJL^{+}17] Jaejun Yoo, Younghoon Jung, Mikyoung Lim, Jong Chul Ye, and Abdul Wahab. A joint sparse recovery framework for accurate reconstruction of inclusions in elastic media. SIAM Journal on Imaging Sciences, 10(3):1104–1138, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Yashtini:2016:FRN

[YK16] Maryam Yashtini and Sung Ha Kang. A fast relaxed normal two split method and an effective weighted TV approach for Euler’s elastica image inpainting. SIAM Journal on Imaging Sciences, 9(4):1552–1581, ????

2016. CODEN SJISBI. ISSN 1936-4954.

Yang:2021:DGM

[YL21] Xu Yang and Zhi-Yong Liu.

A doubly graduated method for inference in Markov random field. SIAM Journal on Imaging Sciences, 14(3):1354–1373, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Yu:2023:TSD

[YLH23] Jing Yu, Fang Li, and Xuyue Hu. Two-stage decolorization based on histogram equalization and local variance maximization. SIAM Journal on Imaging Sciences, 16(2): 740–769, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1509333.

Yang:2020:MCI

[YLLX20] Ming Yang, Qilun Luo, Wen Li, and Mingqing Xiao. Multiview clustering of images with tensor rank minimization via nonconvex approach. SIAM Journal on Imaging Sciences, 13(4): 2361–2392, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Yueh:2019:NAV

[YLLY19] Mei-Heng Yueh, Tiexiang Li, Wen-Wei Lin, and Shing-Tung Yau. A novel algorithm for volume-preserving parameterizations of 3-manifolds. SIAM Journal on Imaging Sciences, 12(2):1071–1098, ???? 2019. CODEN SJISBI. ISSN 1936-

Yueh:2020:NEA

[YLLY20] Mei-Heng Yueh, Tiexiang Li, Wen-Wei Lin, and Shing-Tung Yau. A new efficient algorithm for volume-preserving parameterizations of genus-one 3manifolds. SIAM Journal on Imaging Sciences, 13(3):1536– 1564, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Yao:2022:PBI

[YMA22] Dan Yao, Stephen McLaughlin, and Yoann Altmann. Patchbased image restoration using expectation propagation. SIAM Journal on Imaging Sciences, 15(1):192–227, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1427541.

Yin:2008:BIA

[YOGD08] Wotao Yin, Stanley Osher, Donald Goldfarb, and Jerome Darbon. Bregman iterative algorithms for ℓ_{1}-minimization with applications to compressed sensing. SIAM Journal on Imaging Sciences, 1(1):143–168, ???? 2008. CODEN SJISBI. ISSN 1936-4954.

Yang:2017:ADM

[YPC17] Lei Yang, Ting Kei Pong, and Xiaojun Chen. Alternating direction method of multipliers for a class of nonconvex and nonsmooth problems with applications to background/foreground extraction. SIAM Journal on Imaging Sciences, 10(1):74–110, ???? 2017.

CODEN SJISBI. ISSN 1936-4954.

Yezzi:2020:AOP

[YSB20] Anthony Yezzi, Ganesh Sundaramoorthi, and Minas Benyamin. Accelerated optimization in the PDE framework formulations for the active contour case. SIAM Journal on Imaging Sciences, 13(4):2029–2062, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Yueh:2023:TFS

[Yue23] Mei-Heng Yueh. Theoretical foundation of the stretch energy minimization for areapreserving simplicial mappings. SIAM Journal on Imaging Sciences, 16(3):1142– 1176, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1505062.

Yang:2023:UIR

[YWW^{+}23] Yan Yang, Yizhou Wang, Jiazhen Wang, Jian Sun, and Zongben Xu. An unrolled implicit regularization network for joint image and sensitivity estimation in parallel MR imaging with convergence guarantee. SIAM Journal on Imaging Sciences, 16(3):1791– 1824, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/22M1502094.

Yang:2013:SWP

[YY13] Haizhao Yang and Lexing Ying.

Synchrosqueezed wave packet transform for 2D mode decomposition. SIAM Journal on Imaging Sciences, 6(4):1979– 2009, ???? 2013. CODEN SJISBI. ISSN 1936-4954.

Yanik:2015:SAI

[YY15] H. Çagri Yanik and Birsen Yaz. Synthetic aperture inversion for statistically nonstationary target and clutter scenes. SIAM Journal on Imaging Sciences, 8 (3):1658–1686, ???? 2015. CO-DEN SJISBI. ISSN 1936-4954.

Yun:2017:AKS

[YY17] Joo Dong Yun and Seungjoon Yang. ADMM in Krylov subspace and its application to total variation restoration of spatially variant blur. SIAM Journal on Imaging Sciences, 10(2): 484–507, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Yonel:2019:GWF

[YY19] Bariscan Yonel and Birsen Yazici. A generalization of Wirtinger flow for exact interferometric inversion. SIAM Journal on Imaging Sciences, 12(4):2119–2164, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Yonel:2022:SEF

[YY22] Bariscan Yonel and Birsen Yazici. A spectral estimation framework for phase retrieval via Bregman divergence minimization. SIAM Journal on Imaging Sciences, 15 (2):491–520, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/20M1388061.

Yang:2009:FAE

[YYZW09] Junfeng Yang, Wotao Yin, Yin Zhang, and Yilun Wang. A fast algorithm for edge-preserving variational multichannel image restoration. SIAM Journal on Imaging Sciences, 2(2):569–592, ???? 2009. CODEN SJISBI. ISSN 1936-4954.

Yin:2018:BRA

[YZL^{+}18] Penghang Yin, Shuai Zhang, Jiancheng Lyu, Stanley Osher, Yingyong Qi, and Jack Xin. BinaryRelax: a relaxation approach for training deep neural networks with quantized weights. SIAM Journal on Imaging Sciences, 11(4):2205– 2223, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Zhang:2010:BNR

[ZBBO10] Xiaoqun Zhang, Martin Burger, Xavier Bresson, and Stanley Osher. Bregmanized nonlocal regularization for deconvolution and sparse reconstruction. SIAM Journal on Imaging Sciences, 3(3):253–276, ???? 2010. CODEN SJISBI. ISSN 1936-4954.

Zhang:2017:NTB

[ZBN17] Xiongjun Zhang, Minru Bai, and Michael K. Ng. Nonconvex-TV based image restoration with impulse noise removal. SIAM Journal on Imaging Sciences, 10(3):1627–1667, ????

2017. CODEN SJISBI. ISSN 1936-4954.

Zacur:2014:LIR

[ZBO14] Ernesto Zacur, Matias Bossa, and Salvador Olmos. Leftinvariant Riemannian geodesics on spatial transformation groups. SIAM Journal on Imaging Sciences, 7(3):1503–1557, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Zhao:2022:RTC

[ZBSZ22] Xueying Zhao, Minru Bai, Defeng Sun, and Libin Zheng. Robust tensor completion: Equivalent surrogates, error bounds, and algorithms. SIAM Journal on Imaging Sciences, 15 (2):625–669, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1429539.

Zhu:2012:IDU

[ZC12] Wei Zhu and Tony Chan. Image denoising using mean curvature of image surface. SIAM Journal on Imaging Sciences, 5 (1):1–32, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Zhang:2015:TFO

[ZC15] Jianping Zhang and Ke Chen. A total fractional-order variation model for image restoration with nonhomogeneous boundary conditions and its numerical solution. SIAM Journal on Imaging Sciences, 8(4):2487– 2518, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Zhang:2020:OPV

[ZC20] Daoping Zhang and Ke Chen.

3D orientation-preserving variational models for accurate image registration. SIAM Journal on Imaging Sciences, 13(3): 1653–1691, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Zhu:2022:PGQ

[ZCL22] Zhipeng Zhu, Gary P. T. Choi, and Lok Ming Lui. Parallelizable global quasi-conformal parameterization of multiply connected surfaces via partial welding. SIAM Journal on Imaging Sciences, 15(4): 1765–1807, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1466323.

Zhang:2018:CCM

[ZCO18] Kewei Zhang, Elaine Crooks, and Antonio Orlando. Compensated convexity methods for approximations and interpolations of sampled functions in Euclidean spaces: Applications to contour lines, sparse data, and inpainting. SIAM Journal on Imaging Sciences, 11(4): 2368–2428, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Zhang:2022:UFD

[ZCZL22] Daoping Zhang, Gary P. T.

Choi, Jianping Zhang, and Lok Ming Lui. A unifying framework for n-dimensional quasi-conformal mappings. SIAM Journal on Imaging Sciences, 15(2):960–988, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1457497.

Zhan:2016:CIR

[ZD16] Ruohan Zhan and Bin Dong. CT image reconstruction by spatial-Radon domain datadriven tight frame regularization. SIAM Journal on Imaging Sciences, 9(3):1063–1083, ???? 2016. CODEN SJISBI. ISSN 1936-4954.

Zhang:2018:RJS

[ZDL18] Haimiao Zhang, Bin Dong, and Baodong Liu. A reweighted joint spatial-Radon domain CT image reconstruction model for metal artifact reduction. SIAM Journal on Imaging Sciences, 11(1):707–733, ???? 2018. CO-DEN SJISBI. ISSN 1936-4954.

Zhang:2023:CFM

[ZE23] Kai Zhang and Alireza Entezari. Convolutional forward models for X-ray computed tomography. SIAM Journal on Imaging Sciences, 16(4): 1953–1977, ???? 2023. CO-DEN SJISBI. ISSN 1936-4954.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1464191.

Zangerl:2021:MFW

[ZH21] Gerhard Zangerl and Markus Haltmeier. Multiscale factorization of the wave equation with application to compressed sensing photoacoustic tomography. SIAM Journal on Imaging Sciences, 14(2):558–579, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Zhuang:2016:DAS

[Zhu16] Xiaosheng Zhuang. Digital affine shear transforms: Fast realization and applications in image/video processing. SIAM Journal on Imaging Sciences, 9 (3):1437–1466, ???? 2016. CO-DEN SJISBI. ISSN 1936-4954.

Zhang:2022:NMD

[ZHW22] Huayan Zhang, Zhishuai He, and Xiaochao Wang. A novel mesh denoising method based on relaxed second-order total generalized variation. SIAM Journal on Imaging Sciences, 15(1):1–22, ???? 2022. CO-DEN SJISBI. ISSN 1936-4954.

URL https://epubs.siam.
org/doi/10.1137/21M1397945.

Zou:2021:RSF

[ZJ21] Qing Zou and Mathews Jacob.

Recovery of surfaces and functions in high dimensions: Sampling theory and links to neural networks. SIAM Journal on Imaging Sciences, 14(2):580– 619, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Zhang:2021:DIR

[ZL21] Jianping Zhang and Yanyan Li.

Diffeomorphic image registration with an optimal control relaxation and its implementation. SIAM Journal on Imaging Sciences, 14(4):1890–1931, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Zhu:2018:CCS

[ZLD^{+}18] Zhihui Zhu, Gang Li, Jiajun Ding, Qiuwei Li, and Xiongxiong He. On collaborative compressive sensing systems: The framework, design, and algorithm. SIAM Journal on Imaging Sciences, 11(2):1717–1758, ???? 2018. CODEN SJISBI. ISSN 1936-4954.

Zhang:2019:CTN

[ZN19] Xiongjun Zhang and Michael K. Ng. A corrected tensor nuclear norm minimization method for noisy low-rank tensor completion. SIAM Journal on Imaging Sciences, 12(2):1231–1273, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Zosso:2015:NLR

[ZTO15] Dominique Zosso, Giang Tran, and Stanley J. Osher. Non-local Retinex — a unifying framework and beyond. SIAM Journal on Imaging Sciences, 8(2): 787–826, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Zeune:2017:MSB

[ZvDT^{+}17] Leonie Zeune, Guus van Dalum, Leon W. M. M. Terstappen, Stephan A. van Gils, and Christoph Brune. Multiscale segmentation via Bregman distances and nonlinear spectral analysis. SIAM Journal on Imaging Sciences, 10(1):111– 146, ???? 2017. CODEN SJISBI. ISSN 1936-4954.

Zeng:2019:NLM

[ZWJ19] Chao Zeng, Chunlin Wu, and Rui Jia. Non-Lipschitz models for image restoration with impulse noise removal. SIAM Journal on Imaging Sciences, 12(1):420–458, ???? 2019. CO-DEN SJISBI. ISSN 1936-4954.

Zhao:2014:NCO

[ZWN14] Xi-Le Zhao, Fan Wang, and Michael K. Ng. A new convex optimization model for multiplicative noise and blur removal. SIAM Journal on Imaging Sciences, 7(1):456–475, ???? 2014. CODEN SJISBI. ISSN 1936-4954.

Zhou:2020:BIU

[ZYZL20] Qingping Zhou, Tengchao Yu, Xiaoqun Zhang, and Jinglai Li. Bayesian inference and uncertainty quantification for medical image reconstruction with Poisson data. SIAM Journal on Imaging Sciences, 13(1):29–52, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Zhao:2019:HAP

[ZZ19] Hongkai Zhao and Yimin Zhong. A hybrid adaptive phase space method for reflection traveltime tomography. SIAM Journal on Imaging Sciences, 12(1):28–53, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Zhu:2021:SVR

[ZZ21] Ya-Nan Zhu and Xiaoqun Zhang. A stochastic variance reduced primal dual fixed point method for linearly constrained separable optimization. SIAM Journal on Imaging Sciences, 14(3):1326–1353, ???? 2021. CODEN SJISBI. ISSN 1936-4954.

Zisler:2020:SAF

[ZZPS20] Matthias Zisler, Artjom Zern, Stefania Petra, and Christoph Schnörr. Self-assignment flows for unsupervised data labeling on graphs. SIAM Journal on Imaging Sciences, 13(3):1113– 1156, ???? 2020. CODEN SJISBI. ISSN 1936-4954.